A Complete Bibliography of Publications in Scientometrics: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

04 March 2021 Version 1.06

Title word cross-reference

25 [DR08]. 3 [Egg04a, Egg05b]. ^{th} [DL06, GM09a]. g [Bur09, CB08a, Egg06, Sch08]. H [Egg10, AHK08, AJR09, BI08a, BI08b, BD05b, BM09, Bur09, CB08a, DMPM^{+}08, Egg08b, Egg09a, GP09, Glä06, Glä08b, IP07, IRN07, Lia06, LRR09, Mar09, MHM07, MM08b, RT08, Saa06, Sch07a, Sch09, SKM07, Tol09, Ye09, ZK08]. h(2) [Bur09]. J [BMT09]. N [Egg00a, Bur02b]. n^{th} [Egg07]. Q [CR08, RZ08].

-dimensional [Egg04a, Egg05b]. -factor [BMT09]. -Grams [Egg00a]. -index [AHK08, BI08a, BI08b, BD05b, BM09, CB08a, DMPM^{+}08, Egg06, Egg09a, GP09, Glä06, Glä08b, IP07, IRN07, Lia06, Mar09, MM08b, Saa06, Sch09, SKM07, Tol09, Ye09, ZK08]. -indices [AJR09, Egg08b, LRR09, RT08, Sch07a]. -measures [CR08, RZ08]. -sequence [Egg10]. -years [DR08].

1 [GLT06, TMVAR00]. 11 [GM09a]. 1980s [sKtP04]. 1990 [Ano01b, Ano01c]. 1990s [MVH^{+}05]. 1999 [Wor00c]. 19th [HHWD04].

2 [LGT06]. 2003 [Ano03d, Ley04a, Rou04]. 2005 [Edi05, McC05, Rou05b]. 2007 [Ano08b, Whi08]. 2009 [Ano09]. 21st [MMR05].

4th [Wor00c].

7th [Hin03, Moe00b].

= [GS01].

A. [KKK02]. Abels [BSPKN03a]. absolute [Per08]. absorptive [ACdLPM09]. abstract [AJ08a]. Abstracts [HPF03, ND09, Bla09, MDMBA^{+}06, Pin06]. abuse [CMJSOC09]. Academia [Gu04b]. Academic [LCL05, ADC09b, BI04, Bjö06, BCPTC06, BCGMBSVM07, CGR06, Cho08, Dai07, Flo07, Gu03, Har08, HLT09, IGK07, Jud04, KO08, LZ03, LF03, LCWD09, MHGGVJ08, MSU03, MDTU05, MT07, NSS04, OACS08, PT07, Per08, PD07, RDvT03, SRS^{+}02, TT04, TT03, TTP03, TV06, WWW07, Wu09]. Academics [SHO00, MM05]. accepting [BND08]. Access [Cho08, SH08b, Har09a, KT06, Muk09, ST02a, ST04, SH09]. accessibility [Wu09]. accessible [JLH07]. accomplishments [HSSK07]. according [KK04, Pis09, RRZ06, YGY06]. accounting [vCvCvU08]. accounts [Cam09]. accumulated [Ada05]. achieved [Rig09]. achievements [Cho08]. acquisitions [KM04b]. across [Das04, LT08]. Act [Bal06]. active [Jia08]. activities [Chu01a, Chu01b, DP03, YM08, ZRRBL03]. Activity [DDM00, KGS00, Kum06, LM00, BMJARBC05a, Bal06, Bha04, DD09, Gei05, Glä06, GJT09, Hor01, JCTSM^{+}09, McM09, PMO04, RRGGMS07]. actor [MA07]. actor-actor [MA07]. actor-objectives [MA07]. actors [Jia08]. Acute [CHH04]. addition [Wan07]. additional [GSST06]. additionality [Hay03]. addresses [GMA06]. adequacy [LTL^{+}06]. adjustment [DL06]. administration [KS01]. adopt [YDM^{+}05]. Adrian [ZM01]. adsorption [YS04]. Advanced [BDS06]. advantage [Cri06]. Advantages [BFG02, CJ06]. advent [DL06]. aerosol [XZH08]. aesthetic [AMTE07]. affect [KO08, Lee09, RJDD07]. affecting [LZ02]. affects [KKM07, Lee09]. affiliation [LCWD09]. afford [Wra09a]. Africa [AWG00, Bos09, MCMN02, MP06, MMP09, NBRA^{+}02, OO09, Pou07b, Pou09, PP09, Soo09a, Tij07]. African [IJ04, JI00, Pou05, Soo09b]. after [CM09]. AG [DL06]. Age [BD03, LKGdB01, Sid06, WFI03, AY09, BO05, BS03, Bur02a, LGD02, OCA09]. Ageing [Pol00, vR00]. agenda [Dai09]. agendas [RJDD07]. agents [WD05]. agglomeration [BD05a]. Aggregation [RRZR09]. aging [KO08]. agreement [HH90]. agreements [RJDD07]. Agricultural [TMVAR00, Far02, GKL06]. AICTE [GGJ^{+}05]. AIDS [MCMN02, OO09]. al [Xek05, vR05c]. algorithm [SM06]. algorithmic [WWJLG06]. alignment [GSH01]. all-author [SLI09]. also [ZTG09a]. alternative [DD09, Joh09, NLM01]. alternatives [Tol09]. ambiguous [BSPKN03b]. America [GK09, MC02]. American [Yan09, AC04, Cal06, CRLMRPA08, Dan00, DT07, Jon09, Koe01a, KS01, She08, Tri00]. amnesia [SW01]. among [BPMK06, Dan00, HH90, HR09, Kyv03, PL09, Son03, TNK^{+}06]. amongst [Van01b]. analogy [PBZ02]. analyses [CvLT07, Glä04, Jia08, KBWM07, Kra08, PD07, SLI09, vL06]. Analysis [BMJARBC05a, BGG00, BI00, BPA00, BCZM^{+}09, DCF00, DDM00, Fuj02, GAGPM06, HW02, Ioa06, Kum09, LOK00, LL08, LS05, MS01, MSS06, MH01, NRTD00, RWD^{+}09, RG06, RSJ00, SA00, SHO00, Sig00, SGT^{+}06b, TMVAR00, Ver00, Wor00a, ADC09b, Ack06, AKOK09, AJ08b, AU07, AUQ09, BT06, BI01a, BI04, BIE05, BLZ07, Bay05, Bha04, Bla09, BD06, BND08, BV06, CBVN06, CVV^{+}06, CCHF05, CCHC06, CSW^{+}09, CHH04, CH05, CH07, CHH07, CS02, Dal02, Dan06, DR08, DA08, EP03, Fuj06, GFNG08, GSST06, GJT09, GGP05, Gmü03, Gu04a, GH07, GD03, GD08, GD09b, HH90, HG02, Hay03, He03, HZT05, HMG09, HT09, HCLH04, HYA08, HWD01a, HWD01b, IK02, Jar08, JG02, JM07, KKL06, KU05, Kim01, Kim07, KL09, Koe01a, Kos02, Kre01b, Kre01c, LFAH04]. analysis [LGA06, LLSL07, LJ08, Lee09, LJL08, LK08, LH06, LDF^{+}09, LCLA02, Liu05, cL07, Lo08, LCWD09, LINV^{+}09, MDNL09, MCMN02, MC02, MM02, MSW01, Meh07, MMP06, Mey01, MMAHS08, MEG06, ND02, NLD07, ND09, Nwa06, Old01a, Old01b, ÖSAT08, OO09, OLBB^{+}08, PL09, PSD^{+}07, PA08, PID06, QCW04, QLLR08, QC09, RRS07, RRZ06, RWGA07, Sap07, SM06, SP08, ST08, Som05, TT04, TBK06, TZ08, TG08, TG09, yTXwW03a, yTXwW03b, yTlC05, Tsa08, Tsa09a, Tsa09b, UKKP04, VVPCOL09, Van08a, VGK06, VY08, VDL03, WWW07, WWJLG06, XZH08, YO06, YZ08, YS08, YY08a, YK04, YNT06, YNSS09, YW04, ZQH09, ZL02, ZCY09, ZGHW07, ZTG09b, dMACRVQ^{+}07, vL09, JLC09]. analysis- [BT06]. analysis.A [MB01]. analytical [Ley04b, Per06, Tij07]. analyze [CC09, Gok03]. Analyzing [HH06]. Announcement [Ano04a, Ano08a]. annual [LT08, Bro04]. ant [Wet06]. Antarctic [DR08]. anthrax [BIE05]. anti [CMM04]. anti-pollution [CMM04]. Anticipating [KM07]. Antitrust [MMS05]. applicant [BMD09]. applicants [BD06]. Application [AGM^{+}01, Eto00b, LFAH04, MEG06, She08, YHM05, BDS06, JCTSM^{+}09, RdL02, RT08, SJJ04, ST08, SU09, Tol09, WHC07, YNT06, ZRRB03a, vLM02, vL06]. Applications [ER01, Wol06, Glä08b, Iga08, JSJX02, Lia06, LpCY06, Rei09, Ste09b]. Applied [AET09, BMD07, RDvT03, SLI09]. Applying [NCM09, GMA06, Vin02b]. Approach [Glä00a, Glä00b, LM00, Nal01a, SM00, AJR09, Bla09, BM09, CL06, DG04, ER04, Fra02, GG09a, GdL02, GSB02, Glä06, HLSA07, JMB06, Joh09, KP08, KH08, LAFP04, Lee09, MH07, MMAHS08, ML05, MBG01, PFL01, Pri03, SGT06a, Sch07b, TMKM09, TK09, TV06, WBB04, Zsi08, dMACRVQ^{+}07].

Approaches [Sæt00, AJ08a, Ley04b, SLI09]. appropriate [TVvL02]. approval [BD06]. Approved [AGG00a]. approximation [YM08]. archived [VKVF09]. Archives [Fuj06]. Archiving [KL09]. Area [DCF00, GK00, OACS08, PBG^{+}05, BT06, MB06, SU09, VVPCOL09, WL03, ZL02, vCvCvU08, PSD^{+}07]. Areas [BR00, BPA00, IJ04, IMLSSC09, Sma06, WHC07, YDM^{+}05, YNT06]. arena [JRS06]. arising [Bur07a]. Art [CVV^{+}06, Lar08b]. Arteriosus} [HCLH04]. Article [MBR^{+}08, AJOP04, HBK^{+}08, VH01, Yit02, LFB^{+}06]. Article-count [MBR^{+}08]. Articles [Cah00b, Ing00, SLB00, ADC08, Bal09, BCZM^{+}09, Cho08, DF07, GDD03, Kos07, KT06, LPC05, LT08, LT09, MBC06, NC09, PRdKB07, RvHMJ05, Sol07, SKMI06, Wal06]. artificial [CCDJJF04, PFL01]. arXiv [DF07]. Ashok [San00]. Asia [Ste09d, TS02]. Asian [GM07, HR09, Kim07, SA07]. Aspects [DW01, Wor00b, Bur07b, CA09, Izs06, MM08b, OLBB^{+}08]. Assessing [CvPdlPR06, GR06a, HCTC03, MP06, NLM01, PM09, Ski06, Uzu04, HFG^{+}05, Hul07, Sch09]. Assessment [Cal06, DCB^{+}00, GGJ^{+}05, Har09a, Moe08, TBvL00, XZH08, ZNK01, AMTE07, BFG02, But08, Coc05, Gar03, GTSD09, GKB^{+}09, NN08, PBG^{+}05, PGFRD09, Pou05, Pou09, PP09, RRR07, Sch07b, Smi08, VSL09]. assessments [LFB^{+}06]. assist [VM06]. assisted [VY08, KS05]. Associate [VM06]. associated [TBH^{+}02, ZWW09]. Association [Goe05, BMD07, HH06, DT07]. Astronomers [Tri00, SFGRP09]. astronomy [LL03]. Asymmetric [ER01, CMM04]. Atomic [KKS^{+}07]. attempt [Vin01]. attention [Fra02, vDH05]. Attraction [LCL05, vR05a]. attributes [Pin06]. attributions [YNSS09]. August [Wor00c, Ano03d]. Australasian [ST02b]. Australian [But02, JMB06]. Austrian [SGBB03]. Author [Ano07a, Kre04, yTXwW03a, Wal06, Whi01a, Whi01b, CGV07, Cho08, GT04b, GDTS06, Har09b, Jar08, LF09, MDNL09, Per01, PS02, Saa06, SLI09, SGT06a, TG06, WBB04, WWJLG06, YA09, YNT06, YNSS09]. Author-centered [Whi01a, Whi01b]. Authored [Lan01, Abt07b]. authoritative [ERS06, MG04]. authorities [HS06]. authority [dB04]. Authors [GP00, Lan01, BRSYB04, BGS01a, BGS01b, BBG07, CS02, Gor07, HWD01a, HWD01b, Kre01b, Kre01c, KGK02, LT09, Mar09, Nwa06, Ski09, Sma04b, Sol07, TG04]. Authorship [Eto00a, Eto02, Far02, BK02, GT01, GS01, Glä01, GT04a, GS05, GÖ09, ISS09, IMLSSC09, KKKM05, Kao09a, Kao09b, Kim02, Kre04, KKK07, LF09, LTL^{+}06, OLVLH09, SG06, ST04, Uzu04, WWP^{+}05, YNT06]. authorships [HME07, HLSA07, ST02a]. automated [LH06]. automatically [Whi01a, Whi01b]. Autopoietic [SF00]. Average [SMYS05, YYL04]. avoidance [GS03b]. Award [McC05, Rou04, Rou05b, Whi08]. awardees [BSPKN03a]. awards [BSPKN03a]. awareness [KKM07]. axiomatic [Mar09].

B. [San00]. backbone [BKB05]. Background [KS05]. backlink [QCW04]. bad [CM09]. Balbi [ZM01]. Barrier [MHGGVJ08]. barriers [BL07]. base [DL06, MH07, MP06, RdL02]. Based [Sch01, ADC08, BMT09, BD08, BM09, BDZZ07a, Cho08, CBVS07, Dai09, Dan06, Egg07, GL05, GJT09, IK02, Jia08, KKSK06, LAFP04, LSGZMM^{+}08, LF09, Mar09, MMP06, MM08b, MAVQHS^{+}04, NLM01, ND09, PHL05, PK09, Pou07b, RB06, SLP06, SL05, TG09, TV06, Uzu02, VY06, ZGL09, vPdlPvZ08, vR05b, XCS03]. baseline [IGK07]. bases [TK09]. Basic [BD05e, HZT05, MHSE04, Egg07, GSB02, GM09b, MHL06, MB01, MG04, RDvT03, Vin01]. Bayer [DL06]. Bayesian [KP08]. Be [Hud07, AJOP04, Bur05a, KS09, LVCS03, LINV^{+}09, VL01]. Beauties [Bur05a, vR04]. become [GST03, MG04]. Becoming [Ley00, PR09, ZTG09a]. beginning [CHH04]. behavior [Cal06, NK07]. Behaviour [Hav01, WD01a, HHWD04, Kyv03]. beholder [CMU09]. Belgian [VCP00]. Belgium [RDvT03, RJDD07]. bells [BSPKN03b]. Belver [Chu01a, Chu01b, Moy01a, Moy01b, Sma01a, Sma01b]. Benchmarking [AVGC^{+}06, Smi08, TVvL02, Pou09, SLP06]. benchmarks [TG09]. bene [De 05]. benefits [GS03b, vL06]. Berlin [Hav01, KLK01b]. Better [GST03, CB08a, LFB^{+}06, Wra09a, vdWFM^{+}09]. between [AGG00a, BKM03, Bur06a, CvLT07, CR08, CB03, Dan00, DV03, Egg99, Egg08a, Eto00b, FJL03, GR06a, GBZAEC^{+}07, HH06, Har09b, HE09a, Jan00, Kos07, KT08, KKM07, Kre01b, Kre01c, Ley07, LTMW03, LHHWD06, Mey00a, Mey00b, MB04, Mey07, Nis02, PH09, Per08, PZ03, RRS07, RvLB^{+}02, She08, SS05, TS02, Tho01, TM03, Tsa09a, TBvL00, VMD07, Vin08, YO06, Ye07, YY08a, YNT06, YNSS09, ZGBI09, vdB00]. Betweenness [RZ08]. Beyond [AHK08, But08, GST03, Ran09, Uzu06]. Bhabha [KKS^{+}07]. Bias [LOK00, SV01, Van09, SFG09]. biased [PRdKB07]. biases [VMT^{+}01]. Biblio [ZM01]. Biblio- [ZM01]. Bibliographic [AJ08a, BH05, ER02b, FRJ08, HCC03, HV03, Kre04, KM07, VDL^{+}02, VDL03]. bibliographies [MG04]. Bibliography [Moy01a, Moy01b, Sch01]. Bibliometric [AOS00, AC04, AGG00a, AU07, AWG00, BMJARBC05b, BT06, CHH04, CH05, CH07, CCDJJF04, Dal02, Dan00, Eto00b, Glä00a, Glä00b, Hay03, HCLH04, HYA08, JI00, LDF^{+}09, LR09a, Luw00, MCMN02, MC02, MP00, Moe00a, Ned06, Noy01, PM06, QC09, RSJ00, SA00, Som05, TJM00, Van08a, WD01a, WWW07, Wor00a, ADP08, ADC09b, AGS07, BBV^{+}08, BMFT05, BMFG03, BM09, BN01, BNVL06, CSGGB09, CCHC06, CHH07, DG04, Fre04, GM02, GST03, GDP03, GT04b, GSST06, Glä08b, GGP05, Gu04a, GM07, GH07, HZT05, JLH07, Jar05, JRC09, JM07, KU05, Kim01, Kim07, Koe01a, Kra08, LSGZMM^{+}08, Lau03, LVCS03, Lew01, LCLA02, LFB^{+}06, Mar09, MVH^{+}05, MMP09, MBG01, Muk09, Mus06, Ned08, NC09, NSS04, NCM09, Nwa06, ÖSAT08, PRPFC06, PGD04, Pon02, RWGA07, RCE^{+}09, Saa06, Sap07]. bibliometric [Sch07b, SW01, SS05, TMKM09, TK09, TG06, Tsa08, VVM^{+}03, VM06, VWJ09, WBB04, WWJLG06, XZH08, XCS03, ZQH09, vL06, vL09, vR05a, vR06]. bibliometric-based [BM09]. bibliometrical [TG09]. Bibliometrics [Ack05, HW01, Ive00, Wor00c, vR01, AUQ09, GDTS06, God06, PBI02, Wei05, Whi01a, Whi01b, Uzu02]. bibliometrics-scientometrics [PBI02]. Big [CRLMR04]. bigger [vdWFM^{+}09]. binary [CR08]. bio [CvLT07]. bio-pharmaceutical [CvLT07]. biochemistry [HZT05]. biodefense [MWD02]. bioinformatics [GJT09, GG08, MMP09, PM06]. biological [QC09, RWD^{+}09]. Biology [Lee03, MG05, MG00, MSGGRR08, Pri03, RRGGMS06, RRGGMS07, CDFB05, CCDJJF04, HME07, HZT05, KSB03, Rei09, PH09]. Biomedical [Ton00, DLM06, GR06a, HR09, Lew02a, Lew02b, LP04, Nwa06, PA08, RV08, TI02]. Biomedicine [MSGGRR08, Pri03, RRGGMS06, RRGGMS07]. bionanotechnology [RM07]. Biotechnology [AGP07, AP04, BGG00, Kim07, LdLM00, PC05, SPN00, CBO^{+}03, Dal02, DL06, FH06, Hem06, MH07, McM09, MP06, RdL02, RJDD07, VMD07]. biotechnology- [FH06]. bioweapon [MWD02]. Blind [LOK00]. board [CGR06, Nis02, BD05c]. body [YMB09]. BOF [DG04]. BOF-key [DG04]. Bond [Lew04]. bonding [RR06]. Book [Kha05, OB05, Rou06]. books [Lew04]. boomerang [Lar02]. both [HG02]. Bottom [CC08]. Bottom-up [CC08]. boundaries [Bjö06]. Bradford [BRV00, vUS00]. brain [Lau03]. Brazil [GLT06, GKB^{+}09, LGT06, WXB09, Zan02]. Brazilian [ABN^{+}09, DMPM^{+}08, DR01, FJL03, LJFD01, LC02, LL03, MBC06, MMP06, PFE00, Per08, PMO04, RPFD07, VLdM04]. breadth [Pon02]. breakthroughs [KM07]. Bridges [CR08]. Bridging [ZRRB03a]. brief [vPdlPvZ08]. Bringing [Gor05b]. British [Bon05a]. brokerage [HB07]. Brought [Mar01]. Buckland [Egg04b]. build [VDL^{+}02]. building [MAVQHS^{+}04]. Bulgarian [KT01]. Bulletin [He03]. bureaucracy [Coc09]. Burrell [Egg01a, Egg01b]. Business [RRS07, BT06, Hem06, KS01, Lom08, Nis02, Saa06, Vau04, VW04, VY06, VGK06]. business-related [Saa06]. butterflies [LH06].

C [Chu01a, Chu01b, Moy01a, Moy01b, San00]. C. [PKK09]. CAB [KS05]. calculating [SMP04]. calculation [Jar08]. Calibrating [AGJ08]. CAMEOs [Whi01a, Whi01b]. Can [Lau03, LVCS03, Mey00a, GR06a, KS09, LINV^{+}09, OO09, ZK08]. Canadian [LGA06, Nel06, VKG07]. cancer [LMJS06]. candidates [LLR04]. cannot [VL01]. capabilities [VDC^{+}06]. capacity [ACdLPM09]. capita [Ye07]. capital [BH05, LMJS06, Ski09]. capitals [YZ08]. Capturing [DL06]. cardiovascular [Jar01]. care [GR06a]. Career [DCB^{+}00]. Caribbean [CRLMRPA08, MC02]. Carlos [IMLSSC09]. Case [AOS00, CH00, sKtP04, MG00, TMVAR00, vdB00, AY09, BI04, Bha04, But02, CC08, CMP03, CYKK09, CvPdlPR06, CIM07, DG04, Fra03, GAGPM06, Gor07, GÖ09, GBGCRFdMA09, Gup06, HE09a, HG08, HVG06, JM07, KLC^{+}08, KNM07, KKM07, LLSL07, LC02, LL03, LLR04, LC06, LZ07b, LRR09, LR09b, Mai06, Mey06b, MA07, RDvT03, RRMSG02, RV08, RJDD07, RB06, SS09a, SMYS05, SKMI06, SMB^{+}07, SS05, TG06, Uzu04, VH01, VKG07, Wag05, WHC07, WH07, YMB09, ZL02, ZWZM04, vCvCvU08, vdBH06, MSRRGG02]. Catalan [AUQ09]. Catalonia [RSJ00, VL01]. catch [JLC09]. categories [GBZAEC^{+}07, LR09b, LINV^{+}09, MAVQHS^{+}04, SMYS05, Vin01]. categorisations [HFG^{+}05]. Categorization [SMB^{+}07]. categorizing [BPM^{+}09]. Category [SMP04, BMFG03, LRR09]. causes [GOS08]. cautionary [PBZ02]. CD [CIM07]. Cell [Lee03, HCTC03, Lar08a, LH06, LDF^{+}09]. cells [Has05, KNM07]. censoring [BO05]. center [GS03b]. centered [Whi01a, Whi01b]. Central [BK06, Bos09, LJFD01, MCMN02]. centrality [LF09, RZ08]. Centre [KLK01b, KKS^{+}07]. centred [TZ08]. Centres [Goe05, Pri03, GK09]. centric [YS08]. Century [Hav01, Tri00, DL06, HHWD04, MMR05]. ceramics [SMB^{+}07]. chain [OLC07]. challenges [He03, HW03, RK04]. chance [GST03, Hud07, Vin02a]. change [Bha04, ERW05, Jap07, LGS03, Ley02, ÖSAT08, Soo09b, Uzu06, YK04]. Changes [Fuj06, Prp07, WM04, DL06, GDM08, RKM01, Vin02a]. Changing [Kyv03, GS07, ZGBI09]. Characterised [BRV00]. Characteristics [CYKK09, Tri00, vLM05, CL06, DA08, Glä01, GS07, IK02, ID04, LR09a, MN03]. characterization [ER02b, Egg05a, Mar09]. Characterizations [Whi01a, Whi01b]. Characterizing [BR00, BKM03, HB07, SKT09, Vin09, OCA09]. charting [GSB02]. charts [GTSD09]. Chemical [MS01, ND09, SHO00, CDR06, HL09, YHM05]. chemicals [LGS03]. Chemistry [KGS00, Kum09, MG00, VM06, BMJARBC05a, BK06, BBK09, BSPKN03b, GS01, HME07, Kum06, ND09, Tsa09a, vR06, KKS^{+}07]. Chile [PZ03]. Chilean [DSBH09, KKM07]. China [Gu03, AB01, Bha04, CCHC06, CCHC07, Gar02, GH05, GM07, GD03, HZT05, HL09, He09b, HVG06, Jon09, KG05, LZ05, LZ03, LF03, LZ02, LHHWD06, LpCY06, MDNL09, MHL06, Moe02, TT03, WWP^{+}05, WWW07, WPZ^{+}04, XCS03, ZTG09a, ZWW09]. Chinese [GG09b, GH07, GG08, He03, HG08, HL09, JZCZ02, JRS06, JT08, KLK01a, LKGdB01, LGD02, LLR04, tLtC07, LZ07b, MG05, MDNL09, QCW04, RR02, VY06, Wu09, YZ08, ZL07, ZTG09b]. Chinese-Indian-German [KLK01a]. Chinese-language [LZ07b]. choice [RJDD07]. circuit [MF09]. circulation [ZQH09]. citable [CM09, Fra08]. Citation [Abt00a, AET09, AQ03, ABZMD^{+}07, Ano01b, Ano01c, CG06, CIM07, DCF00, Egg00b, Eto00a, GKL06, Hal09, Hav01, Ing00, KKSK06, Kim02, KK04, Kos02, KT08, Lan01, LLJ09, Ley02, Ley04b, Ley07, MSW01, Moe02, NSS04, OW00, PL08, PRdKB07, Pis09, RVB^{+}01, RvHMJ05, Sæt00, SIMM08, Sch01, SHO00, Sig00, SU09, SM00, SMP04, SMB^{+}07, SG09, SG01b, TMVAR00, Tsa09b, TBvL00, VMT^{+}01, Ver00, WM04, YS04, ZL02, ZWW09, Ada05, AGM07, Aks03, AS04, AVDT04, AJ08b, AUQ09, BLZ07, Bas06, BO05, BN06, BD06, BBG07, Bur01a, Bur02b, Bur02a, BV06, CSW^{+}09, Chu01a, Chu01b, CHH07, CRLMR04, CRLMRPA08, CS02, Egg01c, ER02b, Egg07, EP03, Eto02, FA07, GL05, GM03, Glä04, Glä06, GJT09, Gmü03, GDD03, Hen09, HCC03, IK02, ISS09]. citation [KKS^{+}07, KBWM07, Kra08, Kra07, Kre01b, Kre01c, Lar02, LLSL07, LT08, LT09, LH08, Liu05, MDNL09, MC09, MH07, Meh07, Mey01, MB01, NK07, NLM01, Nel06, NLD07, ND09, Nis02, OLBB^{+}08, OLVLH09, PL09, SV01, SM06, Sch06, SLI09, Sch08, SGT06a, SP08, SKM07, SR05, Ski06, Ski09, Sma04a, SRS^{+}02, SKMI06, Soo09b, SH09, SL05, TV06, yTXwW03a, yTXwW03b, yTlC05, Van01b, VS08, VM06, VDL03, Wet06, Whi04, Woh09, Yan09, YAH08, YR01, YL07, YW04, ZGL09, ZCY09, ZK08, ZL07, ZRRB03a, ZRRB05, vLM05, KS05]. Citation-Assisted [KS05]. citation-based [NLM01, TV06]. citation-weighted [MH07]. Citations [BR00, CGR06, Hud07, Mey00a, Mey00b, Ran09, SLB00, SF00, WPZ^{+}04, ACF09, AGJC09, AR06, ACdLPM09, BM03, CB09, CG06, CM09, DF07, Dre03, GT04a, GT04b, GS05, GDTS06, HFG^{+}05, Har09b, Har08, HLSA07, JLC09, KT06, KT08, KBWM07, Lew04, MC09, Med06, MN03, PH09, Per01, PS02, Pod05, PBZ02, RCE^{+}09, SG06, SGT06a, TG06, TM03, Wal06, Wan07, WBB04, BK06]. Cited [Bas06, Ioa06, SF00, AJOP04, AS04, GST03, GM03, Kos07, LT08, LT09, Per06, PA08, Sma04b, TVvL02, VDL03, YW07, ZWZM04]. citedness [Vin02a]. cities [BH05]. Citing [SF00, Kap04, Per06]. civil [Igi02]. class [Cam09]. classes [MAVQHS^{+}04]. classical [RRZR09]. Classics [Ano01b, Ano01c, Yan09]. classification [CL06, GS03a, LFAH04, LP04, LR09b, TG08]. classifying [Vin01]. clinical [OMD01, RCE^{+}09, YW04]. closed [FPGBD03]. closer [Gor05b, MB01]. clustering [AJ08a, SG03, ZCY09]. clusters [MBPKT03, Sch06]. Co [BK02, DCF00, ER02b, Gmü03, GÖ09, OLVLH09, PBZ02, Ste01, WWJLG06, Bur06a, Bur07b, CvLT07, GSD09, GT01, GS01, Glä01, GT04a, GS05, HME07, HLSA07, ISS09, Jon09, Kap04, Kre04, KKK07, Kre01b, Kre01c, LJ08, Liu05, LF09, LTL^{+}06, MDNL09, MLNS08, PID06, QLLR08, SM06, SS08, SLI09, SG06, ST02a, ST04, SG03, SBSCOR06, yTXwW03a, yTXwW03b, VY06, VKG07, VY08, WWP^{+}05, YNT06, YNSS09, ZCY09, ZRRB03a, ZRRBL03, vdBH06, BK06]. Co-author [WWJLG06, YNSS09]. Co-authorship [BK02, GÖ09, OLVLH09, GT01, GS01, Glä01, GT04a, GS05, ISS09, Kre04, KKK07, LF09, LTL^{+}06, SG06, ST04, WWP^{+}05, YNT06]. co-authorships [HME07, HLSA07, ST02a]. Co-Citation [DCF00, ER02b, Gmü03, Kre01b, Kre01c, Liu05, MDNL09, SM06, SLI09, yTXwW03a, yTXwW03b, ZCY09, ZRRB03a]. Co-citations [PBZ02, BK06]. co-concentration [Bur06a, Bur07b]. co-existing [Kap04]. co-item [ZRRB03a]. co-link [QLLR08, VY06, VY08]. co-linked [VKG07]. co-location [ZRRBL03]. co-occurrences [vdBH06]. Co-operation [Ste01, SBSCOR06]. co-outlinks [GSD09]. co-publication [Jon09, MLNS08]. co-publications [CvLT07, SS08]. co-sitations [PBZ02]. co-usage [PID06]. co-word [LJ08, SG03]. coauthored [GD03]. coauthorship [YK04]. COBIB [Pet08]. cocitation [Jar08, MAVQHS^{+}04]. code [LL08]. Coefficient [Hu07]. Coexistence [MHGGVJ08]. Cognitive [MH01, Jar01, Van01a]. cohesiveness [MLR05]. cohort [DT07]. coin [PS02]. COLCIENCIAS [AGG00a]. Cold [Ack06]. Colima [Moe00b]. collaborate [BL07]. Collaboration [BK00, BA01, Bea01, DW01, Fre04, GDD03, Hav01, KLK01a, KT01, LKGdB01, LTL^{+}06, MH01, Soo09a, Tho00, WD01a, ZBO00, AU07, BL07, Bos09, EL09, GP01, GdL02, GD03, He09b, HHV03, HKL08, Jap07, Jia08, JT08, KKKM05, Kim05, Kim06, KP08, KA04, LGA06, LV09, LC02, LZ02, MG05, MSRRGG02, MB04, NM08, PGD04, Sch04, Soo09b, Wag05, WWP^{+}05, YO06, YNT06, YNSS09, ZGBI09, Soo09b]. collaborations [Boy09, OLC07]. Collaborative [LGD02, dB04, LF09, Rig09, SSM07, Soo09a]. collaborator [KKK02]. collection [BI01a]. collective [Fra02]. colleges [Gmü03]. COLLNET [KLK01b, KLK01a]. Colombian [AGG00a, AGG00b]. colonial [DGAVK03]. colonialism [Bos09]. colons [LH05]. column [BMD07]. Combinatorial [MG00]. Combining [BNVL06, GGP05, VVM^{+}03]. combustion [AVDT04]. coming [BS03]. Comment [Hal09, LLJ09, SG09, LCL05]. Comments [Ano04b, Egg01a, Egg01b, Gu03, Xek05, vR05c, Rou02]. commercial [VW04]. committee [BD05c, BD06]. common [AJ08a]. Commonalities [MB04]. communal [Ski06]. Communication [LK00, SSM07, vdB00, Bal09, Bro04, CSW^{+}09, Dai09, HHV03, Koe01b, Koe01c, LAL07, PL09, San01a, San01b, SS05, WBB04]. Communities [Mey00a, CA09, Jia08, RB06, Ski06]. Community [DFRG02, GT01, Req05, ZV09]. companie [HD07]. companies [CC09, FG06, Hem06, HCC03]. company [Hud07]. Comparative [BCGMBSVM07, Egg09a, KKL06, LOK00, Pou09, YK04, YNT06, CDFB05, GLT06, GJT09, GM04, HR09, LGA06, OLBB^{+}08, PMO04, Pri03, SLI09, Ski06, SB09, WXB09]. compare [AB01, BCK06]. compared [GM07]. Comparing [Pis09, Rig09, VY06, Per06]. Comparison [Cah00a, GH05, GG08, OW00, Pod05, SLP06, vR06, AJ08a, BMT09, BI08b, Cri06, Eto02, FJL03, HL09, HLT09, Jar05, Kim07, KB08, KT08, Kum06, LT08, MDTU05, Mey06a, PHL05, Prp05, SL05, yTlC05, Tsa09a, TLL^{+}09, VM06, VG09, vCvCvU08]. Comparisons [GLM^{+}08, Arc02, GSH01, VMT^{+}01, ZRRB03b]. Competition [BGG00, BS01, Eto00a, GSM05, RdL00, Van01b, VY06]. competitive [HHWD04, MMLA01a, MMLA01b, VY08]. competitiveness [CLH07, GSM05, MNQ08, Ygl03, YY08a]. complementary [DD09]. Complex [Wen01]. Composite [Vin06, ML05]. composition [Nis02, YMB09]. compound [LAFP04]. compounds [Ban06]. Computer [LTMW03, LKGdB01, WXB09, ABN^{+}09, GM04, HG08, KG05, MNQ08]. COMS [PD07]. concatenation [Glä08a]. Concentration [ER01, Ioa06, VLdM04, Bur06a, Bur07b, ER02c, Fra05, GT01]. Concept [Sch06, BNVL06, Tsa09b]. conceptsto [AGM^{+}01]. Conceptual [vR05a, Lee09]. concerning [RR06]. concise [GDTS06]. Condensed [GD08]. Conference [GM09a, Hin03, Moe00b]. conferences [ABGAMD^{+}09]. conjecture [ERS06]. connection [YNSS09]. connectivity [YDM^{+}05]. consequences [AL09, Gar04, Wei05]. consequencesfor [VMT^{+}01]. consideration [Kim06]. considerations [HH90, Kra08, Llu05, TT03]. considered [VL01]. considering [YNT06]. consolidated [RRMSG02]. Consolidation [RSJ00, MSRRGG02, MSGGRR08, RRGGMS06]. Constructing [HCC03, NR03]. construction [GG09a]. Constructions [Lia06]. consumer [Saa06]. contemporary [Zsi08]. Content [BI00, KBWM07, VY08, BIE05, BND08, CIM07, Fuj06, VGK06]. Contents [Ano01b, Ano01c]. Context [GW01, vUS00, AJ08a, Kim06, Noy01, NCM09, PID06]. contexts [Ing02, LH06, Sch06]. Contingencies [AL09]. contingent [LH08]. continuant [Gor07]. continue [AJOP04]. Continuity [WD01a]. continuous [CM09]. contribute [CG06]. Contribution [HG08, Ton00, TI02, Vin09, ADC09a, Bon05a, CCHC07, CRLMR04, LL03, LPC05, Sap07, Tij07]. contributions [HLT09, JJ04a, LZ05, RV08, Wra04, ZV09]. contributors [Joh09]. control [CvPdlPR06, YYL05, YY08b]. controversial [Ack05]. Controversies [Sæt00, LH06]. Convexity [RRZR09]. Cooperation [AET09, Gus00, Ste09b, ZBO00, BGS01a, BGS01b, CvLT07, GLM^{+}07, Nag02, Nag03, PZ03, vL09]. Copenhagen [Wor00c]. copying [Wet06]. Core [CCHF05, FG06, BS01, BD05d, Bur03, Chr04, ER02c, GJT09, Gor05b, KU05]. Cornelius [Moe00b]. corporate [Dai09, GMA07, VD06]. corporations [BM03, LpCY06]. corpus [ADC08]. corpus-based [ADC08]. Correcting [ZRRB03b]. Correction [GTS05, SH21]. corrections [Sch08]. correlate [Ada05]. correlated [Dai07]. Correlation [HE09a, SS05, Vin08, YY08a]. Correspondence [Ano03a, Ano06a, AU07, VWJ09]. cost [GS03b]. costs [WFI03]. Council [AET09, AGP07, RRGGMS06]. Councils [Rin00, Pri03]. count [LLSL07, MBR^{+}08]. Counting [GL05, GLM^{+}08, KBWM07, Lar08b, SLI09, SGT06a]. countrie [SH08b]. Countries [SH09, AC04, BH05, Bon05b, BDZZ07a, Cal06, CBVS07, DGAVK03, Das04, Gok03, GM07, He09b, HR09, Ing02, Kim06, KS01, KKM07, LHHWD06, Mus06, Pou09, SIMM08, SA07, Vin08, WXB09]. Country [ZRRB03b, Arc02, BI04, Bas06, Egg99, Kao09a, Kum06, ST08, YA09]. Counts [Abt00a, Lan01, Ada05, LLSL07, QCW04]. coupling [AJ08a, ER02b, HCC03, KM07, Lo08]. Coverage [dMACRVQ^{+}07, Egg04c, VMT^{+}01]. covered [CG06]. covering [VBBF08, YA07]. created [VGK06]. Creation [BMT09]. creationists [CA09]. Creative [Hem06, HB07, HSSK07, LVCS03]. Creativity [Sol06]. creators [CS02]. credit [TG04]. Crick [AJOP04]. crime [Wal06]. crime-psychology [Wal06]. criteria [BND08, Fuj06, Ye07]. criterion [MM08b]. Critical [Ben03, Wor00b, GS01]. Croatian [AJ08b, Gol02, KK04, Luk06]. Cross [SG06, Arc02, Gar03, OALL07, RM07, ZGL09, ZRRB05, Kum06]. cross-citation [ZGL09]. cross-country [Arc02, Kum06]. cross-disciplinary [RM07]. cross-field [ZRRB05]. Cross-national [SG06, Gar03]. cross-scale [ZRRB05]. cross-sectional [OALL07]. crossing [Bjö06]. Crown [GDD03]. CSI [JZCZ02]. CSIC [AGP07]. CSIR [Gup06]. CSTPC [WPZ^{+}04]. CubaCiencias [RvHMJ05]. Cuban [RvHMJ05, TMVAR00, TMVAR00]. cultural [Inö03]. cumulative [CJ06, Egg07]. currency [OCA09]. Current [Vin09, Jar01, YHM05, Ano01b, Ano01c, CEM06]. Curriculum [DCB^{+}00]. curve [Woh09]. Curves [BRV00]. cycle [LV09, SK01b]. cycles [SK09]. cystic [YMB09]. Czech [Van08a, Van08b].

D [Edi05, McC05]. Danish [Bal06, Fre04, Wor00b]. Data [BI01a, FPGBD03, LS05, BO05, Bay05, BDS06, Bur02a, Bur05b, Bur07a, CBVS07, EP03, GMA07, GG09b, GSST06, LLSL07, MB01, Mus06, Old01a, Old01b, Saa06, ST08, SL05, Tsa09a, VY06, ZL02, vZvPdlPH06]. Database [AWG00, LAB^{+}09, NSS04, Sig00, AJ08b, BPM^{+}09, GD09a, Luk06, MBR^{+}08, ND02, SKMI06, SMB^{+}07, ZRRB03b]. databases [AVGC^{+}06, BH05, CCHC06, CIM07, HW03, HV03, MSW01, MMP06, OMD01, RvHMJ05, TSLCJC09]. DDT [YX08]. dealing [WWJLG06]. decade [SSG06]. Decades [Hav01, SG01a]. Decay [Ing00]. decisions [BD05c, HH06]. decisive [GL05]. decline [GDP03]. declining [LZA08]. Decomposing [GDP03]. decomposition [Ley04b]. decreases [CM09, MG04]. Decreasing [WD01b]. deeply [Bon04]. deeply-felt [Bon04]. Defenders [CA09]. define [ER02c]. Defining [Bur03, Ste09a, Vin01]. definition [ER02c, Eto08]. degree [BE04, vZvPdlPH06]. delay [Ama08, RVB^{+}01, SK01b, YA09, YRL03, YYL04, YYL05, YGY06, YY08b]. delays [YWY05, YGL06]. delineation [LZ07a, Ley08, ZRRB03b]. Delivered [Moe00b]. Delivery [VCP00]. Demand [SLB00]. demographers [VH01]. demon [Jas09]. Denmark [Wor00c, And01]. dense [SM06]. Department [Kra08, Per08]. Departmental [MP00, LTWM05a, LTWM05b]. Departments [LTMW03, Kra08, Ned08, SGBB03, TT04, TSLCJC09]. departure [Gol02]. dependence [SGT^{+}06b, TNK^{+}06]. Dependency [DSBH09, Mey07]. dependent [Bur07b]. depends [Abt07c]. depth [LLSL07]. Derek [Ano08b, Ano09, Edi05, Ley04a, McC05, Rou04, Rou05b, Whi08]. derived [ABGAMD^{+}09]. Dermatological [BRSYB04]. dermatology [SG01b]. describing [Bla09]. description [CS02]. Design [YY08b]. designed [GS03a]. detailed [Boy09]. detect [TK09]. detecting [TLL^{+}09]. detection [CC08]. determinant [RRGGMS07]. Determinants [LCWD09, DT07, Fre04, MSM06, RRS07]. developed [DGAVK03]. Developing [Gu03, Gus00, LZ03, VMD07, Gok03, KKM07, LL08, VZV^{+}03, YDM^{+}05]. Development [JZCZ02, vLM02, BKK^{+}08, CB03, Dan06, DL06, GGJ^{+}05, Hul07, MML^{+}04, Mus06, Pon02, WD05, Zan02]. diabetes [SS05, SASA08]. diachronous [Glä04]. diagnostic [Fuj06]. dibenzo [PRPFC06]. Did [Wra09a]. differ [HPF03]. difference [AY09, HR09, Kos07, LINV^{+}09, TNK^{+}06]. Differences [BMFG03, Kim02, YNT06, ADC09a, ADC09b, Gan09, Goe02, McM09, MB04, Prp05, Mey00b]. different [BK06, BCK06, CIM07, GLM^{+}08, Hu07, IP07, Lew02b, Pod05, SLI09, SFG09, SMYS05, TNK^{+}06, TG09, TSLCJC09, YK04, ZGL09]. differential [BMFT05]. differentials [Prp02]. Differentiating [Bla09]. Differentiation [vR00, vdB00, BD08]. Difficulties [He03]. diffusion [BPMK06, CH04, CEM06, Fra03, Kor01, MSS06, PSD^{+}07]. digital [MSS06, YS08]. dimension [Egg05b, Hu07]. dimensional [Egg04a, Egg05b]. dimensionality [Glä09, Pri03]. dimensions [CX09, YDM^{+}05]. dioxins [PRPFC06]. Direct [OALL07, Lee09]. directed [RZ08]. directions [Gei05]. disaster [CM05]. Disciplinarity [DW01]. Disciplinary [Fre04, TT03, CL06, JD05, RM07, TT04]. discipline [CRLMR04, CEM06, Dan06, Joh09]. disciplines [HHWD04, KT08, PA08, SL05]. disclosing [SKM07]. Discontinuity [WD01a]. Discourse [Sæt00]. discoveries [Cam09]. discovery [DL06]. discrepancy [Per08]. discussions [HPF03]. display [JLC09]. disproportionate [Egg05b]. dissemination [MF09]. dissertation [LZA08]. Distance [Eto00b, Kre04]. distances [BK02]. distinct [CA09, OO09]. distinguish [Har09b]. distinguished [DT07]. distributed [RMTM07, WL03]. Distribution [BRV00, Egg00a, Egg00b, Bur01a, Bur02b, Egg01c, ER02a, Egg07, Fuj02, JSJX02, LF09, NBR01, SJJ04, SH09, Wol06, YM08, YL07, vLM05, Xek05]. distributional [GdL02]. Distributions [BRV00, ABN^{+}09, Bur06a, Van01b, ZRRB03a]. diversity [KSB03, San09]. divide [MSS06]. divided [CR08]. Division [KKS^{+}07]. DNA [AJOP04]. Do [Abt00a, BND08, MFG07, RJDD07, SFG09, Soo09b, VZV^{+}03, ACYYA06, ACAYY07, ACdLPM09, BL07, HPF03, JLC09, Mey07, MM08a, PL08, YDM^{+}05, ZIJSCMGG07]. doctoral [BE04, BD06]. Document [Sig00, VCP00, AJ08a, KBL08, SB09]. document-document [AJ08a]. documentary [LFAH04]. Documents [CG06, CVV^{+}06, Chr04, Gup06, PA08]. Does [Aru01, dB04, BD05b, CGC^{+}00, DF07, FA07, GT04a, KO08, Ski09, AB01, LINV^{+}09]. domain [Bjö06, KL09, LJ08, MVH^{+}05, MMAHS08, TMKM09]. domains [DV03, Fuj02, MAVQHS^{+}04, TK09, VDC^{+}06, YK04]. domestic [FJL03, GH05, JJ04a]. Domesticity [GS05]. dominance [VKVF09]. dominant [Joh09]. done [LINV^{+}09]. Double [GS01]. down [Ley04b]. downloads [DF07]. Dr. [Moe00b, San00]. drain [Lau03]. driven [LV09, Mey06b]. Driving [PFE00]. drug [DL06]. Ductus [HCLH04]. due [DLM06, ISS09]. Durfee [AHK08]. During [KGS00, CH05, Fuj06, GP02, GD03, GD09b, HYA08, Nwa06, ORSCSB00, PV00, QC09, SRS^{+}02, Uzu06, YX08]. Dutch [Ver00, vLM02]. Dynamic [Vin02a, GDM08, RB06, WL03]. dynamical [CA09, CC08]. Dynamics [Req05, Bra09, Fuj02, Gu04b, HHWD04, LV09, Ley02, Ley03, LHHWD06, MA01, RDvT03]. e-COMS [PD07]. E. [DL06]. Early [Ada05, JLH07, Mea01a, Mea01b, Pon02]. earth [GOS08, JJ04a]. East [Ste09d]. ecological [Kra07]. ecology [NC09]. Econometrica [Fra03]. Economic [DT07, SF00, Coz06, GR06a, KZ09, Lee09, Pis09, RDPAC09, WM04, YT09]. Economics [Cah00b, RRS07, SL05, Fra08, KS01, Med06, Ned08, Rou02, RT08, SSG06, SK01a]. economies [PHL05]. economists [DT07, Rou08, RT09, Tol09]. economy [CvPdlPR06, Fra02, Mey06b, WH07]. Ecosystems [Ioa06]. edited [Whi01a, Whi01b]. editor [Egg10, HH06, AJOP05, Ano03a, Ano04b, Ano04c, Egg01a, Egg01b, KKKM06, Xek05]. Editorial [Ama08, Ano01d, Ano01e, Ano03b, Ano04a, Ano08a, Bra04a, Bra04b, MS04, YA09, CG06, CGR06, Nis02, dMR01]. Editors [San00, BND08]. Education [CGR06, OACS08, PSD^{+}07, BD08, ISS09, OLC07, PD07, RK04]. Educational [CGR06, FCB02, GGJ^{+}05, PD07]. Effect [GDD03, AS04, Cri06, Har08, HT06, Lar02, MSRRGG02, MN03, Big08, Bro04, Bon05b]. effective [Per08]. effectiveness [Kim07]. effects [Bal08, BD03, BD05a, Ski06, YT09, ZRRB05, vLM05]. Efficiency [MHL06, JJ04b, Med06, ST08]. effort [GS01]. Egghe [Bur01b, Bur01c, Bur03, Bur09, Sch08]. ego [YS08]. ego-centric [YS08]. Egypt [Far02]. EICSTES [LAFP04]. eight [BIP09]. elaborated [LAFP04]. elasticity [DV03]. election [DT07]. electrical [SGT^{+}06b]. electronic [LZA08, MF09, Muk09]. electronics [FG06, MM02, Ste09d]. element [MF09]. Elementary [CRLMR04]. elicitation [MVH^{+}05]. Elite [DW01, LMJS06, Nag02]. Email [ZL00]. emancipation [Bon04]. emergence [BK06, BKK^{+}08, LAL07]. Emerging [Jon09, MG00, Mey00a, San00, Ben03, Lee08a, Mey07, RRR07, SU09, TMKM09, TK09, VDC^{+}06, WHC07]. Emphasis [NRTD00]. Empirical [BMJARBC05b, LK00, LRR09, Prp00, BCZM^{+}09, Bur06a, GAGPM06, LLSL07, LCWD09, MSGGRR08, Prp05, Wu09, RRGGMS07]. encountering [Egg07]. end [GSB02]. enduring [Bon05a]. energy [Tsa08]. engine [MDNL09]. Engineering [Ben03, DR05, Kao09b, MVH^{+}05, PV00, SHO00, cL07, Lo08, Son03, Tsa09a]. Engineers [DCB^{+}00, Kim02]. engines [AVDT04]. England [HL09]. English [Gu03, CR08, LZ03, VKVF09, WWW07]. English-language [Gu03, LZ03, WWW07]. enhance [OLC07]. Enhancing [AHK08]. enrolments [HHWD04]. Entering [Tri00]. Entrepreneurial [RDvT03]. Entropy [Ley02]. Entry [WD05]. Envelopment [LS05, ST08]. environment [FPGBD03, MMS05]. environmental [ACF09, ERW05, Jap07, Rou08, vdWFM^{+}09]. environments [Hem06, ZL07]. Enzyme [PV00]. epidemics [Ack06]. Epidemiology [KT01, NM08]. episodic [vR05b]. epistemic [dB04, RB06]. epistemologists [Kre01b, Kre01c]. EPO [Cri06]. equation [GM09b, YW04, YS04]. Equipment [PC00]. era [Bar06, Koe01b, Koe01c, MJL^{+}09]. Erdos [BK02].

Erratum [Ano03c]. error [Wet06]. Essays [San00]. Essential [CLH07, CBVS07]. Established [Tri00]. establishment [Fuj06]. estimation [CEM06]. estimators [LFAH04]. ethics [Prp05]. EU [Ing02, MLNS08, RWGA06, SH04, Vin08, ZRRBL03]. Europe [ACF09, BDZZ07a, CB08b, GS07, Mus06, TTP03, ZGBI09]. European [BH05, BRSYB04, Dan00, Gus00, Ing02, Jon09, LdLM00, Ley00, OACS08, OCA09, PC00, PBG^{+}05, PSD^{+}07, Pri03, RWD^{+}09, SIMM08, SSG06, TBH^{+}02, TZ08, TG06, TG08, TG09, UCCM01]. evaluate [CLH07, IRN07, LH08, Llu05, Wan07, YAH08]. Evaluating [tLtC07, OMD01, SGBB03, SM00, ZCY09, JCTSM^{+}09, QMC08, SASA08, Vin06]. Evaluation [BM09, JJ04b, LOK00, AGP07, BN06, CMU09, Cho08, DMPM^{+}08, FCB02, FH06, GAGPM06, GS03a, Gmü03, GSMT09, GKB^{+}09, GH05, GG08, HCC06, KP09, KB06, KBWM07, Kra08, MHL06, PA08, Rou08, SMYS05, ZWZM04, vL06]. evaluations [JD05, SKMI06]. evaluative [PGD04, Vin01]. Everything [Bra09]. Evidence [DT07, LK00, Rou02, Sæt00, VD06, ACF09, Bal08, Chr04, HD07, RRGGMS07, VS08, Pou07a, SK01a]. evolution [BBK09, CJ06, Has05, Iga08, NC09, RWGA07, RRR07, VD06, WHC07, CA09]. evolutionary [CDFB05]. evolving [sKtP04]. examination [GS03b, McM09, Ste09b, VMD07, Wra04, ZGBI09]. examined [BND08]. examiner [ACdLPM09]. Examining [Kre01b, Kre01c]. Example [SLB00, AJR09, CBVN06, He03, LAFP04, Lee08a]. excellence [Bas06, BD06, HBK^{+}09, KKK02, TVvL02, VVM^{+}03, ZRRB05]. excellent [SS09a]. exception [Bon04]. exchange [RVB^{+}01]. Excursion [Ive00]. exemplary [Kra08]. Exercise [Har09a, Pou07b, Smi08]. Exercises [Moe08]. existing [AVGC^{+}06, Kap04]. Expanded [GKL06, AJ08b]. Expansion [LINV^{+}09]. expectations [SH08b, vR01]. expected [Bur05a]. expenditure [Lom08]. expenditures [YM06]. Experience [Gus00, YA09]. experiment [JMB06]. experimentation [ER02a]. Explaining [Jap07]. explanation [Egg99, Egg05b, SSG06]. Exploiting [Lar02]. Exploration [Iga08, GG09b, Mey01, NC09, OACS08, VDC^{+}06]. explorations [BNVL06, RM07]. Exploratory [DCB^{+}00, LM00, CB08a, DFRG02, EP03, GM03, GS07, HME07, Lee09, MG05, Mey06a, QLLR08, QMC08, TT03, TBK06, vdBH06]. explore [KM07]. Exploring [ACF09, BD05a, DD04, LTL^{+}06, MT07, Nag03, Per06, RRGGMS07, Saa06, VVM^{+}03, Vau04, ZV09]. export [GBZAEC^{+}07, SMW^{+}09, SG01b]. exposure [VVPCOL09]. expressed [Pod05]. expression [YYL04]. Extending [BV06]. extension [Ban06, ERS06, GP09]. Extensive [BI00]. extent [Joh09]. External [Tho00]. extinctions [Wet06]. extra [MLNS08]. extra-EU [MLNS08]. Extremadura [GBGCRFdMA09]. eye [CMU09].

Facing [Nal01a]. Factor [SMYS05, ABZMD^{+}07, AGJC09, AL09, BMT09, BFG02, BDZZ07b, BCZM^{+}09, CG06, CGR06, CM09, Dan06, Egg08a, GSMT09, LPC05, Ley04b, LC06, Llu05, MBR^{+}08, Van09, YGY06, dMR01, CEM06, RGD00, SASA08]. factor-analytical [Ley04b]. Factors [Jon09, PFE00, Wan07, BL07, CMU09, Gol02, Inö03, LZ02, Mai06, MFG07, NLM01, Pis09, Rou05a, YWY05, YW07, vLM05, Big08, QCW04, ST02b, SMP04, Vin02b]. facts [SKM07]. Faculties [SL05]. Faculty [TI02, Kyv03, LCWD09, OALL07]. failed [Ack06]. fair [ZRRB03b, Zsi08]. fairness [BD05c, Rei09]. families [Cri06, Ste09a, vPdlPvZ08]. family [DW09, LC06]. far [HPF03]. Fatal [LCL05, vR05a]. Feasibilities [MSW01]. feasible [vL06]. feature [HME07]. Features [AGG00b, Vau04]. feelings [FH06]. fellowship [BD05c, BD06]. Felsenstein [CDFB05]. felt [Bon04]. female [Lew01, MM05]. few [AGJC09]. fibrosis [YMB09]. Field [AOS00, Mey00a, AGP07, AVDT04, AP04, BMJARBC05a, CBVN06, CB09, CGR06, CC08, GKB^{+}09, Gor07, GG08, HLSA07, JLC09, Lee08a, LZ07a, LL08, MSGGRR08, Mey01, Mey07, RRGGMS07, SG01a, SFGRP09, VMD07, ZRRB05]. field-normalisation [ZRRB05]. Fields [Ing00, MH01, Sch00, vdB00, BKK^{+}08, CJ06, Fuj02, GSB02, GS03a, Hu07, HR09, IP07, KS01, Lee08a, ND09, Pod05, PR09, RvLB^{+}02, Wra04]. filed [Ste09b]. films [BM03]. filter [Chr04]. filtering [Sch06]. Financing [RS00]. finite [GMA06]. finite-state [GMA06]. Finland [HT09]. Finnish [RV08]. firm [YM06]. firms [BM03, CvLT07]. First [BNVL06, Egg00b, Lan01, AQ03, Bur01a, CL06, Egg01c, Jar08, LT09, Per01, SLI09]. First-Citation [Egg00b, Bur01a, Egg01c]. First-named [Lan01]. Fish [AB01]. fit [BMJARBC05b, BMJARBC05a]. fitting [Izs06]. Five [ZBO00, Cal06, SIMM08]. Flash [BD05e]. flat [JLC09]. flaws [BMJARBC05b]. Flemish [DG04, Luw00]. Flow [PB09, Bro04, SLP06, VDL03]. Flows [SPN00, BLZ07, Mey02, PK09]. Fluctuation [Dan06]. Fluid [CR08]. Fluorine [Pou09]. focused [PBG^{+}05]. Food [Ama08, AC04, AY09, Cal06, YA09]. foreign [CvPdlPR06, JJ04a, LpCY06, Uzu04]. Forensic [VVPCOL09]. forerunners [JLC09]. Foreword [Bra04b, Gar01a]. Forgetting [Pol00]. formal [MC09]. Formalised [ZNK01]. format [CIM07]. formats [GMA07]. formula [TG04]. found [CVV^{+}06]. Foundation [KLK01a, KLK01b]. foundations [GTSD09]. four [CM09, Kim07, KT08, MN03, Pou09]. fourfold [AGJC09]. fourth [MF09]. Fractal [vR00, Bru03, Egg05b]. fractality [BMJARB^{+}07, BMJARBC05c]. fraction [Egg99]. fractional [Egg99, SGT06a]. Fragment [BRV00]. Fragmentation [LK00]. frame [TVvL02]. France [SLB00, Oku00, RdL02, YO06]. Frankenfoods [LH06]. French [LMJS06]. frequencies [Abt07a]. Frequency [Abt00b, PRdKB07, Ski09, Wol06]. Frömmichen [ZM01]. front [Jar05]. Frontier [BPA00]. fronts [MMAHS08, SW01]. FSG [GMA06]. Fuchs [Bon05a]. fuel [Has05, KNM07]. Full [ABGAMD^{+}09, GGP05, VM06]. Full-text [ABGAMD^{+}09, GGP05]. fullerenes [LAL07]. function [Egg04c, Egg05c, YM08, ZRRB03a, ZRRB03b]. Functional [PD07]. functionality [BN01]. functioning [BMFT05]. functions [Bur06b].

Fundamental [Cah00b]. fundamentalism [Pou07a]. funded [GM09b]. Funding [HBK^{+}09, GSM05, GGJ^{+}05, Goe05, Hor01, JMB06, LTL^{+}06, MMLA01a, MMLA01b, Rig09, San09, Som05]. Further [Lan01, BMFG03]. Fusion [Ack06]. Future [Bea01, Abt07b]. fuzzy [ER02c, HW02, ML05].

G [Rou02]. G7 [He09b]. gaining [vDH05]. gap [Zan02]. Garfield [ERS06, Jas09, Vin02b]. Garg [Ano04b, San00]. Gasset [BMJARB^{+}07]. gastroenterology [CCHC06]. Gatekeeper [BDZZ07b]. gatekeepers [BD05d, BDZZ07a]. Gatekeeping [BZDZ07]. GDP [Vin08, Ye07]. Gender [ADC09b, ABN^{+}09, Goe02, McM09, Prp02, And01, BMFG03, BE04, BMD09, MB06, TBK06, VBBF08]. genealogy [Zsi08]. General [Fuj06, Kra08, MN03, yTlC05, VWJ09]. Generalized [SKM07, Bur05b, DW09]. Generating [Lar02]. generation [GTSD09, SLI09, SG03]. Generational [Prp05]. genes [Bra09]. genetic [cL07, Lo08]. genetically [Lew07]. Genetics [GK00, CB08b, HSSK07, KGK02, Nis02]. genius [Egg07]. genomics [Bra09]. geodesic [Kre04]. geographic [HT09, VDL03, ZRRBL03]. Geographical [Fra05, LZ02, PRdKB07]. geologists [RRMSG02]. Geology [MSRRGG02]. geometric [GG09a]. geostatistical [HMG09]. geostatistics [HMG09, ZGHW07]. German [Bay05, CR08, HHWD04, Hor01, JLH07, KZ09, KLK01a, SS09a, SL05]. Germany [ERW05, LGS03, RdL02]. giant [ZTG09a]. glasses [ZRRB03b]. Global [Gu04a, KKL07, KLK01b, LDF^{+}09, YS08, Dai07, Das04, ERW05, GDM08, Jap07, LW09, Sch07b, VY06]. good [Ste04]. Google [BI08b, KT08, MDNL09, MF09, Smi08]. government [BL07, Coz06, HT09, Ley03, LM03, LM07, YZ08]. governmental [VLdM04]. graduate [LZA08, LLR04, YS08]. graduation [DMPM^{+}08]. Grail [VVM^{+}03]. Grams [Egg00a]. Granada [MM05]. grant [BMD07, MMLA01a, MMLA01b, Rei09, Rig09]. grant-holders [MMLA01a, MMLA01b]. granted [Ski06]. grants [JMB06]. Graph [Ley04b]. Graph- [Ley04b]. graphical [Bur02a]. graphs [BK02, GMA06, YR01]. Grasper [Sha01]. Great [SH08b]. greater [dB04]. Griffith [Chu01a, Chu01b, Moy01a, Moy01b]. ground [LW09]. grounded [Bal08, Pin06]. Group [SA00, ER04, MLR05]. grouping [SM06]. groups [AY09, CBVN06, Hem06, vR06]. Growth [GK00, HHWD04, KGS00, MA01, PT07, SK09, vR00, BS07, Dai07, Egg05b, FCTV04, GKSK02, IK02, LHHWD06, Mea01a, Mea01b, Sma06, YK04]. Gu [LF03]. Guest [MS04]. guided [SG03]. Gupta [San00].

H [Glä08a, KKK02]. H-index [Glä08a]. habits [MB06]. Hacettepe [TI02]. Hargens [Lin91]. headings [Pet08]. Health [CGC^{+}00, NM08, RPFD07, Sæt00, FH06, GR06a, GKB^{+}09, HFG^{+}05, LSGZMM^{+}08, Lew02a, LCLA02, MC02, SS05]. health-related [FH06]. healthcare [LSGZMM^{+}08]. hearted [Bon04]. Heinrich [ZM01]. Helix [BL07, DP03, GS07, Gok03, Ley03, LM03, LM07, MSU03, PHL05, Pri03]. Helmholtz [Goe05]. help [Lau03, ZRRB03b]. Henry [Sma01a, Sma01b]. herbal [tLtC07]. Hermann [Goe05]. Hermeneutics [Nal01b]. Herting [Lin91]. heterogeneous [GSD09, HHV03]. Heuristic [Egg00b, Egg01c]. HGF [Goe05]. Hierarchical [CSGGB09, CC08]. High [Abt00a, PFE00, Roy01, CB03, HCC03, JLH07, NCM09]. high-tech [HCC03]. Higher [OACS08, PSD^{+}07, BD08, DF07, ISS09, OLC07, PD07]. highest [LPC05]. Highly [AJOP04, Bas06, ZWZM04, AS04, But02, Egg05c, GST03, Kos07, LT08, LT09, Sma04b, TVvL02]. highly-rated [But02]. Hirsch [Ban06, BT09, BGS06, Bur07a, Bur09, CBVS07, ER06, GP09, IRN07, LR09b, Ran09, RT09, SKT09, SKM07, vR06]. Hirsch-index [CBVS07, ER06, vR06]. Hirsch-type [BGS06, LR09b, SKT09]. historical [Bru03, CB08b, RDPAC09]. History [AL09, WPZ^{+}04, OLVLH09, ZCY09, vPdlPvZ08]. hits [MF09]. HIV [OO09]. HIV/AIDS [OO09]. Hoax [Whi04]. Hoechst [DL06]. holders [BE04, MMLA01a, MMLA01b, Rig09]. Holy [VVM^{+}03]. home [Cri06, JT08]. Homeland [Gor05a]. homeopathy [CH05]. homonyms [WWJLG06]. Hong [CCHC07]. Honour [Moe00b, San00]. horizon [GST03]. Horticulture [LRR09]. hot [TLL^{+}09]. Houses [HG02, BD05d]. Howard [Edi05, McC05]. hubs [HS06, Lar08a]. human [CB08b, HSSK07, Ski09]. Humanities [Ned06, Wen01, AVGC^{+}06, AUQ09, Eto08, KKL06, LGA06]. hundred [DR01]. Hungarian [ISS09]. hybrid [JGD08, MSU03]. hydrogen [KNM07, Tsa08]. hype [Mey07]. Hyperlink [YZ08, TT03]. hyperlinks [KKK07, VGK06]. Hypertext [De 00, Tsa09b]. Hypothesis [SG03, AGJ08, GBZAEC^{+}07]. Hypothetical [ABZMD^{+}07].

Iberian [AC04, Cal06]. Iceberg [GBZAEC^{+}07]. Iceland [Lew01]. ICIAM [AET09]. ICT [But08, RG06]. Identification [AVDT04, LFB^{+}06, YL07]. identify [Boy09, CBVN06, Lee08a]. Identifying [But02, Chr04, HSSK07, Joh09, MBPKT03]. identities [Whi04]. Identity [CS02, BN06]. Identity-creators [CS02]. idiosyncrasy [JR06]. idiosyncratic [EL09]. IFPA [SMYS05]. II [AJOP04]. III [IMLSSC09]. illustrated [GTSD09]. illustration [RT09]. image [BN06, CS02]. image-makers [CS02]. immunology [DR01, LL03]. Impact [BA01, CG01, CEM06, Ing00, KM04b, LSGZMM^{+}08, Mai06, Pis09, QCW04, RvLvR02, SASA08, Vin02b, Wei05, Ada05, AGM07, ABZMD^{+}07, AGJC09, AJ08b, AL09, AJR09, BFG02, BDZZ07b, BCZM^{+}09, CB09, CG06, CGR06, CM09, CDFB05, CMU09, CSW^{+}09, Coz06, Dan06, DW09, Egg08a, Fre04, GG09a, GGJ^{+}05, GS01, GM02, Glä06, GJT09, Glä09, GSMT09, HBK^{+}08, Ing02, ISS09, IGK07, KKS^{+}07, KP08, LZA08, Lar08a, LV09, LVCS03, LPC05, LC06, Llu05, LCWD09, LFB^{+}06, MBR^{+}08, MC09, MB06, NLM01, Ned08, PM09, PRdKB07, Pod05, Pou05, PKK09, RRZR09, Rou05a, Van09, Vin09, Wal06, WPZ^{+}04, YWY05, YGY06, YW07, YW04, dMR01, vLM02, vLM05, Big08, RGD00, ST02b, SMP04, SMYS05]. Impacts [JT08, TV06, ZL07].

Impala [VCP00]. Impetus [KLK01a]. implementation [SM06]. Implementing [JD05]. Implications [LZA08, Koe01a, MT07, SKB08]. imply [BT09, BA07]. Import [GBZAEC^{+}07]. importance [YNSS09, vDH05]. Important [Abt00a]. improvement [Lew02a]. Improving [BN01, PKL02, SH08b]. IMS [AET09]. IMU [AET09]. in-depth [LLSL07]. Inadvertent [Wei05]. inbound [ST04]. inbound/outbound [ST04]. incentive [MHGGVJ08]. incentives [BGS07, SS09a]. including [YNSS09]. inclusion [WWJLG06]. increase [AGJC09]. Increasing [Per08, VKVF09]. Incremental [ISS09]. Index [Ano01f, Ano01g, Ano06b, Ano07a, Ano07b, CLH07, Kum06, SMB^{+}07, Vin09, YO06, AHK08, Ban06, BI08a, BI08b, BT09, BD05b, BM09, BGS06, BDZZ07b, Bur07a, CB08a, CBVS07, DMPM^{+}08, ER06, Egg06, Egg09a, GP09, Glä06, Glä08a, Glä08b, Hir05, IP07, IRN07, Lia06, Mar09, MHM07, MM08b, Ran09, Saa06, Sch08, Sch09, SKM07, SMYS05, Tol09, Woh09, Ye09, ZK08, vPdlPvZ08, vR06, CG06, CIM07, GKL06, Hav01, KKSK06, KK04, KT08, Ley04b, Ley07, Moe02, OW00, PL08, Sig00, VMT^{+}01, WM04, ZWW09]. Indexed [Ste01, ABZMD^{+}07]. indexes [Cho08, OLBB^{+}08, Per08, VS08, RvHMJ05]. Indexing [Ban06, SV01]. India [AB01, Aru01, BA01, Bha04, Gar02, GP02, Goe02, GD03, Gup06, GD08, GD09a, GD09b, KGS00, KG05, Nag02, PC05, RKM01, RK04]. Indian [KLK01a, AU07, BK00, BA07, GKL06, HG02, JG02, KT01, LK08]. indicate [ACYYA06, ACdLPM09]. Indicator [Sol06, AY09, Bas06, BDZZ07a, GSM05, GDP03, Hor01, KB08, Ley03, LH08, LRS06, SS08, VSL09, YAH08, vPdlPvZ08]. indicator-based [BDZZ07a]. Indicators [Ack06, AOS00, BGG00, BI00, CBVS07, GL01, Gus00, JZCZ02, Ley02, Moe00a, RdL00, AS04, AMTE07, AU07, BMJARBC05a, BDS06, BD08, BMFG03, BSPKN03a, BSPKN03b, CMJSOC09, GT04b, GTSD09, GGP05, GSH01, JLH07, JRC09, Kum09, LSGZMM^{+}08, LJL08, Lom08, MSU03, MT07, MBG01, OACS08, PHL05, PGD04, PA08, PID06, RRZR09, RdL02, RCE^{+}09, Rou05a, SH04, SS05, TG06, TG09, Tho01, Tij07, Vin01, Vin03, Vin06, Vin08, WD05, YGL06, ZIJSCMGG07, ZGL09, ZRRB03b, vR06, Hin03]. indices [AJR09, Dai07, DW09, Egg08b, GG09a, LRR09, LR09b, ML05, PM09, RT08, Sch07a, SKT09]. Indirect [AR06, Lee09]. individual [AJR09, CB08a, Hir05, MFG07, SB09, ZGL09]. individuals [Ned08]. Industrial [AET09, Ben03, BA07, Kao09b, CLH07, Fre04, sKtP04, TV06, VDC^{+}06]. industrialised [DGAVK03]. industries [AMTE07, CCHF05, MM02]. industry [BL07, CvLT07, CGV07, GS03b, KM04b, Ley03, LM03, LM07, LTL^{+}06, MHGGVJ08, McM09, OLC07, Ste09d]. industry-science [CGV07]. industry-university [OLC07]. industy [CDR06]. inequality [CB03, YS08]. inflate [GT04a, ZK08]. inflated [LLSL07]. Inflationary [PGD04]. inflow [Son03]. inflow/outflow [Son03]. Influence [AJR09, Luk06, ABZMD^{+}07, BL07, Big08, BMD09, BCZM^{+}09, CC09, Egg04b, Fox04, GT04b, Hem06, Igi02, Inö03, Joh09, RCE^{+}09, Sch08, TG06, Tsa09b, YWY05, YGY06, YGL06, vCvCvU08, Ama08]. influenced [PRdKB07, YA09]. influences [Chu01a, Chu01b, ÖSAT08]. influential [HBK^{+}08, VH01]. infometric [BI01a]. informal [MC09]. Information [CG01, DCF00, Gu04b, GH05, Koe01a, Lar02, Lee08a, LT09, MH01, PGFRD09, Yan09, vUS00, Ack06, AVDT04, AY09, BT06, CMP03, Eto03, GDTS06, JGD08, Koe01b, Koe01c, KT06, LFAH04, Ley03, LL08, LCWD09, MDNL09, Mea01a, Mea01b, MF09, Old01a, Old01b, PD07, PSD^{+}07, SLB00, Sap07, Uzu04, Vau04, VW04, VM06, WD05, YY08a]. Informed [Moe08]. Informetric [BRV00, HW03, ND02, BI08a, Bur05b, Bur06a, ER06, Gu04b, KS02, Rou02, Wol06]. Informetrics [HW01, Moe00b, BIP09, Bur07b, Egg04a, Egg05b, GM09a, Uzu02]. Infrastructure [Ver00]. Ingwersen [Edi05, Rou05b]. inlinks [LTMW03, TT04]. Innovation [AMTE07, CH00, SPN00, ZIJSCMGG07, CDR06, CLH07, Dai09, FG06, JM07, KKM07, MMS05, MSM06, MS05, Mey02, Mey07, Mus06, PHL05, PB09, RMTM07, SP08, Uzu06, VDC^{+}06, VL01, WH07, YM08, vdP07]. innovations [Bru03]. innovative [KP08]. Inometric [TGJd01]. input [Eto03, MB01, VSL09]. inserted [ACdLPM09]. insight [ZRRBL03]. Insights [Ste09d]. insignificant [Egg07]. Institut [SLB00]. Institute [JG02, AET09, Lee03]. institutes [CR08, Coc05, KZ09, Lee08b, TG08, TG09, Vin06, GDD03]. Institutional [BA01, CH00, DMPM^{+}08, Hen09, ÖSAT08, AJR09, Bay05, GMA06, Gar03, GTSD09, Hu07, KB08, LGT06, MMAHS08]. institutionalization [Sch07b]. institutions [AMBB^{+}05, Bal06, BMT09, BI04, BD08, BPMK06, ISS09, Jud04, MM08b, MM08c, Nag02, PMO04, Uzu02, YMB09]. instruments [Rig09]. integrated [GM09b]. Integrating [GL01]. Integration [CB08b, vdB00, MSGGRR08, RRGGMS07]. integrators [Mey06a]. Intellectual [Chu01a, Chu01b, CBO^{+}03, DR08, DCF00, LK00, BKM03, Dai09, Liu05, PL08, Pon02, TK09]. intelligence [VY08]. intended [Gar04]. intensity [VMD07]. Inter [ST08, YO06, BS07, Dai07, sKtP04, LZ02]. inter-and [BS07]. inter-correlated [Dai07]. Inter-country [ST08]. inter-industrial [sKtP04]. inter-regional [LZ02]. Inter-sectoral [YO06]. Interaction [Dan00, Jan00, Jia08, Mey06b]. Interactions [Mey00a, DP03, OALL07, VZV^{+}03]. interactive [BN01]. Interdisciplinarity [LT09, Son03, BS03, Ley07, MBG01, PCRP07, Sch04]. Interdisciplinary [Eto03, KLK01b, BMT09, Fuj02, Pon02, PR09, RVB^{+}01, RvLvR02, SG01a, SW01]. interests [BCK06]. interface [BM03]. interfaces [Ley07]. Interlinking [TS02, LTWM05a, LTWM05b, MBPKT03]. intermediaries [GT01]. internal [AVDT04, LK08, yTlC05, VWJ09]. International [BK00, BA01, Gan09, GM09a, GOS08, GSH01, He09b, Hin03, Ing00, MMP06, Moe00b, NLD07, RR02, Sap07, Ste01, VVPCOL09, YS08, ZBO00, BGB03, BSPKN03a, BCGMBSVM07, CCHC07, DR05, DA08, FJL03, Gar03, GS01, Glä01, GdL02, GH05, GD09a, HLT09, ISS09, Jap07, JJ04a, JRS06, JT08, Kim05, Kim06, KS01, Kor01, LJFD01, LC02, LTWM05a, LTWM05b, MBC06, Nis02, RCWE07, SS08, ST02a, SKMI06, SG01b, Ste09b, SBSCOR06, TT04, TVvL02, VMT^{+}01, Wag05, ZRRB03b, AET09]. Internationalisation [RDPAC09, BN06]. internationality [BCPTC06, GS05, Kao09b, Uzu04]. Internationalization [DZ03, WHC07, ERW05, HL09, WWW07]. internet [TGJd01, vR01, Bar06, CSW^{+}09, DSBH09, MSS06, YDM^{+}05]. interpretation [JJ04b]. Intra [MP00, MLNS08]. Intra-Departmental [MP00]. Intra-EU [MLNS08]. intricacies [PS02]. Introducing [Vin09]. Introduction [Ano01h, Hin03, LM07, McC01, Moe01, Noy06, Oku00, SF00]. introductions [HPF03]. invariant [KS09, LV09]. invasions [QC09]. inventing [Mey06a]. inventions [Bal06]. Inventive [MDTU05, Bha04, Ste09c]. inventor [CGV07]. inventor-author [CGV07]. inventors [BGS07, MSM06, MSU03]. investigate [MSM06]. Investigating [GT01]. investigation [BM09, SASA08, SH09, Ye09, YR01]. investment [LpCY06, She08, VD06]. invisible [Gmü03]. IP [MT07]. Iran [MHSE04, MMR05, YHM05]. Iranian [AKOK09, HME07, HE09a, OW00]. Ireland [RT08]. Irreproducibility [Flo07]. ISI [Bas06, Bay05, CCHC06, CIM07, IP07, JR06, MMP06, RvHMJ05, SMP04, YGL06, ZRRB03b]. Islamic [MHSE04]. Israel [BI04, ZGBI09]. Israeli [Gok03]. ISSI [ABGAMD^{+}09, Ano03d, Dav03]. Issue [Ano01a, Wor00b, LM07]. Issued [Dai07, LAFP04]. Issues [vZvPdlPH06, vdP07, HG02, RRZR09]. Italian [ADP08, ADC09b, BGS07, BDS06, RRZ06, Tuz05, UC03, Wra09b]. Italy [CDR06, MSS06]. item [BMFT05, Egg04c, ZRRB03a]. items [BV06, CM09, Egg05c, Fra08, Per06].

JAERI [YT09]. Jain [San00]. James [Lew04]. Japan [HL09, cL07, YO06]. Japanese [FG06, Hay03, HT06, LL08, MM02, NSS04, SGT^{+}06b]. JCR [LRR09, MMP06]. John [Ano08b, Ano09, Edi05, Ley04a, SSMV06]. Joint [PZ03, GT01]. Journal [Ano01a, BRV06, BN06, BDZZ07a, CG06, DCF00, GM02, Ioa06, Kra07, Ley02, MA07, Rou08, SM00, SMP04, TMVAR00, yTXwW03b, Ack06, AQ03, AGJC09, AL09, ADC08, BMT09, BGB03, BD05d, BCZM^{+}09, But08, CB09, CM09, CMU09, EP03, Gar03, Glä09, HH90, Har09b, Kor01, Kos07, LR09a, Liu05, LRR09, LINV^{+}09, McC05, NLM01, PH09, Saa06, SG01a, Van09, Wal06, Whi08, Yit02, YGY06, dMACRVQ^{+}07, vCvCvU08, vLM02, vLM05, ABZMD^{+}07, Ley04b, Pis09, SIMM08, SASA08, SG01b, CR08, Yan09]. journal-based [BMT09]. Journals [AGJC09, Dan00, Eto00a, Sig00, SM00, yTlC05, vdB00, ABZMD^{+}07, ABGAMD^{+}09, Ama08, AJ08b, BS01, BGS01a, BGS01b, BD05d, BGS06, BDZZ07b, BZDZ07, BCPTC06, But02, CB09, CG06, CGR06, CM05, CRLMRPA08, CEM06, DA08, DLM06, FCB02, Fra08, GG09b, Gu03, He03, HL09, Igi02, JJ04a, KLC^{+}08, Kao09a, Kao09b, KKL06, KK04, LPC05, Lew02b, LP04, LZ07a, Ley08, LZ03, LF03, Llu05, LS05, MA01, Mai06, MBR^{+}08, MBC06, MMP06, MN03, MM08a, Muk09, NBR01, NLM01, Nel06, Nis02, PL08, PL09, Pis09, Pou05, RR02, Rou02, RCWE07, Saa06, SIMM08, SFG09, SRS^{+}02, SMP04, SH09, SG01b, SK01a, TM03, Tsa09a, UC03, Uzu04, Vin09, VWJ09, Wal06, WWW07, Wu09, YW07, YW04, ZGL09, ZL07, dMR01]. judged [LVCS03]. judgment [Mai06, vR06]. judgments [Moe08]. June [Ano01c]. just [DD04, LFB^{+}06].

K. [San00]. Karl [ZM01]. Katherine [Ano08b, Whi08]. keep [Hud07]. Key [JRS06, CCHF05, DG04]. keywords [HLSA07]. Kharagpur [JG02]. kinetic [YS04]. King [Bon03]. KIST [LLSL07]. Klaus [Bon05a]. know [Mey07]. Knowledge [Lar08a, LAL07, Mey06a, PL09, Pon02, SPN00, SP08, WH07, AP04, BPMK06, CH04, CMP03, DL06, GFNG08, Gan09, Gu04a, Gu04b, GBZAEC^{+}07, Has05, Hem06, JJ04b, sKtP04, MVH^{+}05, Mey02, Mey06b, MP06, PHL05, PK09, PB09, Prp07, RdL02, RMTM07, RDvT03, RVB^{+}01, RvLB^{+}02, SLP06, Son03, SG01b, Ste04, Ste09c, SK09, TNK^{+}06, TGJd01, YS08, ZRRBL03]. knowledge-based [PHL05]. knowledge-driven [Mey06b]. known [Hud07]. Kocher [Rou02]. Kong [CCHC07]. Korea [CH00, CMP03, CYKK09, Kap04, Kim01, Kim07, KP08, KBL08, LOK00, LJ08, LZ05, cL07, OLC07, PHL05, PL08, SP08, SLP06, Son03, YY08a]. Korean [Kim02, sKtP04, Kim05, Lee08b, PL08]. Koshland [LR09a]. Kosmulski [Bur09, DW09]. Kosmulski-indices [DW09]. Kuwaiti [AQ03].

labeled [CG06]. Laboratory [LM00]. Labs [JRS06, Coc09]. Lack [KS02, Rou02]. lag [DLM06]. Lagergren [YS04]. Lancet [Kos07]. landscape [ZGBI09, ZRRB03b]. Language [ADC08, LZ07b, VMT^{+}01, Gu03, LZ03, WWW07]. languages [AUQ09]. Large [BM03, Ver00, KB06, MAVQHS^{+}04, SGT^{+}06b]. largest [LPC05]. Laser [GP00, GP01, Gar02, GP02]. Last [Hav01]. late [GST03, Gok03]. late-developing [Gok03]. latecomers [JLC09]. latent [BMD09, SKM07]. later [DR01]. Latin [CRLMRPA08, GK09, KS01, MC02]. Lattice [RB06, ER02b]. Lattice-based [RB06]. Lattices [YR01]. Laureates [Cam09, DFRG02, KKKM05]. Law [CJ06, Pet08, Egg05a, JSJX02, KSB03, vUS00]. laws [BMJARBC05b, Rou02, SK01a]. lead [DF07]. leadership [Hem06, KB08, SH04]. leading [AMBB^{+}05, CM05, Pis09, SG01a]. learn [OO09]. learned [SGBB03]. least [AGJC09, CM09, DGAVK03]. lectures [UKKP04]. Legal [VVPCOL09]. legislation [IGK07]. length [Yit02]. Lessons [Koe01b, Koe01c, SGBB03, Gok03]. Letter [AJOP05, Ano03a, Ano04b, Ano04c, Egg10, KKKM06, Xek05, Egg01a, Egg01b]. Level [BPA00, ACYYA06, Bas06, BMFG03, CB08a, GLT06, LMJS06, LP04, NCM09, SB09]. Levels [De 00, Saa06]. Leydesdorff [Ley04a, Rou04]. Liang [Bur07a]. Liberman [RR06]. libraries [BT06]. Library [LT09, KT06, LR09b, Sap07, YY08a]. licensability [Lee08b]. Life [Fox04, Hav01, LPC05, Luw00, Bal09, Jon09, LV09, Lee08b]. lifespan [BIP09]. light [Bon04]. light-hearted [Bon04]. like [Gok03]. Likelihood [ST02a, ST04]. likes [Sid06]. limitations [BFG02, Ran09]. limits [AVGC^{+}06]. Limnology [MBC06]. linear [CSGGB09, Egg05b]. Linguistic [TTP03, TT03]. Linguistics [NLM01]. link [BI04, Bjö06, CYKK09, KKM07, QLLR08, TZ08, VY06, VY08]. linkage [PL09, VDL^{+}02]. Linkages [CH00, MG00, TBvL00, BA07, GH07, RRR07]. linked [Jud04, VKG07]. Linking [VDL^{+}02, BIE05, Gei05, Jud04, Lom08, TT03]. Links [VW04, CGV07, Mey06a, TBK06]. LIS [DA08, KL09, Muk09]. listed [MSM06, SG01b]. Lists [ZL00, BSPKN03a, BSPKN03b]. literacy [PD07]. literary [AUQ09]. Literature [Aru01, GK00, HW01, Ioa06, Ton00, TJM00, Ack05, Ack06, GM03, GSST06, GJT09, GOS08, Gor05a, Gor05b, GM07, GKSK02, Gup06, HW02, Joh09, KKSK06, KU05, MCMN02, OO09, PM06, PA08, PKL02, RWGA07, RWD^{+}09, Tij07, TI02, yTXwW03a, yTXwW03b, TM03, Tsa08, YRL03, ZWW09, vCvCvU08]. literatures [He03]. Little [Koe01a]. Liu [vR05c]. LNCS [HG08]. Local [HT09]. Locating [Jia08]. location [ZRRBL03]. Loet [Ley04a, Rou04]. logic [HWD01a]. logistic [CEM06]. long [CEM06, LGS03]. long-term [CEM06]. longitudinal [HB07, PT07]. look [BND08, MB01, VS08]. losing [LW09]. Lost [BBG07]. Lotka [Bur06b, Egg04c, Egg05a, Pet08]. Lotkaian [Egg09b]. Low [KKM07, ACdLPM09]. LUC [RS00].

M [San00]. machinery [SGT^{+}06b]. machines [RRR07]. macro [Aks03, GT04b, GLT06, GSH01]. macro-indicators [GSH01]. macro-level [GLT06]. made [Whi01a, Whi01b]. Madeira [Wet06]. Madrid [IMLSSC09]. magic [Cro04]. Magnet [Gok03]. Magnetic [PC00]. Mailing [ZL00]. Mainland [TT03, ZWW09]. mainstream [NBRA^{+}02, PZ03]. Major [LZ02, BPMK06, BD05d, GM07, LLR04, Uzu04]. make [LINV^{+}09]. makers [CS02]. makes [HBK^{+}08, VH01]. Making [Gus00, DG04, VM06]. Management [Dan00, Eto02, Moe00a, PFE00, Pon02, DFRG02, Gan09, Gu04a, Gu04b, JM07, KLC^{+}08, KP09, MT07, OLC07]. managers [OALL07]. Mandelbrot [Izs06]. manipulated [KS09]. manipulation [YW07]. manufacturers [SGT^{+}06b]. Manufacturing [Ben03, sKtP04]. manuscript [Lin91]. manuscripts [BND08, YA09]. map [HCC03, MVH^{+}05]. Mapping [BLZ07, Bha04, BvdS06, BPMK06, BKB05, BBK09, GK09, HG02, HHV03, LJ08, Ley07, MP00, SW01, SA07, YMB09, ZV09, vdBH06, AJ08a, CC08, Iga08, JGD08, Jar05, LFAH04, LDF^{+}09, Noy01, NCM09, PFL01, PR09, Pou07b, Pou09, SB09]. Maps [Cah00a, OACS08, Boy09, BN01, BNVL06, KB06, MAVQHS^{+}04, PFL01]. Marginal [WD01b]. Markers [DCF00]. Market [LdLM00, CC09, Dai07, Dai09, Tho01, VY06]. Markov [BMD09]. marks [Bal09]. Martin [Rou02]. mass [Lew07]. material [CG06, RJDD07, YT09]. Materials [Ben03, MB06, BPM^{+}09, MBR^{+}08, SMB^{+}07]. Mathematical [AET09, MM08b, YRL03, DZ03, Egg08a, Eto08, Glä06, HWD01a, Ye09]. Mathematics [AET09, Aru01, BSPKN03a, DF07, ZL07]. matrix [Lia06, SLI09]. matter [GD08, MFG07]. Matthew [BMJARB^{+}07, Big08, BS01, Bon05b, Bro04]. Matthias [Rou02]. May [Ano01b]. McCain [Ano08b, Whi08]. mean [CB09]. meaning [LH06]. Meanings [Mar01, NDNZ01]. Measure [De 00, Gra01, ER04, GR06a, Glä06, LF09, MS05, MM08a, NR03, PRdKB07, RT09, Ste09a]. measured [AJ08b, MF09]. Measurement [VCP00, ADP08, Gei05]. measures [BFG02, CR08, Coz06, DW09, GLM^{+}07, GdL02, GM02, Lar08a, Nis02, RvLvR02, RZ08, SB09, TGJd01, ZRRB05, vLM02]. Measuring [BCPTC06, Bur06a, CGV07, CB03, Hul07, Kim06, LH06, Mey06b, Moe02, PCRP07, RvLB^{+}02, RCE^{+}09, TV06, YW04, LTL^{+}06, vZvPdlPH06, vdP07, PR09]. mechanical [Kim02]. Mechanics [CR08]. Medal [Ano08b, Ano09, Edi05, Ley04a]. media [Dai07, Dai09, Lew07]. Median [AY09, Rou05a]. mediated [TT03]. medical [ABZMD^{+}07, AKOK09, Kos07, OMD01, OALL07, RMTM07, RWD^{+}09, AQ03]. Medicine [PH09, Sch00, VVPCOL09, Bal09, BSPKN03b, CCDJJF04, DD09, KU05, LC06, tLtC07, Rei09, yTlC05, VWJ09, TI02]. Medicine-Related [Sch00]. MEDLINE [Luk06, XCS03, CCHC06, KT01]. MEDLINE-based [XCS03]. meeting [MDMBA^{+}06]. members [CGR06, OALL07]. Memorial [Sti04, Ano01a]. memories [vR05b]. mental [GKB^{+}09]. Merck [DL06]. mergers [KM04b]. merging [GFNG08]. merit [BSPKN03a, BSPKN03b]. Merton [Bon03, Col04, Cro04, Fox04, Gar04, Sma04a, Ste04, Sti04]. Mertonian [Har04]. meso [TG06]. meso-indicators [TG06]. Messages [ZL00]. Meta [NRTD00]. Meta-Analysis [NRTD00]. Metallurgy [BPM^{+}09, SMB^{+}07]. method [BPM^{+}09, CDFB05, KS09, MBPKT03, SLP06, SMP04, WWJLG06, vdBH06]. Methodical [GL01]. Methodological [GTSD09, GL01, Gmü03, Pri03, vR05a]. methodologies [MM08a]. Methodology [Eto00b, HSSK07, JCTSM^{+}09, ADP08, AGM07, ADC08, BMT09, CGV07, CC08, MM08c]. Methods [Arc02, BN06, SF00, AGM^{+}01, BI01a, BDS06, Bur02a, GL05, GLM^{+}08, HYA08, Jar05, Lau03, MN03, TLL^{+}09, TGJd01, YHM05, ZGHW07, vR05a]. Metric [PSD^{+}07, DMPM^{+}08, dMACRVQ^{+}07]. Metrical [Nal01b]. Metrics [De 00, Gei05, Kos01, KKL07]. metropolitan [MML^{+}04]. Mexican [CRLMR04, LCLA02, SFGRP09]. México [MLR05]. Michel [Ano09]. Micro [BPA00, BMFG03]. Microbiological [SA00]. Microbiology [AOS00]. microdata [BD08]. microscopic [BI04]. migration [LK08]. millennium [GSB02]. mine [GSD09]. Mini [Bjö06]. mining [FPGBD03]. minority [AUQ09]. misconduct [BND08]. misuse [CMJSOC09]. mobility [HLSA07, JT08]. Model [AGS07, BRV00, Pri03, BMJARBC05c, BMD07, Coc05, CEM06, DLM06, ER06, FCTV04, Glä04, Gok03, GM09b, Izs06, LVCS03, LLSL07, Nag03, San09, SK01b, Van01b, YRL03, dMR01, BMJARBC05a]. Modeling [EL09, GKSK02, Ygl03, BKK^{+}08, BMD09, GM09b, SR05, YM06]. Modelling [BO05, Bur02a, CMM04, Egg08b, GK00, ID04, Bur01a, Bur05b, YW04]. Models [NK07, BMD09, CSGGB09, Ye09]. Modern [Roy01, DL06, Req05]. modification [QLLR08]. modified [Lew07, SMP04]. Moed [ERS06]. Molecular [Lee03, MG05, HZT05, Jon09, RM07, RRR07]. Monarch [LH06]. money [San09]. Monitoring [DV03, SSG06, Ned06]. Moroccan [BM09]. Most [Ioa06, DGAVK03, ER02a, LT09, Ley08, Tol09]. Most-Cited [Ioa06]. most-recent-reference [ER02a]. motivation [Kap04]. Motivational [Gol02]. Motivations [KT06, BGS07]. motors [RM07]. movies [FH06]. Moving [But08]. MS [LS05]. Multi [DW01, Lan01, Glä09, IMLSSC09, LMJS06, NR03, PFL01]. Multi-Authored [Lan01]. multi-authorship [IMLSSC09]. multi-dimensionality [Glä09]. Multi-Disciplinarity [DW01]. multi-level [LMJS06]. multi-objective [NR03]. multi-self-organizing-maps [PFL01]. Multidimensional [DDM00, GP09]. Multidisciplinarity [Sch04, BS07, HB07]. Multilateral [Gus00]. Multilevel [MMAHS08]. Multinational [Ste01, Abt07a, BM03, Egg99]. multinationality [GdL02]. Multinationals [Ver00]. multiple [Chr04]. Multivariate [FCB02, EP03, SJJ04, Xek05, ZGHW07]. museums [GK09]. music [ND02]. mutual [Ley03]. Mystery [Nal01a]. Myth [Ing00]. Myths [Bal08].

Nagpaul [San00]. Nalimov [Ano01b, Ano01c, Ano01a, CG01, Gar01b, Gra01, Roy01, SZ01]. Name [LLR04, Pet08, Sol07]. named [Lan01]. Naming [Sch06]. nano [BZDZ07, Eto03, HB07, Mey01, Mey06a]. nano-S&T [HB07]. nano-science [Mey01, Mey06a]. nano-technology [Eto03, Mey01]. nano-titled [BZDZ07]. Nanobiotechnology [TMKM09]. Nanoscale [Pou07b]. nanoscience [Ley08, LZ07b, Sch04]. nanosciences [BLZ07]. nanostructured [Lar08a]. nanotechnology [CBVN06, HSSK07, HM03, Hul07, Iga08, KKL07, KBL08, Ley08, Mey07, Sch04, Sch07b, TMKM09, WHC07, LZ07a]. nanotubes [KM07, LAL07]. National [CH00, Glä01, GS07, Goe05, LTWM05a, LTWM05b, Pri03, RV08, AS04, Bay05, BSPKN03a, BSPKN03b, CYKK09, Gar03, GDP03, HT06, JRS06, KB08, Kor01, LJ08, MMP06, ML05, SG06, She08, SFG09, SS05, TT04, VMT^{+}01, VDC^{+}06, WM04, vL09, MLR05, Uzu06]. nations [Kim07, Ygl03, YY08a]. Natural [Luw00, Wen01, LGA06, Ski09, SMW^{+}09]. nature [RMTM07, TM03]. Nazi [CM05]. Nazi-power [CM05]. necessarily [vdWFM^{+}09]. need [Har09b, PGD04]. Negative [Bal08, FH06]. Neglected [HH90]. neighboring [SIMM08]. neighbors [SS09b]. Nelson [Tsa09b]. Neo [Bos09, CA09, DGAVK03]. Neo-colonial [DGAVK03]. Neo-colonialism [Bos09]. Neo-creationists [CA09]. Neoplasms [KT01]. nepotism [SH08a, SH21]. ness [Ski06]. Netherlands [PHL05, Rin00]. Network [CvLT07, KLK01b, CA09, CC09, HB07, KP08, LAFP04, Lar02, Lar08a, LMJS06, MEG06, RRZ06, SM06, YS08, ZGL09]. Networks [GG09b, MP00, RMTM07, CR08, CCDJJF04, DSBH09, ER02b, EL09, FRJ08, GSD09, GDD03, GÖ09, HHV03, HS06, HKL08, KNM07, Kre04, KKK07, LGA06, LF09, Nag02, OLVLH09, PFL01, Rig09, RZ08, SKT09, SKM07, YR01, YK04, YNT06, YNSS09, ZV09, vR05b]. neural [CC09, CCDJJF04, LAFP04, PFL01]. neurology [YW04]. Neuroscience [XCS03, BGS01a, BGS01b, GDP03, SW01]. neutral [GSMT09]. never [GST03]. nevertheless [Bon04]. newsgroups [KSB03]. Newsletter [Ano03d, Dav03]. Nigeria [Nwa06]. No [BSPKN03b]. No-bells [BSPKN03b]. Nobel [UKKP04, Cam09, KKKM05, KM04a]. Nobelists [BSPKN03b]. noise [HD07]. Non [GP00, ADP08, ABZMD^{+}07, BV06, CVV^{+}06, ERS06, Fra08, HD07, MML^{+}04, MSS06, Mey06a, Old01a, Old01b, RRMSG02, Vin08]. non-authoritative [ERS06]. non-citable [Fra08]. non-consolidated [RRMSG02]. non-EU [Vin08]. non-indexed [ABZMD^{+}07]. non-inventing [Mey06a]. non-metropolitan [MML^{+}04]. non-patent [CVV^{+}06, HD07]. non-profit [MSS06]. non-source [BV06]. non-spatial-data [Old01a, Old01b]. nonparametric [BDS06]. Nordic [Wor00c]. norm [And01]. normalisation [ZRRB05]. normalized [GTSD09]. Normative [Cro04, Sma04a]. North [Ing02, Jon09]. Norway [IGK07, KNM07]. Norwegian [AOS00, Ive00, SA00]. Nota [De 05]. Note [RdL00, CDR06, MM08a, YM08]. noun [Sch06]. Novel [Mey00a, Bla09, ER04, Fra02, Mey01, Ste09c, WWJLG06]. Nuclear [Ack06]. Number [Ano01b, Ano01c, Nal01b, Abt07c, LPC05, MN03, PH09, Pod05, RR06, YM08]. numbers [BK02, CM09, LH05].

obesity [AGS07]. Obituary [Ano03e, Bon03, SV09]. Objective [NN08, NR03]. objectives [MA07]. Observations [IK02, Bur03, Rig09, vR01]. observed [YL07]. obsolescence [Bur02b, Egg04b]. obstacles [BGS07]. Obtain [RdL00, Bas06]. obtained [CB08b]. Occasion [Moe00b]. occurrences [vdBH06]. Ocean [Das04, DR05, ZQH09]. oceanography [LL03]. OECD [Cri06, YY08a]. off [VZV^{+}03]. OLAP [HV03]. old [AJOP04, OCA09]. One [BMFG03, DR01]. online [Ama08, MSW01, Whi01a, Whi01b]. only [Abt07c, GST03]. Ootprint [SHO00]. Open [Cho08, Har09a, JRS06, LOK00, SH08b, CvPdlPR06, KT06, Muk09, SH09, WH07]. Operation [Kao09a, Ste01, SBSCOR06]. Operational [Eto00a, Eto02, Fuj06]. Ophthalmology [DW01, UCCM01, ZWW09]. opinion [CBO^{+}03, MMLA01a, MMLA01b]. Opportunities [HW03, Wra09a]. Optics [TK09]. optometry [ZWW09]. OR/MS [LS05]. order [GBGCRFdMA09, LLR04]. Ordering [VCP00]. ordinary [Egg07]. Organic [KGS00, Kum09, Kum06]. Organisations [RDPAC09, MFG07]. organisms [Lew07]. Organization [ZL00]. Organizational [CA09, LMJS06, Hem06]. organizations [HE09a, Lo08, OALL07]. organizing [PFL01]. Oriented [ZL00, Goe05, HCC06, MMP06]. origin [And01, BE04, LCWD09, PRdKB07, YA09]. origins [God06]. Ortega [BMJARB^{+}07]. other [BI08a, GM07, Kim07, Mar09, PS02, RRZR09, WXB09]. outbound [ST04]. Outcomes [ZWW09, HFG^{+}05]. outflow [Son03]. outflows [SFGRP09]. outlets [Har08]. outlinks [GSD09, ID04]. outlook [WPZ^{+}04]. Outperform [SS09b]. Output [AGG00a, De 00, RS00, AMBB^{+}05, AHK08, AVGC^{+}06, Bay05, BM03, Eto03, GYY02, GD08, GD09a, GD09b, Hir05, IK02, JG02, Lar08a, MJL^{+}09, MDTU05, MMR05, NBRA^{+}02, Per08, QMC08, RG06, Rou02, San09, She08, SL05, SK01a, VVPCOL09, WM04, XCS03, ZTG09b]. Output-Financing [RS00]. outputs [But08, Gei05, GH05]. overall [ADC09a, vCvCvU08]. overlapping [RB06]. overlaps [Lar02]. overseas [ST02a]. overview [CCDJJF04, FCTV04, Gar03, NBRA^{+}02]. own [ER04]. own-group [ER04].

P [GMA06]. P-FSG [GMA06]. P. [San00]. Pacific [TS02]. page [ID04]. PageRank [FRJ08]. pages [BIE05]. pain [RWD^{+}09]. Pair [Moe00b]. panel [JLC09]. Panorama [AWG00]. Paper [Ano04b, Big08, Bur01b, CB09, Gu03, Xek05]. Papers [Abt00a, BK00, Ioa06, Lan01, NSS04, Sig00, Ste01, TBvL00, WPZ^{+}04, Abt07a, Abt07b, Abt07c, AJOP04, AS04, Ama08, BS07, Egg07, Fuj02, GBGCRFdMA09, GD03, Hin03, Hu07, Lar02, Lew02b, PK09, PH09, Sma04b, TVvL02, YA09, ZWZM04]. paradigm [KM07]. Paradox [She08, HVG06, RRZR09]. Parameter [YL07]. parametric [ADP08, YM06]. parametric-bibliometric [ADP08]. parametrized [GMA06]. Paretian [BT09]. Parliament [HG02]. parsing [Sch06]. Part [AJOP04, GLT06, LGT06, SHO00, TMVAR00, ZRRB03a]. participating [GS03b]. particle [CRLMR04]. Partnership [YO06]. party [Hor01]. PASCAL [AWG00]. passive [MF09]. Past [Bea01, ZBO00, YNSS09]. Patent [BGG00, CLH07, DDM00, GH07, HCLH04, Ive00, Lee08b, LpCY06, cL07, Lo08, Mey00a, Mey00b, Mey01, MB01, RdL00, VD06, Ver00, ACF09, AR06, ACYYA06, ACAYY07, ACdLPM09, BGS07, BM03, Bha04, CVV^{+}06, CCHF05, CC09, Cri06, Dai07, Dai09, GG09b, Gup06, HD07, HCC03, Iga08, JLC09, LGS03, LFAH04, LLSL07, Lee09, LL08, MS05, PA08, RdL02, SLP06, Ste09a, Ste09b, SGT^{+}06b, Wan07, vPdlPvZ08, vZvPdlPH06]. patent-based [SLP06]. Patent-bibliometric [GH07]. Patent-Bibliometrics [Ive00]. patent-related [Dai07]. Patenting [BGG00, Ive00, Van08b, Bal06, BGS07, Bal08, IGK07, Lom08, MM02, McM09, Mey06a, MFG07, PMO04, Ste09d]. patentometric [PB09]. Patents [BM03, GM03, Ste09c, TBvL00, Arc02, Bru03, CMM04, Cri06, Dai07, HD07, HM03, JLC09, Lee08b, LZ07a, Ley08, MSM06, MSU03, MT07, PK09, RDPAC09, VMD07, VDL^{+}02, VDL03]. path [GM09b, Mey07]. path-dependency [Mey07]. Paths [DCB^{+}00, Bjö06]. pattern [SSM07]. Patterns [CMP03, EP03, Eto00a, JI00, JJ04a, KT01, LT08, PK09, SBSCOR06, TT04, YO06, AQ03, AVDT04, AU07, BT09, BGS01a, BGS01b, BZDZ07, Bru03, CRLMR04, CRLMRPA08, Dai07, Eto02, Far02, Jon09, Kim02, sKtP04, LT09, LGD02, LLR04, MLNS08, Nel06, Nwa06, Sch04, Soo09b, SH09, TTP03, Uzu04, VDL03, vL09]. Paulo [LPC05]. pay [FA07, VZV^{+}03]. PBRF [Smi08]. Peer [CGC^{+}00, Rei09, ABGAMD^{+}09, ADC08, BD05c, BD06, BMD07, BMD09, DGAVK03, Igi02, JMB06, Mai06, SH08a, VWJ09, vR06, SH21]. peer-review [SH08a, SH21]. peer-reviewed [ABGAMD^{+}09, ADC08, DGAVK03, Igi02]. peers [Mey06a]. Peking [ZWZM04]. pentachlorophenol [HYA08, YAH08]. People [GD03]. perceived [BL07]. percentage [MG04]. percentile [Rou05a]. perception [CMU09]. perceptions [Lew02b, SKB08, SASA08]. Percolation [BRV00]. Performance [BA01, CGC^{+}00, Coz06, Egg09b, PFE00, VCP00, AGP07, BFG02, Cal06, CC09, Coc05, Coc09, DMPM^{+}08, GAGPM06, GTSD09, GBFM09, GM04, GG08, HT06, Hu07, HR09, KP09, KP08, KO08, LMJS06, LV09, Lee03, LJL08, tLtC07, MSRRGG02, MSGGRR08, Moe02, NR03, Ned06, Ned08, PM09, RRGGMS06, RRGGMS07, SMYS05, Tho01, Tij07, VMT^{+}01, VDC^{+}06, ZRRB05]. Performances [CH00, NN08]. performer [Kos02]. Period [KK04, ORSCSB00, PFE00, Bon05a, CH05, HYA08, LR09a, QC09, Uzu06, YX08]. periodical [DLM06, YYL04, YYL05]. periods [LGS03]. peripheral [Gor05b, Kim06]. periphery [GDD03, NBR01]. Persistent [SH08a, SH21]. Personal [Pet08, Fuj02, LMJS06]. personality [HBK^{+}08]. personnel [WFI03]. Perspective [BI01b, PC00, CCHF05, Chu01a, Chu01b, CB08b, HB07, LW09, RRZ06, RDPAC09, Ste04]. Perspectives [MS01, BCPTC06, Tij07]. persuasion [Whi04]. Peter [SSMV06, Edi05, Rou05b, Ano09]. pharmaceutical [CvLT07, CC09, KM04b]. PhD [VBBF08, YA07]. phenomenon [Kap04]. philosophers [Kre01b, Kre01c]. Philosophical [Nal01a]. Philosophy [GW01, Nal01b, Gei05, Kre01b, Kre01c]. phrases [HW02, Sch06]. Phylogenies [CDFB05]. Physical [Abt00b, BMJARBC05a]. physicists [IK02, Kim02]. Physics [BPA00, RSJ00, Sig00, AGM^{+}01, Bal09, Bon05a, BSPKN03b, CM05, CRLMR04, GD08, GD09b, HME07, HWD01b, Kim01, RvLvR02, Tsa09a, KM04a, UKKP04, PH09]. pilot [BDZZ07a, GGP05, UKKP04]. Pioneers [ZM01]. place [CS02, Zsi08]. Plans [Kim07]. Platform [MG00]. PLS [GM09b]. Point [SMYS05, WD01a]. points [WD05]. Poland [Sap07]. policies [CMJSOC09, KKM07, dMR01]. Policy [Gus00, Jan00, Ned08, Rin00, Sæt00, DG04, GDTS06, JM07, Noy01, NCM09, Sch07b, Uzu06, VVM^{+}03, VM06, YHM05]. Polish [SZ01]. Political [LK00, CMJSOC09, Nis02, SKB08]. politico [WM04]. politico-economic [WM04]. pollution [CMM04]. Polymer [SMP04, BPM^{+}09, SMB^{+}07]. Polywater [Ack05, Ack06]. poorly [Kos07]. Population [AGG00b, BKK^{+}08, GK00, KGK02, RRZ06]. portal [PD07]. Porter [Ygl03]. portrait [SSMV06, SSM07]. Portuguese [MFG07]. posed [Bur01c]. posets [BK06]. positioning [BD08, SH09]. positions [VY06]. Positive [Egg04a, FH06]. Possible [Gra01, Nal01b, BCK06, GOS08]. Post [DMPM^{+}08, BD06, LR09a, MJL^{+}09, VWJ09]. post-doctoral [BD06]. Post-graduation [DMPM^{+}08]. post-Koshland [LR09a]. post-publication [VWJ09]. post-Soviet [MJL^{+}09]. posting [Ama08]. Potential [ZRRBL03, Boy09, Per06]. Power [Rou02, SK01a, CM05]. PPI [YO06]. practical [Izs06]. practice [Cro04, DSBH09, Ski06, Vin02b, AQ03]. practices [MT07]. practise [Egg06]. Prathap [AJR09]. Pre [LR09a]. Pre-Koshland [LR09a]. Precision [Lin91]. Predict [CGC^{+}00, Ski09]. Predicting [Wal06, Sma06, HBK^{+}08]. prediction [JRC09, LT08]. predictive [BD05c]. Preface [Ano04d, Ano05, GR06b, Glä08c, GM09a, Kre01a, OGYR06, VG06]. preference [ER04, SG06]. preliminary [BLZ07, ÖSAT08, Ste09b]. presence [LH05]. Present [Bea01]. presentations [ABGAMD^{+}09]. Presidency [DT07]. President [Ano03d]. pressure [HHWD04]. prestige [CMU09, RRGGMS07]. previously [BD06]. Price [Ano08b, Ano09, Edi05, Ley04a, McC05, Rou04, Rou05b, Whi08, Dai09, FCTV04]. Priest [Roy01]. primary [Lo08]. Principle [WD01b]. Principles [AQ03]. Prior [CVV^{+}06]. priority [Ben03, Ste04]. private [OALL07, Som05, VD06]. Prize [KM04a]. Probabilistic [GW01, YO06]. Probabilities [Egg07]. problem [ER04, Egg04b]. Problems [GL01, Bur01c, Vin02b, vR05a]. Proceeding [Sig00]. Proceedings [GSST06, Wor00c]. process [BMD09, Bur05b, GAGPM06, Lin91, YRL03, YYL05]. processes [PGFRD09]. processing [Wol06]. proclamation [Bon04]. Produce [Abt00a]. Producers [ZBO00, YS08]. Production [ORSCSB00, SPN00, AC04, BRSYB04, BMFG03, CMP03, CvPdlPR06, GSMT09, HE09a, HWD01a, HWD01b, Inö03, JJ04b, KBL08, NM08, PRPFC06, PB09, Prp07, RDvT03, WXB09]. productive [AP04, Egg05c, Saa06]. productivities [LTMW03]. Productivity [DA08, HB07, MB06, PC00, Prp00, SA00, Uzu02, ADP08, ADC09a, ADC09b, Bar06, BD03, BD05a, BDS06, DD04, Egg09b, Gan09, GLM^{+}07, Goe02, GKB^{+}09, HCTC03, HVG06, HLT09, JT08, KKKM05, KM04b, KS01, Kre04, KGK02, LVCS03, LZ02, LCWD09, MHGGVJ08, Nwa06, Prp02, Prp07, RRMSG02, SFGRP09, SMW^{+}09, TM03, VMD07, WFI03, YNSS09]. productivity-export [SMW^{+}09]. Prof. [SSMV06]. professional [BMFG03, DSBH09, Prp05, YS08]. professionalization [Wra09a]. professionals [BT06, OALL07]. professors [Wra09b, VM06]. Profile [Luw00, PV00, BD08, DR05, GKL06, PC05, Req05, TG09, WFI03]. Profiles [LM00, ZBO00, BMT09, GS07, LGT06, MG05, TG08, VM06]. Profiling [AGM07, CGV07, KL09, PKL02]. profit [MSS06]. profitability [VD06]. Program [Gok03]. programme [Goe05]. Programmes [Gus00, Hay03]. Programs [DMPM^{+}08]. progress [DSBH09]. project [Eto03, LAFP04]. Projects [AGG00a, CGC^{+}00]. Prolific [GP00, Tol09]. promotions [JRC09]. Proof [ERS06]. propensity [SGT^{+}06b, Van08b]. Properties [LR09b, EL09, RRZR09]. property [CBO^{+}03, Dai09]. property-based [Dai09]. Proposal [BR00, LOK00, AMTE07, Bur03, ER02c]. propositions [Mey07]. prospectively [LVCS03]. Prospects [RK04]. Provided [KLK01a, PGFRD09]. providing [SSG06]. provincial [YZ08]. Provision [vUS00, Ste04]. proximity [LZ02]. proxy [Ste09b]. pseudo [Nag03]. pseudo-regression [Nag03]. Psychiatric [DMPM^{+}08, LJFD01, MN03]. Psychiatry [Fuj06, Ing02, ZNK01, CSGGB09, FJL03, JLH07]. psychological [BBV^{+}08, KBWM07]. Psychology [CGR06, Goe02, HBK^{+}08, Kra08, Llu05, Wal06]. Public [NM08, RPFD07, RDPAC09, AP04, Bal06, BD05a, Coc05, Coc09, CBO^{+}03, KKM07, Lee08b, MH07, MFG07, OALL07, Ste04, VD06]. public/private [VD06]. Publication [BGS01a, BGS01b, Cal06, CRLMR04, CRLMRPA08, DD09, DLM06, GLM^{+}07, Har08, JI00, Ley00, Moe00a, Prp00, ZNK01, Abt07c, Ack06, Ama08, AHK08, BMT09, BT09, BND08, CBVN06, GL05, GLM^{+}08, Glä06, GJT09, Goe02, GD09a, Jon09, KKKM05, Kim02, KS01, Kra08, Lar08a, Lar08b, LH08, MLNS08, MB06, OMD01, She08, Soo09a, SL05, TG08, Uzu02, Van01b, VM06, VWJ09, YAH08, YYL04, YWY05, YYL05, YGY06, YGL06, YY08b, vR05b]. Publications [HM03, KZ09, MDMBA^{+}06, OW00, Pol00, ABGAMD^{+}09, BD06, BPM^{+}09, CvLT07, CCHC07, CR08, Dai07, Egg99, FJL03, Fra03, GT01, Goe05, Gor05a, GSH01, GYY02, HG08, KKS^{+}07, Kim01, KK04, KBWM07, LJFD01, LGD02, LLR04, LZ07b, Luk06, MG05, Moy01a, Moy01b, ÖSAT08, PKK09, PZ03, RRS07, RWGA06, SS08, Sch09, SMW^{+}09, Soo09b, Ste09c, TG04, VKVF09, Van08a, Vin06, ZL02]. publicly [GM09b]. publish [ACAYY07]. Published [Sig00, Bla09, BD06, PH09, SRS^{+}02, Wal06]. publisher [DF07]. Publishing [DW01, RCWE07, BD05d, DGAVK03, DLM06, Igi02, Kyv03, YRL03]. PubMed [VKVF09]. Puerto [ORSCSB00, SBSCOR06]. pulsing [BMJARB^{+}07]. purposes [BI01a, GS03a]. put [CS02]. pyramid [Zan06].

Qualitative [SH04]. Quality [CMU09, PBG^{+}05, RSJ00, ZNK01, ACF09, HE09a, Hud07, JLH07, Kos02, LFAH04, Lew01, Moe08, NLM01, PH09, Pin06, PGFRD09, Rig09, RT09, SV01, San09, SS08, SKMI06]. quality-scores [PH09]. quantify [Hir05]. Quantifying [GS03b]. Quantitative [FH06, KB06, VVPCOL09, BND08, BS03, HL09, LJL08, SH04, Ye07]. quantity [HE09a, Lew01, Moe08]. question [Bal09]. questions [HG02]. Quotation [Wet06].

R&D [AP04, CMJSOC09, Coz06, Dai09, DP03, GGJ^{+}05, Hay03, KP08, KNM07, LJ08, Lom08, ST08, VD06, YM06]. R. [Fox04]. race [SH04]. radiology [MDMBA^{+}06]. RAE [LTMW03]. raised [HG02]. Raman [PKK09]. Ramaseshan [SSM07]. Randić [Egg10]. random [BPM^{+}09]. rank [CB09, JSJX02]. rank-size [JSJX02]. ranked [BSPKN03a, BSPKN03b]. Ranking [KLC^{+}08, TSLCJC09, BD05b, Flo07, LT08, LCL05, Mar09, MM08b, MM08c, vR05a]. rankings [BCGMBSVM07, GL05, KH08, Mar09, Ned08, YGY06]. rate [Abt07c, Bur07a, YW07, YS04]. rated [But02]. rates [Hen09, Jud04, Kra07, KKK07, SG01b]. ratings [LTMW03, Uzu02]. ratio [Fra08, MS05]. Rational [RT08, Sol06, BK02]. rationality [Fra02]. RC [BMD07]. reactions [Mea01a, Mea01b, YS04]. reader [JMB06]. realization [MF09]. really [BD05b]. reasons [AJOP04, BND08]. Recipient [McC05, Rou04, Rou05b, Whi08]. recipients [BD05c]. reciprocating [AVDT04]. Recognition [LC02, JLH07, SH08b]. recommendations [HH06]. Reconsidering [FCTV04, SASA08]. Recorded [Sig00]. recovery [CM09]. recursive [WWJLG06]. Redistributed [BPM^{+}09]. reduced [DF07]. reductionism [BK06]. referee [HH06]. referees [BND08, HH90]. Reference [Abt00b, vR05b, ER02a, Kra08, LINV^{+}09, ND09, TVvL02, ZRRB03a, vdBH06]. Reference-based [vR05b]. References [Gup06, AY09, Big08, CVV^{+}06, ERS06, GS05, HD07, Kap04, KBWM07, MC09, MG04, PBI02, SG06, UKKP04, VDL^{+}02, Wu09]. referencing [ZRRB03a]. referencing-structure [ZRRB03a]. Reflect [Moe00a]. Reflected [Hav01, Ste01, CM05, Gar03, GD09a, PBI02, WWP^{+}05, WM04]. Reflection [KKK07]. Reflections [Bea01, BCPTC06, YHM05]. reflects [ADC08]. Reform [Eto00a]. reforms [MC02]. regard [GSH01]. region [AMBB^{+}05, ACdLPM09]. Regional [DP03, PC00, ZIJSCMGG07, ZTG09b, AMBB^{+}05, ABN^{+}09, ACYYA06, LZ02, Mai06, VL01]. regions [ACAYY07, LV09, Tuz05, ZRRBL03]. registered [Lee08b]. regression [Nag03]. regularities [KM04a]. reinforcement [Egg04a]. Rejecting [Cam09, BND08]. rejection [BD06]. Related [Sch00, Sig00, BK06, CHH04, CHH07, Dai07, FH06, Glä08b, IMLSSC09, KNM07, MSU03, ND09, Saa06, ZQH09]. relatedness [MA07]. relating [San09]. Relation [Abt00b, Eto00a, KBL08, Yit02, Egg99, Egg05b, Egg08a, Egg09b, Gor05a, VD06, WM04]. relational [BMJARBC05a, GSB02, GTSD09]. Relations [She08, Vin03, CGV07, Glä01, GM03, GK09, Ley03, LM03, LM07, MHGGVJ08]. Relationship [SMW^{+}09, CB03, GR06a, Kre01b, Kre01c, LTMW03, Nis02, PH09, Tho01, TM03, VMD07, Ye07]. relationships [LF09, PB09]. Relative [ER01, Hu07, GTSD09, HLT09, LPC05, Lew02b, PGD04, RRZR09, Vin03]. Relativity [ZRRB05]. Relevance [TBvL00, Chr04]. relevant [JD05, VDC^{+}06, Vin09]. reliability [Bur02a, ZCY09, BD05c, Rei09]. reliable [OMD01]. Remarks [RR06, Bur01b]. remediation [HYA08]. Reminiscences [Gar01b]. Renaissance [Wra09b]. renewals [vPdlPvZ08]. Renormalized [RGD00]. Reply [Ano04c, KS02, vR05c]. Report [AET09, SRS^{+}02, Ley04b]. reporting [Lew07, MN03]. reports [MB01, ABZMD^{+}07, Ley02, Pis09, SIMM08, SM00, SMP04, SG01b]. Reports-indexed [ABZMD^{+}07]. repositories [LZA08]. Representation [SIMM08]. representative [Kre01b, Kre01c]. represented [SG01a]. representing [ZCY09]. reprint [Dre03]. Republic [MHSE04, GD03, RT08, Van08b]. reputation [DD04]. requests [Dre03]. Rescher [WD01b]. Research [AGP07, Ama08, AGG00b, Aru01, Bar06, BPA00, BD05e, CvLT07, CGC^{+}00, Coc09, Dan00, De 00, Eto00a, Gei05, Goe05, GDD03, Gra01, Gus00, Har09a, HCC06, Ing00, KKS^{+}07, Kao09a, KGS00, KLK01b, KKK02, LOK00, Luw00, MG00, MML^{+}04, Moe08, Mus06, MH01, NRTD00, Oku00, PBG^{+}05, PKL02, PV00, Pri03, RRMSG02, Rin00, RDPAC09, San09, SA00, Tho00, TMVAR00, TBvL00, WD01b, ADP08, ADC09a, ADC09b, Ack05, AY09, AU07, ABN^{+}09, AB01, Bal06, BMT09, BBV^{+}08, Bay05, dB04, BD05a, BFG02, BPMK06, BD05c, BD06, Bos09, BM09, BBK09, Bra09, BS07, BPM^{+}09, CBVN06, Cal06, CHH04, CH05, CH07, CHH07, Coc05, CBO^{+}03, DZ03, Das04, DR05, DG04, ERW05, Eto02, FCB02, Fre04, Fuj02, Fuj06, GAGPM06, GR06a, GSM05, Gar02, GP02]. research [GKL06, GLM^{+}07, GT01, GM02, GSB02, GS07, GTSD09, GBFM09, Gor07, GS03b, Gu04a, GM04, GH05, GG08, GM09b, GBGCRFdMA09, GD08, GD09b, HFG^{+}05, HT06, HZT05, HHV03, HSSK07, Hem06, Hir05, HCTC03, Hor01, Hu07, HR09, IRN07, Ing02, IJ04, IK02, IMLSSC09, Jap07, Jar01, Jar05, JMB06, JG02, JRS06, JT08, Jon09, JM07, KKSK06, KP09, KZ09, Kim06, KL09, KS02, KM04b, Kos02, KKL07, KG05, Kum06, Kum09, KO08, LAB^{+}09, Lar08a, LSGZMM^{+}08, Lee03, Lee08a, Lee08b, LGT06, Lew02a, LP04, LTMW03, LH08, LDF^{+}09, LZ02, LCLA02, tLtC07, Llu05, cL07, Lo08, LCWD09, MNQ08, MHGGVJ08, MJL^{+}09, MSRRGG02, MSGGRR08, MHL06, MDTU05, MMAHS08, Moe02, MM08a, MMP09, MFG07, Muk09, NR03, Ned06, Ned08, OMD01, ÖSAT08, OLVLH09, PM09, PRdKB07, PC05, Per08]. research [PA08, PR09, Pou07b, Pou09, Prp07, QC09, RKM01, RDvT03, Req05, RRGGMS06, RRGGMS07, Rig09, RV08, RvLvR02, RRZ06, RWGA06, RWD^{+}09, RJDD07, Rou02, RT09, SSM07, SS08, SS09a, Sch04, Sch07b, SW01, SA07, She08, SKB08, SU09, Smi08, SMYS05, SKMI06, SMB^{+}07, SMW^{+}09, Som05, SASA08, Son03, SSG06, SK01a, TMKM09, TK09, TG08, TG09, TVvL02, TV06, Tij07, VMT^{+}01, VDC^{+}06, Van08a, VSL09, VM06, Vin06, Vin08, WD05, WL03, XZH08, YHM05, YMB09, YA09, YX08, YAH08, YY08a, YW04, ZQH09, ZL02, ZCY09, ZWZM04, Zsi08, ZV09, vL06, vR06, vdBH06, vdWFM^{+}09]. research-oriented [HCC06]. Researcher [Bas06, AJR09, PCRP07, Prp05, QMC08, SASA08, YNSS09]. Researchers [AGG00a, DW01, Hav01, Lew02b, Ton00, And01, BCK06, BL07, JLH07, JT08, LMJS06, Lew01, MSRRGG02, Mey06a, TSLCJC09, Woh09]. resistance [BK02]. resisting [Cam09]. Resonance [PC00]. resource [PD07, Rou08]. Resources [Cho08, BIE05, CSW^{+}09]. respect [Eto08]. respectively [Saa06]. Respiratory [CHH04]. Response [JCTSM^{+}09, Gor05a, LF03]. responsibility [GBGCRFdMA09]. Restructuring [HT06]. resulting [MDMBA^{+}06]. Results [BI00, KLK01a, MMLA01a, MMLA01b, Bra09, CB08b, Flo07, GLM^{+}08, HSSK07, Jas09]. Retirement [Moe00b]. Retrieval [DCF00, Lar02]. retrograde [SW01]. returning [JT08]. Returns [WD01b, GR06a]. Reveal [Aru01, BA07, PL08]. Revealed [Wor00b, EP03]. reverse [GM03]. Review [CGC^{+}00, Kha05, LOK00, Ned06, OB05, Rou06, BI01a, BD05c, BD06, BMD07, BND08, BMD09, GDTS06, JMB06, Lin91, PKL02, Rei09, SH08a, VWJ09, YS04, ZV09, SH21]. reviewed [ABGAMD^{+}09, ADC08, DGAVK03, Igi02]. reviewer [PH09]. Reviewers [LOK00]. reviewing [BND08]. Reviews [Bro04]. revisited [EL09, Lin91, Sch06]. Reward [Whi04]. rhythm [LRS06]. Rico [ORSCSB00, SBSCOR06]. right [BO05]. rise [CCHC06, Meh07]. risks [Lew07]. Robert [Bon03, Col04, Gar04, Sma04a, Ste04, Sti04]. robot [LJ08]. robust [BDS06, TG04]. Role [LdLM00, Rin00, Ver00, AGJC09, CM09, GDTS06, Goe05, JRS06, PGD04, SH08b, VWJ09, WBB04, ZGL09]. roles [YNT06, ZV09]. ROM [CIM07]. round [LTV05]. Rousseau [Ano04c, Bur03]. routes [ACYYA06]. Row [BMD07]. Row-column [BMD07]. rubber [SMW^{+}09]. Russia [MML^{+}04, WM04]. Russian [MMLA01a, MMLA01b, MJL^{+}09, Roy01, ZNK01].

S [San00]. S&T [BI00, Gus00, HG02, HB07, HVG06, ML05, PL08, RKM01, RK04, RRR07]. Saint [BMJARB^{+}07]. salaries [Wra09b]. same [CB09, LFB^{+}06, Sol07]. sampling [BPM^{+}09, Bur06b, Egg05a]. São [LPC05]. scale [ZRRB05]. Scaling [HHWD04, IP07]. Scandinavia [Glä00a, Glä00b]. Scandinavian [Ing00, KH08]. scare [BIE05]. scattering [Egg04b]. scenario [Das04]. scenes [FH06]. scheme [GS03a]. schemes [VDL^{+}02]. Schering [DL06]. Scholar [BI08b, KT08, MDNL09, Smi08]. Scholarly [Bal09, LAB^{+}09, Muk09, San01a, San01b, ZGBI09, BPMK06, CSW^{+}09, CX09, CYKK09, DD04, Gan09, Koe01b, Koe01c, MA01, MB04, MM05, NLM01, Uzu04, VS08, YY08a]. Scholars [JI00, DFRG02, Saa06]. school [OALL07, VM06]. SCI [SMB^{+}07, AJ08b, Big08, BPM^{+}09, CRLMRPA08, Ley02, MBR^{+}08, SFG09, SM00, SKMI06, Ste01]. SCI-expanded [AJ08b]. SciELO [MMP06]. Science [AMBB^{+}05, AWG00, BK06, BA01, Cah00a, CMJSOC09, CG01, CRLMR04, CBVS07, Das04, Eto02, GFNG08, GP00, Glä00a, Glä00b, GLT06, GDD03, Gra01, GH05, GD03, Hav01, He03, Hin03, Ing00, Jan00, LGT06, LT09, LTMW03, LKGdB01, MG00, MB06, Mey00a, MH01, NBRA^{+}02, ORSCSB00, SZ01, SMP04, Ste01, Tho00, TBvL00, Uzu06, VDL03, WXB09, Wen01, Yan09, ZL00, ZBO00, vR00, vdB00, AJ08a, AC04, AGM^{+}01, ABN^{+}09, AB01, BMJARB^{+}07, BKM03, BM03, BvdS06, BS01, BKB05, Boy09, BSPKN03a, BSPKN03b, BD05d, BDZZ07b, BS07, BN01, Cal06, CM05, CGV07, CC08, DGAVK03, DR08, DV03, DSBH09, Eto08, Fra02, GP01, Gei05, GSB02, GDP03, GS03a, GDTS06, GOS08, Gol02, GK09, GM04, GH07, GYY02, GKSK02, GD09a, Har04, HG08, HS06, JGD08]. science [Jas09, Jon09, Kim05, KB06, KNM07, Koe01a, Kos01, KT06, KT08, KBWM07, KKM07, Kre01b, Kre01c, KG05, Lar08a, LSGZMM^{+}08, LL03, Ley02, LZ05, LZ07a, LW09, LRS06, LHHWD06, MNQ08, MDNL09, Mai06, MBR^{+}08, MML^{+}04, MH07, MHSE04, Mey01, MDTU05, Mey06a, Mey06b, MHM07, MG04, Nag03, NBR01, Nis02, Noy01, NCM09, ÖSAT08, OLVLH09, PB09, Pod05, PFL01, PID06, PR09, Pou07a, PP09, Prp02, RvLB^{+}02, RRZ06, RWGA07, Sap07, Sch07b, SH04, Ski09, Sma06, SMB^{+}07, Son03, Soo09b, SGT^{+}06b, TBK06, TV06, Uzu04, VVM^{+}03, VZV^{+}03, VMD07, VSL09, VLdM04, VDL^{+}02, Wag05, Wei05, YHM05, YT09, ZRRB03b, ZRRBL03, ZWW09, Zsi08, vDH05, vLM02, vL06, vL09, vR04, CG06, CIM07, GKL06, Hav01, KKSK06, KK04, KT08, Ley04b, Ley07, Moe02]. Science [OW00, PL08, Sæt00, Sig00, SMB^{+}07, TSLCJC09, VMT^{+}01, VG09, WM04, ZWW09, LRS06, LR09a, TMVAR00]. science-dependence [SGT^{+}06b]. Science-Technology [MG00, BM03, Mey06b, VZV^{+}03, ZRRBL03]. Sciences [Abt00b, CGC^{+}00, JI00, JJ04a, LPC05, Luw00, Ned06, PFE00, RPFD07, Ste01, SL05, Abt07a, AKOK09, And01, AVGC^{+}06, BK06, Bal09, BS03, Eto08, Far02, GÖ09, JD05, KKL06, Kra07, LGA06, LCLA02, MEG06, YA07, ZTG09a]. Scientific [AGP07, AC04, BR00, BK00, Bea01, HVG06, HWD01a, HWD01b, Ioa06, KGK02, LKGdB01, MNQ08, MLR05, Mey00b, MMR05, NM08, ORSCSB00, OW00, PRPFC06, PC00, PA08, Pol00, PZ03, RRGGMS06, SLB00, SA00, SM00, Sol06, SKMI06, SK01b, VDC^{+}06, WD01b, WXB09, WWP^{+}05, WPZ^{+}04, Zan02, vdB00, Abt07c, Ack05, Ack06, AVGC^{+}06, ADC08, BMT09, Bal09, Ban06, BCK06, BRSYB04, BKK^{+}08, BM03, Bla09, BD03, BD06, BND08, Bro04, CJ06, CC08, CCHC07, Coc05, CEM06, Cro04, DW09, ER02c, Eto08, FCTV04, Fra02, Fra03, GG09a, GG09b, GLM^{+}07, Gei05, Glä01, GdL02, GM03, GSMT09, GKB^{+}09, GSH01, GBZAEC^{+}07, HHWD04, HE09a, He03, HL09, He09b, Hir05, HBK^{+}09, HKL08, IP07, IRN07, IJ04, Inö03, Jia08, JCTSM^{+}09, JRS06, JT08, KK04, Kor01, LZA08]. scientific [Lar02, LVCS03, LJL08, LZ02, Llu05, LINV^{+}09, LAL07, MA01, MHGGVJ08, MLNS08, MMP06, MM08a, MM08b, MM08c, MAVQHS^{+}04, NBRA^{+}02, NN08, ÖSAT08, OCA09, PM09, PK09, PGD04, PMO04, Pod05, Prp07, RR02, Rig09, RG06, RvHMJ05, SS08, Ski06, SBSCOR06, TNK^{+}06, TVvL02, Tuz05, VVPCOL09, VH01, VVM^{+}03, Van01b, Vin08, WL03, WFI03, WM04, WBB04, YO06, YW07, ZL02, ZTG09b, Zsi08, dMR01]. scientifique [SLB00]. Scientist [RRGGMS06, SKB08, HLSA07, HBK^{+}09, LMJS06, MSGGRR08, MFG07, SB09]. Scientists [DCB^{+}00, Prp00, ADC09a, Abt07c, BMFG03, BD05b, CM05, CBO^{+}03, FG06, Gol02, HB07, JRC09, Kap04, Kra08, LK08, MSRRGG02, RRMSG02, RRGGMS07, Soo09a, Van01b, Wra04, Wra09a, YDM^{+}05, Zan06]. Scientometric [BMJARBC05a, Eto08, GKL06, GL01, JZCZ02, MS01, Moe00b, Rin00, SSMV06, UKKP04, ZGHW07, BMJARBC05c, Coc05, DA08, GS03a, GD03, HG02, JG02, KKSK06, Kum09, Lee03, MMP06, PP09, RWGA06, SSM07, Tsa09a, Vin03, Vin06, Vin08, YHM05, Ye07, ZGBI09, Zsi08]. Scientometrics [BGB03, GP00, Gar02, GP02, GM09a, Gra01, GW01, HW01, Jan00, KG05, Oku00, San00, Sch00, Sch02, SF00, ZM01, CL06, Har09a, LAB^{+}09, Lee08a, LM07, LCL05, PBI02, Pou07a, Pou07b, Uzu04, Vin01, Ano01a, BGB03, Gar03, HKL08, McC05, SG01a, Sch01, Uzu02, Whi08]. scope [vPdlPvZ08]. scope-year [vPdlPvZ08]. Scopus [BI08b, GD09a, Pis09, TSLCJC09, VG09, dMACRVQ^{+}07]. score [AHK08, Egg99]. scores [PH09]. Screen [Sæt00]. scripts [NLD07]. Search [BI00, Cah00b, Gmü03, MDNL09, MB01, VVM^{+}03]. second [GSB02]. secondary [YRL03]. sectional [OALL07]. sections [ND09, VWJ09]. sector [CMP03, DR05, MSS06, OALL07, RG06]. Sectoral [LGT06, GS07, YO06]. sectors [AP04, RdL02]. Security [Lee08a, Gor05a]. Seen [Sæt00, GKL06]. Seismology [WL03]. Selected [Ing00, Ama08, BH05, Hin03, HR09, IJ04]. Selecting [BD06]. Selection [BD05c]. selective [IK02]. Self [Bon04, HLSA07, Jud04, ZL00, ZK08, Aks03, AGJC09, CG06, CM09, FA07, GT04a, GT04b, GDTS06, Har09b, Hen09, Kra07, Med06, PS02, PFL01, Sch08, SGT06a, TG06, WBB04, YW07]. Self-citation [ZK08, Aks03, FA07, Hen09, Kra07, Sch08, SGT06a]. Self-citations [HLSA07, AGJC09, CG06, CM09, GT04a, GT04b, GDTS06, Har09b, Med06, PS02, SGT06a, TG06, WBB04]. self-cited [YW07]. Self-emancipation [Bon04]. self-linked [Jud04]. Self-linking [Jud04]. Self-Organization [ZL00]. semantic [MEG06]. semi [YM06]. semi-parametric [YM06]. Semiconductor [TJM00, GM07, yTXwW03a, yTXwW03b, TM03]. seminal [KBL08]. senior [BBG07]. Separating [Sol07]. sequence [Egg10, Lia06]. sequences [Egg09a]. series [Bro04, Bru03, LRR09]. set [HW02, ML05, Rou05a]. sets [ER02c, MBPKT03]. setting [Ben03]. several [RCE^{+}09]. Severe [CHH04]. sex [ADC09a]. Shadows [ZBO00]. Shan [Xek05]. Shanghai [Flo07]. shaping [Cro04]. share [Egg05c, GT04a]. sharing [TGJd01]. Short [SS05]. shortcomings [MSW01]. shorter [MG04]. shortest [Bjö06]. Should [CB09]. shoulders [Sma04a]. side [PS02]. Sighted [LOK00]. signaling [vDH05]. Signals [vDH05]. signatures [GBGCRFdMA09]. Significant [LCLA02, TNK^{+}06]. significantly [PRdKB07]. signs [BND08]. similar [BSPKN03a]. similarities [LHHWD06, Prp05]. Similarity [SB09, AJ08a]. Simonton [LVCS03]. Simple [Bur02a]. Simpson [RRZR09]. simulation [YYL05, YY08b]. since [WD01a]. Singapore [Lee03, WH07]. Single [Ley00, Abt07b, PM09, Rig09, Sch09]. single-authored [Abt07b]. sitations [PBZ02]. site [LTWM05a, LTWM05b]. sites [HT09, PBG^{+}05, TS02]. Sitkis [SM06]. Sivaraj [SSM07]. Six [Wag05, PR09]. Size [BRV00, LdLM00, SA00, BD05a, Egg05c, JSJX02]. sleep [RWGA06, RWGA07]. Sleeping [Bur05a, vR04]. slips [SR05]. Small [Tho00, Bjö06, CA09, CvPdlPR06, LC06, Mey06b, WH07]. SMEs [LdLM00]. snapshot [RWGA06, RWD^{+}09]. Social [BE04, Ing00, Ley04b, Ley07, MH01, Ned06, Ste01, SL05, AVGC^{+}06, BS07, Eto08, GÖ09, GYY02, GKSK02, HBK^{+}08, JD05, Kap04, KKL06, LGA06, LMJS06, MSGGRR08, MSS06, MEG06, NBR01, ÖSAT08, PB09, RRGGMS07, RRZ06, SKB08, Van01a, YS08, ZTG09a, vL06, And01]. Society [Koe01a, Yan09, NSS04, Prp07]. Socio [MP00, MH01, YT09, Coz06, GR06a]. Socio-Bibliometric [MP00]. Socio-Cognitive [MH01]. Socio-economic [YT09, Coz06, GR06a]. socioecological [San01a, San01b]. Sociological [GL01]. sociology [Har04]. sociometric [ER04]. Software [MVH^{+}05, SM06]. Soil [MHM07]. Sokal [Whi04]. solar [Lar08a]. Solla [Ano08b, Ano09, Edi05, Ley04a, McC05, Rou04, Rou05b, Whi08]. Solution [Egg04b]. Solutions [Mar01]. Some [AGG00b, Bra09, Bur01b, Izs06, Jan00, Kra08, LK00, Llu05, OLBB^{+}08, Rig09, Tri00, vR01, CM05, Glä08b, JJ04a, Mar09, Nel06, Sid06, SSG06, Vin01, Bur07a, Mey07]. Source [BR00, BV06, Egg04c, GMA07, GSST06, Gor05a, Kim02, Per06, VW04, ZL02]. source-item [Egg04c]. Sources [KT08, AVDT04, Egg05c, PBI02, SP08, Ste09c, VS08, WH07]. South [YS08, CYKK09, IJ04, JI00, KBL08, Lee08b, MP06, MMP09, PHL05, Pou05, Pou07b, Pou09, Soo09a, Soo09b]. Soviet [MJL^{+}09]. Space [DCF00, Pri03, Bjö06, Bla09, CYKK09, vPdlPvZ08]. spaces [BKM03, MBPKT03]. Spain [IRN07, OLBB^{+}08, PZ03, VBBF08]. span [AJ08b]. Spaniards [CB08b]. Spanish [AGP07, ABZMD^{+}07, AY09, AP04, FCB02, GBFM09, JM07, LSGZMM^{+}08, MSRRGG02, MDMBA^{+}06, OLVLH09, PC00, PBG^{+}05, PSD^{+}07, PGFRD09, PA08, RRS07, RRMSG02, RRGGMS06, RG06, VVPCOL09]. spatial [Old01a, Old01b]. Special [Mey00b]. specialisation [Tuz05, vZvPdlPH06]. specialization [Ste09d]. specialties [LC06]. specialty [RM07, YMB09]. specific [GTSD09, Kap04, LL08]. Speech [Moe00b]. speed [MF09]. spend [Dai09]. spending [Coz06]. spices [SA07]. spillovers [ZRRBL03]. sponsored [tLtC07]. sports [YA07]. spread [Kao09a]. SQL [Wol06]. square [AHK08]. SSCI [CRLMRPA08, Ste01, SL05]. stabilisation [PT07]. staff [KO08]. stage [CHH04, MML^{+}04, Pon02]. standard [GST03, ND09, vR06]. standardisation [Rou02]. standardization [Fuj06, GFNG08, KS02]. Standardizing [GMA07]. standing [Lew02b]. star [ADC09a]. stars [SFGRP09]. State [Wor00b, GMA06, Lar08b, PP09, YYL04]. statement [RRR07]. statement-linkages [RRR07]. States [Hen09, LW09, VD06]. Statistical [He03, AGM^{+}01, GFNG08, HWD01a, HWD01b, MN03]. Statistics [AET09, BGG00, Eto00b, Hal09, LLJ09, RdL00, SG09, Glä08b, LGS03, Ley02, RdL02]. Status [GD09a, GD09b, IJ04, BRV06, KKKM05, Van01b]. steady [YYL04]. stem [HCTC03, LH06, LDF^{+}09]. stems [AJ08a]. step [BMFG03]. still [vL06]. Stimulation [Kim07]. STN [NLD07]. Stochastic [Bur01a, SR05]. stock [Tho01]. strategic [BD08]. Strategies [DW01, LdLM00, Moe00a, CB03, LFB^{+}06, Mus06, Ste09d, Uzu06]. strategy [Dai09]. Stratification [Dan00]. Stream [MT07, Fuj06]. streamline [NLD07]. Strength [CLH07, vL09, Jar08, Ste09a]. strengths [AMBB^{+}05]. stroke [CHH07]. stroke-related [CHH07]. Structural [GM09b, LHHWD06, RKM01, Ley02, TG08, TG09, YW04]. Structure [CGR06, GBFM09, LK00, AJOP04, BMJARB^{+}07, BvdS06, BBK09, CYKK09, DR08, Has05, HKL08, Jar01, sKtP04, Lar08a, Liu05, MMAHS08, PL08, Pon02, SV01, SLP06, SU09, Son03, Van01a, Vin08, YS08, YR01, ZRRB03a]. Structures [LKGdB01, MH01, GFNG08, LGD02, PL09]. STS [Van01a].

Studies [GL01, Rin00, ZL00, vdB00, AUQ09, BND08, GL05, HW03, LZA08, Liu05, PL09, PGD04, SP08, Wag05, ZRRB03a]. Study [AOS00, Dan00, DCB^{+}00, Egg00b, JI00, LK00, Prp00, TMVAR00, TJM00, vdB00, Aks03, AC04, AGM^{+}01, AY09, AGS07, ADC08, BIP09, BBV^{+}08, Bea01, Ben03, BMFT05, Bha04, BDZZ07a, BCGMBSVM07, Bur06a, But02, CSGGB09, CC08, CB08a, DFRG02, Egg01c, Egg09a, GP01, GM03, GLT06, GS07, GGP05, GM04, GM07, Gup06, HME07, HE09a, HD07, HG08, Hen09, HR09, HCC03, JM07, KKSK06, KLC^{+}08, KNM07, Koe01b, Koe01c, LMJS06, Lee03, Lee08b, LL03, Lew01, LLR04, LC06, LRR09, MG05, MSRRGG02, MSGGRR08, MSU03, MN03, MMP09, Muk09, MA07, OALL07, PM06, PT07, PRPFC06, PB09, PMO04, Pon02, Pri03, QLLR08, RDvT03, RV08, SLI09, Ski06, SMYS05, SKMI06, SMB^{+}07, SB09, TT03, UKKP04, VBBF08, VWJ09, WXB09, Whi04, Wu09, XCS03, YA07, YMB09, YYL05, ZL02]. study [ZWZM04, vdBH06, IMLSSC09]. Studying [Kor01, Lau03]. style [HPF03]. sub [Fuj02, SM06, vCvCvU08]. sub-area [vCvCvU08]. sub-domains [Fuj02]. sub-grouping [SM06]. Subfield [GTSD09, Vin02b]. Subfield-specific [GTSD09]. subfields [BPM^{+}09, GS03a, Jon09, SMB^{+}07]. subject [ER02c, GBZAEC^{+}07, SMYS05, ZCY09]. subjects [LT08]. submitted [Ama08]. subsequent [Wal06]. substantive [QLLR08]. success [MS05]. Successive [Sch07a, AJR09, Egg08b, RT08]. summarise [CB09]. Superiority [Hu07]. supervision [RT09]. supervisors [LLR04]. supply [OLC07]. support [DG04, Hem06]. Surface [JCTSM^{+}09]. surfactants [BMJARBC05a]. surgery [yTlC05]. surnames [LK08]. surprising [Jas09]. survey [BGS07, MMLA01a, MMLA01b, MSU03, MFG07, RWGA06, Rou08]. Surveying [MSM06]. Surviving [CM09]. Sustainability [SS09a]. Sustained [PKK09]. Sutter [Rou02]. Swarming [Bru03]. Swedish [DP03, GDP03, SPN00]. symbols [Sch06]. Symmetric [ER01]. synchronous [Glä04, UKKP04]. Syndrome [CHH04]. Syndrome-related [CHH04]. Synthetic [Kum09, Kum06]. System [Ley00, SPN00, SM00, VCP00, ADC09b, AP04, Egg05c, GM09b, HT06, HS06, HVG06, JRS06, LSGZMM^{+}08, LZ05, LW09, MC02, Rig09, San01a, San01b, SH09, VL01, WD05, Wei05, YY08b, vLM02]. systematic [Bur06b, HD07]. Systems [CH00, MH01, SF00, Tho00, Wen01, ZIJSCMGG07, AGM^{+}01, Egg09b, Jon09, LCWD09, Mey02, MDTU05, Old01a, Old01b, PHL05, VDC^{+}06, vL09].

Table [SF00]. Taiwan [CCHF05, CH00, CHH07, HCC03, HCC06, KP09, LC06, cL07, SLP06, TT03]. Taiwanese [KLC^{+}08]. taken [Ski06]. taken-for-granted-ness [Ski06]. Taking [Kim06]. tale [CMJSOC09]. taxonomies [RB06]. Team [MSGGRR08, MSRRGG02]. teams [RRMSG02, RRGGMS06, Ski09]. tech [HCC03]. Technical [Luw00, WPZ^{+}04, MB04]. technique [MAVQHS^{+}04, SLB00]. techniques [MVH^{+}05]. techno [KNM07]. techno-science [KNM07]. Technological [CLH07, Ste09d, TBvL00, AGP07, Bha04, KM07, LGS03, LV09, Lee09, LL08, PK09, PB09, RG06, Ste09a, VDC^{+}06, VMD07, Ygl03, Zan02, vZvPdlPH06]. technologies [Ben03, CCHF05, JLC09, TNK^{+}06]. Technology [Das04, GP00, GH05, GD03, Hin03, Jan00, JG02, LdLM00, MG00, Mey00a, ORSCSB00, Ver00, ZL00, vdB00, ACF09, AC04, AMBB^{+}05, AGM^{+}01, BKM03, BM03, Cal06, CMJSOC09, CvPdlPR06, CB03, DV03, Eto03, GFNG08, GP01, Gei05, GH07, GD09a, JM07, KNM07, Kos01, LJ08, LZ07b, Mey01, Mey06a, Mey06b, MEG06, PB09, PFL01, PP09, SH04, Tho01, Uzu06, VZV^{+}03, VMD07, VSL09, VDL^{+}02, WHC07, ZRRBL03, Yan09]. Ted [Tsa09b]. telecommunication [CMP03]. Tell [Mey00a, ZIJSCMGG07]. Ten [Bon05b, BH05, IMLSSC09]. tendencies [KM04a]. tenured [KO08]. term [CEM06]. Terminology [MS01]. terms [Egg05a, LZ07a, Ley08]. terrorism [Gor05a, Gor05b, Gor07]. test [AGJ08, LVCS03]. Testing [JRC09, Pet08, dMR01, Izs06]. tetrachloro [PRPFC06]. tetrachloro-dibenzo-dioxins [PRPFC06]. tetrad [GDM08]. text [ABGAMD^{+}09, GGP05]. th [Bur02b]. th-citation [Bur02b]. Thai [SRS^{+}02]. Thailand [SMYS05, SKMI06]. Their [JRS06, KKKM05, Rin00, AGJC09, CL06, CIM07, CS02, Har08, JRC09, Kra08, LT08, LTMW03, Mey06a, MT07, PGFRD09, Sma04b, SKB08, TG09, TG04, TM03, WWW07, YMB09, YNSS09, YW07]. theJournal [Koe01a]. theme [LL08]. Theoretical [Bur03, GK00, BCZM^{+}09, KGK02, YNT06]. Theory [ER01, ER02a, Egg06, Bur02a, ML05, Pin06, Sma04a]. thermometers [Pis09]. theses [LZA08, VBBF08, YA07]. thesis [LZA08]. thick [CS02]. thin [BM03]. think [Sma04b]. Third [Hor01, MT07]. Third-Stream [MT07]. thorium [KKSK06]. those [LPC05]. Thought [KB08]. Thoughts [Jan00, Bur07a]. Threaded [ZL00]. threat [Pou07a]. Three [Sæt00, ABGAMD^{+}09, HL09, Hen09, Jon09, LLR04, YGL06]. threshold [MMR05]. ties [Jon09, YMB09]. Time [Bur07b, AJ08b, Bru03, Fox04, GST03, LV09, PR09, vPdlPvZ08]. Time-dependent [Bur07b]. times [CM09]. title [LH05, Yit02]. titled [BZDZ07]. titles [Bal09]. Today [Aru01]. tool [LDF^{+}09, Llu05, NSS04, SM06]. tools [VVM^{+}03]. Top [Ley04b, LpCY06, ZBO00, BDZZ07a, CL06, KK04, SSG06, Yan09]. Top-down [Ley04b]. Topic [Wor00a]. topical [LM07]. topics [Ban06, BI08a, TLL^{+}09, vdBH06]. topology [Jia08]. total [AHK08, KBL08, QLLR08]. tourism [BBV^{+}08]. Traces [CVV^{+}06]. Tracing [CH04, LGS03, Mey02, ZGL09]. track [KZ09]. Tracking [KNM07, RRR07, Sma06, Wor00b]. trade [RT09]. trademark [Dai09]. traditional [AMTE07, CMJSOC09]. training [YS08]. transfer [AP04, RvLB^{+}02, RJDD07]. transfractality [BMJARBC05c, BMJARB^{+}07]. Transient [Gor07]. transition [Bal09, MM05]. transitional [Prp07]. transnational [Nag03]. transport [MM02]. Transtheoretical [AGS07]. treatments [HCLH04]. Treaty [Ste09b]. Treaty- [Ste09b]. trend [LDF^{+}09, SMB^{+}07]. Trends [DDM00, GOS08, KU05, Kum06, MM02, MJL^{+}09, PV00, San00, YX08, KL09, Kyv03, LDF^{+}09, LZ07b, OMD01, Tij07, XZH08]. Triad [GDM08]. triadic [Cri06, Ste09a]. trial [JMB06]. trials [RCE^{+}09]. Triple [BL07, DP03, GS07, Gok03, Ley03, LM03, LM07, MSU03, PHL05, Pri03]. Triple-Helix [BL07]. tropical [KU05]. Trustees [BD05c]. tsunami [CH07]. Turkey [GOS08, Gok03, GÖ09, GYY02, ÖSAT08, Uzu06]. Turkish [IK02, Ton00, YA07]. twelve [GSB02, Wal06]. Twentieth [Hav01]. Two [Bur01c, Hav01, Rin00, Sæt00, SG01a, Tho00, AJ08a, BCZM^{+}09, CA09, Jar05, MDTU05, Pis09, Pri03, SGBB03, SLI09, TSLCJC09, Wu09]. type [BGS06, Kim02, LR09b, SKT09]. Types [Sig00, BCZM^{+}09, Soo09b]. typical [VG09].

U.K. [RdL02]. U.S. [BPMK06, ERW05, Lee08b, PMO04]. UK [CMU09, Har09a, LTMW03, MT07, Moe08, SHO00]. Ukraine [KKL06]. uncitedness [Egg08a, vLM05]. uncovered [LTL^{+}06]. underlying [DL06, Jon09]. Understanding [LF03, vLM05]. unexpected [Jas09]. UNICAMP [PB09]. unification [GMA06]. unified [BMJARBC05c, BMJARBC05a]. Union [AET09, BH05, BRSYB04, Ley00, RWD^{+}09, TBH^{+}02, TZ08, UCCM01]. United [Hen09, LW09, VD06]. universal [YYL04]. universalism [And01]. Universe [NDNZ01, Sha01]. universitie [ADP08, Bal06, PBG^{+}05]. universities [AKOK09, ACAYY07, BA07, BDS06, BDZZ07a, BCGMBSVM07, CL06, Flo07, GBFM09, Hem06, Hen09, Hor01, HCC06, KP09, KO08, LLR04, LCL05, MNQ08, MML^{+}04, MT07, RRS07, RDvT03, ST02b, TG06, VKG07, VLdM04, vR05a]. University [IMLSSC09, MLR05, PSD^{+}07, TS02, TI02, ACYYA06, BGS07, Bal08, GS03b, HHWD04, KH08, Kra08, Kyv03, Ley03, LM03, LM07, LTWM05a, LTWM05b, LTL^{+}06, MHGGVJ08, MSRRGG02, MSU03, OLC07, PGFRD09, QCW04, RRMSG02, SGBB03, TBH^{+}02, TZ08, TSLCJC09, VVPCOL09, VG09, YY08a, ZWZM04, GBGCRFdMA09, LPC05, MM05, WFI03, ZWZM04]. university-centred [TZ08]. university-industry [LTL^{+}06, MHGGVJ08]. university-industry-government [Ley03, LM07]. university-related [MSU03]. Untapped [BR00]. update [AY09, Ley08]. upon [CC09, DGAVK03, SL05, TNK^{+}06, Wei05]. upsetting [LZ05]. urban [Liu05]. URL [KT06]. Uruguay [VM06]. USA [CMM04, CB08b, MM02, SLP06]. usage [BvdS06, PID06]. Use [LH08, VM06, YAH08, ADC08, Bal09, BFG02, Bur05b, Bur06b, CB09, CMJSOC09, Llu05, MVH^{+}05, MM08a, TTP03]. used [PBI02]. Usefulness [IRN07]. User [PID06]. users [Lew02b, PD07]. Using [AWG00, BR00, Bas06, Bay05, Boy09, CLH07, CC09, CS02, De 00, DG04, DCB^{+}00, GSD09, Gus00, HFG^{+}05, KM07, LAFP04, NLD07, OLC07, PFL01, RdL00, RdL02, Sch09, VDL^{+}02, Arc02, BMD09, BM09, CBVN06, CL06, ER02c, Egg05b, Fuj02, GLM^{+}08, GT01, He03, HW02, HW03, HCC03, KL09, LGS03, LLSL07, Lee08a, LJ08, LCL05, LL08, LINV^{+}09, MSW01, Moe02, PHL05, PL09, Saa06, SMYS05, TSLCJC09, YO06, ZL02, ZGHW07, vdBH06].

USPTO [Cri06, PA08]. utility [LAB^{+}09]. utilization [CCDJJF04]. Utilizing [Old01a, Old01b].

V [Ano01b, Ano01c, CG01, Gar01b, Gra01, PKK09]. V. [Ano01b, Ano01c, CG01, Gra01]. Validation [AOS00, NRTD00]. validity [BD05c, BM09, KBWM07, Rei09]. valuation [LH08]. Value [JM07, CC09, Lee09, MT07, Ste09b, vPdlPvZ08]. valued [RZ08]. values [PGD04, YM06]. Vanishing [BBG07]. vanity [Zsi08]. variables [Lee09, Pin06]. variation [Jar08]. Variations [CIM07, TT04]. varies [DLM06]. various [Abt07a]. varying [Dai07]. Vassily [Gar01b, Roy01]. Vector [Pri03]. Venezuelan [Req05]. verge [SU09]. versions [CIM07]. versus [BDZZ07b, Fra08, Fuj02, KKK07, MC09, MMP06, Per01, Ran09, TSLCJC09, Wal06]. via [GT01, Iga08, RRR07]. vicinity [VMD07]. Vienna [WFI03]. View [WD01a, BS03, GS01, Gu04b, NBR01, Pri03, PBZ02, SW01]. Vinkler [Ano09]. Virtual [KLK01b]. Visibility [GBGCRFdMA09, Ing00, Ing02, KA04, AJR09, BBG07, Fre04, IMLSSC09, Kre04, KKK07, LJFD01, PSD^{+}07, Pou05, RR02, SG01b, UC03]. Vision [RRR07]. visual [OACS08, ZWW09]. Visualization [Nel06, DV03]. Visualizing [Liu05, Nag02, Bla09]. Vita [DCB^{+}00]. volatility [CMM04, MM02]. Volume [Ano01f, Ano01g, Ano07a, Ano07b]. volumes [SMW^{+}09]. vs [CA09, JJ04a, LMJS06, MLNS08, Mus06, RDvT03, RRMSG02, Tri00, VY06, Ygl03].

W [Ano08b]. wage [CB03]. Walk [LTV05]. Walk-round [LTV05]. war [Igi02, Luk06]. Waring [Xek05, Bur05b, SJJ04]. Watson [AJOP04]. Way [SZ01]. weak [HHWD04, Mey06a, YMB09]. weakness [vL09]. web [GK09, HT09, LFAH04, OACS08, PBG^{+}05, VY08, VS08, Wu09, BI00, BI01a, BIE05, BIP09, Bjö06, Bro04, CX09, CYKK09, CIM07, FPGBD03, HS06, ID04, Jud04, Kre04, KA04, KKK07, LTWM05a, LTWM05b, MBPKT03, OCA09, PID06, QCW04, Sch02, ST02b, TT04, TS02, TT03, TTP03, TBK06, TSLCJC09, VY06, VG09, ZL02]. Webometric [AKOK09, HT09, MC02, Pri03]. webometrics [BI01b, LAFP04]. webs [PT07]. website [Vau04]. websites [PSD^{+}07, PGFRD09, QCW04, TBH^{+}02, VW04, YZ08, VGK06, VKG07]. weight [SGT06a]. Weighted [HS06, MH07]. Weighting [Pin06]. Welfare [Wor00a, Wor00b]. Western [HR09, Jon09, LHHWD06, TTP03]. Where [MDMBA^{+}06]. Which [BI08b, MDMBA^{+}06, ACAYY07]. White [Edi05, McC05]. Who [MDMBA^{+}06, RPFD07, RPFD07, vL06]. Wide [Bro04, KH08, NN08]. widening [Zan02]. WIFs [LTMW03]. win [Ano09, Edi05]. wins [Ano08b, Ley04a]. within [BI04, Bur06a, CvLT07, Coc05, Coc09, GT01, KM04b, Rig09, TSLCJC09, VMD07]. Wolf [RR06]. women [GBGCRFdMA09, LL03]. word [LJ08, SG03, vdBH06]. word-reference [vdBH06]. words [LH05, LH06]. work [BD05b, QMC08]. working [OALL07]. works [Tsa09b]. Workshop [Wor00c]. World [BD05e, KKSK06, KH08, NBR01, Bjö06, BDZZ07a, CL06, CX09, Flo07, GSB02, GDM08, GKSK02, Ing02, KKL06, LZ05, LW09, LCL05, MNQ08, NN08, TI02, XZH08, Bro04, Ton00]. World-wide [KH08, NN08]. worldwide [JLH07, Tij07]. WoS [BI08b]. writer [Har08]. writes [UCCM01]. Wyllie [SSMV06].

XML [ZL02].

year [BIP09, Hen09, vPdlPvZ08]. years [AQ03, AGJC09, Bon05b, CM09, DR08, DR01, NC09, RWGA07, Sid06, VBBF08, YA07]. Yinian [LF03]. young [Gol02, HBK^{+}09, SFGRP09, Wra04, Wra09a, Prp00, Tri00]. Yugoslavia [Igi02]. Yule [RRZR09].

Zealand [GDD03, Smi08, HD07]. zero [LLSL07, YM06]. zero-inflated [LLSL07]. Zewail [KKK02]. Zipf [Izs06, JSJX02, KSB03]. Zitt [AGJ08, Ano09]. zoom [AGJ08].

References

Arunachalam:2001:FSR

[AB01] Subbiah Arunachalam and Jayashree Balaji. Fish science research in China: How does it compare with fish research in India? Scientometrics, 52(1):13–28, September 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

                 tronic). URL http://link.springer.com/article/10.
                 1023/A%3A1012742826813.

Aleixandre-Benavent:2009:FTP

[ABGAMD^{+}09] Rafael Aleixandre-Benavent, Gregorio González-Alcaide, Alberto Miguel-Dasit, Carolina Navarro-Molina, and Juan Carlos Valderrama-Zurián. Full-text publications in peer-reviewed journals derived from presentations at three ISSI conferences. Scientometrics, 80(2):407–418, August 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

                 1588-2861 (electronic). URL http://link.springer.com/
                 article/10.1007/s11192-008-2066-z.

Arruda:2009:BCS

[ABN^{+}09] Denis Arruda, Fábio Bezerra, Vânia Almeida Neris, Patricia Rocha De Toro, and Jacques Wainera. Brazilian computer science research: Gender and regional distributions. Scientometrics, 79(3):651–665, June 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1944-0.

Abt:2000:DIP

[Abt00a] Helmut A. Abt. Do important papers produce high citation counts? Scientometrics, 48(1):65–70, June 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005680318379.

Abt:2000:RFR

[Abt00b] Helmut A. Abt. The reference–frequency relation in the physical sciences. Scientometrics, 49(3):443–451, November 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1010589623808.

Abt:2007:FMP

[Abt07a] Helmut A. Abt. The frequencies of multinational papers in various sciences. Scientometrics, 72(1):105–115, July 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-1686-z.

Abt:2007:FSA

[Abt07b] Helmut A. Abt. The future of single-authored papers. Scientometrics, 73(3):353–358, December 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1822-9.

Abt:2007:PRS

[Abt07c] Helmut A. Abt. The publication rate of scientific papers depends only on the number of scientists. Scientometrics, 73(3):281–288, December 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1807-8.

Aleixandre-Benavent:2007:HIN

[ABZMD^{+}07] Rafael Aleixandre-Benavent, Juan Carlos Valderrama Zurián, Alberto Miguel-Dasit, Adolfo Alonso Arroyo, and Miguel Castellano Gómez. Hypothetical influence of nonindexed Spanish medical journals on the impact factor of the Journal Citation Reports-indexed journals. Scientometrics, 70(1):53–66, January 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-0104-x.

Alfaraz:2004:BSF

[AC04] P. H. Alfaraz and Amalia Mirta Calviño. Bibliometric study on food science and technology: Scientific production in Iberian–American countries (1991–2000). Scientometrics, 61(1):89–102, September 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000037365.53469.91.

Azagra-Caro:2007:WRD

[ACAYY07] Joaquín M. Azagra-Caro, Fragiskos Archontakis, and Alfredo Yegros-Yegros. In which regions do universities patent and publish more? Scientometrics, 70(2):251–266, February 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-0202-9.

Azagra-Caro:2009:WDP

[ACdLPM09] Joaquín M. Azagra-Caro, Ignacio Fernández de Lucio, François Perruchas, and Pauline Mattsson. What do patent examiner inserted citations indicate for a region with low absorptive capacity? Scientometrics, 80(2):441–455, August 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2081-0.

Acosta:2009:EQE

[ACF09] Manuel Acosta, Daniel Coronado, and Ana Fernández. Exploring the quality of environmental technology in Europe: evidence from patent citations. Scientometrics, 80(1):131– 152, July 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2057-0.

Ackermann:2005:BCS

[Ack05] Eric Ackermann. Bibliometrics of a controversial scientific literature: Polywater research, 1962–1974. Scientometrics, 63(2):189–208, April 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0210-6.

Ackermann:2006:IFI

[Ack06] Eric Ackermann. Indicators of failed information epidemics in the scientific journal literature: A publication analysis of Polywater and Cold Nuclear Fusion. Scientometrics, 66 (3):451–466, March 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0033-0.

Azagra-Caro:2006:WDU

[ACYYA06] Joaquín M. Azagra-Caro, Alfredo Yegros-Yegros, and Fragiskos Archontakis. What do university patent routes indicate at regional level? Scientometrics, 66(1):219–230, January 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0015-2.

Adams:2005:ECC

[Ada05] Jonathan Adams. Early citation counts correlate with accumulated impact. Scientometrics, 63(3):567–581, June 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-005-0228-9.

Argamon:2008:LUR

[ADC08] Shlomo Argamon, Jeff Dodick, and Paul Chase. Language use reflects scientific methodology: A corpus-based study of peer-reviewed journal articles. Scientometrics, 75(2):203– 238, May 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1768-y.

Abramo:2009:CSS

[ADC09a] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Alessandro Caprasecca. The contribution of star scientists to overall sex differences in research productivity. Scientometrics, 81 (1):137–156, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-2131-7.

Abramo:2009:GDR

[ADC09b] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Alessandro Caprasecca. Gender differences in research productivity: A bibliometric analysis of the Italian academic system. Scientometrics, 79(3):517–539, June 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-2046-8.

Abramo:2008:MIU

[ADP08] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Fabio Pugini. The measurement of Italian universities’ research productivity by a non parametric-bibliometric methodology. Scientometrics, 76(2):225–244, August 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1942-2.

Adler:2009:CS

[AET09] Robert Adler, John Ewing, and Peter Taylor. Citation statistics: A report from the International Mathematical Union (IMU) in cooperation with the International Council of Industrial and Applied Mathematics (ICIAM) and the Institute of Mathematical Statistics (IMS). Statistical Science, 24(1):1–14, February 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1255009002. See com-

ments [Hal09, SG09, LLJ09].

Anduckia:2000:BOC

[AGG00a] Juan Carlos Anduckia, Julián Gómez, and Yuri Jack Gómez. Bibliometric output from Colombian researchers with approved projects by COLCIENCIAS between 1983 and 1994. Scientometrics, 48(1):3–25, June 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005680900632.

Anduckia:2000:SFC

[AGG00b] Juan Carlos Anduckia, Julián Gómez, and Yuri Jack Gómez. Some features of Colombian research population (1983–1994). Scientometrics, 48(3):285–305, July 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005684303869.

Adams:2008:CZT

[AGJ08] Jonathan Adams, Karen Gurney, and Louise Jackson. Calibrating the zoom — a test of Zitt’s hypothesis. Scientometrics, 75(1):81–95, April 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-1832-7.

Andrade:2009:JIT

[AGJC09] Antonia Andrade, Raúl González-Jonte, and Juan Miguel Campanario. Journals that increase their impact factor at least fourfold in a few years: The role of journal selfcitations. Scientometrics, 80(2):515–528, August 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2085-9.

Amaral:2001:ASP

[AGM^{+}01] Luis A. N. Amaral, P. Gopikrishnan, Kaushik Matia, Vasiliki Plerou, and H. E. Stanley. Application of statistical physics methods and conceptsto the study of science & technology systems. Scientometrics, 51(1):9–36, April 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1010556426328.

Adams:2007:PCI

[AGM07] Jonathan Adams, Karen Gurney, and Stuart Marshall. Profiling citation impact: A new methodology. Scientometrics, 72(2):325–344, August 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-1696-x.

Albert:2007:STP

[AGP07] Armando Albert, Begoña Granadino, and Luis M. Plaza. Scientific and technological performance evaluation of the Spanish council for scientific research (CSIC) in the field of biotechnology. Scientometrics, 70(1):41–51, January 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-0103-y.

Andres:2007:TMO

[AGS07] Ana Andrés, Juana Gómez, and Carmina Saldaña. The transtheoretical model and obesity: A bibliometric study. Scientometrics, 73(3):289–301, December 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1692-1.

Anderson:2008:BDS

[AHK08] Thomas R. Anderson, Robin K. S. Hankin, and Peter D. Killworth. Beyond the Durfee square: Enhancing the h-index to score total publication output. Scientometrics, 76(3):577–588, September 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-007-2071-2.pdf.

Ahlgren:2008:BCC

[AJ08a] Per Ahlgren and Bo Jarneving. Bibliographic coupling, common abstract stems and clustering: A comparison of two document-document similarity approaches in the context of science mapping. Scientometrics, 76(2):273–290, August 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1935-1.

Andreis:2008:ICJ

[AJ08b] Mladen Andreis and Maja Jokić. An impact of Croatian journals measured by citation analysis from SCI-expanded database in time span 1975–2001. Scientometrics, 75(2): 263–288, May 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1858-x.

Ahmed:2004:HCO

[AJOP04] Tanzila Ahmed, Ben Johnson, Charles Oppenheim, and Catherine Peck. Highly cited old papers and the reasons why they continue to be cited. Part II. The 1953 Watson and Crick article on the structure of DNA. Scientometrics, 61(2):147–156, October 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000041645.60907.57.

Ahmed:2005:LE

[AJOP05] Tanzila Ahmed, Ben Johnson, Charles Oppenheim, and Catherine Peck. Letter to the Editor. Scientometrics, 63(3): 631, June 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-005-0232-0.

Arencibia-Jorge:2009:IIR

[AJR09] Ricardo Arencibia-Jorge and Ronald Rousseau. Influence of individual researchers’ visibility on institutional impact: an example of Prathap’s approach to successive hindices. Scientometrics, 79(3):507–516, June 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-2025-0.

Aminpour:2009:WAI

[AKOK09] Farzaneh Aminpour, Payam Kabiri, Zahra Otroj, and Abbas Ali Keshtkar. Webometric analysis of Iranian universities of medical sciences. Scientometrics, 80(1):253–264, July 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2059-y.

Aksnes:2003:MSS

[Aks03] Dag W. Aksnes. A macro study of self-citation. Scientometrics, 56(2):235–246, January 2003. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1021919228368.

Archambault:2009:HJI

[AL09] Éric Archambault and Vincent Larivière. History of the journal impact factor: Contingencies and consequences. Scientometrics, 79(3):635–649, June 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-2036-x.

Amat:2008:EPD

[Ama08] Carlos B. Amat. Editorial and publication delay of papers submitted to 14 selected food research journals. Influence of online posting. Scientometrics, 74(3):379–389, March 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1823-8.

Altvater-Mackensen:2005:STR

[AMBB^{+}05] Nicole Altvater-Mackensen, Gregor Balicki, Lucie Bestakowa, Bianca Bocatius, Johannes Braun, Lars Brehmer, Verena Brune, Kirstina Eigemeier, Füsun Erdem, Ralf Fritscher, Anne Jacobs, Bernd Klingsporn, Marcin Kosinski, Julia Kuntze, Ju-Ra Lee, and et al. Science and technology in the region: The output of regional science and technology, its strengths and its leading institutions. Scientometrics, 63(3):463–529, June 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0225-z.

Alcaide-Marzal:2007:IAT

[AMTE07] Jorge Alcaide-Marzal and Enrique Tortajada-Esparza. Innovation assessment in traditional industries. A proposal of aesthetic innovation indicators. Scientometrics, 72(1):33– 57, July 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1708-x.

Andersen:2001:NUS

[And01] Heine Andersen. The norm of universalism in sciences. Social origin and gender of researchers in Denmark. Scientometrics, 50(2):255–272, February 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010521606702.

Anonymous:2001:ANM

[Ano01a] Anonymous. About the Nalimov Memorial issue of the journal Scientometrics. Scientometrics, 52(2):107–109, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1017986917073.

Anonymous:2001:CCVa

[Ano01b] Anonymous. Citation classics of V. V. Nalimov 1. Current Contents, Number 21, May 21, 1990. Scientometrics, 52(2): 171–174, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017951403870.

Anonymous:2001:CCVb

[Ano01c] Anonymous. Citation classics of V. V. Nalimov 2. Current Contents, Number 24, June 11, 1990. Scientometrics, 52 (2):175–177, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017903520708.

Anonymous:2001:Ea

[Ano01d] Anonymous. Editorial. Scientometrics, 50(1):3–4, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1017396817090.

Anonymous:2001:Eb

[Ano01e] Anonymous. Editorial. Scientometrics, 52(2):101–105, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1017944300235.

Anonymous:2001:IVa

[Ano01f] Anonymous. Index to volume 51. Scientometrics, 51(3):639– 642, July 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1017329225967.

Anonymous:2001:IVb

[Ano01g] Anonymous. Index to volume 52. Scientometrics, 52(3): 539–542, November 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017343321242.

Anonymous:2001:I

[Ano01h] Anonymous. Introduction. Scientometrics, 51(3):465–467, July 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1019634800743.

Anonymous:2003:CLE

[Ano03a] Anonymous. Correspondence: Letter to the Editor. Scientometrics, 57(1):141–144, May 2003. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1023631805730.

Anonymous:2003:Eb

[Ano03b] Anonymous. Editorial. Scientometrics, 58(1):3, September 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1025471322573.

Anonymous:2003:Ea

[Ano03c] Anonymous. Erratum. Scientometrics, 57(1):145–146, May 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1023683822569.

Anonymous:2003:NIP

[Ano03d] Anonymous. Newsletter from the ISSI President, August 2003. Scientometrics, 58(2):469–471, October 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1026205214647.

Anonymous:2003:O

[Ano03e] Anonymous. Obituary. Scientometrics, 57(1):5–6, May 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1023618602096.

Anonymous:2004:EA

[Ano04a] Anonymous. Editorial announcement. Scientometrics, 60(1):3, May 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/B%3ASCIE.
0000027300.55755.2d.

Anonymous:2004:LEC

[Ano04b] Anonymous. Letter to the Editor: Comments on a paper of Garg. Scientometrics, 59(2):277–278, February 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3ASCIE.0000018533.22139.9c.

Anonymous:2004:LER

[Ano04c] Anonymous. Letter to the Editor: Reply to Rousseau. Scientometrics, 59(2):279, February 2004. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1023/B%3ASCIE.0000018534.58677.b9.

Anonymous:2004:P

[Ano04d] Anonymous. Preface. Scientometrics, 60(3):281–282, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/B%3ASCIE.
0000034373.49824.3d.

Anonymous:2005:P

[Ano05] Anonymous. Preface. Scientometrics, 62(1):115–116, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-005-0006-8.

Anonymous:2006:C

[Ano06a] Anonymous. Correspondence. Scientometrics, 67(2):327– 333, May 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0104-2.

Anonymous:2006:I

[Ano06b] Anonymous. Index. Scientometrics, 66(3):619–626, March 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0031-2.

Anonymous:2007:AIV

[Ano07a] Anonymous. Author index to volume 73. Scientometrics, 73(3):361–362, December 2007. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-007-0002-2.

Anonymous:2007:IV

[Ano07b] Anonymous. Index to volume 73. Scientometrics, 73(3): 359–360, December 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-007-
0001-3.

Anonymous:2008:EA

[Ano08a] Anonymous. Editorial announcement. Scientometrics, 74 (1):3, January 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-008-
0191-3.

Anonymous:2008:KWM

[Ano08b] Anonymous. Katherine W. McCain wins the 2007 Derek John de Solla Price Medal. Scientometrics, 74(1):5–6, January 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-008-0192-2.

Anonymous:2009:PVM

[Ano09] Anonymous. Péter Vinkler and Michel Zitt win the 2009 Derek John de Solla Price Medal. Scientometrics, 81(1):1–5, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0095-x.

Aksnes:2000:VBI

[AOS00] Dag W. Aksnes, Terje Bruen Olsen, and Per O. Seglen. Validation of bibliometric indicators in the field of microbiology: A Norwegian case study. Scientometrics, 49(1):7–22, August 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005653006993.

Armando:2004:TKS

[AP04] Albert Armando and Luis M. Plaza. The transfer of knowledge from the Spanish public R&D system to the productive sectors in the field of biotechnology. Scientometrics, 59(1):3–14, January 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000013295.
88415.16.

Al-Qallaf:2003:CPK

[AQ03] Charlene L. Al-Qallaf. Citation patterns in the Kuwaiti journal Medical Principles and Practice: The first 12 years, 1989–2000. Scientometrics, 56(3):369–382, March 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1022330920925.

Atallah:2006:IPC

[AR06] Gamal Atallah and Gabriel Rodríguez. Indirect patent citations. Scientometrics, 67(3):437–465, June 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1556/Scient.67.2006.3.7.

Archambault:2002:MUP

[Arc02] Éric Archambault. Methods for using patents in crosscountry comparisons. Scientometrics, 54(1):15–30, April 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1015654903218.

Arunachalam:2001:MRI

[Aru01] Subbiah Arunachalam. Mathematics research in India today: What does the literature reveal? Scientometrics, 52 (2):235–259, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017915823434.

Aksnes:2004:EHC

[AS04] Dag W. Aksnes and Gunnar Sivertsen. The effect of highly cited papers on national citation indicators. Scientometrics, 59(2):213–224, February 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000018529.58334.eb.

Anuradha:2007:BII

[AU07] K. T. Anuradha and Shalini R. Urs. Bibliometric indicators of Indian research collaboration patterns: A correspondence analysis. Scientometrics, 71(2):179–189, May 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1657-4.

Ardanuy:2009:CAC

[AUQ09] Jordi Ardanuy, Cristóbal Urbano, and Lluís Quintana. A citation analysis of Catalan literary studies (1974–2003): Towards a bibliometrics of humanities studies in minority languages. Scientometrics, 81(2):347–366, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-008-2143-3.

Aleixandre:2004:IIS

[AVDT04] Rafael Aleixandre, Juan Carlos Valderrama, José María Desantes, and Antonio J. Torregrosa. Identification of information sources and citation patterns in the field of reciprocating internal combustion engines. Scientometrics, 59 (3):321–336, March 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000018537.
65848.2f.

Archambault:2006:BSO

[AVGC^{+}06] Éric Archambault, Étienne Vignola-Gagné, Grégoire Côté, Vincent Larivière, and Yves Gingrasb. Benchmarking scientific output in the social sciences and humanities: The limits of existing databases. Scientometrics, 68(3):329– 342, September 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0115-z.

Arvanitis:2000:SAB

[AWG00] Rigas Arvanitis, Roland Waast, and Jacques Gaillard. Science in Africa: A bibliometric panorama using PASCAL database. Scientometrics, 47(3):457–473, March 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005615816165.

Amat:2009:MAD

[AY09] C. B. Amat and A. Yegros Yegros. Median age difference of references as indicator of information update of research groups: A case study in Spanish food research. Scientometrics, 78(3):447–465, March 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1993-4.

Basu:2001:ICS

[BA01] Aparna Basu and Ritu Aggarwal. International collaboration in science in India and its impact on institutional performance. Scientometrics, 52(3):379–394, November 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1014239731175.

Bhattacharya:2007:ILI

[BA07] Sujit Bhattacharya and Praveen Arora. Industrial linkages in Indian universities: What they reveal and what they imply? Scientometrics, 70(2):277–300, February 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0204-7.

Baldini:2006:AIP

[Bal06] Nicola Baldini. The act on inventions at public research institutions: Danish universities’ patenting activity. Scientometrics, 69(2):387–407, November 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0159-0.

Baldini:2008:NEU

[Bal08] Nicola Baldini. Negative effects of university patenting: Myths and grounded evidence. Scientometrics, 75(2):289– 311, May 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1865-y.

Ball:2009:SCT

[Bal09] Rafael Ball. Scholarly communication in transition: The use of question marks in the titles of scientific articles in medicine, life sciences and physics 1966–2005. Scientometrics, 79(3):667–679, June 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-1984-5.

Banks:2006:EHI

[Ban06] Michael G. Banks. An extension of the Hirsch index: Indexing scientific topics and compounds. Scientometrics, 69 (1):161–168, October 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0146-5.

Barjak:2006:RPI

[Bar06] Franz Barjak. Research productivity in the Internet era. Scientometrics, 68(3):343–360, September 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0116-y.

Basu:2006:UIH

[Bas06] Aparna Basu. Using ISI’s ‘highly cited researchers’ to obtain a country level indicator of citation excellence. Scientometrics, 68(3):361–375, September 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0117-x.

Bayers:2005:UID

[Bay05] Nancy K. Bayers. Using ISI data in the analysis of German national and institutional research output. Scientometrics, 62(1):155–163, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
005-0010-z.

Buehring:2007:LCV

[BBG07] Gertrude Case Buehring, Jessica E. Buehring, and Patrick D. Gerard. Lost in citation: Vanishing visibility of senior authors. Scientometrics, 72(3):459–468, September 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1762-4.

Boyack:2009:MSE

[BBK09] Kevin W. Boyack, Katy Börner, and Richard Klavans. Mapping the structure and evolution of chemistry research. Scientometrics, 79(1):45–60, April 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0403-5.

Barrios:2008:BSP

[BBV^{+}08] Maite Barrios, Angel Borrego, Andreu Vilaginés, Candela Ollé, and Marta Somoza. A bibliometric study of psychological research on tourism. Scientometrics, 77(3):453–467, December 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1952-0.

Buela-Casal:2007:CSI

[BCGMBSVM07] Gualberto Buela-Casal, Olga Gutiérrez-Martínez, María Paz Bermúdez-Sánchez, and Oscar Vadillo-Muñoz. Comparative study of international academic rankings of universities. Scientometrics, 71(3):349–365, June 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1653-8.

Batista:2006:IPC

[BCK06] Pablo D. Batista, Mônica G. Campiteli, and Osame Kinouchi. Is it possible to compare researchers with different scientific interests? Scientometrics, 68(1):179–189, July 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0090-4.

Buela-Casal:2006:MIR

[BCPTC06] Gualberto Buela-Casal, Pandelis Perakakis, Michael Taylor, and Purificación Checa. Measuring internationality: Reflections and perspectives on academic journals. Scientometrics, 67(1):45–65, April 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0050-z.

Buela-Casal:2009:AIT

[BCZM^{+}09] Gualberto Buela-Casal, Izabela Zych, Ana Medina, María I. Viedma Del Jesus, Susana Lozano, and Gloria Torres. Analysis of the influence of the two types of the journal articles; theoretical and empirical on the impact factor of a journal. Scientometrics, 80(1):265–282, July 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-1715-6.

Bonaccorsi:2003:AES

[BD03] Andrea Bonaccorsi and Cinzia Daraio. Age effects in scientific productivity. Scientometrics, 58(1):49–90, September 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1025427507552.

Bonaccorsi:2005:ESA

[BD05a] Andrea Bonaccorsi and Cinzia Daraio. Exploring size and agglomeration effects on public research productivity. Scientometrics, 63(1):87–120, March 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0205-3.

Bornmann:2005:DIR

[BD05b] Lutz Bornmann and Hans-Dieter Daniel. Does the hindex for ranking of scientists really work? Scientometrics, 65(3):391–392, December 2005. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-005-0281-4.

Bornmann:2005:SRF

[BD05c] Lutz Bornmann and Hans-Dieter Daniel. Selection of research fellowship recipients by committee peer review. Reliability, fairness and predictive validity of Board of Trustees’s decisions. Scientometrics, 63(2):297–320, April 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0214-2.

Braun:2005:JGM

[BD05d] Tibor Braun and Ildikó Dióspatonyi. The journal gatekeepers of major publishing houses of core science journals. Scientometrics, 64(2):113–120, August 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0243-x.

Braun:2005:WFB

[BD05e] Tibor Braun and Ildikó Dióspatonyi. World flash on basic research. Scientometrics, 62(3):297–319, March 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0023-7.

Bornmann:2006:SSE

[BD06] Lutz Bornmann and Hans-Dieter Daniel. Selecting scientific excellence through committee peer review — a citation analysis of publications previously published to approval or rejection of post-doctoral research fellowship applicants. Scientometrics, 68(3):427–440, September 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0121-1.

Bonaccorsi:2008:DSP

[BD08] Andrea Bonaccorsi and Cinzia Daraio. The differentiation of the strategic profile of higher education institutions. New positioning indicators based on microdata. Scientometrics, 74(1):15–37, January 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-0101-8.

Bonaccorsi:2006:AIP

[BDS06] Andrea Bonaccorsi, Cinzia Daraio, and Léopold Simar. Advanced indicators of productivity of universities. An application of robust nonparametric methods to Italian data. Scientometrics, 66(2):389–410, February 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0028-x.

Braun:2007:JGI

[BDZZ07a] Tibor Braun, Ildikó Dióspatonyi, Erika Zádor, and Sándor Zsindely. Journal gatekeepers indicator-based top universities of the world, of Europe and of 29 countries — a pilot study. Scientometrics, 71(2):155–178, May 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1843-4.

Braun:2007:GIV

[BDZZ07b] Tibor Braun, Ildikó Dióspatonyi, Sándor Zsindely, and Erika Zádor. Gatekeeper index versus impact factor of science journals. Scientometrics, 71(3):541–543, June 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-1844-3.

Bornmann:2004:SOG

[BE04] Lutz Bornmann and Jürgen Enders. Social origin and gender of doctoral degree holders. Scientometrics, 61(1): 19–41, September 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000037360.
59788.fb.

Beaver:2001:RSC

[Bea01] Donald Deb. Beaver. Reflections on scientific collaboration (and its study): Past, present, and future. Scientometrics, 52(3):365–377, November 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014254214337.

Bengisu:2003:CET

[Ben03] Murat Bengisu. Critical and emerging technologies in materials, manufacturing, and industrial engineering: A study for priority setting. Scientometrics, 58(3):473–487, November 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000006875.61813.f6.

Bordons:2002:ALU

[BFG02] Maria Bordons, M. T. Fernández, and Isabel Gómez. Advantages and limitations in the use of impact factor measures for the assessment of research performance. Scientometrics, 53(2):195–206, February 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014800407876.

Bharvi:2003:SIJ

[BGB03] Dutt Bharvi, K. C. Garg, and Anita Bali. Scientometrics of the international journal Scientometrics. Scientometrics, 56(1):81–93, January 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1021950607895.

Banerjee:2000:PSI

[BGG00] Parthasarathi Banerjee, B. M. Gupta, and K. C. Garg. Patent statistics as indicators of competition an analysis of patenting in biotechnology. Scientometrics, 47(1):95–116, January 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005669810018.

Braun:2001:PCPa

[BGS01a] Tibor Braun, Wolfgang Glänzel, and András Schubert. Publication and cooperation patterns of the authors of neuroscience journals. Scientometrics, 50(3):499–510, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1012774206340.

Braun:2001:PCPb

[BGS01b] Tibor Braun, Wolfgang Glänzel, and András Schubert. Publication and cooperation patterns of the authors of neuroscience journals. Scientometrics, 51(3):499–510, July 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1019643002560.

Braun:2006:HTI

[BGS06] Tibor Braun, Wolfgang Glänzel, and András Schubert. A Hirsch-type index for journals. Scientometrics, 69(1): 169–173, October 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0147-4.

Baldini:2007:PPS

[BGS07] Nicola Baldini, Rosa Grimaldi, and Maurizio Sobrero. To patent or not to patent? A survey of Italian inventors on motivations, incentives, and obstacles to university patenting. Scientometrics, 70(2):333–354, February 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0206-5.

Bartol:2005:CCT

[BH05] Tomaz Bartol and Marjan Hocevar. The capital cities of the ten new European Union countries in selected bibliographic databases. Scientometrics, 65(2):173–187, November 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0266-3.

Bhattacharya:2004:MIA

[Bha04] Sujit Bhattacharya. Mapping inventive activity and technological change through patent analysis: A case study of India and China. Scientometrics, 61(3):361–381, November 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000045115.23375.50.

Bar-Ilan:2000:RES

[BI00] Judit Bar-Ilan. Results of an extensive search for “S&T Indicators” on the Web: A content analysis. Scientometrics, 49 (2):257–277, October 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010569119265.

Bar-Ilan:2001:DCM

[BI01a] Judit Bar-Ilan. Data collection methods on the Web for infometric purposes — a review and analysis. Scientometrics, 50(1):7–32, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005682102768.

Bjorneborn:2001:PW

[BI01b] Lennart Björneborn and Peter Ingwersen. Perspective of webometrics. Scientometrics, 50(1):65–82, January 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005642218907.

Bar-Ilan:2004:MLA

[BI04] Judit Bar-Ilan. A microscopic link analysis of academic institutions within a country — the case of Israel. Scientometrics, 59(3):391–403, March 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000018540.33706.c1.

Bar-Ilan:2008:IIO

[BI08a] Judit Bar-Ilan. The h-index of h-index and of other informetric topics. Scientometrics, 75(3):591–605, June 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-1880-z.

Bar-Ilan:2008:WIC

[BI08b] Judit Bar-Ilan. Which h-index? — a comparison of WoS, Scopus and Google Scholar. Scientometrics, 74(2):257–271, February 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-0216-y.

Bar-Ilan:2005:ASW

[BIE05] Judit Bar-Ilan and Ana Echermane. The anthrax scare and the Web: A content analysis of Web pages linking to resources on anthrax. Scientometrics, 63(3):443–462, June 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0224-0.

Biglu:2008:IRP

[Big08] Mohammad Hossein Biglu. The influence of references per paper in the SCI to Impact Factors and the Matthew Effect. Scientometrics, 74(3):453–470, March 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1815-8.

Bar-Ilan:2009:LIW

[BIP09] Judit Bar-Ilan and Bluma C. Peritz. The lifespan of “informetrics” on the Web: An eight year study (1998– 2006). Scientometrics, 79(1):7–25, April 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0401-7.

Bjorneborn:2006:MSW

[Bjö06] Lennart Björneborn. ‘Mini small worlds’ of shortest link paths crossing domain boundaries in an academic Web space. Scientometrics, 68(3):395–414, September 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0119-8.

Basu:2000:ICI

[BK00] Aparna Basu and B. S. Vinu Kumar. International collaboration in Indian scientific papers. Scientometrics, 48(3):381– 402, July 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005692505687.

Balaban:2002:CAR

[BK02] Alexandru T. Balaban and Douglas J. Klein. Co-authorship, rational Erdős numbers, and resistance distances in graphs. Scientometrics, 55(1):59–70, September 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016098803527.

Balaban:2006:CCS

[BK06] Alexandru T. Balaban and Douglas J. Klein. Is chemistry ‘the central science’? How are different sciences related? Co-citations, reductionism, emergence, and posets. Scientometrics, 69(3):615–637, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0173-2.

Boyack:2005:MBS

[BKB05] Kevin W. Boyack, Richard Klavans, and Katy Börner. Mapping the backbone of science. Scientometrics, 64(3): 351–374, August 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0255-6.

Bettencourt:2008:PME

[BKK^{+}08] Luís M. A. Bettencourt, David I. Kaiser, Jasleen Kaur, Carlos Castillo-Chávez, and David E. Wojick. Population modeling of the emergence and development of scientific fields. Scientometrics, 75(3):495–518, June 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1888-4.

Bhattacharya:2003:CIS

[BKM03] Sujit Bhattacharya, Hildrun Kretschmer, and Martin Meyer. Characterizing intellectual spaces between science and technology. Scientometrics, 58(2):369–390, October 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1026244828759.

Belkhodja:2007:THC

[BL07] Omar Belkhodja and Réjean Landry. “The triple-helix collaboration: Why do researchers collaborate with industry and the government? What are the factors that influence the perceived barriers?”. Scientometrics, 70(2):301–332, February 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-0205-6.

Blatt:2009:DDV

[Bla09] Eli M. Blatt. Differentiating, describing, and visualizing scientific space: A novel approach to the analysis of published scientific abstracts. Scientometrics, 80(2):385–406, August 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2070-3.

Bassecoulard:2007:MNC

[BLZ07] Elise Bassecoulard, Alain Lelu, and Michel Zitt. Mapping nanosciences by citation flows: A preliminary analysis. Scientometrics, 70(3):859–880, March 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0315-1.

Bhattacharya:2003:LFS

[BM03] Sujit Bhattacharya and Martin Meyer. Large firms and the science-technology interface patents, patent citations, and scientific output of multinational corporations in thin films. Scientometrics, 58(2):265–279, October 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1026284510104.

Bouabid:2009:EMR

[BM09] Hamid Bouabid and Ben R. Martin. Evaluation of Moroccan research using a bibliometric-based approach: investigation of the validity of the h-index. Scientometrics, 78(2):203–217, February 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-2005-4.

Bornmann:2007:RCR

[BMD07] Lutz Bornmann, Rüdiger Mutz, and Hans-Dieter Daniel. Row-column (RC) association model applied to grant peer review. Scientometrics, 73(2):139–147, November 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-006-1797-y.

Bornmann:2009:IAG

[BMD09] Lutz Bornmann, Rüdiger Mutz, and Hans-Dieter Daniel. The influence of the applicants’ gender on the modeling of a peer review process by using latent Markov models. Scientometrics, 81(2):407–411, November 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2189-2.

Bordons:2003:OSF

[BMFG03] María Bordons, Fernanda Morillo, M. Teresa Fernández, and Isabel Gómez. One step further in the production of bibliometric indicators at the micro level: Differences by gender and professional category of scientists. Scientometrics, 57 (2):159–173, June 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1024181400646.

Benito:2005:BSD

[BMFT05] Juana Gómez Benito, María Dolores Hidalgo Montesinos, Georgina Guilera Ferré, and Macarena Moreno Torrente. A bibliometric study of differential item functioning. Scientometrics, 64(1):3–16, July 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
005-0234-y.

Bailon-Moreno:2007:PSS

[BMJARB^{+}07] Rafael Bailón-Moreno, Encarnación Jurado-Alameda, Rosario Ruiz-Baños, Jean Pierre Courtial, and Evaristo Jiménez-Contreras. The pulsing structure of science: Ortega y Gasset, Saint Matthew, fractality and transfractality. Scientometrics, 71(1):3–24, April 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-1600-8.

Bailon-Moreno:2005:AFP

[BMJARBC05a] R. Bailón-Moreno, E. Jurado-Alameda, R. Ruiz-Baños, and J. P. Courtial. Analysis of the field of physical chemistry of surfactants with the Unified Scientometric Model. Fit of relational and activity indicators. Scientometrics, 63 (2):259–276, April 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0212-4.

Bailon-Moreno:2005:BLE

[BMJARBC05b] R. Bailón-Moreno, E. Jurado-Alameda, R. Ruiz-Baños, and J. P. Courtial. Bibliometric laws: Empirical flaws of fit. Scientometrics, 63(2):209–229, April 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0211-5.

Bailon-Moreno:2005:USM

[BMJARBC05c] R. Bailón-Moreno, E. Jurado-Alameda, R. Ruiz-Baños, and J. P. Courtial. The unified scientometric model. Fractality and transfractality. Scientometrics, 63(2):231–257, April 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0221-3.

Ball:2009:CJB

[BMT09] Rafael Ball, Bernhard Mittermaier, and Dirk Tunger. Creation of journal-based publication profiles of scientific institutions — a methodology for the interdisciplinary comparison of scientific research based on the J-factor. Scientometrics, 81(2):381–392, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-2120-5.

Buter:2001:IFI

[BN01] Renald K. Buter and Ed C. M. Noyons. Improving the functionality of interactive bibliometric science maps. Scientometrics, 51(1):55–68, April 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010560527236.

Bonnevie-Nebelong:2006:JCI

[BN06] Ellen Bonnevie-Nebelong. Journal citation identity, journal citation image and internationalisation: Methods for journal evaluation. Scientometrics, 66(2):411–424, February 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0029-9.

Bornmann:2008:DER

[BND08] Lutz Bornmann, Irina Nast, and Hans-Dieter Daniel. Do editors and referees look for signs of scientific misconduct when reviewing manuscripts? A quantitative content analysis of studies that examined review criteria and reasons for accepting and rejecting manuscripts for publication. Scientometrics, 77(3):415–432, December 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1950-2.

Buter:2006:CCM

[BNVL06] R. K. Buter, E. C. M. Noyons, M. Van Mackelenbergh, and T. Laine. Combining concept maps and bibliometric maps: First explorations. Scientometrics, 66(2):377–387, February 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0027-y.

BasultoSantos:2005:MCA

[BO05] Jesús Basulto Santos and Francisco Javier Ortega Irizo. Modelling citation age data with right censoring. Scientometrics, 62(3):329–342, March 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
005-0025-5.

Bonitz:2003:ORK

[Bon03] Manfred Bonitz. Obituary: Robert King Merton (1910– 2003). Scientometrics, 57(2):151–152, June 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1024146632506.

Bonitz:2004:SEP

[Bon04] Manfred Bonitz. Self-emancipation proclamation and a light-hearted but nevertheless deeply-felt exception. Scientometrics, 60(1):19–24, May 2004. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/B%
3ASCIE.0000027303.64409.9b.

Bonitz:2005:KFE

[Bon05a] Manfred Bonitz. Klaus Fuchs — the enduring contribution to physics from his British period. Scientometrics, 62 (3):343–350, March 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0026-4.

Bonitz:2005:TYM

[Bon05b] Manfred Bonitz. Ten years Matthew effect for countries. Scientometrics, 64(3):375–379, August 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0256-5.

Boshoff:2009:NCR

[Bos09] Nelius Boshoff. Neo-colonialism and research collaboration in central Africa. Scientometrics, 81(2):413–434, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2211-8.

Boyack:2009:UDM

[Boy09] Kevin W. Boyack. Using detailed maps of science to identify potential collaborations. Scientometrics, 79(1):27–44, April 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0402-6.

Bhattacharya:2000:IFA

[BPA00] Sujit Bhattacharya, Chandra Pal, and Jagdish Arora. Inside the frontier areas of research in physics: A micro level analysis. Scientometrics, 47(1):131–142, January 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005626027764.

Buranathiti:2009:RRS

[BPM^{+}09] T. Buranathiti, N. Premkamolnetr, T. Markpin, P. Ratchatahirun, W. Yochai, and N. Sombatsompop. Redistributed random sampling method for categorizing materials research publications from SCI database: Metallurgy and polymer subfields. Scientometrics, 78(1):13–21, January 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-1764-x.

Borner:2006:MDS

[BPMK06] Katy Börner, Shashikant Penumarthy, Mark Meiss, and Weimao Ke. Mapping the diffusion of scholarly knowledge among major U.S. research institutions. Scientometrics, 68(3):415–426, September 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
006-0120-2.

Balaban:2000:PUU

[BR00] Alexandru T. Balaban and Milan Randić. Proposal for using an untapped source of citations characterizing scientific areas. Scientometrics, 49(3):517–521, November 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010597925625.

Braun:2004:E

[Bra04a] T. Braun. Editorial. Scientometrics, 61(2):1, October 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/B%3ASCIE.
0000041644.59602.f3.

Braun:2004:EF

[Bra04b] Tibor Braun. Editorial foreword. Scientometrics, 60 (1):9, May 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/B%3ASCIE.
0000027301.70522.4c.

Braam:2009:EAG

[Bra09] Robert Braam. Everything about genes: Some results on the dynamics of genomics research. Scientometrics, 79(1):61– 77, April 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0404-4.

Brown:2004:MEB

[Bro04] Cecelia Brown. The Matthew Effect of the Annual Reviews series and the flow of scientific communication through the World Wide Web. Scientometrics, 60(1):25–30, May 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/B%3ASCIE.0000027304.80068.0c.

Belinchon:2004:DSP

[BRSYB04] Isabel Belinchón, José Manuel Ramos, Evaristo Sánchez-Yus, and Isabel Betlloch. Dermatological scientific production from European Union authors (1987–2000). Scientometrics, 61(2):271–281, October 2004. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/B%
3ASCIE.0000041652.90675.d6.

Brunk:2003:SIF

[Bru03] Gregory G. Brunk. Swarming of innovations, fractal patterns, and the historical time series of US patents. Scientometrics, 56(1):61–80, January 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1021998523825.

Bogaert:2000:PMI

[BRV00] Jan Bogaert, Ronald Rousseau, and Piet Van Hecke. Percolation as a model for informetric distributions: Fragment size distribution characterised by Bradford curves. Scientometrics, 47(2):195–206, February 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005678707987.

Bollen:2006:JS

[BRV06] Johan Bollen, Marko A. Rodriquez, and Herbert Van de Sompel. Journal status. Scientometrics, 69(3):669–687, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-006-0176-z.pdf.

Bonitz:2001:CSM

[BS01] Manfred Bonitz and Andrea Scharnhorst. Competition in science and the Matthew core journals. Scientometrics, 51 (1):37–54, April 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010508510398.

Braun:2003:QVC

[BS03] Tibor Braun and András Schubert. A quantitative view on the coming of age of interdisciplinarity in the sciences 1980– 1999. Scientometrics, 58(1):183–189, September 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1025439910278.

Braun:2007:GRI

[BS07] Tibor Braun and András Schubert. The growth of research on inter-and multidisciplinarity in science and social science papers, 1975–2006. Scientometrics, 73(3):345–351, December 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1933-3.

Braun:2003:AAS

[BSPKN03a] Tibor Braun, Zsuzsa Szabadi-Peresztegi, and Éva Kovács-Németh. About Abels and similar international awards for ranked lists of awardees as science indicators of national merit in mathematics. Scientometrics, 56(2):161–168, January 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1021911026551.

Braun:2003:NBA

[BSPKN03b] Tibor Braun, Zsuzsa Szabadi-Peresztegi, and Éva Kovács-Németh. No-bells for ambiguous lists of ranked Nobelists as science indicators of national merit in physics, chemistry and medicine, 1901–2001. Scientometrics, 56(1):3–28, January 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1021998006078.

Ball:2006:BAN

[BT06] Rafael Ball and Dirk Tunger. Bibliometric analysisa new business area for information professionals in libraries? Scientometrics, 66(3):561–577, March 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0041-0.

Barcza:2009:PPP

[BT09] Krisztina Barcza and András Telcs. Paretian publication patterns imply Paretian Hirsch index. Scientometrics, 81(2):513–519, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
008-2175-8.

Burrel:2001:SMF

[Bur01a] Quentin L. Burrel. Stochastic modelling of the first-citation distribution. Scientometrics, 52(1):3–12, September 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1012751509975.

Burrell:2001:SRP

[Bur01b] Quentin L. Burrell. Some remarks on a paper by Egghe. Scientometrics, 50(2):345–350, February 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010586026266. See [Egg99].

Burrell:2001:TPP

[Bur01c] Quentin L. Burrell. Two problems posed by Egghe. Scientometrics, 50(3):539–544, January 2001. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%

3A1010571017605. See comments [Egg01a, Egg01b].

Burrell:2002:MCA

[Bur02a] Quentin L. Burrell. Modelling citation age data: Simple graphical methods from reliability theory. Scientometrics, 55(2):273–285, August 2002. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1019671808921.

Burrell:2002:CDO

[Bur02b] Quentin L. Burrell. The n th-citation distribution and obsolescence. Scientometrics, 53(3):309–323, March 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014816911511.

Burrell:2003:DCT

[Bur03] Quentin L. Burrell. Defining a core: Theoretical observations on the Egghe–Rousseau proposal. Scientometrics, 57(1):75–92, May 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023623603913.

Burrell:2005:SBE

[Bur05a] Quentin L. Burrell. Are “sleeping beauties” to be expected? Scientometrics, 65(3):381–389, December 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-005-0280-5.

Burrell:2005:UGW

[Bur05b] Quentin L. Burrell. The use of the generalized Waring process in modelling informetric data. Scientometrics, 64(3): 247–270, August 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0250-y.

Burrell:2006:MCW

[Bur06a] Quentin L. Burrell. Measuring concentration within and coconcentration between informetric distributions: An empirical study. Scientometrics, 68(3):441–456, September 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-006-0122-0.

Burrell:2006:ULF

[Bur06b] Quentin L. Burrell. The use of Lotka functions and systematic sampling. Scientometrics, 67(2):323–325, May 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-006-0103-3.

Burrell:2007:HIH

[Bur07a] Quentin L. Burrell. Hirsch index or Hirsch rate? Some thoughts arising from Liang’s data. Scientometrics, 73(1): 19–28, October 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-1774-5.

Burrell:2007:TDA

[Bur07b] Quentin L. Burrell. Time-dependent aspects of coconcentration in informetrics. Scientometrics, 73(2):161– 174, November 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1688-x.

Burrell:2009:HEK

[Bur09] Quentin L. Burrell. On Hirsch’s h, Egghe’s g and Kosmulski’s h(2). Scientometrics, 79(1):79–91, April 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0405-3.

Butler:2002:IHR

[But02] Linda Butler. Identifying “highly-rated” journals — an Australian case study. Scientometrics, 53(2):207–227, February 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1014852424715.

Butler:2008:IAM

[But08] Linda Butler. ICT assessment: Moving beyond journal outputs. Scientometrics, 74(1):39–55, January 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-0102-7.

Butler:2006:ECA

[BV06] Linda Butler and Martijn S. Visser. Extending citation analysis to non-source items. Scientometrics, 66(2):327–343, February 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0024-1.

Bollen:2006:MSS

[BvdS06] Johan Bollen and Herbert van de Sompel. Mapping the structure of science through usage. Scientometrics, 69(2): 227–258, November 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0151-8.

Braun:2007:GPN

[BZDZ07] Tibor Braun, Sándor Zsindely, Ildikó Dióspatonyi, and Erika Zádor. Gatekeeping patterns in nano-titled journals. Scientometrics, 70(3):651–667, March 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0306-2.

Cantu:2009:ODA

[CA09] Anselmo Garcia Cantú and Marcel Ausloos. Organizational and dynamical aspects of a small network with two distinct communities: Neo-creationists vs. Evolution defenders. Scientometrics, 80(2):457–472, August 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2065-0.

Cahlik:2000:CMS

[Cah00a] Tomas Cahlik. Comparison of the maps of science. Scientometrics, 49(3):373–387, November 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010581421990.

Cahlik:2000:SFA

[Cah00b] Tomas Cahlik. Search for fundamental articles in economics. Scientometrics, 49(3):389–402, November 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1010533506061.

Calvino:2006:ARP

[Cal06] Amalia Mirta Calviño. Assessment of research performance in food science and technology: Publication behavior of five Iberian–American countries (1992–2003). Scientometrics, 69(1):103–116, October 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0141-x.

Campanario:2009:RRN

[Cam09] Juan Miguel Campanario. Rejecting and resisting Nobel class discoveries: accounts by Nobel laureates. Scientometrics, 81(2):549–565, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2141-5.

Cozzens:2003:MRB

[CB03] Susan E. Cozzens and Kamau Bobb. Measuring the relationship between high technology development strategies and wage inequality. Scientometrics, 58(2):351–368, October 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1026292711921.

Costas:2008:IBT

[CB08a] Rodrigo Costas and María Bordons. Is g-index better than h-index? An exploratory study at the individual level. Scientometrics, 77(2):267–288, November 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1997-0.

Curras:2008:IEH

[CB08b] Emilia Currás and Enrique Wulff Barreiro. Integration in Europe of human genetics results obtained by Spaniards in the USA: A historical perspective. Scientometrics, 75 (3):473–493, June 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1861-2.

Calver:2009:SWU

[CB09] Michael C. Calver and J. Stuart Bradley. Should we use the mean citations per paper to summarise a journal’s impact or to rank journals in the same field? Scientometrics, 81(3): 611–615, December 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-2229-y.

Coutinho:2003:IPP

[CBO^{+}03] Marilia Coutinho, Elizabeth Balbachevsky, Denilde Oliveira Holzhacker, Diogo Da Costa Patrao, Ricardo Nicoliello Zorzetto Vencio, Rodrigo Luiz Medeiros Da Silva, Márcio Luís Gomes Lucatelli, Lucimara Flávio Dos Reis, and Maria Angélica Marin. Intellectual property and public research in biotechnology: the scientists opinion. Scientometrics, 58(3):641–656, November 2003. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/B%
3ASCIE.0000006885.67651.9e.

Calero:2006:HIR

[CBVN06] Clara Calero, Renald Buter, Cecilia Cabello Valdés, and Ed Noyons. How to identify research groups using publication analysis: an example in the field of nanotechnology. Scientometrics, 66(2):365–376, February 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0026-z.

Csajbok:2007:HIC

[CBVS07] Edit Csajbók, Anna Berhidi, Lívia Vasas, and András Schubert. Hirsch-index for countries based on essential science indicators data. Scientometrics, 73(1):91–117, October 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1859-9.

Chavalarias:2008:BSF

[CC08] David Chavalarias and Jean-Philippe Cointet. Bottom-up scientific field detection for dynamical and hierarchical science mapping, methodology and case study. Scientometrics, 75(1):37–50, April 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1825-6.

Chen:2009:UNN

[CC09] Yu-Shan Chen and Ke-Chiun Chang. Using neural network to analyze the influence of the patent performance upon the market value of the US pharmaceutical companies. Scientometrics, 80(3):637–655, September 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-2095-2.

Claude:2004:BOU

[CCDJJF04] Robert Claude, Arreto Charles-Daniel, Azerad Jean, and Gaudy Jean-François. Bibliometric overview of the utilization of artificial neural networks in medicine and biology. Scientometrics, 59(1):117–130, January 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000013302.59845.34.

Chen:2006:RCG

[CCHC06] Tzeng-Ji Chen, Yu-Chun Chen, Shinn-Jang Hwang, and Li-Fang Chou. The rise of China in gastroenterology? A bibliometric analysis of ISI and Medline databases. Scientometrics, 69(3):539–549, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0177-y.

Chen:2007:CHK

[CCHC07] Tzeng-Ji Chen, Yu-Chun Chen, Shinn-Jang Hwang, and Li-Fang Chou. The contribution of Hong Kong to China’s international scientific publications. Scientometrics, 71(1): 151–154, April 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1654-7.

Chen:2005:CTK

[CCHF05] Dar-Zen Chen, Han-Wen Chang, Mu-Hsuan Huang, and Feng-Cheng Fu. Core technologies and key industries in Taiwan from 1978 to 2002: A perspective from patent analysis. Scientometrics, 64(1):31–53, July 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0236-9.

Carvalho:2005:IFP

[CDFB05] Priscilla Carvalho, José Alexandre Felizola Diniz-Filho, and Luis Mauricio Bini. The impact of Felsenstein’s “phylogenies and the comparative method” on evolutionary biology. Scientometrics, 62(1):53–66, January 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0003-y.

Ceci:2006:NIC

[CDR06] Anna Ceci, Mario De Marchi, and Maurizio Rocchi. A note on innovation in the chemical industy in Italy. Scientometrics, 69(3):607–614, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0172-3.

Contreras:2006:CIF

[CEM06] Claudia Contreras, Gonzalo Edwards, and Alejandra Mizala. The Current Impact Factor and the long-term impact of scientific journals by discipline: A logistic diffusion model estimation. Scientometrics, 69(3):689–696, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0165-2.

Cherny:2001:IVV

[CG01] Arkady I. Cherny and Ruggero S. Gilyarevsky. The impact of V. V. Nalimov on information science. Scientometrics, 52 (2):159–163, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017995118891.

Campanario:2006:JSC

[CG06] Juan Miguel Campanario and Lidia González. Journal selfcitations that contribute to the impact factor: Documents labeled “editorial material” in journals covered by the Science Citation Index. Scientometrics, 69(2):365–386, November 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0158-1.

Claveria:2000:DPR

[CGC^{+}00] L. Erik Clavería, Eliseo Guallar, Jordi Camí, José Conde, Roberto Pastor, José R. Ricoy, Eduardo Rodríguez-Farré, Fernando Ruiz-Palomo, and Emilio Muñoz. Does peer review predict the performance of research projects in health sciences? Scientometrics, 47(1):11–23, January 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005609624130.

Campanario:2006:SIF

[CGR06] Juan Miguel Campanario, Lidia González, and Cristina Rodríguez. Structure of the impact factor of academic journals in the field of education and educational psychology: Citations from editorial board members. Scientometrics, 69 (1):37–56, October 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0137-6.

Cassiman:2007:MIS

[CGV07] Bruno Cassiman, Patrick Glenisson, and Bart Van Looy. Measuring industry-science links through inventor-author relations: A profiling methodology. Scientometrics, 70(2): 379–391, February 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-0208-3.

Choung:2000:NSI

[CH00] Jae-Yong Choung and Hye-Ran Hwang. National systems of innovation: Institutional linkages and performances in the case of Korea and Taiwan. Scientometrics, 48(3):413– 426, July 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005696606595.

Chen:2004:TKD

[CH04] Chaomei Chen and Diana Hicks. Tracing knowledge diffusion. Scientometrics, 59(2):199–211, February 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000018528.59913.48.

Chiu:2005:BAH

[CH05] Wen-Ta Chiu and Yuh-Shan Ho. Bibliometric analysis of homeopathy research during the period of 1991 to 2003. Scientometrics, 63(1):3–23, March 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0201-7.

Chiu:2007:BAT

[CH07] Wen-Ta Chiu and Yuh-Shan Ho. Bibliometric analysis of tsunami research. Scientometrics, 73(1):3–17, October 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-1523-1.

Chiu:2004:BAS

[CHH04] Wen-Ta Chiu, Jing-Shan Huang, and Yuh-Shan Ho. Bibliometric analysis of severe acute respiratory syndromerelated research in the beginning stage. Scientometrics, 61(1):69–77, September 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/content/pdf/10.1023/B%3ASCIE.
0000037363.49623.28.pdf.

Chuang:2007:BCA

[CHH07] Kun-Yang Chuang, Ya-Li Huang, and Yuh-Shan Ho. A bibliometric and citation analysis of stroke-related research in Taiwan. Scientometrics, 72(2):201–212, August 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1721-0.

Cho:2008:NEI

[Cho08] Soo-Ryun Cho. New evaluation indexes for articles and authors’ academic achievements based on open access resources. Scientometrics, 77(1):91–112, October 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1834-5.

Christoffersen:2004:ICD

[Chr04] Mikkel Christoffersen. Identifying core documents with a multiple evidence relevance filter. Scientometrics, 61(3): 385–394, November 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000045116.
51534.5c.

Chu:2001:IAIa

[Chu01a] Heting Chu. Intellectual activities and influences of Belver C. Griffith: A citation perspective. Scientometrics, 50(3): 481–488, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1012770105431.

Chu:2001:IAIb

[Chu01b] Heting Chu. Intellectual activities and influences of Belver C. Griffith: A citation perspective. Scientometrics, 51 (3):481–488, July 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1019638901652.

Costas:2007:VCF

[CIM07] Rodrigo Costas and Isabel Iribarren-Maestro. Variations in content and format of ISI databases in their different versions: The case of the Science Citation Index in CD–ROM and the Web of Science. Scientometrics, 72(2):167–183, August 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1589-z.

Cahlik:2006:LCA

[CJ06] Tomás Cahlík and Marcel Jirina. Law of cumulative advantages in the evolution of scientific fields. Scientometrics, 66 (3):441–449, March 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0032-1.

Cheng:2006:FAC

[CL06] Ying Cheng and Nian Cai Liu. A first approach to the classification of the top 500 world universities by their disciplinary characteristics using scientometrics. Scientometrics, 68(1):135–150, July 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0087-z.

Lo:2007:PAG

[cL07] Szu chia Lo. Patent analysis of genetic engineering research in Japan, Korea and Taiwan. Scientometrics, 70(1): 183–200, January 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-0111-y.

Chen:2007:UEP

[CLH07] Dar-Zen Chen, Wen-Yau Cathy Lin, and Mu-Hsuan Huang. Using essential patent index and essential technological strength to evaluate industrial technological innovation competitiveness. Scientometrics, 71(1):101–116, April 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-1655-6.

Cardona:2005:DNP

[CM05] Manuel Cardona and Werner Marx. The disaster of the nazipower in science as reflected by some leading journals and scientists in physics. Scientometrics, 64(3):313–324, August 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0253-8.

Campanario:2009:SBT

[CM09] Juan Miguel Campanario and Antonio Molina. Surviving bad times: The role of citations, self-citations and numbers of citable items in recovery of the journal impact factor after at least four years of continuous decreases. Scientometrics, 81(3):859–864, December 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-008-2257-7.

Castro-Martinez:2009:STP

[CMJSOC09] Elena Castro-Martínez, Fernando Jiménez-Sáez, and Francisco Javier Ortega-Colomer. Science and technology policies: A tale of political use, misuse and abuse of traditional R&D indicators. Scientometrics, 80(3):827–844, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-2132-6.

Chan:2004:MAV

[CMM04] Felix Chan, Dora Marinova, and Michael McAleer. Modelling the asymmetric volatility of anti-pollution patents in the USA. Scientometrics, 59(2):179–197, February 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/B%3ASCIE.0000018527.22276.10.

Choung:2003:PKP

[CMP03] Jae-Yong Choung, Hong-Ghi Min, and Myeong-Cheol Park. Patterns of knowledge production: The case of information and telecommunication sector in Korea. Scientometrics, 58(1):115–128, September 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1025431608461.

Catling:2009:QEB

[CMU09] Jonathan C. Catling, Victoria L. Mason, and Dominic Upton. Quality is in the eye of the beholder? An evaluation of impact factors and perception of journal prestige in the UK. Scientometrics, 81(2):333–345, November 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-2124-1.

Coccia:2005:SMA

[Coc05] Mario Coccia. A scientometric model for the assessment of scientific research performance within public institutes. Scientometrics, 65(3):307–321, December 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0276-1.

Coccia:2009:RPB

[Coc09] Mario Coccia. Research performance and bureaucracy within public research labs. Scientometrics, 79(1):93–107, April 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0406-2.

Cole:2004:RKM

[Col04] Jonathan R. Cole. Robert K. Merton, 1910–2003. Scientometrics, 60(1):37–40, May 2004. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/B%
3ASCIE.0000027305.61662.68.

Cozzarin:2006:PMS

[Coz06] Brian P. Cozzarin. Performance measures for the socioeconomic impact of government spending on R&D. Scientometrics, 68(1):41–71, July 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
006-0083-3.

Chen:2008:QMB

[CR08] Lixin Chen and Ronald Rousseau. Q-measures for binary divided networks: Bridges between German and English institutes in publications of the Journal of Fluid Mechanics. Scientometrics, 74(1):57–69, January 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-0103-6.

Criscuolo:2006:HAE

[Cri06] Paola Criscuolo. The ‘home advantage’ effect and patent families. A comparison of OECD triadic patents, the USPTO and the EPO. Scientometrics, 66(1):23–41, January 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0003-6.

Collazo-Reyes:2004:PCP

[CRLMR04] Francisco Collazo-Reyes, Ma. Elena Luna-Morales, and Jane M. Russell. Publication and citation patterns of the Mexican contribution to a “big science” discipline: Elementary particle physics. Scientometrics, 60(2):131–143, June 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000027676.41604.e9.

Collazo-Reyes:2008:PCP

[CRLMRPA08] F. Collazo-Reyes, M. E. Luna-Morales, J. M. Russell, and M. A. Pérez-Angón. Publication and citation patterns of Latin American & Caribbean journals in the SCI and SSCI from 1995 to 2004. Scientometrics, 75(1):145–161, April 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1841-6.

Cronin:2004:NSS

[Cro04] Blaise Cronin. Normative shaping of scientific practice: The magic of Merton. Scientometrics, 60(1):41–46, May 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/B%3ASCIE.0000027306.30138.49.

Cronin:2002:ICI

[CS02] Blaise Cronin and Debora Shaw. Identity-creators and image-makers: Using citation analysis and thick description to put authors in their place. Scientometrics, 54(1):31–49, April 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1015628320056.

Cervantes:2009:HLM

[CSGGB09] Víctor H. Cervantes, Ana Cristina Santana, Georgina Guilera, and Juana Gómez-Benito. Hierarchical linear models in psychiatry: A bibliometric study. Scientometrics, 80(3): 797–808, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-2121-4.

Chen:2009:IIR

[CSW^{+}09] Chuanfu Chen, Kai Sun, Gang Wu, Qiong Tang, Jian Qin, Kuei Chiu, Yushuang Fu, Xiaofang Wang, and Jing Liu. The impact of Internet resources on scholarly communication: A citation analysis. Scientometrics, 81(2):459–474, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2180-y.

Calero:2007:RCW

[CvLT07] Clara Calero, Thed N. van Leeuwen, and Robert J. W. Tijssen. Research cooperation within the bio-pharmaceutical industry: Network analyses of co-publications within and between firms. Scientometrics, 71(1):87–99, April 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-007-1650-y.pdf.

Cincera:2006:AFC

[CvPdlPR06] Michele Cincera, Bruno van Pottelsberghe de la Potterie, and Veugelers Reinhilde. Assessing the foreign control of production of technology: The case of a small open economy. Scientometrics, 66(3):493–512, March 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0036-x.

Callaert:2006:TPA

[CVV^{+}06] Julie Callaert, Bart Van Looy, Arnold Verbeek, Koenraad Debackere, and Bart Thijs. Traces of prior art: An analysis of non-patent references found in patent documents. Scientometrics, 69(1):3–20, October 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0135-8.

Chu:2009:WDS

[CX09] Heting Chu and Chen Xu. Web 2.0 and its dimensions in the scholarly world. Scientometrics, 80(3):717–729, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2103-y.

Chung:2009:CLS

[CYKK09] Young Mee Chung, So Young Yu, Yong Kwang Kim, and Su Yeon Kim. Characteristics and link structure of a national scholarly Web space: The case of South Korea. Scientometrics, 80(3):595–612, September 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2091-y.

Davarpanah:2008:SAI

[DA08] M. R. Davarpanah and S. Aslekia. A scientometric analysis of international LIS journals: Productivity and characteristics. Scientometrics, 77(1):21–39, October 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1803-z.

Daizadeh:2007:IUP

[Dai07] Iraj Daizadeh. Issued US patents, patent-related global academic and media publications, and the US market indices are inter-correlated, with varying growth patterns. Scientometrics, 73(1):29–36, October 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1749-1.

Daizadeh:2009:IPB

[Dai09] Iraj Daizadeh. An intellectual property-based corporate strategy: An R&D spend, patent, trademark, media communication, and market price innovation agenda. Scientometrics, 80(3):731–746, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2105-9.

Dalpe:2002:BAB

[Dal02] Robert Dalpé. Bibliometric analysis of biotechnology. Scientometrics, 55(2):189–213, August 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1019663607103.

Danell:2000:SAJ

[Dan00] Richard Danell. Stratification among journals in management research: A bibliometric study of interaction between European and American journals. Scientometrics, 49(1): 23–38, August 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005605123831.

Dang:2006:FAD

[Dan06] Yaru Dang. Fluctuation analysis of discipline development based on impact factor. Scientometrics, 67(2):175–186, May 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0092-2.

Dastidar:2004:OST

[Das04] Prabir G. Dastidar. Ocean science & technology research across the countries: A global scenario. Scientometrics, 59 (1):15–27, January 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000013296.
33677.ef.

Davis:2003:IN

[Dav03] Mari Davis. ISSI newsletter. Scientometrics, 57(1):147–149, May 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1023635906639.

Beaver:2004:DCR

[dB04] Donald deB. Beaver. Does collaborative research have greater epistemic authority? Scientometrics, 60(3):399–408, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000034382.85360.cd.

Dietz:2000:UCV

[DCB^{+}00] James S. Dietz, Ivan Chompalov, Barry Bozeman, Eliesh O’Neil Lane, and Jongwon Park. Using the curriculum vita to study the career paths of scientists and engineers: An exploratory assessment. Scientometrics, 49(3):419– 442, November 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010537606969.

Ding:2000:JMI

[DCF00] Ying Ding, Gobinda G. Chowdhury, and Schubert Foo. Journal as markers of intellectual space: Journal cocitation analysis of information retrieval area, 1987–1997. Scientometrics, 47(1):55–73, January 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005665709109.

Dewett:2004:ESR

[DD04] Todd Dewett and Angelo S. Denisi. Exploring scholarly reputation: It’s more than just productivity. Scientometrics, 60(2):249–272, June 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000027796.
55585.61.

Danell:2009:PAC

[DD09] Jenny-Ann Brodin Danell and Rickard Danell. Publication activity in complementary and alternative medicine. Scientometrics, 80(2):539–551, August 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2078-8.

Dore:2000:MAT

[DDM00] Jean-Christophe Doré, Christian Dutheuil, and Jean-François Miquel. Multidimensional analysis of trends in patent activity. Scientometrics, 47(3):475–492, March 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1005667800235.

DeBra:2000:UHM

[De 00] Paul De Bra. Using hypertext metrics to measure research output levels. Scientometrics, 47(2):227–236, February 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005682808896.

DeMarchi:2005:NB

[De 05] Mario De Marchi. Nota bene. Scientometrics, 63(2): 403–405, April 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0219-x.

Davis:2007:DAL

[DF07] Philip M. Davis and Michael J. Fromerth. Does the arXiv lead to higher citations and reduced publisher downloads for mathematics articles? Scientometrics, 71(2):203–215, May 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1661-8.

Duncan:2002:CSE

[DFRG02] W. Jack Duncan, Eric W. Ford, Matthew D. Rousculp, and Peter M. Ginter. Community of scholars: An exploratory study of management laureates. Scientometrics, 55(3):395– 409, November 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1020414829761.

Debackere:2004:UBA

[DG04] Koenraad Debackere and Wolfgang Glänzel. Using a bibliometric approach to support research policy making: The case of the Flemish BOF-key. Scientometrics, 59(2):253– 276, February 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000018532.
70146.02.

Dahdouh-Guebas:2003:NCS

[DGAVK03] Farid Dahdouh-Guebas, J. Ahimbisibwe, Rita Van Moll, and Nico Koedam. Neo-colonial science by the most industrialised upon the least developed countries in peer-reviewed publishing. Scientometrics, 56(3):329–343, March 2003.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1022374703178.

Dominguez-Lacasa:2006:CCK

[DL06] Iciar Dominguez-Lacasa. Capturing the changes in the knowledge base underlying drug discovery and development in the 20^{th} century and the adjustment of Bayer, Hoechst, Schering AG and E. Merck to the advent of modern biotechnology. Scientometrics, 66(2):345–364, February 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0025-0.

Dong:2006:PLB

[DLM06] Peng Dong, Marie Loh, and Adrian Mondry. Publication lag in biomedical journals varies due to the periodical’s publishing model. Scientometrics, 69(2):271–286, November 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0148-3.

deMoya-Anegon:2007:CAS

[dMACRVQ^{+}07] Félix de Moya-Anegón, Zaida Chinchilla-Rodríguez, Benjamín Vargas-Quesada, Elena Corera-Álvarez, Francisco José Muñoz-Fernández, Antonio González-Molina, and Victor Herrero-Solana. Coverage analysis of Scopus: A journal metric approach. Scientometrics, 73(1):53–78, October 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1681-4.

DaLuz:2008:IIP

[DMPM^{+}08] Mariana Pires Da Luz, Carla Marques-Portella, Mauro Mendlowicz, Sonia Gleiser, Evandro Silva Freire Coutinho, and Ivan Figueira. Institutional h-index: The performance of a new metric in the evaluation of Brazilian psychiatric post-graduation programs. Scientometrics, 77(2): 361–368, November 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1964-9.

deMarchi:2001:EPS

[dMR01] Mario de Marchi and Maurizio Rocchi. The editorial policies of scientific journals: Testing an impact factor model. Scientometrics, 51(2):395–404, June 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1012705818635.

Danell:2003:RRA

[DP03] Rickard Danell and Olle Persson. Regional R&D activities and interactions in the Swedish triple helix. Scientometrics, 58(2):203–218, October 2003. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1026228425125.

DosSantos:2001:BIO

[DR01] Neusa F. Dos Santos and Vivian M. Rumjanek. Brazilian immunology: One hundred years later. Scientometrics, 50 (3):405–418, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010502513971.

Dastidar:2005:ERO

[DR05] Prabir G. Dastidar and S. Ramachandran. Engineering research in ocean sector: An international profile. Scientometrics, 65(2):199–213, November 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0268-1.

Dastidar:2008:ISA

[DR08] Prabir G. Dastidar and S. Ramachandran. Intellectual structure of Antarctic science: A 25-years analysis. Scientometrics, 77(3):389–414, December 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1947-x.

Drenth:2003:MRR

[Dre03] Joost P. H. Drenth. More reprint requests, more citations? Scientometrics, 56(2):283–286, January 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1021975313347.

Duque:2009:IPP

[DSBH09] Ricardo B. Duque, Wesley M. Shrum, Omar Barriga, Jr., and Guillermo Henríquez. Internet practice and professional networks in Chilean science: Dependency or progress? Scientometrics, 81(1):239–263, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-009-2125-0.

Diamond:2007:DEP

[DT07] Arthur M. Diamond,, Jr. and Robert J. Toth. The determinants of election to the Presidency of the American Economic Association: Evidence from a cohort of distinguished 1950’s economists. Scientometrics, 73(2):131–137, November 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-1747-8.

Deleus:2003:MEB

[DV03] Filip Deleus and Marc M. Van Hulle. Monitoring elasticity between science and technology domains and its visualization. Scientometrics, 56(1):147–160, January 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1021958909713.

Davis:2001:ERO

[DW01] Mari Davis and Concepción S. Wilson. Elite researchers in ophthalmology: Aspects of publishing strategies, collaboration and multi-disciplinarity. Scientometrics, 52(3): 395–410, November 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014291715245.

Deineko:2009:NFS

[DW09] Vladimir G. Deineko and Gerhard J. Woeginger. A new family of scientific impact measures: The generalized Kosmulski-indices. Scientometrics, 80(3):819–826, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-2130-0.

Dang:2003:IMR

[DZ03] Yaru Dang and Wenling Zhang. Internationalization of mathematical research. Scientometrics, 58(3):559– 570, November 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000006880.
38873.dd.

Editorial:2005:PIH

[Edi05] Editorial. Peter Ingwersen and Howard D. White win the 2005 Derek John de Solla Price Medal. Scientometrics, 65(3):265–266, December 2005. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-005-0271-6.

Egghe:1999:ERB

[Egg99] L. Egghe. An explanation of the relation between the fraction of multinational publications and the fractional score of a country. Scientometrics, 45(2):291–310, June 1999. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/BF02458438. See remarks [Bur01b].

Egghe:2000:DG

[Egg00a] Leo Egghe. The distribution of N-grams. Scientometrics, 47 (2):237–252, February 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005634925734.

Egghe:2000:HSF

[Egg00b] Leo Egghe. A heuristic study of the first-citation distribution. Scientometrics, 48(3):345–359, July 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005688404778.

Egghe:2001:CLEa

[Egg01a] Leo Egghe. Comments on the “Letter to the Editor” by Burrell. Scientometrics, 50(2):351, February 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1010538110336. See [Bur01c].

Egghe:2001:CLEb

[Egg01b] Leo Egghe. Comments on the “Letter to the Editor” by Burrell. Scientometrics, 50(3):551–552, January 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1010523101676. See [Bur01c].

Egghe:2001:HSF

[Egg01c] Leo Egghe. A heuristic study of the first-citation distribution. Scientometrics, 50(2):363, February 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1010594228083.

Egghe:2004:PRD

[Egg04a] Leo Egghe. Positive reinforcement and 3-dimensional informetrics. Scientometrics, 60(3):497–509, August 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000034390.96418.bf.

Egghe:2004:SPB

[Egg04b] Leo Egghe. Solution of a problem of Buckland on the influence of obsolescence on scattering. Scientometrics, 59(2): 225–232, February 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000018530.
54281.68.

Egghe:2004:SIC

[Egg04c] Leo Egghe. The source-item coverage of the Lotka function. Scientometrics, 61(1):103–115, September 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000037366.83414.09.

Egghe:2005:CLL

[Egg05a] Leo Egghe. A characterization of the law of Lotka in terms of sampling. Scientometrics, 62(3):321–328, March 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0024-6.

Egghe:2005:EDG

[Egg05b] Leo Egghe. An explanation of disproportionate growth using linear 3-dimensional informetrics and its relation with the fractal dimension. Scientometrics, 63(2):277–296, April 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0213-3.

Egghe:2005:SIH

[Egg05c] Leo Egghe. The share of items of highly productive sources as a function of the size of the system. Scientometrics, 65(3): 275–291, December 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0274-3.

Egghe:2006:TPI

[Egg06] Leo Egghe. Theory and practise of the g-index. Scientometrics, 69(1):131–152, October 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
006-0144-7.

Egghe:2007:PEG

[Egg07] Leo Egghe. Probabilities for encountering genius, basic, ordinary or insignificant papers based on the cumulative n^{th} citation distribution. Scientometrics, 70(1):167–181, January 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-0110-z.

Egghe:2008:MRB

[Egg08a] Leo Egghe. The mathematical relation between the impact factor and the uncitedness factor. Scientometrics, 76 (1):117–123, July 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1902-x.

Egghe:2008:MSI

[Egg08b] Leo Egghe. Modelling successive h-indices. Scientometrics, 77(3):377–387, December 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-1968-5.

Egghe:2009:CSI

[Egg09a] Leo Egghe. Comparative study of h-index sequences. Scientometrics, 81(2):311–320, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2170-0.

Egghe:2009:PRP

[Egg09b] Leo Egghe. Performance and its relation with productivity in Lotkaian systems. Scientometrics, 81(2):567– 585, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-2226-1.

Egghe:2010:LER

[Egg10] Leo Egghe. Letter to the editor: On Randić’s Hsequence. Scientometrics, 84(3):795–797, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-009-0110-2. See [Ran09].

Elmacioglu:2009:MIP

[EL09] Ergin Elmacioglu and Dongwon Lee. Modeling idiosyncratic properties of collaboration networks revisited. Scientometrics, 80(1):195–216, July 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-2047-7.

Enachescu:2003:PJC

[EP03] Cornelia Enachescu and Tiberiu Postelnicu. Patterns in journal citation data revealed by exploratory multivariate analysis. Scientometrics, 56(1):43–59, January 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1021946506987.

Egghe:2001:SAT

[ER01] Leo Egghe and Ronald Rousseau. Symmetric and asymmetric theory of relative concentration and applications. Scientometrics, 52(2):261–290, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1017967807504.

Egghe:2002:TEM

[ER02a] Leo Egghe and I. K. R. Ravichandra Rao. Theory and experimentation on the most-recent-reference distribution. Scientometrics, 53(3):371–387, March 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014825113328.

Egghe:2002:CCB

[ER02b] Leo Egghe and Ronald Rousseau. Co-citation, bibliographic coupling and a characterization of lattice citation networks. Scientometrics, 55(3):349–361, November 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1020458612014.

Egghe:2002:PDC

[ER02c] Leo Egghe and Ronald Rousseau. A proposal to define a core of a scientific subject: A definition using concentration and fuzzy sets. Scientometrics, 54(1):51–62, April 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1015680304126.

Egghe:2004:HMO

[ER04] Leo Egghe and Ronald Rousseau. How to measure owngroup preference? A novel approach to a sociometric problem. Scientometrics, 59(2):233–252, February 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000018531.24886.00.

Egghe:2006:IMH

[ER06] Leo Egghe and Ronald Rousseau. An informetric model for the Hirsch-index. Scientometrics, 69(1):121–129, October 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0143-8.

Egghe:2006:PCM

[ERS06] Leo Egghe, I. K. Ravichandra Rao, and Bibhuti Bhusan Sahoo. Proof of a conjecture of Moed and Garfield on authoritative references and extension to non-authoritative references. Scientometrics, 66(3):537–549, March 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0039-7. See [MG04].

Engels:2005:RIG

[ERW05] Anita Engels, Tina Ruschenburg, and Peter Weingart. Recent internationalization of global environmental change research in Germany and the U.S. Scientometrics, 62(1): 67–85, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0004-x.

Eto:2000:ACP

[Eto00a] Hajime Eto. Authorship and citation patterns in operational research journals in relation to competition and reform. Scientometrics, 47(1):25–42, January 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005661608200.

Eto:2000:BDB

[Eto00b] Hajime Eto. Bibliometric distance between methodology and application in statistics. Scientometrics, 48(1):85–97, June 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005684419288.

Eto:2002:ACP

[Eto02] Hajime Eto. Authorship and citation patterns in management science in comparison with operational research. Scientometrics, 53(3):337–349, March 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014821012420.

Eto:2003:III

[Eto03] Hajime Eto. Interdisciplinary information input and output of a nano-technology project. Scientometrics, 58(1):

5–33, September 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025423406643.

Eto:2008:SDS

[Eto08] Hajime Eto. Scientometric definition of science: In what respect is the humanities more scientific than mathematical and social sciences? Scientometrics, 76(1):23–42, July 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-1883-9.

Fowler:2007:DSC

[FA07] James H. Fowler and Dag W. Aksnes. Does self-citation pay? Scientometrics, 72(3):427–437, September 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1777-2.

Farahat:2002:APA

[Far02] Hashem Farahat. Authorship patterns in agricultural sciences in Egypt. Scientometrics, 55(2):157–170, August 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1019659506195.

Fernandez-Cano:2002:MES

[FCB02] Antonio Fernández-Cano and Ángel Bueno. Multivariate evaluation of Spanish educational research journals. Scientometrics, 55(1):87–102, September 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016003104436.

Fernandez-Cano:2004:RPM

[FCTV04] Antonio Fernández-Cano, Manuel Torralbo, and Mónica Vallejo. Reconsidering Price’s model of scientific growth: An overview. Scientometrics, 61(3):301–321, November 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/B%3ASCIE.0000045112.11562.11.

Furukawa:2006:CSI

[FG06] Ryuzo Furukawa and Akira Goto. Core scientists and innovation in Japanese electronics companies. Scientometrics, 68(2):227–240, August 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
006-0109-x.

Fujimaki:2006:QEP

[FH06] Kiyoko Fujimaki and Rockbill Haklak. Quantitative evaluation of positive or negative feelings for biotechnologyor health-related scenes in movies. Scientometrics, 68(2): 213–226, August 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0108-y.

Figueira:2003:CBD

[FJL03] Ivan Figueira, Raphael Jacques, and Jacqueline Leta. A comparison between domestic and international publications in Brazilian psychiatry. Scientometrics, 56(3):317–327, March 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1022322719108.

Florian:2007:IRS

[Flo07] Razvan V. Florian. Irreproducibility of the results of the Shanghai academic ranking of world universities. Scientometrics, 72(1):25–32, July 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-1712-1.

Fox:2004:RKM

[Fox04] Mary Frank Fox. R. K. Merton — life time of influence. Scientometrics, 60(1):47–50, May 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000027307.67775.01.

Faba-Perez:2003:DMC

[FPGBD03] Cristina Faba-Pérez, Vicente P. Guerrero-Bote, and Félix De Moya-Anegón. Data mining in a closed Web environment. Scientometrics, 58(3):623–640, November 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000006884.08036.73.

Franck:2002:SEA

[Fra02] Georg Franck. The scientific economy of attention: A novel approach to the collective rationality of science. Scientometrics, 55(1):3–26, September 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016059402618.

Franses:2003:DSP

[Fra03] Philip Hans Franses. The diffusion of scientific publications: The case of econometrica, 1987. Scientometrics, 56 (1):29–42, January 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1021994422916.

Frandsen:2005:GC

[Fra05] Tove Faber Frandsen. Geographical concentration. Scientometrics, 63(1):69–83, March 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
005-0204-4.

Frandsen:2008:RCV

[Fra08] Tove Faber Frandsen. On the ratio of citable versus noncitable items in economics journals. Scientometrics, 74(3): 439–451, March 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1697-9.

Frederiksen:2004:DDB

[Fre04] Lars Frode Frederiksen. Disciplinary determinants of bibliometric impact in Danish industrial research: Collaboration and visibility. Scientometrics, 61(2):253–270, October 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000041651.26664.14.

Fiala:2008:PBN

[FRJ08] Dalibor Fiala, François Rousselot, and Karel Jezek. PageRank for bibliographic networks. Scientometrics, 76(1):135– 158, July 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1908-4.

Fujigaki:2002:ADR

[Fuj02] Yuko Fujigaki. Analysis on dynamics of research subdomains in interdisciplinary fields: Analysis using personal distribution versus papers. Scientometrics, 54(1):63–74, April 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1015632520964.

Fujigaki:2006:CRS

[Fuj06] Yuko Fujigaki. Changes in the research stream by standardization: A content analysis of the archives of general psychiatry during the establishment of operational diagnostic criteria. Scientometrics, 68(2):203–212, August 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0162-5.

Garcia-Aracil:2006:AEP

[GAGPM06] Adela García-Aracil, Antonio Gutiérrez Gracia, and Marián Pérez-Marín. Analysis of the evaluation process of the research performance: An empirical case. Scientometrics, 67 (2):213–230, May 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0095-z.

Gantman:2009:IDP

[Gan09] Ernesto R. Gantman. International differences of productivity in scholarly management knowledge. Scientometrics, 80(1):153–165, July 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-2054-8.

Garfield:2001:F

[Gar01a] Eugene Garfield. Foreword. Scientometrics, 50(1):5–6, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1005644801160.

Garfield:2001:RVV

[Gar01b] Eugene Garfield. Reminiscences of Vassily V. Nalimov. Scientometrics, 52(2):165–166, October 2001. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1023/A%3A1017947202961.

Garg:2002:SLRa

[Gar02] K. C. Garg. Scientometrics of laser research in India and China. Scientometrics, 55(1):71–85, September 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016050920365.

Garg:2003:OCN

[Gar03] K. C. Garg. An overview of cross-national, national, and institutional assessment as reflected in the international journal Scientometrics. Scientometrics, 56(2):169–199, January 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1021963010621.

Garfield:2004:ICR

[Gar04] Eugene Garfield. The intended consequences of Robert K. Merton. Scientometrics, 60(1):51–61, May 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000027308.27185.30.

Gomez:2009:SRP

[GBFM09] Isabel Gómez, María Bordons, M. Teresa Fernández, and Fernanda Morillo. Structure and research performance of Spanish universities. Scientometrics, 79(1):131–146, April 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0408-0.

Guerrero-Bote:2009:VRW

[GBGCRFdMA09] Vicente P. Guerrero-Bote, Rocío Gómez-Crisóstomo, Luz María Romo-Fernández, and Félix de Moya-Anegón. Visibility and responsibility of women in research papers through the order of signatures: the case of the University of Extremadura, 1990–2005. Scientometrics, 81(1):225–238, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-2122-3.

Guerrero-Bote:2007:IEK

[GBZAEC^{+}07] Vicente P. Guerrero-Bote, Felipe Zapico-Alonso, María Eugenia Espinosa-Calvo, Rocío Gómez-Crisóstomo, and Félix de Moya-Anegón. Import–export of knowledge between scientific subject categories: The Iceberg Hypothesis. Scientometrics, 71(3):423–441, June 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-1682-3.

Gupta:2003:ICP

[GD03] B. M. Gupta and S. M. Dhawan. India’s collaboration with People’s Republic of China in science and technology: A scientometric analysis of coauthored papers during 1994–1999. Scientometrics, 57(1):59–74, May 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1023671519843.

Gupta:2008:CMP

[GD08] B. M. Gupta and S. M. Dhawan. Condensed matter physics: An analysis of India’s research output, 1993– 2001. Scientometrics, 75(1):123–144, April 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1814-9.

Gupta:2009:SIS

[GD09a] B. M. Gupta and S. M. Dhawan. Status of India in science and technology as reflected in its publication output in the Scopus international database, 1996– 2006. Scientometrics, 80(2):473–490, August 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2083-y.

Gupta:2009:SPR

[GD09b] B. M. Gupta and S. M. Dhawan. Status of physics research in India: An analysis of research output during 1993– 2001. Scientometrics, 78(2):295–316, February 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1926-2.

Goldfinch:2003:SPC

[GDD03] Shaun Goldfinch, Tony Dale, and Karl DeRouen, Jr. Science from the periphery: Collaboration, networks and ‘periphery effects’ in the citation of New Zealand Crown Research Institutes articles, 1995–2000. Scientometrics, 57(3):321–337, July 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1025048516769.

Glanzel:2002:DAM

[GdL02] Wolfgang Glänzel and Cornelius de Lange. A distributional approach to multinationality measures of international scientific collaboration. Scientometrics, 54(1):75–89, April 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1015684505035.

Glanzel:2008:TTG

[GDM08] Wolfgang Glänzel, Koenraad Debackere, and Martin Meyer. ‘Triad’ or ‘tetrad’? On global changes in a dynamic world. Scientometrics, 74(1):71–88, January 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-0104-5.

Glanzel:2003:DSN

[GDP03] Wolfgang Glänzel, Rickard Danell, and Olle Persson. The decline of Swedish neuroscience: Decomposing a bibliometric national science indicator. Scientometrics, 57(2):197– 213, June 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1024185601555.

Glanzel:2006:CRR

[GDTS06] Wolfgang Glänzel, Koenraad Debackere, Bart Thijs, and András Schubert. A concise review on the role of author self-citations in information science, bibliometrics and science policy. Scientometrics, 67(2):263–277, May 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0098-9.

Geisler:2005:MSA

[Gei05] Eliezer Geisler. The measurement of scientific activity: Research directions in linking philosophy of science and metrics of science and technology outputs. Scientometrics, 62 (2):269–284, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0020-x.

Gamber:2008:STS

[GFNG08] Thilo Gamber, Monika Friedrich-Nishio, and Hariolf Grupp. Science and technology in standardization: A statistical analysis of merging knowledge structures. Scientometrics, 74(1):89–108, January 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-0105-4.

Guan:2008:CEC

[GG08] Jiancheng Guan and Xia Gao. Comparison and evaluation of Chinese research performance in the field of bioinformatics. Scientometrics, 75(2):357–379, May 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1871-0.

Gagolewski:2009:GAC

[GG09a] Marek Gagolewski and Przemyslaw Grzegorzewski. A geometric approach to the construction of scientific impact indices. Scientometrics, 81(3):617–634, December 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2253-y.

Gao:2009:NSJ

[GG09b] Xia Gao and Jiancheng Guan. Networks of scientific journals: An exploration of Chinese patent data. Scientometrics, 80(1):283–302, July 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-2013-4.

Garg:2005:AIA

[GGJ^{+}05] K. C. Garg, B. M. Gupta, Tabassum Jamal, Santanu Roy, and Suresh Kumar. Assessment of impact of AICTE funding on R&D and educational development. Scientometrics, 65(2):151–160, November 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0264-5.

Glenisson:2005:CFT

[GGP05] Patrick Glenisson, Wolfgang Glänzel, and Olle Persson. Combining full-text analysis and bibliometric indicators. A pilot study. Scientometrics, 63(1):163–180, March 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-005-0208-0.

Guan:2005:CED

[GH05] Jiancheng Guan and Ying He. Comparison and evaluation of domestic and international outputs in information science & technology research of China. Scientometrics, 65(2): 215–244, November 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0269-0.

Guan:2007:PBA

[GH07] Jiancheng Guan and Ying He. Patent-bibliometric analysis on the Chinese science–technology linkages. Scientometrics, 72(3):403–425, September 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1741-1.

Glanzel:2009:CAP

[GJT09] Wolfgang Glänzel, Frizo Janssens, and Bart Thijs. A comparative analysis of publication activity and citation impact based on the core literature in bioinformatics. Scientometrics, 79(1):109–129, April 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0407-1.

Gupta:2000:MGL

[GK00] B. M. Gupta and C. R. Karisiddappa. Modelling the growth of literature in the area of theoretical population genetics. Scientometrics, 49(2):321–355, October 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010577321082.

Gouveia:2009:MWR

[GK09] Fábio C. Gouveia and Eleonora Kurtenbach. Mapping the web relations of science centres and museums from Latin America. Scientometrics, 79(3):491–505, June 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1949-8.

Goncalves:2009:ESP

[GKB^{+}09] Renata R. Gonçalves, Christian Kieling, Rodrigo A. Bressan, Jair J. Mari, and Luis A. Rohde. The evaluation of scientific productivity in Brazil: An assessment of the mental health field. Scientometrics, 80(2):529–537, August 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2077-9.

Garg:2006:SPI

[GKL06] K. C. Garg, Suresh Kumar, and Kashmiri Lal. Scientometric profile of Indian agricultural research as seen through Science Citation Index expanded. Scientometrics, 68(1):151– 166, July 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0088-y.

Gupta:2002:MGW

[GKSK02] B. M. Gupta, Suresh Kumar, S. L. Sangam, and C. R. Karisiddappa. Modeling the growth of world social science literature. Scientometrics, 53(1):161–164, January 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1014844222898.

Glaser:2001:ISI

[GL01] Jochen Gläser and Grit Laudel. Integrating scientometric indicators into sociological studies: Methodical and methodological problems. Scientometrics, 52(3):411–434, November 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1014243832084.

Gauffriau:2005:CMD

[GL05] Marianne Gauffriau and Peder Olesen Larsen. Counting methods are decisive for rankings based on publication and citation studies. Scientometrics, 64(1):85–93, July 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-005-0239-6.

Glanzel:2000:SSBa

[Glä00a] Wolfgang Glänzel. Science in Scandinavia: A bibliometric approach. Scientometrics, 48(2):121–150, September 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1005640604267.

Glanzel:2000:SSBb

[Glä00b] Wolfgang Glänzel. Science in Scandinavia: A bibliometric approach. Scientometrics, 49(2):357, October 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1010559323943.pdf.

Glanzel:2001:NCI

[Glä01] Wolfgang Glänzel. National characteristics in international scientific co-authorship relations. Scientometrics, 51(1):

69–115, April 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010512628145.

Glanzel:2004:TMD

[Glä04] Wolfgang Glänzel. Towards a model for diachronous and synchronous citation analyses. Scientometrics, 60(3): 511–522, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000034391.
06240.2a.

Glanzel:2006:IMA

[Glä06] Wolfgang Glänzel. On the h-index — a mathematical approach to a new measure of publication activity and citation impact. Scientometrics, 67(2):315–321, May 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0102-4.

Glanzel:2008:HIC

[Glä08a] Wolfgang Glänzel. H-index concatenation. Scientometrics, 77(2):369–372, November 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
008-2104-x.

Glanzel:2008:SNB

[Glä08b] Wolfgang Glänzel. On some new bibliometric applications of statistics related to the h-index. Scientometrics, 77(1): 187–196, October 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1989-0.

Glanzel:2008:P

[Glä08c] Wolfgang Glänzel. Preface. Scientometrics, 74(1):13–14, January 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-008-0111-6.

Glanzel:2009:MDJ

[Glä09] Wolfgang Glänzel. The multi-dimensionality of journal impact. Scientometrics, 78(2):355–374, February 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2166-9.

Gauffriau:2007:PCP

[GLM^{+}07] Marianne Gauffriau, Peder Olesen Larsen, Isabelle Maye, Anne Roulin-Perriard, and Markus von Ins. Publication, cooperation and productivity measures in scientific research. Scientometrics, 73(2):175–214, November 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1800-2.

Gauffriau:2008:CRP

[GLM^{+}08] Marianne Gauffriau, Peder Olesen Larsen, Isabelle Maye, Anne Roulin-Perriard, and Markus von Ins. Comparisons of results of publication counting using different methods. Scientometrics, 77(1):147–176, October 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1934-2.

Glanzel:2006:SBP

[GLT06] Wolfgang Glänzel, Jacqueline Leta, and Bart Thijs. Science in Brazil. Part 1: A macro-level comparative study. Scientometrics, 67(1):67–86, April 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
006-0055-7.

Glanzel:2002:JIM

[GM02] Wolfgang Glänzel and Henk F. Moed. Journal impact measures in bibliometric research. Scientometrics, 53(2): 171–193, February 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014848323806.

Glanzel:2003:PCS

[GM03] Wolfgang Glänzel and Martin Meyer. Patents cited in the scientific literature: An exploratory study of ‘reverse’ citation relations. Scientometrics, 58(2):415–428, October 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1026248929668.

Guan:2004:CSR

[GM04] Jiancheng Guan and Nan Ma. A comparative study of research performance in computer science. Scientometrics, 61(3):339–359, November 2004. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/B%
3ASCIE.0000045114.85737.1b.

Guan:2007:BSC

[GM07] Jiancheng Guan and Nan Ma. A bibliometric study of China’s semiconductor literature compared with other major Asian countries. Scientometrics, 70(1):107–124, January 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-0107-7.

Glanzel:2009:PIC

[GM09a] Wolfgang Glänzel and Henk F. Moed. Preface: The 11^{th} International Conference on Scientometrics and Informetrics. Scientometrics, 79(1):5, April 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-009-0400-8.pdf.

Guan:2009:SEM

[GM09b] Jiancheng Guan and Nan Ma. Structural equation model with PLS path modeling for an integrated system of publicly funded basic research. Scientometrics, 81(3):683–698, December 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-2058-7.

Galvez:2006:UIA

[GMA06] Carmen Galvez and Félix Moya-Anegón. The unification of institutional addresses applying parametrized finite-state graphs (P-FSG). Scientometrics, 69(2):323–345, November 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0156-3.

Galvez:2007:SFC

[GMA07] Carmen Galvez and Félix Moya-Anegón. Standardizing formats of corporate source data. Scientometrics, 70(1):3–26, January 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-0101-0.

Gmur:2003:CCA

[Gmü03] Markus Gmür. Co-citation analysis and the search for invisible colleges: A methodological evaluation. Scientometrics, 57(1):27–57, May 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023619503005.

Gossart:2009:CAN

[GÖ09] Cédric Gossart and Müge Özman. Co-authorship networks in social sciences: The case of Turkey. Scientometrics, 78 (2):323–345, February 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1963-x.

Godin:2006:OB

[God06] Benoît Godin. On the origins of bibliometrics. Scientometrics, 68(1):109–133, July 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
006-0086-0.

Goel:2002:GDP

[Goe02] Kamlesh Goel. Gender differences in publication productivity in psychology in India. Scientometrics, 55(2):243–258, August 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1019667708012.

Goebelbecker:2005:RPN

[Goe05] Jürgen Goebelbecker. The role of publications in the new programme oriented funding of the Hermann von Helmholtz Association of National Research Centres (HGF). Scientometrics, 62(1):173–181, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0012-x.

Goktepe:2003:THM

[Gok03] Devrim Goktepe. The triple helix as a model to analyze Israeli magnet program and lessons for late-developing countries like Turkey. Scientometrics, 58(2):219–239, October 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1026280409195.

Golub:2002:MFD

[Gol02] Branka Golub. Motivational factors in departure of young scientists from Croatian science. Scientometrics, 53(3): 429–445, March 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014881231075.

Gordon:2005:HSL

[Gor05a] Avishag Gordon. Homeland security literature in relation to terrorism publications: The source and the response. Scientometrics, 65(1):55–65, October 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0259-2.

Gordon:2005:PTL

[Gor05b] Avishag Gordon. The peripheral terrorism literature: Bringing it closer to the core. Scientometrics, 62(3): 403–414, March 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0030-8.

Gordon:2007:TCA

[Gor07] Avishag Gordon. Transient and continuant authors in a research field: The case of terrorism. Scientometrics, 72 (2):213–224, August 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1714-z.

Gokceoglu:2008:IES

[GOS08] Candan Gokceoglu, Aral I. Okay, and Ebru Sezer. International earth science literature from Turkey — 1970–2005: Trends and possible causes. Scientometrics, 74(3):409–423, March 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1813-x.

Garg:2000:SPN

[GP00] K. C. Garg and P. Padhi. Scientometrics of prolific and non–prolific authors in laser science and technology. Scientometrics, 49(3):359–371, November 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010529405152.

Garg:2001:SCL

[GP01] K. C. Garg and P. Padhi. A study of collaboration in laser science and technology. Scientometrics, 51(2):415–427, June 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1012709919544.

Garg:2002:SLRb

[GP02] K. C. Garg and P. Padhi. Scientometrics of laser research in India during 1970–1994. Scientometrics, 55(2):215–241, August 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1019615723942.

Garcia-Perez:2009:MEH

[GP09] Miguel A. García-Pérez. A multidimensional extension to Hirsch’s h-index. Scientometrics, 81(3):779–785, December 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-2290-1.

Garcia-Romero:2006:ASE

[GR06a] Antonio García-Romero. Assessing the socio-economic returns of biomedical research (i): How can we measure the relationship between research and health care? Scientometrics, 66(2):249–261, February 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0019-y.

Glanzel:2006:P

[GR06b] Wolfgang Glänzel and Ronald Rousseau. Preface. Scientometrics, 68(3):327, September 2006. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-006-0114-0.

Granovsky:2001:IPM

[Gra01] Yuri V. Granovsky. Is it possible to measure science? V. V. Nalimov’s research in scientometrics. Scientometrics, 52 (2):127–150, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017991017982.

Glanzel:2001:DED

[GS01] Wolfgang Glänzel and András Schubert. Double effort = double impact? A critical view at international coauthorship in chemistry. Scientometrics, 50(2):199–214, February 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1010561321723.

Glanzel:2003:NCS

[GS03a] Wolfgang Glänzel and András Schubert. A new classification scheme of science fields and subfields designed for scientometric evaluation purposes. Scientometrics, 56(3): 357–367, March 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022378804087.

Gray:2003:QBP

[GS03b] Denis O. Gray and Harm-Jan Steenhuis. Quantifying the benefits of participating in an industry university research center: An examination of research cost avoidance. Scientometrics, 58(2):281–300, October 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1026236626942.

Glanzel:2005:DIC

[GS05] Wolfgang Glänzel and András Schubert. Domesticity and internationality in co-authorship, references and citations. Scientometrics, 65(3):323–342, December 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0277-0.

Glanzel:2007:NRP

[GS07] Wolfgang Glänzel and Balázs Schlemmer. National research profiles in a changing Europe (1983–2003) an exploratory study of sectoral characteristics in the triple helix. Scientometrics, 70(2):267–275, February 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0203-8.

Glanzel:2002:RCA

[GSB02] Wolfgang Glänzel, András Schubert, and Tibor Braun. A relational charting approach to the world of basic research in twelve science fields at the end of the second millennium. Scientometrics, 55(3):335–348, November 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1020406627944.

Garcia-Santiago:2009:UCO

[GSD09] Lola García-Santiago and Felix De Moya-Anegón. Using cooutlinks to mine heterogeneous networks. Scientometrics, 79(3):681–702, June 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1855-0.

Grupp:2001:IAS

[GSH01] Hariolf Grupp, Ulrich Schmoch, and Sybelle Hinze. International alignment and scientific regard as macroindicators for international comparisons of publications. Scientometrics, 51(2):359–380, June 2001. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1012703117727.

Garcia:2005:CFI

[GSM05] Clara Eugenia Garcia and Luis Sanz-Menéndez. Competition for funding as an indicator of research competitiveness. Scientometrics, 64(3):271–300, August 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0251-x.

Gomez-Sancho:2009:ESP

[GSMT09] José María Gómez-Sancho and María Jesús Mancebón-Torrubia. The evaluation of scientific production: Towards a neutral impact factor. Scientometrics, 81(2):435– 458, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-2137-1.

Glanzel:2006:PLA

[GSST06] Wolfgang Glänzel, Balázs Schlemmer, András Schubert, and Bart Thijs. Proceedings literature as additional data source for bibliometric analysis. Scientometrics, 68(3):457– 473, September 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0124-y.

Glanzel:2003:BLT

[GST03] Wolfgang Glänzel, Balázs Schlemmer, and Bart Thijs. Better late than never? On the chance to become highly cited only beyond the standard bibliometric time horizon. Scientometrics, 58(3):571–586, November 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000006881.30700.ea.

Genest:2001:ICW

[GT01] Christian Genest and Carl Thibault. Investigating the concentration within a research community using joint publications and co-authorship via intermediaries. Scientometrics, 51(2):429–440, June 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1012761903614.

Glanzel:2004:DCA

[GT04a] Wolfgang Glänzel and Bart Thijs. Does co-authorship inflate the share of self-citations? Scientometrics, 61(3): 395–404, November 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000045117.
13348.b1.

Glanzel:2004:IAS

[GT04b] Wolfgang Glänzel and Bart Thijs. The influence of author self-citations on bibliometric macro indicators. Scientometrics, 59(3):281–310, March 2004. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/B%
3ASCIE.0000018535.99885.e9.

Glanzel:2005:C

[GTS05] Wolfgang Glänzel, Bart Thijs, and Balázs Schlemmer. Correction. Scientometrics, 63(3):633, June 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-005-0233-z.

Glanzel:2009:SSN

[GTSD09] Wolfgang Glänzel, Bart Thijs, András Schubert, and Koenraad Debackere. Subfield-specific normalized relative indicators and a new generation of relational charts: Methodological foundations illustrated on the assessment of institutional research performance. Scientometrics, 78(1):165–188, January 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2109-5.

Gu:2003:CPD

[Gu03] Yinian Gu. Comments on the paper “Developing Englishlanguage academic journals of China”. Scientometrics, 58 (3):695–696, November 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/accesspage/article/10.1023/B%
3ASCIE.0000006889.42979.f6. See [LZ03].

Gu:2004:GKM

[Gu04a] Yinian Gu. Global knowledge management research: A bibliometric analysis. Scientometrics, 61(2):171–190, October 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000041647.01086.f4.

Gu:2004:IMK

[Gu04b] Yinian Gu. Information management or knowledge management? An informetric view of the dynamics of academia.

Scientometrics, 61(3):285–299, November 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000045111.51946.ba.

Gupta:2006:RLP

[Gup06] V. K. Gupta. References to literature in patent documents: A case study of CSIR in India. Scientometrics, 68(1):29– 40, July 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0082-4.

Gusmao:2000:DUI

[Gus00] Regina Gusmão. Developing and using indicators of multilateral S&T cooperation for policy making: The experience from European research programmes. Scientometrics, 47 (3):493–514, March 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005619917073.

Gurjeva:2001:SCP

[GW01] Lyubov G. Gurjeva and Paul Wouters. Scientometrics in the context of probabilistic philosophy. Scientometrics, 52 (2):111–126, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017939001144.

Gulgoz:2002:TOS

[GYY02] Sami Gülgöz, Ömer A. Yedekçioglu, and Ersin Yurtsever. Turkey’s output in social science publications: 1970– 1999. Scientometrics, 55(1):103–121, September 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016055121274.

Hall:2009:CCS

[Hal09] Peter Gavin Hall. Comment: Citation statistics. Statistical Science, 24(1):25–26, February 2009. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1255009006. See

[Hal09].

Hargens:2004:WMS

[Har04] Lowell L. Hargens. What is Mertonian sociology of science? Scientometrics, 60(1):63–70, May 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000027309.30756.6c.

Harwood:2008:POT

[Har08] Nigel Harwood. Publication outlets and their effect on academic writers’ citations. Scientometrics, 77(2):253– 265, November 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1955-x.

Harnad:2009:OAS

[Har09a] Stevan Harnad. Open access scientometrics and the UK research assessment exercise. Scientometrics, 79(1):147–156, April 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0409-z.

Hartley:2009:NDB

[Har09b] James Hartley. On the need to distinguish between author and journal self-citations. Scientometrics, 81(3):787– 788, December 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-008-
2241-2.

Hassan:2005:EKS

[Has05] Emmanuel Hassan. The evolution of the knowledge structure of fuel cells. Scientometrics, 62(2):223–238, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0016-6.

Havemann:2001:CBB

[Hav01] Frank Havemann. Collaboration behaviour of Berlin life science researchers in the last two decades of the Twentieth century as reflected in the Science Citation Index. Scientometrics, 52(3):435–443, November 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014295816154.

Hayashi:2003:BAA

[Hay03] Takayuki Hayashi. Bibliometric analysis on additionality of Japanese R&D programmes. Scientometrics, 56(3):301–316, March 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1022370602270.

Heinze:2007:CCS

[HB07] Thomas Heinze and Gerrit Bauer. Characterizing creative scientists in nano-S&T: Productivity, multidisciplinarity, and network brokerage in a longitudinal perspective. Scientometrics, 70(3):811–830, March 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0313-3.

Haslam:2008:WMA

[HBK^{+}08] Nick Haslam, Lauren Ban, Leah Kaufmann, Stephen Loughnan, Kim Peters, Jennifer Whelan, and Sam Wilson. What makes an article influential? Predicting impact in social and personality psychology. Scientometrics, 76(1):169–185, July 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1892-8.

Hornbostel:2009:FYS

[HBK^{+}09] Stefan Hornbostel, Susan Böhmer, Bernd Klingsporn, Jörg Neufeld, and Markus von Ins. Funding of young scientist and scientific excellence. Scientometrics, 79(1):171–190, April 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0411-5.

Huang:2003:CPC

[HCC03] Mu-Hsuan Huang, Li-Yun Chiang, and Dar-Zen Chen. Constructing a patent citation map using bibliographic coupling: A study of Taiwan’s high-tech companies. Scientometrics, 58(3):489–506, November 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000006876.29052.bf.

Huang:2006:RER

[HCC06] Mu-Hsuan Huang, Han-Wen Chang, and Dar-Zen Chen. Research evaluation of research-oriented universities in Taiwan from 1993 to 2003. Scientometrics, 67(3):419–435, June 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1556/Scient.67.2006.3.6.

Hsieh:2004:BAP

[HCLH04] Wang-Huu Hsieh, Wen-Ta Chiu, Yee-Shuan Lee, and Yuh-Shan Ho. Bibliometric analysis of patent Ductus Arteriosus treatments. Scientometrics, 60(2):205–215, June 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000027793.12866.58.

Ho:2003:ASC

[HCTC03] Y. S. Ho, C. H. Chiu, T. M. Tseng, and W. T. Chiu. Assessing stem cell research productivity. Scientometrics, 57 (3):369–376, July 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025004701748.

He:2007:ESN

[HD07] Zi-Lin He and Min Deng. The evidence of systematic noise in non-patent references: A study of New Zealand companies’ patents. Scientometrics, 72(1):149–166, July 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-1702-3.

He:2003:DCC

[He03] Tianwei He. Difficulties and challenges of Chinese scientific journals: Statistical analysis of Chinese literatures using Chinese science Bulletin as example. Scientometrics, 57 (1):127–139, May 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023679721660.

Hayati:2009:CBQ

[HE09a] Zouhayr Hayati and Saeideh Ebrahimy. Correlation between quality and quantity in scientific production: A case study of Iranian organizations from 1997 to 2006. Scientometrics, 80(3):625–636, September 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-2094-3.

He:2009:ISC

[He09b] Tianwei He. International scientific collaboration of China with the G7 countries. Scientometrics, 80(3):571– 582, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-2043-y.

Hemlin:2006:CKE

[Hem06] Sven Hemlin. Creative knowledge environments for research groups in biotechnology. The influence of leadership and organizational support in universities and business companies. Scientometrics, 67(1):121–142, April 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0053-9.

Hendrix:2009:ISC

[Hen09] Dean Hendrix. Institutional self-citation rates: A three year study of universities in the United States. Scientometrics, 81(2):321–331, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2160-2.

Hanney:2005:UCC

[HFG^{+}05] Steve Hanney, Iain Frame, Jonathan Grant, Martin Buxton, Tracey Young, and Grant Lewison. Using categorisations of citations when assessing the outcomes from health research. Scientometrics, 65(3):357–379, December 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0279-y.

Haritash:2002:MSI

[HG02] Nirmal Haritash and B. M. Gupta. Mapping of S&T issues in the Indian Parliament: A scientometric analysis of questions raised in both houses of the Parliament. Scientometrics, 54(1):91–102, April 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1015636621873.

He:2008:CCP

[HG08] Ying He and Jiancheng Guan. Contribution of Chinese publications in computer science: A case study on LNCS. Scientometrics, 75(3):519–534, June 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1781-1.

Hargens:1990:NCA

[HH90] L. L. Hargens and J. R. Herting. Neglected considerations in the analysis of agreement among journal referees. Scientometrics, 19(1–2):91–106, July 1990. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/BF02130467. See [HH06] and comments [Lin91].

Hargens:2006:AAB

[HH06] Lowell L. Hargens and Jerald R. Herting. Analyzing the association between referees’ recommendations and editors’ decisions. Scientometrics, 67(1):15–26, April 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-006-0048-6. See [HH90] and comments

[Lin91].

Heimeriks:2003:MCC

[HHV03] Gaston Heimeriks, Marianne Hörlesberger, and Peter Van Den Besselaar. Mapping communication and collaboration in heterogeneous research networks. Scientometrics, 58(2): 391–413, October 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026296812830.

Havemann:2004:GDG

[HHWD04] Frank Havemann, Michael Heinz, and Roland Wagner-Döbler. Growth dynamics of German university enrolments and of scientific disciplines in the 19th century: Scaling behaviour under weak competitive pressure. Scientometrics, 60(3):283–294, August 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000034374.79769.ed.

Hinze:2003:ISP

[Hin03] Sybille Hinze. Introduction to the selected papers of the 7th International Conference on Science and Technology Indicators. Scientometrics, 57(2):155–157, June 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1024129416576.

Hirsch:2005:IQI

[Hir05] Jorge E. Hirsch. An index to quantify an individual’s scientific research output. Proceedings of the National Academy of Sciences of the United States of America, 102(56):16569– 16572, November 15, 2005. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.

pnas.org/content/102/46/16569.abstract.

Hou:2008:SSC

[HKL08] Haiyan Hou, Hildrun Kretschmer, and Zeyuan Liu. The structure of scientific collaboration networks in Scientometrics. Scientometrics, 75(2):189–202, May 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1771-3.

He:2009:ICS

[HL09] Tianwei He and Wei Liu. The internationalization of Chinese scientific journals: A quantitative comparison of three chemical journals from China, England and Japan. Scientometrics, 80(3):583–593, September 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2067-y.

Hellsten:2007:SCC

[HLSA07] Iina Hellsten, Renaud Lambiotte, Andrea Scharnhorst, and Marcel Ausloos. Self-citations, co-authorships and keywords: A new approach to scientists’ field mobility? Scientometrics, 72(3):469–486, September 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-007-1680-5.pdf.

Hung:2009:ICR

[HLT09] Wen-Chi Hung, Ling-Chu Lee, and Min-Hua Tsai. An international comparison of relative contributions to academic productivity. Scientometrics, 81(3):703–718, December 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2210-9.

Hullmann:2003:PPN

[HM03] Angela Hullmann and Martin Meyer. Publications and patents in nanotechnology. Scientometrics, 58(3):507– 527, November 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000006877.
45467.a7.

Harirchi:2007:ESF

[HME07] Goya Harirchi, Göran Melin, and Shapour Etemad. An exploratory study of the feature of Iranian co-authorships in biology, chemistry and physics. Scientometrics, 72(1):11– 24, July 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1693-0.

Hengl:2009:GAG

[HMG09] Tomislav Hengl, Budiman Minasny, and Michael Gould. A geostatistical analysis of geostatistics. Scientometrics, 80 (2):491–514, August 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0073-3.

Hornbostel:2001:TPF

[Hor01] Stefan Hornbostel. Third party funding of German universities. An indicator of research activity? Scientometrics, 50 (3):523–537, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010566916697.

Hartley:2003:AID

[HPF03] James Hartley, James W. Pennebaker, and Claire Fox. Abstracts, introductions and discussions: How far do they differ in style? Scientometrics, 57(3):389–398, July 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1025008802657.

Hu:2009:CSD

[HR09] Xiaojun Hu and Ronald Rousseau. A comparative study of the difference in research performance in biomedical fields among selected Western and Asian countries. Scientometrics, 81(2):475–491, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2202-9.

Horlesberger:2006:WNS

[HS06] Marianne Hörlesberger and Edgar Schiebel. Web networks of the science system: Weighted hubs and authorities. Scientometrics, 66(2):263–278, February 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0020-5.

Heinze:2007:ICR

[HSSK07] Thomas Heinze, Philip Shapira, Jacqueline Senker, and Stefan Kuhlmann. Identifying creative research accomplishments: Methodology and results for nanotechnology and human genetics. Scientometrics, 70(1):125–152, January 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-0108-6.

Hayashi:2006:RJN

[HT06] Takayuki Hayashi and Hiroyuki Tomizawa. Restructuring the Japanese national research system and its effect on performance. Scientometrics, 68(2):241–264, August 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0163-4.

Holmberg:2009:LGW

[HT09] Kim Holmberg and Mike Thelwall. Local government web sites in Finland: A geographic and webometric analysis. Scientometrics, 79(1):157–169, April 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0410-6.

Hu:2007:RSC

[Hu07] Xiaojun Hu. Relative superiority coefficient of papers: A new dimension for institutional research performance in different fields. Scientometrics, 72(3):389–402, September 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-1733-1.

Hudson:2007:KCY

[Hud07] John Hudson. Be known by the company you keep: Citations — quality or chance? Scientometrics, 71(2):231–238, May 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1671-6.

Hullmann:2007:MAD

[Hul07] Angela Hullmann. Measuring and assessing the development of nanotechnology. Scientometrics, 70(3):739–758, March 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-0310-6.

Hudomalj:2003:OBD

[HV03] Emil Hudomalj and Gaj Vidmar. OLAP and bibliographic databases. Scientometrics, 58(3):609–622, November 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/B%3ASCIE.0000006883.28709.d2.

Huang:2006:SPP

[HVG06] Can Huang, Celeste Amorim Varum, and Joaquim Borges Gouveia. Scientific productivity paradox: The case of China’s S&T system. Scientometrics, 69(2):449– 473, November 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0153-6.

Hood:2001:LBS

[HW01] William W. Hood and Concepción S. Wilson. The literature of bibliometrics, scientometrics, and informetrics. Scientometrics, 52(2):291–314, October 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1017919924342.

Hood:2002:AFS

[HW02] William W. Hood and Concepción S. Wilson. Analysis of the fuzzy set literature using phrases. Scientometrics, 54 (1):103–118, April 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1015688605943.

Hood:2003:ISU

[HW03] William W. Hood and Concepción S. Wilson. Informetric studies using databases: Opportunities and challenges. Scientometrics, 58(3):587–608, November 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000006882.47115.c6.

Huber:2001:SPSa

[HWD01a] John C. Huber and Roland Wagner-Döbler. Scientific production: A statistical analysis of authors in mathematical logic. Scientometrics, 50(2):323–337, February 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010581925357.

Huber:2001:SPSb

[HWD01b] John C. Huber and Roland Wagner-Döbler. Scientific production: A statistical analysis of authors in physics, 1800– 1900. Scientometrics, 50(3):437–453, January 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010558714879.

Huang:2008:BAP

[HYA08] Yi Huang, Qing Yang, and Xiaolan Ao. Bibliometric analysis of pentachlorophenol remediation methods during the period of 1994 to 2005. Scientometrics, 77(1):177–186, October 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1899-1.

He:2005:BRB

[HZT05] Tianwei He, Jinglin Zhang, and Lirong Teng. Basic research in biochemistry and molecular biology in China: A bibliometric analysis. Scientometrics, 62(2):249–259, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0018-4.

Isola:2004:MCW

[ID04] Ajiferuke Isola and Wolfram Dietmar. Modelling the characteristics of Web page outlinks. Scientometrics, 59(1):43–62, January 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000013298.22207.2b.

Igami:2008:EEN

[Iga08] Masatsura Igami. Exploration of the evolution of nanotechnology via mapping of patent applications. Scientometrics, 77(2):289–308, November 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1973-8.

Igic:2002:ICW

[Igi02] Rajko Igić. The influence of the civil war in Yugoslavia on publishing in peer-reviewed journals. Scientometrics, 53 (3):447–452, March 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014833315145.

Iversen:2007:BIA

[IGK07] Eric J. Iversen, Magnus Gulbrandsen, and Antje Klitkou. A baseline for the impact of academic patenting legislation in Norway. Scientometrics, 70(2):393–414, February 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0209-2.

Ingwersen:2004:SAR

[IJ04] Peter Ingwersen and Daisy Jacobs. South African research in selected scientific areas: Status 1981–2000. Scientometrics, 59(3):405–423, March 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000018541.29653.f5.

Inonu:2002:OGC

[IK02] Erdal Inönü and L. Kurnaz. Observations on the growth characteristics of the research output of Turkish physicists based on a selective citation analysis. Scientometrics, 55(3): 437–444, November 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1020470914740.

Iribarren-Maestro:2009:MAV

[IMLSSC09] Isabel Iribarren-Maestro, María Luisa Lascurain-Sánchez, and Elías Sanz-Casado. Are multi-authorship and visibility related? Study of ten research areas at Carlos III University of Madrid. Scientometrics, 79(1):191–200, April 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0412-4.

Ingwersen:2000:IVC

[Ing00] Peter Ingwersen. The international visibility and citation impact of Scandinavian research articles in selected social science fields: The decay of a myth. Scientometrics, 49 (1):39–61, August 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005657107901.

Ingwersen:2002:VIR

[Ing02] Peter Ingwersen. Visibility and impact of research in psychiatry for North European countries in EU, US and world contexts. Scientometrics, 54(1):131–144, April 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1015692706852.

Inonu:2003:ICF

[Inö03] Erdal Inönü. The influence of cultural factors on scientific production. Scientometrics, 56(1):137–146, January 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1021906925642.

Ioannidis:2006:CMC

[Ioa06] John P. A. Ioannidis. Concentration of the most-cited papers in the scientific literature: Analysis of journal ecosystems. PLoS One, 1(1):e5:1–e5:7, 12 2006. CODEN POL-

NCL. ISSN 1932-6203. URL http://dx.plos.org/10.
1371/journal.pone.0000005.

Iglesias:2007:SID

[IP07] Juan E. Iglesias and Carlos Pecharromán. Scaling the hindex for different scientific ISI fields. Scientometrics, 73(3): 303–320, December 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1805-x.

Imperial:2007:UHI

[IRN07] Juan Imperial and Alonso Rodríguez-Navarro. Usefulness of Hirsch’s h-index to evaluate scientific research in Spain. Scientometrics, 71(2):271–282, May 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1665-4.

Inzelt:2009:ICI

[ISS09] Annamária Inzelt, András Schubert, and Mihály Schubert. Incremental citation impact due to international co-authorship in Hungarian higher education institutions. Scientometrics, 78(1):37–43, January 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1957-8.

Iversen:2000:EPB

[Ive00] Eric J. Iversen. An excursion into the patent-bibliometrics of Norwegian patenting. Scientometrics, 49(1):63–80, August 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005609224740.

Izsak:2006:SPA

[Izs06] János Izsák. Some practical aspects of fitting and testing the Zipf–Mandelbrot model. Scientometrics, 67(1):107–120, April 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0052-x.

Jansz:2000:STI

[Jan00] Margriet C. N. Jansz. Some thoughts on the interaction between scientometrics and science and technology policy. Scientometrics, 47(2):253–264, February 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005686909804.

Jappe:2007:EIC

[Jap07] Arlette Jappe. Explaining international collaboration in global environmental change research. Scientometrics, 71 (3):367–390, June 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1676-1.

Jarneving:2001:CSC

[Jar01] Bo Jarneving. The cognitive structure of current cardiovascular research. Scientometrics, 50(3):365–389, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1010546312154.

Jarneving:2005:CTB

[Jar05] Bo Jarneving. A comparison of two bibliometric methods for mapping of the research front. Scientometrics, 65(2): 245–263, November 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0270-7.

Jarneving:2008:VCF

[Jar08] Bo Jarneving. A variation of the calculation of the first author cocitation strength in author cocitation analysis. Scientometrics, 77(3):485–504, December 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1943-1.

Jasienski:2009:GDS

[Jas09] Michal Jasienski. Garfield’s demon and “surprising” or “unexpected” results in science. Scientometrics, 78(2):347–353, February 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1979-2.

Jimenez-Contreras:2009:RSM

[JCTSM^{+}09] Evaristo Jiménez-Contreras, Daniel Torres-Salinas, Rafael Bailón Moreno, Rosario Ruiz Baños, and Emilio Delgado López-Cózar. Response surface methodology and its application in evaluating scientific activity. Scientometrics, 79(1):201–218, April 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0413-3.

Jeannin:2005:IRD

[JD05] Philippe Jeannin and Joëlle Devillard. Implementing relevant disciplinary evaluations in the social sciences. Scientometrics, 63(1):121–144, March 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0206-2.

Jeevan:2002:SAR

[JG02] V. K. J. Jeevan and B. M. Gupta. A scientometric analysis of research output from Indian Institute of Technology, Kharagpur. Scientometrics, 53(1):165–168, January 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1014896206968.

Janssens:2008:HMI

[JGD08] Frizo Janssens, Wolfgang Glänzel, and Bart De Moor. A hybrid mapping of information science. Scientometrics, 75 (3):607–631, June 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-2002-7.

Jacobs:2000:BSP

[JI00] Daisy Jacobs and Peter Ingwersen. A bibliometric study of the publication patterns in the sciences of South African scholars 1981–96. Scientometrics, 47(1):75–93, January 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005617825947.

Jiang:2008:LAA

[Jia08] Yichuan Jiang. Locating active actors in the scientific collaboration communities based on interaction topology analyses. Scientometrics, 74(3):471–482, March 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1587-1.

Jesus:2004:PFC

[JJ04a] Rey-Rocha Jesús and Martín-Sempere María José. Patterns of the foreign contributions in some domestic vs. international journals on Earth sciences. Scientometrics, 59(1): 95–115, January 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000013301.
01329.7a.

Jiancheng:2004:EIK

[JJ04b] Guan Jiancheng and Wang Junxia. Evaluation and interpretation of knowledge production efficiency. Scientometrics, 59(1):131–155, January 2004. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/B%
3ASCIE.0000013303.25298.ae.

Jang:2009:HDL

[JLC09] Show-Ling Jang, Shihmin Lo, and Wen Hao Chang. How do latecomers catch up with forerunners? Analysis of patents and patent citations in the field of flat panel display technologies. Scientometrics, 79(3):563–591, June 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-2032-1.

Jacob:2007:ERH

[JLH07] Jürgen Harald Jacob, Siegfried Lehrl, and Andreas Wolfram Henkel. Early recognition of high quality researchers of the German psychiatry by worldwide accessible bibliometric indicators. Scientometrics, 73(2):117–130, November 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-1729-x.

Junquera:2007:VBA

[JM07] Beatriz Junquera and María Mitre. Value of bibliometric analysis for research policy: A case study of Spanish research into innovation and technology management. Scientometrics, 71(3):443–454, June 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1689-9.

Jayasinghe:2006:NRT

[JMB06] Upali W. Jayasinghe, Herbert W. Marsh, and Nigel Bond. A new reader trial approach to peer review in funding research grants: An Australian experiment. Scientometrics, 69(3): 591–606, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0171-4.

Johnston:2009:EIA

[Joh09] Ron Johnston. The extent of influence: An alternative approach to identifying dominant contributors to a discipline’s literature. Scientometrics, 78(3):409–420, March 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-2015-2.

Jonkers:2009:ETF

[Jon09] Koen Jonkers. Emerging ties: Factors underlying China’s co-publication patterns with Western European and North American research systems in three molecular life science subfields. Scientometrics, 80(3):775–795, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-008-2115-7.

Jin:2006:AII

[JR06] Bihui Jin and Ronald Rousseau. Another ISI idiosyncrasy. Scientometrics, 66(3):613–614, March 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-006-0043-y.

Jensen:2009:TBI

[JRC09] Pablo Jensen, Jean-Baptiste Rouquier, and Yves Croissant. Testing bibliometric indicators by their prediction of scientists promotions. Scientometrics, 78(3):467–479, March 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-2014-3.

Jin:2006:KLO

[JRS06] Bihui Jin, Ronald Rousseau, and Xiaoxing Sun. Key labs and open labs in the Chinese scientific research system: Their role in the national and international scientific arena. Scientometrics, 67(1):3–14, April 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0047-7.

Jiang:2002:NRS

[JSJX02] Guohua Jiang, Shi Shan, Lan Jiang, and Xuesong Xu. A new rank-size distribution of Zipf’s Law and its applications. Scientometrics, 54(1):119–130, April 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1015640722782.

Jonkers:2008:CRR

[JT08] Koen Jonkers and Robert Tijssen. Chinese researchers returning home: Impacts of international mobility on research collaboration and scientific productivity. Scientometrics, 77(2):309–333, November 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1971-x.

Judit:2004:SLS

[Jud04] Bar-Ilan Judit. Self-linking and self-linked rates of academic institutions on the Web. Scientometrics, 59(1):29–41, January 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000013297.48237.6f.

Jin:2002:DCS

[JZCZ02] Bihui Jin, Jiangong Zhang, Dingquan Chen, and Xianyou Zhu. Development of the Chinese scientometric indicators (CSI). Scientometrics, 54(1):145–154, April 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1015644823690.

Kretschmer:2004:VCW

[KA04] Hildrun Kretschmer and Isidro F. Aguillo. Visibility of collaboration on the Web. Scientometrics, 61(3):405– 426, November 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000045118.
68430.fd.

Kao:2009:ACS

[Kao09a] Chiang Kao. The authorship and country spread of operation research journals. Scientometrics, 78(3):397–407, March 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-1850-0.

Kao:2009:AII

[Kao09b] Chiang Kao. The authorship and internationality of industrial engineering journals. Scientometrics, 81(1):123–136, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-2093-4.

Kapseon:2004:MCS

[Kap04] Kim Kapseon. The motivation for citing specific references by social scientists in Korea: The phenomenon of co-existing references. Scientometrics, 59(1):79–93, January 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000013300.68224.c6.

Klavans:2006:QEL

[KB06] Richard Klavans and Kevin W. Boyack. Quantitative evaluation of large maps of science. Scientometrics, 68(3):475– 499, September 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0125-x.

Klavans:2008:TLN

[KB08] Richard Klavans and Kevin W. Boyack. Thought leadership: A new indicator for national and institutional comparison. Scientometrics, 75(2):239–250, May 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1854-1.

Kostoff:2008:RSN

[KBL08] Ronald N. Kostoff, Ryan B. Barth, and Clifford G. Y. Lau. Relation of seminal nanotechnology document production to total nanotechnology document production — South Korea. Scientometrics, 76(1):43–67, July 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1891-9.

Krampen:2007:VCC

[KBWM07] Günter Krampen, Ralf Becker, Ute Wahner, and Leo Montada. On the validity of citation counting in science evaluation: Content analyses of references and citations in psychological publications. Scientometrics, 71(2):191–202, May 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1659-2.

Kumar:2005:SCS

[KG05] Suresh Kumar and K. C. Garg. Scientometrics of computer science research in India and China. Scientometrics, 64(2): 121–132, August 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0244-9.

Kumar:2002:SPA

[KGK02] Suresh Kumar, B. M. Gupta, and C. R. Karisiddappa. Scientific productivity of authors in theoretical population genetics. Scientometrics, 53(1):73–93, January 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014883904242.

Karki:2000:AGO

[KGS00] M. M. S. Karki, K. C. Garg, and Praveen Sharma. Activity and growth of organic chemistry research in India during 1971–1989. Scientometrics, 49(2):279–288, October 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1010521203335.

Kivinen:2008:WWU

[KH08] Osmo Kivinen and Juha Hedman. World-wide university rankings: A Scandinavian approach. Scientometrics, 74 (3):391–408, March 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1820-y.

Kharbanda:2005:BR

[Kha05] V. P. Kharbanda. Book review. Scientometrics, 65(1): 145–149, October 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0263-6.

Kim:2001:BAP

[Kim01] Mee-Jean Kim. A bibliometric analysis of physics publications in Korea, 1994–1998. Scientometrics, 50(3):503–521, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1010514932626.

Kim:2002:CPK

[Kim02] Mee-Jean Kim. Citation patterns of Korean physicists and mechanical engineers: Differences by type of publication source and type of authorship. Scientometrics, 55(3):421– 436, November 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1020418930669.

Kim:2005:KSI

[Kim05] Mee-Jean Kim. Korean science and international collaboration, 1995–2000. Scientometrics, 63(2):321–339, April 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-005-0215-1.

Kim:2006:MIR

[Kim06] Ki-Wan Kim. Measuring international research collaboration of peripheral countries: Taking the context into consideration. Scientometrics, 66(2):231–240, February 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0017-0.

Kim:2007:BAE

[Kim07] Mee-Jean Kim. A bibliometric analysis of the effectiveness of Korea’s biotechnology stimulation plans, with a comparison with four other Asian nations. Scientometrics, 72(3): 371–388, September 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-1585-8.

Klaic:2004:CSP

[KK04] Zvjezdana Bencetić Klaić and Branimir Klaić. Croatian scientific publications in top journals according to the Science Citation Index for the 1980–2000 period. Scientometrics, 61(2):221–250, October 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000041650.23093.87.

Kumar:2002:HZR

[KKK02] Vijai Kumar, V. L. Kalyane, and B. S. Kademani. A. H. Zewail: Research collaborator par excellence. Scientometrics, 53(1):113–121, January 2002. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1014888005151.

Kretschmer:2007:RCA

[KKK07] Hildrun Kretschmer, Ute Kretschmer, and Theo Kretschmer. Reflection of co-authorship networks in the Web: Web hyperlinks versus Web visibility rates. Scientometrics, 70(2): 519–540, February 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-0214-5.

Kademani:2005:NLT

[KKKM05] B. S. Kademani, V. L. Kalyane, Vijai Kumar, and Lalit Mohan. Nobel laureates: Their publication productivity, collaboration and authorship status. Scientometrics, 62(2): 261–268, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0019-3.

Kademani:2006:LE

[KKKM06] B. S. Kademani, V. L. Kalyane, Vijai Kumar, and Lalit Mohan. Letter to the Editor. Scientometrics, 66(3):615–617, March 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0044-x.

Kavunenko:2006:CAJ

[KKL06] Lidiya Kavunenko, Vladimir Khorevin, and Katerina Luzan. Comparative analysis of journals on social sciences and humanities in Ukraine and the world. Scientometrics, 66(1):123–132, January 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
006-0009-0.

Kostoff:2007:GNR

[KKL07] Ronald N. Kostoff, Raymond G. Koytcheff, and Clifford G. Y. Lau. Global nanotechnology research metrics. Scientometrics, 70(3):565–601, March 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0303-5.

Krauskopf:2007:LAL

[KKM07] Manuel Krauskopf, Erwin Krauskopf, and Bernardita Méndez. Low awareness of the link between science and innovation affects public policies in developing countries: The Chilean case. Scientometrics, 72(1):93–103, July 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1737-5.

Kademani:2007:RCI

[KKS^{+}07] B. S. Kademani, Vijai Kumar, Ganesh Surwase, Anil Sagar, Lalit Mohan, Anil Kumar, and C. R. Gaderao. Research and citation impact of publications by the Chemistry Division at Bhabha Atomic Research Centre. Scientometrics, 71(1):25– 57, April 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1651-x.

Kademani:2006:WLT

[KKSK06] B. S. Kademani, Vijai Kumar, Anil Sagar, and Anil Kumar. World literature on thorium research: A scientometric study based on Science Citation Index. Scientometrics, 69(2): 347–364, November 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0157-2.

Kim:2009:ART

[KL09] Heejung Kim and Jae Yun Lee. Archiving research trends in LIS domain using profiling analysis. Scientometrics, 80 (1):75–90, July 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1998-z.

Kao:2008:RTM

[KLC^{+}08] Chiang Kao, Hsiou-Wei Lin, San-Lin Chung, Wei-Chi Tsai, Jyh-Shen Chiou, Yen-Liang Chen, Liang-Hsuan Chen, Shih-Chieh Fang, and Hwei-Lan Pao. Ranking Taiwanese management journals: A case study. Scientometrics, 76(1):95– 115, July 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1895-5.

Kretschmer:2001:CIG

[KLK01a] Hildrun Kretschmer, Liang Liming, and Ramesh Kundra. Chinese-Indian-German collaboration results that provided the impetus for the foundation of COLLNET. Scientometrics, 52(3):445–456, November 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014247900224.

Kretschmer:2001:FGI

[KLK01b] Hildrun Kretschmer, Liang Liming, and Ramesh Kundra. Foundation of a global interdisciplinary research network (COLLNET) with Berlin as the virtual centre. Scientometrics, 52(3):531–537, November 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014268505676.

Karazija:2004:NPP

[KM04a] Romualdas Karazija and Alina Momkauskaite. The Nobel Prize in Physics — regularities and tendencies. Scientometrics, 61(2):191–205, October 2004. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/B%
3ASCIE.0000041648.87075.de.

Koenig:2004:IMA

[KM04b] Michael E. D. Koenig and Elizabeth M. Mezick. Impact of mergers & acquisitions on research productivity within the pharmaceutical industry. Scientometrics, 59(1):157–169, January 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000013304.40957.0d.

Kuusi:2007:ATB

[KM07] Osmo Kuusi and Martin Meyer. Anticipating technological breakthroughs: Using bibliographic coupling to explore the nanotubes paradigm. Scientometrics, 70(3):759–777, March 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-0311-5.

Klitkou:2007:TTS

[KNM07] Antje Klitkou, Stian Nygaard, and Martin Meyer. Tracking techno-science networks: A case study of fuel cells and related hydrogen technology R&D in Norway. Scientometrics, 70(2):491–518, February 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0213-6.

Kyvik:2008:DAT

[KO08] Svein Kyvik and Terje Bruen Olsen. Does the aging of tenured academic staff affect the research performance of universities? Scientometrics, 76(3):439–455, September 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1767-z.

Koehler:2001:ISL

[Koe01a] Wallace Koehler. Information science as “little science”:the implications of a bibliometric analysis of theJournal of the American society for information science. Scientometrics, 51(1):117–132, April 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010516712215.

Koenig:2001:LSSa

[Koe01b] Michael E. D. Koenig. Lessons from the study of scholarly communication for the new information era. Scientometrics, 50(3):511–523, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1012726323178.

Koenig:2001:LSSb

[Koe01c] Michael E. D. Koenig. Lessons from the study of scholarly communication for the new information era. Scientometrics, 51(3):511–523, July 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1019695019399.

Kortelainen:2001:SID

[Kor01] Terttu A. M. Kortelainen. Studying the international diffusion of a national scientific journal. Scientometrics, 51 (1):133–146, April 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010568829054.

Kostoff:2001:MST

[Kos01] Ronald N. Kostoff. The metrics of science and technology. Scientometrics, 50(2):353–361, February 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010590111245.

Kostoff:2002:CAR

[Kos02] Ronald N. Kostoff. Citation analysis of research performer quality. Scientometrics, 53(1):49–71, January 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014831920172.

Kostoff:2007:DBH

[Kos07] Ronald N. Kostoff. The difference between highly and poorly cited medical articles in the journal lancet. Scientometrics, 72(3):513–520, September 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1573-7.

Kim:2008:IRC

[KP08] Hyukjoon Kim and Yongtae Park. The impact of R&D collaboration on innovative performance in Korea: A Bayesian network approach. Scientometrics, 75(3):535–554, June 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1857-y.

Kao:2009:ERP

[KP09] Chiang Kao and Hwei-Lan Pao. An evaluation of research performance in management of 168 Taiwan universities. Scientometrics, 78(2):261–277, February 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1906-6.

Krauss:2007:JSC

[Kra07] Jochen Krauss. Journal self-citation rates in ecological sciences. Scientometrics, 73(1):79–89, October 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1727-7.

Krampen:2008:EUD

[Kra08] Günter Krampen. The evaluation of university departments and their scientists: Some general considerations with reference to exemplary bibliometric publication and citation analyses for a department of psychology. Scientometrics, 76(1):3–21, July 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1900-z.

Kretschmer:2001:P

[Kre01a] Hildrun Kretschmer. Preface. Scientometrics, 52(3): 363–364, November 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/accesspage/article/10.1023/A%
3A1014239330510.

Kreuzman:2001:CCAa

[Kre01b] Henry Kreuzman. A co-citation analysis of representative authors in philosophy: Examining the relationship between epistemologists and philosophers of science. Scientometrics, 50(3):525–539, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1012778307249.

Kreuzman:2001:CCAb

[Kre01c] Henry Kreuzman. A co-citation analysis of representative authors in philosophy: Examining the relationship between epistemologists and philosophers of science. Scientometrics, 51(3):525–539, July 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1019647103469.

Kretschmer:2004:APG

[Kre04] Hildrun Kretschmer. Author productivity and geodesic distance in bibliographic co-authorship networks, and visibility on the Web. Scientometrics, 60(3):409–420, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/B%3ASCIE.0000034383.86665.22.

Koljatic:2001:IPP

[KS01] Mladen M. Koljatic and Mónica R. Silva. The international publication productivity of Latin American countries in the economics and business administration fields. Scientometrics, 51(2):381–394, June 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1012753601797.

Kocher:2002:LSI

[KS02] Martin G. Kocher and Matthias Sutter. Lack of standardization in informetric research: Reply. Scientometrics, 55(2):329–331, August 2002. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1019628026668. See [Rou02].

Kostoff:2005:CCA

[KS05] Ronald N. Kostoff and Michael F. Shlesinger. CAB: Citation-Assisted Background. Scientometrics, 62(2):199– 212, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0014-8.

Koczy:2009:IMC

[KS09] László Á. Kóczy and Martin Strobel. The invariant method can be manipulated. Scientometrics, 81(1):291–293, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2134-4.

Kot:2003:ZLD

[KSB03] Mark Kot, Emily Silverman, and Celeste A. Berg. Zipf’s Law and the diversity of biology newsgroups. Scientometrics, 56(2):247–257, January 2003. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1021971212438.

Kundra:2001:CPI

[KT01] Ramesh Kundra and Dimiter Tomov. Collaboration patterns in Indian and Bulgarian epidemiology of neoplasms in Medline for 1966–1999. Scientometrics, 52(3):519– 523, November 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014264403859.

Kousha:2006:MUC

[KT06] Kayvan Kousha and Mike Thelwall. Motivations for URL citations to open access library and information science articles. Scientometrics, 68(3):501–517, September 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0126-9.

Kousha:2008:SGS

[KT08] Kayvan Kousha and Mike Thelwall. Sources of Google Scholar citations outside the Science Citation Index: A comparison between four science disciplines. Scientometrics, 74 (2):273–294, February 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-0217-x.

Keiser:2005:TCL

[KU05] Jennifer Keiser and Jürg Utzinger. Trends in the core literature on tropical medicine: a bibliometric analysis from 1952–2002. Scientometrics, 62(3):351–365, March 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-005-0027-3.

Kumari:2006:TSO

[Kum06] Lalitha Kumari. Trends in synthetic organic chemistry research. Cross-country comparison of activity index. Scientometrics, 67(3):467–476, June 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1556/Scient.67.2006.3.8.

Kumari:2009:SOC

[Kum09] G. Lalitha Kumari. Synthetic organic chemistry research: Analysis by scientometric indicators. Scientometrics, 80(3): 559–570, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1985-4.

Kyvik:2003:CTP

[Kyv03] Svein Kyvik. Changing trends in publishing behaviour among university faculty, 1980–2000. Scientometrics, 58(1):

35–48, September 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025475423482.

Ketzler:2009:PGE

[KZ09] Rolf Ketzler and Klaus F. Zimmermann. Publications: German economic research institutes on track. Scientometrics, 80(1):231–252, July 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-2062-3.

LaRowe:2009:SDU

[LAB^{+}09] Gavin LaRowe, Sumeet Ambre, John Burgoon, Weimao Ke, and Katy Börner. The scholarly database and its utility for scientometrics research. Scientometrics, 79(2):219–234, May 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0414-2.

Lamirel:2004:UCA

[LAFP04] Jean-Charles Lamirel, Shadi Al Shehabi, Claire François, and Xavier Polanco. Using a compound approach based on elaborated neural network for Webometrics: An example issued from the EICSTES project. Scientometrics, 61(3): 427–441, November 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000045119.
88828.ce.

Lucio-Arias:2007:KES

[LAL07] Diana Lucio-Arias and Loet Leydesdorff. Knowledge emergence in scientific communication: from “fullerenes” to “nanotubes”. Scientometrics, 70(3):603–632, March 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11192-007-0304-4.pdf.

Lange:2001:CCM

[Lan01] Lydia L. Lange. Citation counts of multi-authored papers — first-named authors and further authors. Scientometrics, 52(3):457–470, November 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014299917063.

Larsen:2002:ECO

[Lar02] Birger Larsen. Exploiting citation overlaps for information retrieval: Generating a boomerang effect from the network of scientific papers. Scientometrics, 54(2):155–178, June 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1016011326300.

Larsen:2008:KNH

[Lar08a] Katarina Larsen. Knowledge network hubs and measures of research impact, science structure, and publication output in nanostructured solar cell research. Scientometrics, 74(1): 123–142, January 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-0107-2.

Larsen:2008:SAP

[Lar08b] Peder Olesen Larsen. The state of the art in publication counting. Scientometrics, 77(2):235–251, November 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-1991-6.

Laudel:2003:SBD

[Lau03] Grit Laudel. Studying the brain drain: Can bibliometric methods help? Scientometrics, 57(2):215–237, June 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1024137718393.

Leta:2002:RIC

[LC02] Jacqueline Leta and Hernan Chaimovich. Recognition and international collaboration: the Brazilian case. Scientometrics, 53(3):325–335, March 2002. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1014868928349.

Lin:2006:IIF

[LC06] Ming-Hsien Lin and Liang-Kung Chen. The impact of impact factor on small specialties: A case study of family medicine in Taiwan. Scientometrics, 66(3):513–520, March 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0037-9.

Liu:2005:ARW

[LCL05] Nian Cai Liu, Ying Cheng, and Li Liu. Academic ranking of world universities using scientometrics — a comment to the “fatal attraction”. Scientometrics, 64(1):101–109, July 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0241-z.

LiceaDeArenas:2002:SMR

[LCLA02] Judith Licea De Arenas, Heriberta Castaños-Lomnitz, and Judith Arenas Licea. Significant Mexican research in the health sciences: A bibliometric analysis. Scientometrics, 53 (1):39–48, January 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014879803333.

Long:2009:DFR

[LCWD09] Rebecca Long, Aleta Crawford, Michael White, and Kimberly Davis. Determinants of faculty research productivity in information systems: An empirical analysis of the impact of academic origin and academic affiliation. Scientometrics, 78(2):231–260, February 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1990-7.

Li:2009:GSC

[LDF^{+}09] Ling-Li Li, Guohua Ding, Nan Feng, Ming-Huang Wang, and Yuh-Shan Ho. Global stem cell research trend: Bibliometric analysis as a tool for mapping of trends from 1991 to 2006. Scientometrics, 80(1):39–58, July 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-1939-5.

Lemarie:2000:SES

[LdLM00] Stéphane Lemarié, Marie-Angèle de Looze, and Vincent Mangematin. Strategies of European SMEs in biotechnology: The role of size, technology and market. Scientometrics, 47(3):541–560, March 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005624017982.

Lee:2003:SSR

[Lee03] Chu Keong Lee. A scientometric study of the research performance of the Institute of Molecular and Cell Biology in Singapore. Scientometrics, 56(1):95–110, January 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1021902724734.

Lee:2008:HIE

[Lee08a] Woo Hyoung Lee. How to identify emerging research fields using scientometrics: An example in the field of information security. Scientometrics, 76(3):503–525, September 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-1898-2.

Lee:2008:PLL

[Lee08b] Yong-Gil Lee. Patent licensability and life: A study of U.S. patents registered by South Korean public research institutes. Scientometrics, 75(3):463–471, June 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1879-5.

Lee:2009:WAP

[Lee09] Yong-Gil Lee. What affects a patent’s value? An analysis of variables that affect technological, direct economic, and indirect economic value: An exploratory conceptual approach. Scientometrics, 79(3):623–633, June 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-2020-5.

Lewison:2001:QQF

[Lew01] Grant Lewison. The quantity and quality of female researchers: A bibliometric study of Iceland. Scientometrics, 52(1):29–43, September 2001. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1012794810883.

Lewison:2002:BRH

[Lew02a] Grant Lewison. From biomedical research to health improvement. Scientometrics, 54(2):179–192, June 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016005710371.

Lewison:2002:RUP

[Lew02b] Grant Lewison. Researchers” and users” perceptions of the relative standing of biomedical papers in different journals. Scientometrics, 53(2):229–240, February 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014804608785.

Lewison:2004:JBC

[Lew04] Grant Lewison. James bond and citations to his books. Scientometrics, 59(3):311–320, March 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000018536.84255.b1.

Lewison:2007:RRG

[Lew07] Grant Lewison. The reporting of the risks from genetically modified organisms in the mass media, 2002– 2004. Scientometrics, 72(3):439–458, September 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1769-2.

Leydesdorff:2000:EUB

[Ley00] Loet Leydesdorff. Is the European Union becoming a single publication system? Scientometrics, 47(2):265–280, February 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005639026643.

Leydesdorff:2002:ISC

[Ley02] Loet Leydesdorff. Indicators of structural change in the dynamics of science: Entropy statistics of the SCI Journal Citation Reports. Scientometrics, 53(1):131–159, January 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1014892106059.

Leydesdorff:2003:MIU

[Ley03] Loet Leydesdorff. The mutual information of universityindustry-government relations: An indicator of the triple helix dynamics. Scientometrics, 58(2):445–467, October 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1026253130577.

Leydesdorff:2004:LLW

[Ley04a] Loet Leydesdorff. Loet Leydesdorff wins the 2003 Derek John de Solla Price Medal. Scientometrics, 60(3):273– 274, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/B%3ASCIE.
0000034371.06075.20.

Leydesdorff:2004:TDJ

[Ley04b] Loet Leydesdorff. Top-down decomposition of the Journal Citation Report of the social Science Citation Index: Graph- and factor-analytical approaches. Scientometrics, 60(2):159–180, June 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000027678.
31097.e0.

Leydesdorff:2007:MII

[Ley07] Loet Leydesdorff. Mapping interdisciplinarity at the interfaces between the Science Citation Index and the social Science Citation Index. Scientometrics, 71(3):391–405, June 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-007-1694-z.pdf.

Leydesdorff:2008:DNN

[Ley08] Loet Leydesdorff. The delineation of nanoscience and nanotechnology in terms of journals and patents: A most recent update. Scientometrics, 76(1):159–167, July 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1889-3.

Li:2003:UAJ

[LF03] Li Li and Zhang Fenglian. Understanding academic journals of China. Response to yinian gu. Scientometrics, 58(3): 697–700, November 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000006890.
14683.43.

Lu:2009:MAC

[LF09] Haiyang Lu and Yuqiang Feng. A measure of authors’ centrality in co-authorship networks based on the distribution of collaborative relationships. Scientometrics, 81(2): 499–511, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-2173-x.

Lamirel:2004:NCQ

[LFAH04] Jean-Charles Lamirel, Claire François, Shadi Al Shehabi, and Martial Hoffmann. New classification quality estimators for analysis of documentary information: Application to patent analysis and web mapping. Scientometrics, 60 (3):445–562, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000034386.
05278.e8.

Lundberg:2006:IBJ

[LFB^{+}06] Jonas Lundberg, Anette Fransson, Mats Brommels, John Skår, and Inger Lundkvist. Is it better or just the same? Article identification strategies impact bibliometric assessments. Scientometrics, 66(1):183–197, January 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0013-4.

Lariviere:2006:CCN

[LGA06] Vincent Larivière, Yves Gingras, and Éric Archambault. Canadian collaboration networks: A comparative analysis of the natural sciences, social sciences and the humanities. Scientometrics, 68(3):519–533, September 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0127-8.

Liang:2002:CPA

[LGD02] Liming Liang, Yongzheng Guo, and Mari Davis. Collaborative patterns and age structures in Chinese publications. Scientometrics, 54(3):473–489, July 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016042718548.

Lacasa:2003:TTC

[LGS03] Icíar Dominguez Lacasa, Hariolf Grupp, and Ulrich Schmoch. Tracing technological change over long periods in Germany in chemicals using patent statistics. Scientometrics, 57(2):175–195, June 2003. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1024133517484.

Leta:2006:SBP

[LGT06] Jacqueline Leta, Wolfgang Glänzel, and Bart Thijs. Science in Brazil. Part 2: Sectoral and institutional research profiles. Scientometrics, 67(1):87–105, April 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0051-y.

Lewison:2005:WTN

[LH05] Grant Lewison and James Hartley. What’s in a title? Numbers of words and the presence of colons. Scientometrics, 63(2):341–356, April 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0216-0.

Leydesdorff:2006:MMW

[LH06] Loet Leydesdorff and Iina Hellsten. Measuring the meaning of words in contexts: An automated analysis of controversies about ‘monarch butterflies,’ ‘Frankenfoods,’ and ‘stem cells’. Scientometrics, 67(2):231–258, May 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0096-y.

Li:2008:UCP

[LH08] Zhi Li and Yuh-Shan Ho. Use of citation per publication as an indicator to evaluate contingent valuation research. Scientometrics, 75(1):97–110, April 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1838-1.

Liang:2006:SSB

[LHHWD06] Liming Liang, Frank Havemann, Michael Heinz, and Roland Wagner-Döbler. Structural similarities between science growth dynamics in China and in Western countries. Scientometrics, 66(2):311–325, February 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0023-2.

Liang:2006:ISI

[Lia06] Liming Liang. h-index sequence and h-index matrix: Constructions and applications. Scientometrics, 69(1):153–159, October 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0145-6.

Lindsey:1991:PMR

[Lin91] D. Lindsey. Precision in the manuscript review process: Hargens and Herting revisited. Scientometrics, 22 (2):313–325, October 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02020004. See

[HH90].

Lopez-Illescas:2009:ESJ

[LINV^{+}09] Carmen López-Illescas, Ed C. M. Noyons, Martijn S. Visser, Félix De Moya-Anegón, and Henk F. Moed. Expansion of scientific journal categories using reference analysis: How can it be done and does it make a difference? Scientometrics, 79(3):473–490, June 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1975-6.

Liu:2005:VIS

[Liu05] Zao Liu. Visualizing the intellectual structure in urban studies: A journal co-citation analysis (1992–2002). Scientometrics, 62(3):385–402, March 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0029-1.

Lee:2008:MKN

[LJ08] Bangrae Lee and Yong-Il Jeong. Mapping Korea’s national R&D domain of robot technology by using the co-word analysis. Scientometrics, 77(1):3–19, October 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1819-4.

Leta:2001:CIV

[LJFD01] Jacqueline Leta, Raphael Jacques, Ivan Figueira, and Leopoldo De Meis. Central international visibility of Brazilian psychiatric publications from 1981 to 1995. Scientometrics, 50(2):241–254, February 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010569522632.

Lehmann:2008:QAI

[LJL08] Sune Lehmann, Andrew D. Jackson, and Benny E. Lautrup. A quantitative analysis of indicators of scientific performance. Scientometrics, 76(2):369–390, August 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1868-8.

Lin:2000:FIS

[LK00] Yang Lin and Lynda L. Kaid. Fragmentation of the intellectual structure of political communication study: Some empirical evidence. Scientometrics, 47(1):143–164, January 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005678011835.

Lewison:2008:IMI

[LK08] Grant Lewison and Ramesh Kundra. The internal migration of Indian scientists, 1981–2003, from an analysis of surnames. Scientometrics, 75(1):21–35, April 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1842-5.

Liang:2001:ASS

[LKGdB01] Liming Liang, Hildrun Kretschmer, Yongzheng Guo, and Donald deB. Beaver. Age structures of scientific collaboration in Chinese computer science. Scientometrics, 52(3): 471–486, November 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014252017971.

Leta:2003:CWB

[LL03] Jacqueline Leta and Grant Lewison. The contribution of women in Brazilian science: A case study in astronomy, immunology and oceanography. Scientometrics, 57(3):339– 353, July 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1025000600840.

Liu:2008:ADS

[LL08] Chen-Yuan Liu and Shenq-Yih Luo. Analysis of developing a specific technological field using the theme code of Japanese patent information. Scientometrics, 75(1):51–65, April 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1835-4.

Lehmann:2009:CCS

[LLJ09] Sune Lehmann, Benny E. Lautrup, and Andrew D. Jackson. Comment: Citation statistics. Statistical Science, 24(1):17–20, February 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1255009004. See [Hal09].

Liang:2004:NOP

[LLR04] Liming Liang, Junwan Liu, and Ronald Rousseau. Name order patterns of graduate candidates and supervisors in Chinese publications: A case study of three major Chinese universities. Scientometrics, 61(1):3–18, September 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000037359.22151.6f.

Lee:2007:DEA

[LLSL07] Yong-Gil Lee, Jeong-Dong Lee, Yong-Il Song, and Se-Jun Lee. An in-depth empirical analysis of patent citation counts using zero-inflated count data model: The case of KIST. Scientometrics, 70(1):27–39, January 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0102-z.

Lluch:2005:SCU

[Llu05] Julia Osca Lluch. Some considerations on the use of the impact factor of scientific journals as a tool to evaluate research in psychology. Scientometrics, 65(2):189– 197, November 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0267-2.

Laredo:2000:LAP

[LM00] Philippe Larédo and Philippe Mustar. Laboratory activity profiles: An exploratory approach. Scientometrics, 47 (3):515–539, March 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005671901143.

Leydesdorff:2003:THU

[LM03] Loet Leydesdorff and Martin Meyer. The triple helix of university–industry–government relations. Scientometrics, 58(2):191–203, October 2003. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1026276308287.

Leydesdorff:2007:STH

[LM07] Loet Leydesdorff and Martin Meyer. The scientometrics of a triple helix of university-industry-government relations (introduction to the topical issue). Scientometrics, 70(2):

207–222, February 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-0200-y.

Lazega:2006:OVP

[LMJS06] Emmanuel Lazega, Lise Mounier, Marie-Thérèse Jourda, and Rafaël Stofer. Organizational vs. personal social capital in scientists’ performance: A multi-level network study of elite French cancer researchers (1996–1998). Scientometrics, 67(1):27–44, April 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0049-5.

Lo:2008:PCA

[Lo08] Szu-Chia Lo. Patent coupling analysis of primary organizations in genetic engineering research. Scientometrics, 74 (1):143–151, January 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-0110-7.

Lee:2000:BSR

[LOK00] Mushin Lee, Kiyong Om, and Joon Koh. The bias of sighted reviewers in research proposal evaluation: A comparative analysis of blind and open review in Korea. Scientometrics, 48(1):99–116, June 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005636503358.

Lombardo:2008:NIL

[Lom08] Luciano Lombardo. New indicators linking patenting and business R&D expenditure. Scientometrics, 76(2):201–224, August 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1631-1.

Lewison:2004:CBJ

[LP04] Grant Lewison and Guillermo Paraje. The classification of biomedical journals by research level. Scientometrics, 60 (2):145–157, June 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000027677.
79173.b8.

Leta:2005:LSR

[LPC05] Jacqueline Leta, Júlio C. R. Pereira, and Hernan Chaimovich. The life sciences — the relative contribution of the University of São Paulo to the highest impact factor journals and to those with the largest number of articles, 1980 to 1999. Scientometrics, 63(3):599–616, June 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0230-2.

Liu:2006:PAT

[LpCY06] Yun Liu, Guo ping Cheng, and Yu Yang. Patent applications of the top 500 foreign investment corporations in China. Scientometrics, 68(1):167–177, July 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0089-x.

Liang:2009:BCJ

[LR09a] Liming Liang and Ronald Rousseau. Bibliometric characteristics of the journal Science: Pre-Koshland, Koshland and post-Koshland period. Scientometrics, 80(2):359–372, August 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2086-8.

Liu:2009:PHT

[LR09b] Yuxian Liu and Ronald Rousseau. Properties of Hirsch-type indices: the case of library classification categories. Scientometrics, 79(2):235–248, May 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
009-0415-1.

Liu:2009:ESJ

[LRR09] Yuxian Liu, I. K. Ravichandra Rao, and Ronald Rousseau. Empirical series of journal h-indices: The JCR category horticulture as a case study. Scientometrics, 80(1):59–74, July 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-2026-z.

Liang:2006:RIS

[LRS06] Liming Liang, Ronald Rousseau, and Fei Shi. A rhythm indicator for science and the rhythm of Science. Scientometrics, 68(3):535–544, September 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0128-7.

Lozano:2005:DEA

[LS05] S. Lozano and J. L. Salmerón. Data envelopment analysis of OR/MS journals. Scientometrics, 64(2):133–150, August 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0245-8.

Lascurain-Sanchez:2008:IHS

[LSGZMM^{+}08] M. Luisa Lascurain-Sánchez, Carlos García-Zorita, Carmen Martín-Moreno, Carlos Suárez-Balseiro, and Elías Sanz-Casado. Impact of health science research on the Spanish health system, based on bibliometric and healthcare indicators. Scientometrics, 77(1):131–146, October 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1922-6.

Levitt:2008:PAC

[LT08] Jonathan M. Levitt and Mike Thelwall. Patterns of annual citation of highly cited articles and the prediction of their citation ranking: A comparison across subjects. Scientometrics, 77(1):41–60, October 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1946-y.

Levitt:2009:MHC

[LT09] Jonathan M. Levitt and Mike Thelwall. The most highly cited library and information science articles: Interdisciplinarity, first authors and citation patterns. Scientometrics, 78(1):45–67, January 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-1927-1.

Lundberg:2006:CUE

[LTL^{+}06] Jonas Lundberg, Göran Tomson, Inger Lundkvist, John Skår, and Mats Brommels. Collaboration uncovered: Exploring the adequacy of measuring university-industry collaboration through co-authorship and funding. Scientometrics, 69(3):575–589, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0170-5.

Li:2003:RBW

[LTMW03] Xuemei Li, Mike Thelwall, Peter Musgrove, and David Wilkinson. The relationship between the WIFs or inlinks of computer science departments in UK and their RAE ratings or research productivities in 2001. Scientometrics, 57 (2):239–255, June 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1024189702463.

Leydesdorff:2005:WR

[LTV05] Loet Leydesdorff, Virginia Trimnle, and Péter Vinkler. Walk-round. Scientometrics, 63(2):407–419, April 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-005-0220-4.

Li:2005:NIUa

[LTWM05a] Xuemei Li, Mike Thelwall, David Wilkinson, and Peter Musgrove. National and international university departmental Web site interlinking. Scientometrics, 64(2): 151–185, August 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0246-7.

Li:2005:NIUb

[LTWM05b] Xuemei Li, Mike Thelwall, David Wilkinson, and Peter Musgrove. National and international university departmental Web site interlinking. Scientometrics, 64(2): 187–208, August 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0247-6.

Lukenda:2006:IWC

[Luk06] Josip Lukenda. Influence of the 1991–1995 war on Croatian publications in the MEDLINE database. Scientometrics, 69 (1):21–36, October 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0136-7.

Luwel:2000:BPF

[Luw00] Marc Luwel. A bibliometric profile of Flemish research in natural, life and technical sciences. Scientometrics, 47(2): 281–302, February 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005691010713.

Lecocq:2009:ICT

[LV09] Catherine Lecocq and Bart Van Looy. The impact of collaboration on the technological performance of regions: time invariant or driven by life cycle dynamics? Scientometrics, 80(3):845–865, September 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-2158-4.

Lee:2003:CSI

[LVCS03] John D. Lee, Kim J. Vicente, Andrea Cassano, and Anna Shearer. Can scientific impact be judged prospectively? A bibliometric test of Simonton’s model of creative productivity. Scientometrics, 56(2):223–232, January 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1021967111530.

Leydesdorff:2009:USL

[LW09] Loet Leydesdorff and Caroline Wagner. Is the United States losing ground in science? A global perspective on the world science system. Scientometrics, 78(1):23–36, January 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-008-1830-4.

Liang:2002:MFA

[LZ02] Liming Liang and Ling Zhu. Major factors affecting China’s inter-regional research collaboration: Regional scientific productivity and geographical proximity. Scientometrics, 55 (2):287–316, August 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1019623925759.

Li:2003:DEL

[LZ03] Li Li and Fenglian Zhang. Developing English-language academic journals of China. Scientometrics, 57(1):119–125, May 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1023627704822. See comments

[Gu03].

Leydesdorff:2005:CCK

[LZ05] Loet Leydesdorff and Ping Zhou. Are the contributions of China and Korea upsetting the world system of science? Scientometrics, 63(3):617–630, June 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0231-1.

Leydesdorff:2007:NFS

[LZ07a] Loet Leydesdorff and Ping Zhou. Nanotechnology as a field of science: Its delineation in terms of journals and patents. Scientometrics, 70(3):693–713, March 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-007-0308-0.pdf.

Lin:2007:LTN

[LZ07b] Min-Wei Lin and Jingjing Zhang. Language trends in nanoscience and technology: The case of Chinese-language publications. Scientometrics, 70(3):555–564, March 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-0302-6.

Lariviere:2008:DSI

[LZA08] Vincent Larivière, Alesia Zuccala, and Éric Archambault. The declining scientific impact of theses: Implications for electronic thesis and dissertation repositories and graduate studies. Scientometrics, 74(1):109–121, January 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-0106-3.

Mabe:2001:GDS

[MA01] Michael Mabe and Mayur Amin. Growth dynamics of scholarly and scientific journals. Scientometrics, 51(1):147–162, April 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1010520913124.

Munteanu:2007:JRA

[MA07] Radu Munteanu and Marin Apetroae. Journal relatedness: An actor-actor and actor-objectives case study. Scientometrics, 73(2):215–230, November 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1735-7.

Maier:2006:IFP

[Mai06] Gunther Maier. Impact factors and peer judgment: The case of regional science journals. Scientometrics, 69(3): 651–667, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0175-0.

Markova:2001:BNM

[Mar01] Elena V. Markova. He brought new meanings and new solutions. Scientometrics, 52(2):151–158, October 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1017943102052.

Marchant:2009:ACR

[Mar09] Thierry Marchant. An axiomatic characterization of the ranking based on the h-index and some other bibliometric rankings of authors. Scientometrics, 80(2):325–342, August 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2075-y.

Moya-Anegon:2004:NTB

[MAVQHS^{+}04] Félix Moya-Anegón, Benjamín Vargas-Quesada, Victor Herrero-Solana, Zaida Chinchilla-Rodríguez, Elena Corera-Álvarez, and Francisco J. Munoz-Fernández. A new technique for building maps of large scientific domains based on the cocitation of classes and categories. Scientometrics, 61(1):129–145, September 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000037368.31217.34.

Michel:2001:PCA

[MB01] Jacques Michel and Bernd Bettels. Patent citation analysis.a closer look at the basic input data from patent search reports. Scientometrics, 51(1):185–201, April 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010577030871.

Meyer:2004:CDB

[MB04] Martin Meyer and Sujit Bhattacharya. Commonalities and differences between scholarly and technical collaboration. Scientometrics, 61(3):443–456, November 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000045120.04489.80.

Mauleon:2006:PIP

[MB06] Elba Mauleón and María Bordons. Productivity, impact and publication habits by gender in the area of materials science. Scientometrics, 66(1):199–218, January 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0014-3.

Melo:2006:BAI

[MBC06] Adriano S. Melo, Luis Mauricio Bini, and Priscilla Carvalho. Brazilian articles in international journals on limnology. Scientometrics, 67(2):187–199, May 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0093-1.

Morillo:2001:AIT

[MBG01] Fernanda Morillo, María Bordons, and Isabel Gómez. An approach to interdisciplinarity through bibliometric indicators. Scientometrics, 51(1):203–222, April 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010529114941.

Musgrove:2003:MIC

[MBPKT03] Peter B. Musgrove, Ray Binns, Teresa Page-Kennedy, and Mike Thelwall. A method for identifying clusters in sets of interlinking Web spaces. Scientometrics, 58(3):657– 672, November 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000006886.
37828.4a.

Markpin:2008:ACI

[MBR^{+}08] T. Markpin, B. Boonradsamee, K. Ruksinsut, W. Yochai, N. Premkamolnetr, P. Ratchatahirun, and N. Sombatsompop. Article-count impact factor of materials science journals in SCI database. Scientometrics, 75(2):251–261, May 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1860-3.

Macias-Chapula:2002:BWA

[MC02] César A. Macías-Chapula. Bibliometric and webometric analysis of health system reforms in Latin America and the Caribbean. Scientometrics, 53(3):407–427, March 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1014829214237.

Marx:2009:CIO

[MC09] Werner Marx and Manuel Cardona. The citation impact outside references — formal versus informal citations. Scientometrics, 80(1):1–21, July 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-008-1824-2.pdf.

McCain:2001:I

[McC01] Katherine W. McCain. Introduction. Scientometrics, 50 (3):465–467, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1012766004523.

McCain:2005:HDW

[McC05] Catherine McCain. Howard D. White: Recipient of the 2005 Derek de Solla Price Award of the journal Scientometrics. Scientometrics, 65(3):271–273, December 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0273-4.

McMillan:2009:GDP

[McM09] G. Steven McMillan. Gender differences in patenting activity: An examination of the US biotechnology industry. Scientometrics, 80(3):683–691, September 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2101-0.

Macias-Chapula:2002:BAA

[MCMN02] Cesar A. Macías-Chapula and Acacia Mijangos-Nolasco. Bibliometric analysis of AIDS literature in central Africa. Scientometrics, 54(2):309–317, June 2002. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1016074230843.

Miguel-Dasit:2006:PRS

[MDMBA^{+}06] A. Miguel-Dasit, L. Martí-Bonmatí, R. Aleixandre, P. Sanfeliu, and J. C. Valderrama. Publications resulting from Spanish radiology meeting abstracts: Which, where and who. Scientometrics, 66(3):467–480, March 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0045-9.

Ma:2009:ACC

[MDNL09] Ruimin Ma, Qiangbin Dai, Chaoqun Ni, and Xuelu Li. An author co-citation analysis of information science in China with Chinese Google Scholar search engine, 2004– 2006. Scientometrics, 81(1):33–46, October 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-2063-x.

Meyer:2005:IOA

[MDTU05] Martin Meyer, Mariette Du Plessis, Tania Tukeva, and Jann-Timour Utecht. Inventive output of academic research: A comparison of two science systems. Scientometrics, 63(1):145–161, March 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
005-0207-1.

Meadows:2001:ERIa

[Mea01a] Jack Meadows. Early reactions to information growth. Scientometrics, 50(3):553–561, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1012782408157.

Meadows:2001:ERIb

[Mea01b] Jack Meadows. Early reactions to information growth. Scientometrics, 51(3):553–561, July 2001. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1019651204378.

Medoff:2006:ESC

[Med06] Marshall H. Medoff. The efficiency of self-citations in economics. Scientometrics, 69(1):69–84, October 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0139-4.

Mika:2006:AST

[MEG06] Peter Mika, Tom Elfring, and Peter Groenewegen. Application of semantic technology for social network analysis in the sciences. Scientometrics, 68(1):3–27, July 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0081-5.

Meho:2007:RRC

[Meh07] Lokman I. Meho. The rise and rise of citation analysis. Physics World, 20(1):32–36, January 2007. CO-DEN PHWOEW. ISSN 0953-8585. URL http://

physicsworldarchive.iop.org/full/pwa-pdf/20/1/phwv20i1a33.
pdf.

Meyer:2000:PCN

[Mey00a] Martin Meyer. Patent citations in a novel field of technology — what can they tell about interactions between emerging communities of science and technology? Scientometrics, 48(2):151–178, September 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005692621105.

Meyer:2000:WSA

[Mey00b] Martin Meyer. What is special about patent citations? Differences between scientific and patent citations. Scientometrics, 49(1):93–123, August 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005613325648.

Meyer:2001:PCA

[Mey01] Martin S. Meyer. Patent citation analysis in a novel field of technology:an exploration of nano-science and nanotechnology. Scientometrics, 51(1):163–183, April 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010572914033.

Meyer:2002:TKF

[Mey02] Martin Meyer. Tracing knowledge flows in innovation systems. Scientometrics, 54(2):193–212, June 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016057727209.

Meyer:2006:KIW

[Mey06a] Martin Meyer. Knowledge integrators or weak links? An exploratory comparison of patenting researchers with their non-inventing peers in nano-science and technology. Scientometrics, 68(3):545–560, September 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0129-6.

Meyer:2006:MST

[Mey06b] Martin Meyer. Measuring science-technology interaction in the knowledge-driven economy: The case of a small economy. Scientometrics, 66(2):425–439, February 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0018-z.

Meyer:2007:WDW

[Mey07] Martin Meyer. What do we know about innovation in nanotechnology? Some propositions about an emerging field between hype and path-dependency. Scientometrics, 70(3): 779–810, March 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-0312-4.

Mojzes:2009:SDI

[MF09] Imre Mojzes and Zoltán B. Farkas. The speed of dissemination of information about the realization of the fourth passive electronic circuit element measured by Google hits. Scientometrics, 81(3):699–702, December 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2206-5.

Moutinho:2007:DIF

[MFG07] Paula Susana Figueiredo Moutinho, Margarida Fontes, and Manuel Mira Godinho. Do individual factors matter? A survey of scientists’ patenting in Portuguese public research organisations. Scientometrics, 70(2):355–377, February 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-0207-4.

Malo:2000:STL

[MG00] Stéphane Malo and Aldo Geuna. Science-technology linkages in an emerging research platform: The case of combinatorial chemistry and biology. Scientometrics, 47(2):303–321, February 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005643127551.

Moed:2004:BSP

[MG04] Henk F. Moed and Eugene Garfield. In basic science the percentage of “authoritative” references decreases as bibliographies become shorter. Scientometrics, 60(3):295–303, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000034375.39385.84. See

proof of conjecture [ERS06].

Ma:2005:ESC

[MG05] Nan Ma and Jiancheng Guan. An exploratory study on collaboration profiles of Chinese publications in molecular biology. Scientometrics, 65(3):343–355, December 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0278-z.

Mutschke:2001:CCS

[MH01] Peter Mutschke and Anabel Quan Haase. Collaboration and cognitive structures in social science research fields. Towards socio-cognitive analysis in information systems. Scientometrics, 52(3):487–502, November 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014256102041.

McMillan:2007:PSB

[MH07] G. Steven McMillan and Robert D. Hamilton III. The public science base of US biotechnology: A citation-weighted approach. Scientometrics, 72(1):3–10, July 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1701-4.

Manjarres-Henriquez:2008:CUI

[MHGGVJ08] Liney Manjarrés-Henríquez, Antonio Gutiérrez-Gracia, and Jaider Vega-Jurado. Coexistence of university-industry relations and academic research: Barrier to or incentive for scientific productivity. Scientometrics, 76(3):561–576, September 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1877-7.

Meng:2006:EEB

[MHL06] Wei Meng, Zhenhua Hu, and Wenbin Liu. Efficiency evaluation of basic research in China. Scientometrics, 69(1): 85–101, October 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0140-y.

Minasny:2007:SSI

[MHM07] Budiman Minasny, Alfred E. Hartemink, and Alex McBratney. Soil science and the h index. Scientometrics, 73(3): 257–264, December 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1811-z.

Mehrdad:2004:BSI

[MHSE04] Morteza Mehrdad, Akbar Heydari, Mohammad Nabi Sarbolouki, and Shapour Etemad. Basic science in the Islamic Republic of Iran. Scientometrics, 61(1):79– 88, September 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/B%3ASCIE.
0000037364.65282.4d.

Markusova:2009:TRR

[MJL^{+}09] Valentina A. Markusova, Margriet Jansz, Alexandr N. Libkind, Ilya Libkind, and Alexander Varshavsky. Trends in Russian research output in post-Soviet era. Scientometrics, 79(2):249–260, May 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0416-0.

Moon:2005:FST

[ML05] Hye-Seon Moon and Jeong-Dong Lee. A fuzzy set theory approach to national composite S&T indices. Scientometrics, 64(1):67–83, July 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0238-7.

Mattsson:2008:IEV

[MLNS08] Pauline Mattsson, Patrice Laget, Anna Nilsson, and Carl-Johan Sundberg. Intra-EU vs. extra-EU scientific copublication patterns in EU. Scientometrics, 75(3):555–574, June 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1793-x.

Mariana:2005:SGC

[MLR05] Lima Mariana, Sofía Liberman, and Jane M. Russell. Scientific group cohesiveness at the National University of México. Scientometrics, 64(1):55–66, July 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0237-8.

Marinova:2002:TVJ

[MM02] Dora Marinova and Michael McAleer. Trends and volatility in Japanese patenting in the USA: An analysis of the electronics and transport industries. Scientometrics, 55(2): 171–187, August 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1019611623033.

Munoz-Munoz:2005:STF

[MM05] Ana M. Muñoz-Muñoz. The scholarly transition of female academics at the University of Granada (1975– 1990). Scientometrics, 64(3):325–350, August 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0254-7.

Moghaddam:2008:HDW

[MM08a] Golnessa Galyani Moghaddam and Mostafa Moballeghi. How do we measure the use of scientific journals? A note on research methodologies. Scientometrics, 76(1):125–133, July 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1901-y.

Molinari:2008:MAN

[MM08b] Alain Molinari and Jean-François Molinari. Mathematical aspects of a new criterion for ranking scientific institutions based on the h-index. Scientometrics, 75(2):339–356, May 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1872-z.

Molinari:2008:NMR

[MM08c] Jean-François Molinari and Alain Molinari. A new methodology for ranking scientific institutions. Scientometrics, 75 (1):163–174, April 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1853-2.

Miguel:2008:NAI

[MMAHS08] Sandra Miguel, Félix Moya-Anegón, and Víctor Herrero-Solana. A new approach to institutional domain analysis: Multilevel research fronts structure. Scientometrics, 74 (3):331–344, March 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1586-2.

Markusova:2004:RNM

[MML^{+}04] Valentina A. Markusova, Vladimir A. Minin, Alexandr N. Libkind, C. N. Margriet Jansz, Michel Zitt, and Elise Bassecoulard-Zitt. Research in non-metropolitan universities as a new stage of science development in Russia. Scientometrics, 60(3):365–383, August 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000034380.12874.cc.

Markusova:2001:RGHa

[MMLA01a] Valentina Markusova, Valdimir Minin, Alexandr Libkind, and Michael Arapov. Russian grant-holders opinion on competitive funding: Results of a survey. Scientometrics, 50(3): 541–551, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1012730424087.

Markusova:2001:RGHb

[MMLA01b] Valentina Markusova, Valdimir Minin, Alexandr Libkind, and Michael Arapov. Russian grant-holders opinion on competitive funding: Results of a survey. Scientometrics, 51 (3):541–551, July 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1019699120307.

Meneghini:2006:IVN

[MMP06] Rogerio Meneghini, Rogerio Mugnaini, and Abel L. Packer. International versus national oriented Brazilian scientific journals. A scientometric analysis based on SciELO and JCR–ISI databases. Scientometrics, 69(3):529–538, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0168-z.

Molatudi:2009:BSB

[MMP09] Mohohlo Molatudi, Neo Molotja, and Anastassios Pouris. A bibliometric study of bioinformatics research in South Africa. Scientometrics, 81(1):47–59, October 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-2048-6.

Moin:2005:SOI

[MMR05] Mostafa Moin, Maryam Mahmoudi, and Nima Rezaei. Scientific output of Iran at the threshold of the 21st century. Scientometrics, 62(2):239–248, January 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0017-5.

Marinova:2005:AEI

[MMS05] Dora Marinova, Michael McAleer, and Daniel Slottje. Antitrust environment and innovation. Scientometrics, 64(3): 301–311, August 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0252-9.

Miettunen:2003:ESM

[MN03] Jouko Miettunen and Pentti Nieminen. The effect of statistical methods and study reporting characteristics on the number of citations: A study of four general psychiatric journals. Scientometrics, 57(3):377–388, July 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1025056718587.

Ma:2008:SRC

[MNQ08] Ruimin Ma, Chaoqun Ni, and Junping Qiu. Scientific research competitiveness of world universities in computer science. Scientometrics, 76(2):245–260, August 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1913-7.

Moed:2000:BIR

[Moe00a] Henk F. Moed. Bibliometric indicators reflect publication and management strategies. Scientometrics, 47(2):323–346, February 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005695111622.

Moed:2000:SDI

[Moe00b] Henk F. Moed. Speech delivered at the 7th International Conference on Scientometric and Informetrics in Colima (1999) in the honour of Dr. Cornelius Le Pair on the occasion of his retirement. Scientometrics, 47(2):449–450, February 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1005659615256.

Moed:2001:I

[Moe01] Henk F. Moed. Introduction. Scientometrics, 51(1):5–8, April 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1010504109490.

Moed:2002:MCR

[Moe02] Henk F. Moed. Measuring China’s research performance using the Science Citation Index. Scientometrics, 53(3): 281–296, March 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014812810602.

Moed:2008:URA

[Moe08] Henk F. Moed. UK research assessment exercises: Informed judgments on research quality or quantity? Scientometrics, 74(1):153–161, January 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
008-0108-1.

Moyer:2001:BPBa

[Moy01a] Laura Joy Moyer. Bibliography of publications of Belver C. Griffith. Scientometrics, 50(3):469–479, January 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1012718121361.

Moyer:2001:BPBb

[Moy01b] Laura Joy Moyer. Bibliography of publications of Belver C. Griffith. Scientometrics, 51(3):469–479, July 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1019686817581.

Mahlck:2000:SBM

[MP00] Paula Mählck and Olle Persson. Socio-bibliometric mapping of intra-departmental networks. Scientometrics, 49 (1):81–91, August 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005661208810.

Molatudi:2006:AKB

[MP06] Mohohlo Molatudi and Anastassios Pouris. Assessing the knowledge base for biotechnology in South Africa. Scientometrics, 68(1):97–108, July 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
006-0085-1.

Marshakova-Shaikevich:2001:SPA

[MS01] Irina Marshakova-Shaikevich. Scientometric perspectives of the analysis of chemical terminology. Scientometrics, 52(2): 323–336, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017924025251.

McCain:2004:GE

[MS04] Katherine W. McCain and William E. Snizek. Guest editorial. Scientometrics, 60(1):11–18, May 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000027302.85082.5e.

McAleer:2005:NMI

[MS05] Michael McAleer and Daniel Slottje. A new measure of innovation: The patent success ratio. Scientometrics, 63 (3):421–429, June 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0222-2.

Martin-Sempere:2008:TCS

[MSGGRR08] M. José Martín-Sempere, Belén Garzón-García, and Jesús Rey-Rocha. Team consolidation, social integration and scientists’ research performance: An empirical study in the biology and biomedicine field. Scientometrics, 76(3):457– 482, September 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1866-x.

Mattes:2006:SIL

[MSM06] Eugen Mattes, Michael C. Stacey, and Dora Marinova. Surveying inventors listed on patents to investigate determinants of innovation. Scientometrics, 69(3):475–498, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0164-3.

Martin-Sempere:2002:ETC

[MSRRGG02] María José Martín-Sempere, Jesús Rey-Rocha, and Belén Garzón-García. The effect of team consolidation on research collaboration and performance of scientists. Case study of Spanish university researchers in geology. Scientometrics, 55(3):377–394, November 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1020462712923.

Martinelli:2006:AID

[MSS06] Maurizio Martinelli, Irma Serrecchia, and Michela Serrecchia. Analysis of the Internet diffusion in the non-profit sector: the social digital divide in Italy. Scientometrics, 66 (1):155–170, January 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0011-6.

Meyer:2003:THT

[MSU03] Martin Meyer, Tatiana Siniläinen, and Jan Timm Utecht. Towards hybrid triple helix indicators: A study of university-related patents and a survey of academic inventors. Scientometrics, 58(2):321–350, October 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1026240727851.

Marx:2001:CAU

[MSW01] Werner Marx, Hermann Schier, and Michael Wanitschek. Citation analysis using online databases: Feasibilities and shortcomings. Scientometrics, 52(1):59–82, September 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1012798911792.

Meyer:2007:EVA

[MT07] Martin S. Meyer and Puay Tang. Exploring the “value” of academic patents: IP management practices in UK universities and their implications for third-stream indicators. Scientometrics, 70(2):415–440, February 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0210-9.

Mukherjee:2009:SRL

[Muk09] Bhaskar Mukherjee. Scholarly research in LIS open access electronic journals: A bibliometric study. Scientometrics, 80(1):167–194, July 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-2055-2.

Must:2006:NCE

[Mus06] Ülle Must. “New”countries in Europe — research, development and innovation strategies vs bibliometric data. Scientometrics, 66(2):241–248, February 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0016-1.

McCain:2005:UBK

[MVH^{+}05] Katherine W. McCain, June M. Verner, Gregory W. Hislop, William Evanco, and Vera Cole. The use of bibliometric and knowledge elicitation techniques to map a knowledge domain: Software engineering in the 1990s. Scientometrics, 65 (1):131–144, October 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0262-7.

Markusova:2002:BB

[MWD02] Valentina Markusova, Concepción S. Wilson, and Mari Davis. From bioweapon to biodefense. Scientometrics, 53 (1):21–38, January 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1014827819263.
pdf.

Nagpaul:2002:VCN

[Nag02] P. S. Nagpaul. Visualizing cooperation networks of elite institutions in India. Scientometrics, 54(2):213–228, June 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1016036711279.

Nagpaul:2003:EPR

[Nag03] P. S. Nagpaul. Exploring a pseudo-regression model of transnational cooperation in science. Scientometrics, 56 (3):403–416, March 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022335021834.

Nalimov:2001:FMP

[Nal01a] Vassily Nalimov. Facing the mystery: A philosophical approach. Scientometrics, 52(2):179–184, October 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1017955504778.

Nalimov:2001:PNH

[Nal01b] Vassily V. Nalimov. Philosophy of number: How metrical hermeneutics is possible. Scientometrics, 52(2):185–192, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1017907621616.

Narvaez-Berthelemot:2001:WDS

[NBR01] Nora Narvaez-Berthelemot and Jane M. Russell. World distribution of social science journals:a view from the periphery. Scientometrics, 51(1):223–239, April 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010581131779.

Narvaez-Berthelemot:2002:SAO

[NBRA^{+}02] Nora Narváez-Berthelemot, Jane M. Russell, Rigas Arvanitis, Roland Waast, and Jacques Gaillard. Science in Africa: An overview of mainstream scientific output. Scientometrics, 54(2):229–241, June 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016033528117.

Neff:2009:YAB

[NC09] Mark William Neff and Elizabeth A. Corley. 35 years and 160,000 articles: A bibliometric exploration of the evolution of ecology. Scientometrics, 80(3):657–682, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2099-3.

Noyons:2009:ABM

[NCM09] Ed C. M. Noyons and Clara Calero-Medina. Applying bibliometric mapping in a high level science policy context. Scientometrics, 79(2):261–275, May 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0417-z.

Nelson:2002:IAM

[ND02] Michael Nelson and J. Stephen Downie. Informetric analysis of a music database. Scientometrics, 54(2):243–255, June 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1016013912188.

Neuhaus:2009:NRS

[ND09] Christoph Neuhaus and Hans-Dieter Daniel. A new reference standard for citation analysis in chemistry and related fields based on the sections of chemical abstracts. Scientometrics, 78(2):219–229, February 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-2007-2.

Nalimov:2001:UM

[NDNZ01] Vassily V. Nalimov, Jeanna Drogalina-Nalimov, and Konstantin Zuyev. The universe of meanings. Scientometrics, 52 (2):345–360, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017928126160.

Nederhof:2006:BMR

[Ned06] Anton J. Nederhof. Bibliometric monitoring of research performance in the social sciences and the humanities: A review. Scientometrics, 66(1):81–100, January 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0007-2.

Nederhof:2008:PIB

[Ned08] Anton J. Nederhof. Policy impact of bibliometric rankings of research performance of departments and individuals in economics. Scientometrics, 74(1):163–174, January 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-0109-0.

Nelson:2006:VCP

[Nel06] Michael J. Nelson. Visualization of citation patterns of some Canadian journals. Scientometrics, 67(2):279–289, May 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0099-8.

Nisonger:2002:RBI

[Nis02] Thomas E. Nisonger. The relationship between international editorial board composition and citation measures in political science, business, and genetics journals. Scientometrics, 54(2):257–268, June 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016065929026.

Nadarajah:2007:MCB

[NK07] Saralees Nadarajah and Samuel Kotz. Models for citation behavior. Scientometrics, 72(2):291–305, August 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1717-9.

Neuhaus:2007:USS

[NLD07] Christoph Neuhaus, Andreas Litscher, and Hans-Dieter Daniel. Using scripts to streamline citation analysis on STN international. Scientometrics, 71(1):145–150, April 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-1656-5.

Nederhof:2001:AQS

[NLM01] Anthony J. Nederhof, Marc Luwel, and Henk F. Moed. Assessing the quality of scholarly journals in linguistics:an alternative to citation-based journal impact factors. Scientometrics, 51(1):241–265, April 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010533232688.

Navarro:2008:SPC

[NM08] Albert Navarro and Miguel Martin. Scientific production and collaboration in epidemiology and public health, 1997– 2002. Scientometrics, 76(2):291–313, August 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1931-5.

Nicolini:2008:OAS

[NN08] Claudio Nicolini and Fabrizio Nozza. Objective assessment of scientific performances world-wide. Scientometrics, 76(3): 527–541, September 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1786-9.

Noyons:2001:BMS

[Noy01] Ed Noyons. Bibliometric mapping of science in a policy context. Scientometrics, 50(1):83–98, January 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005694202977.

Noyons:2006:I

[Noy06] Ed C. M. Noyons. Introduction. Scientometrics, 66(1):5–9, January 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0001-8.

Nagpaul:2003:CMO

[NR03] P. S. Nagpaul and Santanu Roy. Constructing a multiobjective measure of research performance. Scientometrics, 56(3):383–402, March 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022382904996.

Nanda:2000:EVR

[NRTD00] Santosh K. Nanda, Ariel L. Rivas, William M. Trochim, and J. David Deshler. Emphasis on validation in research: A meta-analysis. Scientometrics, 48(1):45–64, June 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005628301541.

Negishi:2004:CDJ

[NSS04] Masamitsu Negishi, Yuan Sun, and Kunihiro Shigi. Citation database for Japanese papers: A new bibliometric tool for Japanese academic society. Scientometrics, 60(3):333–351, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000034378.38698.b2.

Nwagwu:2006:BAP

[Nwa06] Williams Nwagwu. A bibliometric analysis of productivity patterns of biomedical authors of Nigeria during 1967– 2002. Scientometrics, 69(2):259–269, November 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0152-7.

Ortega:2008:MAW

[OACS08] Jose Luis Ortega, Isidro Aguillo, Viv Cothey, and Andrea Scharnhorst. Maps of the academic web in the European higher education area — an exploration of visual web indicators. Scientometrics, 74(2):295–308, February 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-0218-9.

Ouimet:2007:DIM

[OALL07] Mathieu Ouimet, Nabil Amara, Réjean Landry, and John Lavis. Direct interactions medical school faculty members have with professionals and managers working in public and private sector organizations: A cross-sectional study. Scientometrics, 72(2):307–323, August 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1731-y.

Okubo:2005:BR

[OB05] Yoshiko Okubo and Tibor Braun. Book review. Scientometrics, 63(1):181–188, March 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
005-0209-z.

Ortega:2009:HOW

[OCA09] José Luis Ortega, Viv Cothey, and Isidro F. Aguillo. How old is the Web? Characterizing the age and the currency of the European scientific Web. Scientometrics, 81(1):295–309, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2149-x.

Okubo:2006:P

[OGYR06] Yoshiko Okubo, Wolfgang Glänzel, Yasuhiro Yamashita, and Ronald Rousseau. Preface. Scientometrics, 68(2): 193–202, August 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0106-0.

Okubo:2000:ISR

[Oku00] Yoshiko Okubo. An introduction to scientometrics research in France. Scientometrics, 47(3):451–455, March 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005663732094.

Osca-Lluch:2008:SAC

[OLBB^{+}08] Julia Osca-Lluch, Pedro Blesa, José Manuel Barrueco, Elena Velasco, and Thomas Krichel. Some aspects of citation indexes in Spain: A comparative analysis. Scientometrics, 75 (2):313–318, May 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1867-9.

Om:2007:USC

[OLC07] Kiyong Om, Jungmann Lee, and Juno Chang. Using supply chain management to enhance industry-university collaborations in IT higher education in Korea. Scientometrics, 71 (3):455–471, June 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1690-3.

Old:2001:USIa

[Old01a] L. John Old. Utilizing spatial information systems for nonspatial-data analysis. Scientometrics, 50(3):563–571, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1012734524996.

Old:2001:USIb

[Old01b] L. John Old. Utilizing spatial information systems for non-spatial-data analysis. Scientometrics, 51(3):563–571, July 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1019603321216.

Osca-Lluch:2009:CAC

[OLVLH09] Julia Osca-Lluch, Elena Velasco, Mayte López, and Julia Haba. Co-authorship and citation networks in Spanish history of science research. Scientometrics, 80(2):373–383, August 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2089-5.

Ojasoo:2001:EPT

[OMD01] Tiiu Ojasoo, Hervé Maisonneuve, and Jean-Christophe Doré. Evaluating publication trends in clinical research: How reliable are medical databases? Scientometrics, 50(3): 391–404, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010598313062.

Onyancha:2009:HAA

[OO09] Omwoyo Bosire Onyancha and Dennis N. Ocholla. Is HIV/ AIDS in Africa distinct? What can we learn from an analysis of the literature? Scientometrics, 79(2):277–296, May 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-009-0418-y.pdf.

Ortiz-Rivera:2000:SPP

[ORSCSB00] Laurie A. Ortiz-Rivera, Elías Sanz-Casado, and Carlos A. Suárez-Balseiro. Scientific production in Puerto Rico in science and technology during the period 1990 to 1998. Scientometrics, 49(3):403–418, November 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010585522899.

Onder:2008:ICS

[ÖSAT08]  Çetin Önder, Mehmet Sevkli, Taner Altinok, and Cen-

giz Tavukçuoğlu. Institutional change and scientific research: A preliminary bibliometric analysis of institutional influences on Turkey’s recent social science publications. Scientometrics, 76(3):543–560, September 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1878-6.

Osareh:2000:CIS

[OW00] Farideh Osareh and Concepción S. Wilson. A comparison of Iranian scientific publications in the Science Citation Index: 1985–1989 and 1990–1994. Scientometrics, 48(3):427–442, July 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005648723433.

Plaza:2008:SLC

[PA08] Luis M. Plaza and Armando Albert. Scientific literature cited in USPTO patent documents as indicators for the evaluation and analysis of Spanish scientific research in biomedical disciplines. Scientometrics, 76(3):429–438, September 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1763-3.

Pereira:2009:FSR

[PB09] César Antonio Pereira and Rogério Eduardo Rodrigues Bazi. Flow and social relationships of knowledge in science, technology and innovation: A patentometric study of UNICAMP’s technological production. Scientometrics, 81 (1):61–72, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-2096-1.

Pinto:2005:QAS

[PBG^{+}05] Maria Pinto, José Luis Alonso Berrocal, José Antonio Cordón García, Viviana Fernández Marcial, Carlos García Figuerola, Javier García Marco, Camarero Carmen Gómez, and Rodríguez Ángel F. Zazo. Quality assessment of Spanish universities’ web sites focused on the European research area. Scientometrics, 65(1):67–93, October 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0260-9.

Peritz:2002:SUB

[PBI02] Bluma C. Peritz and Judit Bar-Ilan. The sources used by bibliometrics-scientometrics as reflected in references. Scientometrics, 54(2):269–284, June 2002. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1016018013096.

Prime:2002:CCC

[PBZ02] Camille Prime, Elise Bassecoulard, and Michel Zitt. Cocitations and co-sitations: A cautionary view on an analogy. Scientometrics, 54(2):291–308, June 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016022214005.

Pestana:2000:SSP

[PC00] Angel Pestaña and Sebastián Cerdán. Spanish scientific productivity and equipment in magnetic resonance from a regional and European perspective. Scientometrics, 49(2): 215–231, October 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010594718356.

Patra:2005:BRP

[PC05] Swapan Kumar Patra and Prakash Chand. Biotechnology research profile of India. Scientometrics, 63(3):583–597, June 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0229-8.

Porter:2007:MRI

[PCRP07] Alan L. Porter, Alex S. Cohen, J. David Roessner, and Marty Perreault. Measuring researcher interdisciplinarity. Scientometrics, 72(1):117–147, July 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1700-5.

Pinto:2007:ERI

[PD07] Maria Pinto and Anne-Vinciane Doucet. An educational resource for information literacy in higher education: Functional and users analyses of the e-COMS academic portal. Scientometrics, 72(2):225–252, August 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1725-9.

Persson:2001:AAC

[Per01] Olle Persson. All author citations versus first author citations. Scientometrics, 50(2):339–344, February 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010534009428.

Persson:2006:EAP

[Per06] Olle Persson. Exploring the analytical potential of comparing citing and cited source items. Scientometrics, 68(3): 561–572, September 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0130-0.

PereiraDeAraujo:2008:IDB

[Per08] AntÔnio F. Pereira De Araújo. Increasing discrepancy between absolute and effective indexes of research output in a Brazilian academic department. Scientometrics, 74(3): 425–437, March 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1817-6.

Petek:2008:PNH

[Pet08] Marija Petek. Personal name headings in COBIB: Testing Lotka’s law. Scientometrics, 75(1):175–188, April 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1829-2.

Pereira:2000:DFH

[PFE00] Júlio Cesar Rodrigues Pereira, André Luiz Fischer, and Maria Mercedes Loureiro Escuder. Driving factors of high performance in Brazilian management sciences for the 1981– 1995 period. Scientometrics, 49(2):307–319, October 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1010525304244.

Polanco:2001:UAN

[PFL01] Xavier Polanco, Claire François, and Jean-Charles Lamirel. Using artificial neural networks for mapping of science and technology: A multi-self-organizing-maps approach. Scientometrics, 51(1):267–292, April 2001. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1010537316758.

Persson:2004:IBV

[PGD04] Olle Persson, Wolfgang Glänzel, and Rickard Danell. Inflationary bibliometric values: The role of scientific collaboration and the need for relative indicators in evaluative studies. Scientometrics, 60(3):421–432, August 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000034384.35498.7d.

Pinto:2009:IPS

[PGFRD09] María Pinto, David Guerrero, Andrés Fernández-Ramos, and Anne-Vinciane Doucet. Information provided by Spanish university websites on their assessment and quality processes. Scientometrics, 81(1):265–289, October 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2123-7.

Patterson:2009:RBR

[PH09] Michael S. Patterson and Simon Harris. The relationship between reviewers’ quality-scores and number of citations for papers published in the journal Physics in Medicine and Biology from 2003–2005. Scientometrics, 80(2):343–349, August 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2064-1.

Park:2005:CKB

[PHL05] Han Woo Park, Heung Deug Hong, and Loet Leydesdorff. A comparison of the knowledge-based innovation systems in the economies of South Korea and the Netherlands using triple helix indicators. Scientometrics, 65(1):3–27, October 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0257-4.

Polanco:2006:USI

[PID06] Xavier Polanco, Roche Ivana, and Besagni Dominique. User science indicators in the Web context and co-usage analysis. Scientometrics, 66(1):171–182, January 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0012-5.

Pinto:2006:GTA

[Pin06] Maria Pinto. A grounded theory on abstracts quality: Weighting variables and attributes. Scientometrics, 69(2):

213–226, November 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0150-9.

Pislyakov:2009:CTT

[Pis09] Vladimir Pislyakov. Comparing two “thermometers”: Impact factors of 20 leading economic journals according to Journal Citation Reports and Scopus. Scientometrics, 79 (3):541–550, June 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-2016-1.

Park:2009:PST

[PK09] Hyun Woo Park and Jay Kang. Patterns of scientific and technological knowledge flows based on scientific papers and patents. Scientometrics, 81(3):811–820, December 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-008-2224-3.

Prakasan:2009:SIP

[PKK09] E. R. Prakasan, V. L. Kalyane, and Vijai Kumar. Sustained impact of publications of C. V. Raman. Scientometrics, 78 (1):77–97, January 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-1937-7.

Porter:2002:RPI

[PKL02] Alan L. Porter, Alisa Kongthon, and Jye-Chyi (JC) Lu. Research profiling: Improving the literature review. Scientometrics, 53(3):351–370, March 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014873029258.

Park:2008:KJS

[PL08] Han Woo Park and Loet Leydesdorff. Korean journals in the Science Citation Index: What do they reveal about the intellectual structure of S&T in Korea? Scientometrics, 75 (3):439–462, June 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1862-1.

Park:2009:KLS

[PL09] Han Woo Park and Loet Leydesdorff. Knowledge linkage structures in communication studies using citation analysis among communication journals. Scientometrics, 81(1): 157–175, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-2119-y.

Patra:2006:BSB

[PM06] Swapan Kumar Patra and Saroj Mishra. Bibliometric study of bioinformatics literature. Scientometrics, 67(3):477–489, June 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1556/Scient.67.2006.3.9.

Panaretos:2009:ASR

[PM09] John Panaretos and Chrisovaladis Malesios. Assessing scientific research performance and impact with single indices. Scientometrics, 81(3):635–670, December 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2174-9.

Pinheiro-Machado:2004:CSP

[PMO04] Rita Pinheiro-Machado and P. L. Oliveira. A comparative study of patenting activity in U.S. and Brazilian scientific institutions. Scientometrics, 61(3):323–338, November 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000045113.27221.bf.

Podlubny:2005:CSI

[Pod05] Igor Podlubny. Comparison of scientific impact expressed by the number of citations in different fields of science. Scientometrics, 64(1):95–99, July 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0240-0.

Pollman:2000:FAS

[Pol00] Thijs Pollman. Forgetting and the ageing of scientific publications. Scientometrics, 47(1):43–54, January 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005613725039.

Ponzi:2002:ISI

[Pon02] Leonard J. Ponzi. The intellectual structure and interdisciplinary breadth of knowledge management: A bibliometric study of its early stage of development. Scientometrics, 55 (2):259–272, August 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1019619824850.

Pouris:2005:AIV

[Pou05] Anastassios Pouris. An assessment of the impact and visibility of South African journals. Scientometrics, 62(2): 213–222, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0015-7.

Pouris:2007:FTS

[Pou07a] Anastassios Pouris. Is fundamentalism a threat to science? Evidence from scientometrics. Scientometrics, 71(2):329– 338, May 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1673-4.

Pouris:2007:NRS

[Pou07b] Anastassios Pouris. Nanoscale research in South Africa: A mapping exercise based on scientometrics. Scientometrics, 70(3):541–553, March 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-0301-7.

Pouris:2009:FRS

[Pou09] Anastassios Pouris. Fluorine research in South Africa and four benchmarking countries: Comparative mapping and assessment. Scientometrics, 78(1):131–143, January 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-1904-8.

Pouris:2009:SST

[PP09] Anastassios Pouris and Anthipi Pouris. The state of science and technology in Africa (2000–2004): A scientometric assessment. Scientometrics, 79(2):297–309, May 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0419-x.

Porter:2009:SBM

[PR09] Alan L. Porter and Ismael Rafols. Is science becoming more interdisciplinary? Measuring and mapping six research fields over time. Scientometrics, 81(3):719–745, December 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2197-2.

Pasterkamp:2007:CFB

[PRdKB07] Gerard Pasterkamp, Joris I. Rotmans, Dominique V. P. de Kleijn, and Cornelius Borst. Citation frequency: A biased measure of research impact significantly influenced by the geographical origin of research articles. Scientometrics, 70(1):153–165, January 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0109-5.

Priego:2003:VSM

[Pri03] José Luis Ortega Priego. A vector space model as a methodological approach to the triple helix dimensionality: A comparative study of biology and biomedicine centres of two European National Research Councils from a Webometric view. Scientometrics, 58(2):429–443, October 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1026201013738.

Prpic:2000:PPY

[Prp00] Katarina Prpić. The publication productivity of Young scientists: An empirical study. Scientometrics, 49(3):453– 490, November 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010541707878.

Prpic:2002:GPD

[Prp02] Katarina Prpić. Gender and productivity differentials in science. Scientometrics, 55(1):27–58, September 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016046819457.

Prpic:2005:GSD

[Prp05] Katarina Prpić. Generational similarities and differences in researchers’ professional ethics: An empirical comparison. Scientometrics, 62(1):27–51, January 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0002-z.

Prpic:2007:CSK

[Prp07] Katarina Prpić. Changes of scientific knowledge production and research productivity in a transitional society. Scientometrics, 72(3):487–511, September 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1760-6.

Pena-Rey:2006:SPT

[PRPFC06] Isabel Peña-Rey, Napoleón Pérez-Farinós, and Pedro Marset Campos. Scientific production on tetrachloro-dibenzo-dioxins: A bibliometric study. Scientometrics, 69(3):639– 650, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0174-1.

Pichappan:2002:OSC

[PS02] P. Pichappan and S. Sarasvady. The other side of the coin: The intricacies of author self-citations. Scientometrics, 54 (2):285–290, June 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1016070029935.

Pinto:2007:MAI

[PSD^{+}07] María Pinto, Dora Sales, Anne-Vinciane Doucet, Andrés Fernández-Ramos, and David Guerrero. Metric analysis of the information visibility and diffusion about the European Higher Education Area on Spanish university websites. Scientometrics, 72(2):345–370, August 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1766-0.

Payne:2007:LSA

[PT07] Nigel Payne and Mike Thelwall. A longitudinal study of academic webs: Growth and stabilisation. Scientometrics, 71(3):523–539, June 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1695-y.

Prasad:2000:TPE

[PV00] Alka Prasad and Srinivasan Visalakshi. Trends and profile in enzyme engineering research during 1971–98. Scientometrics, 48(1):27–44, June 2000. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1005676217471.

Presmanes:2003:SCB

[PZ03] Beatriz Presmanes and Ernesto Zumelzu. Scientific cooperation between Chile and Spain: Joint mainstream publications (1991–2000). Scientometrics, 58(3):547–558, November 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000006879.96909.ef.

Qiu:2009:BAB

[QC09] Hao Qiu and Yi-Feng Chen. Bibliometric analysis of biological invasions research during the period of 1991 to 2007. Scientometrics, 81(3):601–610, December 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2207-4.

Qiu:2004:ABC

[QCW04] Junping Qiu, Jingquan Chen, and Zhi Wang. An analysis of backlink counts and Web Impact Factors for Chinese university websites. Scientometrics, 60(3):463–473, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000034387.76981.83.

Qiu:2008:ESS

[QLLR08] Junping Qiu, Yejun Li, Jiang Li, and Quane Ren. An exploratory study on substantive co-link analysis: A modification to total co-link analysis. Scientometrics, 76(2):327–341, August 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1921-7.

Qiu:2008:NEW

[QMC08] Junping Qiu, Ruimin Ma, and Ni Cheng. New exploratory work of evaluating a researcher’s output. Scientometrics, 77(2):335–344, November 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-1782-8.

Randic:2009:CVL

[Ran09] Milan Randić. Citations versus limitations of citations: beyond Hirsch index. Scientometrics, 80(3):809–818, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2128-2. See letter [Egg10].

Roth:2006:LBD

[RB06] Camille Roth and Paul Bourgine. Lattice-based dynamic and overlapping taxonomies: The case of epistemic communities. Scientometrics, 69(2):429–447, November 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0161-6.

Romero:2009:MIC

[RCE^{+}09] Antonio García Romero, José Navarrete Cortés, Cristina Escudero, Juan Antonio Fernández López, and Juan Antonio Chaichío Moreno. Measuring the influence of clinical trials citations on several bibliometric indicators. Scientometrics, 80(3):747–760, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2108-6.

Royle:2007:PIJ

[RCWE07] Jo Royle, Louisa Coles, Dorothy Williams, and Paul Evans. Publishing in international journals. Scientometrics, 71 (1):59–86, April 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1648-5.

Ramani:2000:NUP

[RdL00] Shyama V. Ramani and Marie-Angele de Looze. A note on using patent statistics to obtain competition indicators. Scientometrics, 49(3):511–515, November 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010545908787.

Ramani:2002:UPS

[RdL02] Shyama V. Ramani and Marie-Angele de Looze. Using patent statistics as knowledge base indicators in the biotechnology sectors: An application to France, Germany and the U.K. Scientometrics, 54(3):319–346, July 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016026314914.

Romero-De-Pablos:2009:IPP

[RDPAC09] Ana Romero-De-Pablos and Joaquín M. Azagra-Caro. Internationalisation of patents by public research organisations from a historical and an economic perspective. Scientometrics, 79(2):329–340, May 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
009-0421-3.

Ranga:2003:EUD

[RDvT03] Liana Marina Ranga, Koenraad Debackere, and Nick von Tunzelmann. Entrepreneurial universities and the dynamics of academic knowledge production: A case study of basic vs. applied research in Belgium. Scientometrics, 58(2):301–320, October 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1026288611013.

Reinhart:2009:PRG

[Rei09] Martin Reinhart. Peer review of grant applications in biology and medicine. Reliability, fairness, and validity. Scientometrics, 81(3):789–809, December 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2220-7.

Requena:2005:DMV

[Req05] Jaime Requena. Dynamics of the modern Venezuelan research community profile. Scientometrics, 65(1):95–130, October 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0261-8.

Rojo:2006:ASS

[RG06] Raquel Rojo and Isabel Gómez. Analysis of the Spanish scientific and technological output in the ICT sector. Scientometrics, 66(1):101–121, January 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0008-1.

Ramirez:2000:RIF

[RGD00] Ana M. Ramírez, Esther O. García, and J. Antonio Del Río. Renormalized Impact Factor. Scientometrics, 47(1):3–9, January 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005600807292.

Rigby:2009:CSQ

[Rig09] John Rigby. Comparing the scientific quality achieved by funding instruments for single grant holders and for collaborative networks within a research system: Some observations. Scientometrics, 78(1):145–164, January 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1970-y.

Rinia:2000:SST

[Rin00] Ed J. Rinia. Scientometric studies and their role in research policy of two research councils in The Netherlands. Scientometrics, 47(2):363–378, February 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005699312530.

Rodriguez:2007:DMT

[RJDD07] Victor Rodriguez, Frizo Janssens, Koenraad Debackere, and Bart De Moor. Do material transfer agreements affect the choice of research agendas? The case of biotechnology in Belgium. Scientometrics, 71(2):239–269, May 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1666-3.

Rai:2004:SEI

[RK04] L. P. Rai and Naresh Kumar. S&T education in India: Prospects and challenges. Scientometrics, 61(2):157–169, October 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000041646.78833.37.

Rai:2001:SCS

[RKM01] L. P. Rai, Naresh Kumar, and S. Madan. Structural changes in S&T research in India. Scientometrics, 50(2):313–321, February 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1010529908519.

Rafols:2007:HCD

[RM07] Ismael Rafols and Martin Meyer. How cross-disciplinary is bionanotechnology? Explorations in the specialty of molecular motors. Scientometrics, 70(3):633–650, March 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-0305-3.

Ramlogan:2007:NKD

[RMTM07] Ronnie Ramlogan, Andrea Mina, Gindo Tampubolon, and J. Stanley Metcalfe. Networks of knowledge: The distributed nature of medical innovation. Scientometrics, 70 (2):459–489, February 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-0212-7.

Rousseau:2002:LSI

[Rou02] Ronald Rousseau. Lack of standardisation in informetric research. Comments on “Power laws of research output. Evidence for journals of economics” by Matthias Sutter and Martin G. Kocher. Scientometrics, 55(2):317–327, August 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1019675909829. See [SK01a] and

reply [KS02].

Rousseau:2004:LLR

[Rou04] Ronald Rousseau. Loet Leydesdorff: Recipient of the 2003 Derek de Solla Price Award. Scientometrics, 60(3):275–277, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000034372.77778.6e.

Rousseau:2005:MPI

[Rou05a] Ronald Rousseau. Median and percentile impact factors: A set of new indicators. Scientometrics, 63(3):431–441, June 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0223-1.

Rousseau:2005:PIR

[Rou05b] Ronald Rousseau. Peter Ingwersen: Recipient of the 2005 Derek de Solla Price Award. Scientometrics, 65(3):267– 269, December 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0272-5.

Rousseau:2006:BR

[Rou06] Ronald Rousseau. Book review. Scientometrics, 67(1): 165–166, April 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-006-
0046-8.

Rousseau:2008:JEE

[Rou08] Sandra Rousseau. Journal evaluation by environmental and resource economists: A survey. Scientometrics, 77(2): 223–233, November 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1980-9.

Roy:2001:VNM

[Roy01] Rustum Roy. Vassily Nalimov — modern Russian High Priest. Scientometrics, 52(2):167–169, October 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1017999319799.

RodriguesPereira:2007:WWW

[RPFD07] Júlio Cesar Rodrigues Pereira, Juliana Parreira Vasconcellos, Lucilla Furusawa, and Augusto De Moura Barbati. Who’s who and what’s what in Brazilian public health sciences. Scientometrics, 73(1):37–52, October 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1787-8.

Ren:2002:IVC

[RR02] Shengli Ren and Ronald Rousseau. International visibility of Chinese scientific journals. Scientometrics, 53(3):389–405, March 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1014877130166.

Rousseau:2006:RCL

[RR06] Ronald Rousseau and Sandra Rousseau. Remarks concerning the Liberman–Wolf bonding number. Scientometrics, 67(2):167–173, May 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0091-3.

Rey-Rocha:2006:SPC

[RRGGMS06] Jesús Rey-Rocha, Belén Garzón-García, and M. José Martín-Sempere. Scientists’ performance and consolidation of research teams in biology and biomedicine at the Spanish council for scientific research. Scientometrics, 69(2): 183–212, November 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0149-2.

Rey-Rocha:2007:ESI

[RRGGMS07] Jesús Rey-Rocha, Belén Garzón-García, and M. José Martín-Sempere. Exploring social integration as a determinant of research activity, performance and prestige of scientists. Empirical evidence in the biology and biomedicine field. Scientometrics, 72(1):59–80, July 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1703-2.

Rey-Rocha:2002:RPS

[RRMSG02] Jesús Rey-Rocha, María José Martín-Sempere, and Belén Garzón. Research productivity of scientists in consolidated vs. non-consolidated teams: The case of Spanish university geologists. Scientometrics, 55(1):137–156, September 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1016059222182.

Robinson:2007:TEN

[RRR07] Douglas K. R. Robinson, Martin Ruivenkamp, and Arie Rip. Tracking the evolution of new and emerging S&T via statement-linkages: Vision assessment in molecular machines. Scientometrics, 70(3):831–858, March 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-007-0314-2.pdf.

Ramos:2007:ADE

[RRS07] Raúl Ramos, Vicente Royuela, and Jordi Suriñach. An analysis of the determinants in economics and business publications by Spanish universities between 1994 and 2004. Scientometrics, 71(1):117–144, April 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1652-9.

Rivellini:2006:SNI

[RRZ06] Giulia Rivellini, Ester Rizzi, and Susanna Zaccarin. The science network in Italian population research: An analysis according to the social network perspective. Scientometrics, 67(3):407–418, June 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1556/Scient.67.2006.3.5.

Ramanana-Rahary:2009:APR

[RRZR09] Suzy Ramanana-Rahary, Michel Zitt, and Ronald Rousseau. Aggregation properties of relative impact and other classical indicators: Convexity issues and the Yule–Simpson paradox. Scientometrics, 79(2):311–327, May 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0420-4.

Rousseau:2000:OFL

[RS00] Ronald Rousseau and Mark Smeyers. Output-financing at LUC. Scientometrics, 47(2):379–387, February 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005651429368.

Rovira:2000:BAP

[RSJ00] Lluís Rovira, Pau Senra, and David Jou. Bibliometric analysis of physics in Catalonia: Towards quality consolidation? Scientometrics, 49(2):233–256, October 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010517102426.

Ruane:2008:RSI

[RT08] Frances Ruane and Richard S. J. Tol. Rational (successive) h-indices: An application to economics in the Republic of Ireland. Scientometrics, 75(2):395–405, May 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1869-7.

Ruane:2009:HMQ

[RT09] Frances P. Ruane and Richard S. J. Tol. A Hirsch measure for the quality of research supervision, and an illustration with trade economists. Scientometrics, 80(3):613– 624, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-2092-x.

Riikonen:2008:NRC

[RV08] Pentti Riikonen and Mauno Vihinen. National research contributions: A case study on Finnish biomedical research. Scientometrics, 77(2):207–222, November 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1962-y.

Rinia:2001:CDI

[RVB^{+}01] Ed J. Rinia, Thed. N. Van Leeuwen, Eppo E. W. Bruins, Hendrik G. Van Vuren, and Anthony F. J. Van Raan. Citation delay in interdisciplinary knowledge exchange. Scientometrics, 51(1):293–309, April 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010589300829.

Ruiz:2005:CSA

[RvHMJ05] Juan A. Araujo Ruiz, Guido van Hooydonk, Raul G. Torricella Morales, and Ricardo Arencibia Jorge. Cuban scientific articles in ISI Citation Indexes and CubaCiencias databases (1988–2003). Scientometrics, 65(2):161– 171, November 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0265-4.

Rinia:2002:MKT

[RvLB^{+}02] Ed J. Rinia, Thed N. van Leeuwen, Eppo E. W. Bruins, Hendrik G. van Vuren, and Anthony F. J. van Raan. Measuring knowledge transfer between fields of science. Scientometrics, 54(3):347–362, July 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016078331752.

Rinia:2002:IMI

[RvLvR02] Ed J. Rinia, Thed N. van Leeuwen, and Anthony F. J. van Raan. Impact measures of interdisciplinary research in physics. Scientometrics, 53(2):241–248, February 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014856625623.

Robert:2009:AMB

[RWD^{+}09] Claude Robert, Concepción S. Wilson, Stéphane Donnadieu, Jean-François Gaudy, and Charles-Daniel Arreto. Analysis of the medical and biological pain research literature in the European Union: A 2006 snapshot. Scientometrics, 80(3):693–716, September 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2102-z.

Robert:2006:SEP

[RWGA06] Claude Robert, Concepción S. Wilson, Jean-François Gaudy, and Charles-Daniel Arreto. A snapshot of EU publications in sleep research: A scientometric survey. Scientometrics, 67(3):385–405, June 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1556/Scient.
67.2006.3.4.

Robert:2007:ESS

[RWGA07] Claude Robert, Concepción S. Wilson, Jean-François Gaudy, and Charles-Daniel Arreto. The evolution of the sleep science literature over 30 years: A bibliometric analysis. Scientometrics, 73(2):231–256, November 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1780-2.

Rousseau:2008:BCM

[RZ08] Ronald Rousseau and Lin Zhang. Betweenness centrality and Q-measures in directed valued networks. Scientometrics, 75(3):575–590, June 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-1772-2.

Seglen:2000:SPG

[SA00] Per O. Seglen and Dag W. Aksnes. Scientific productivity and group size: A bibliometric analysis of Norwegian microbiological research. Scientometrics, 49(1):125–143, August 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005665309719.

Senthilkumaran:2007:MSR

[SA07] P. Senthilkumaran and A. Amudhavalli. Mapping of spices research in Asian countries. Scientometrics, 73(2):149– 159, November 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-1738-9.

Saad:2006:EIA

[Saa06] Gad Saad. Exploring the h-index at the author and journal levels using bibliometric data of productive consumer scholars and business-related journals respectively. Scientometrics, 69(1):117–120, October 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0142-9.

Saetnan:2000:SSS

[Sæt00] Ann Rudinow Sætnan. To screen or not to screen? Science discourse in two health policy controversies, as seen through three approaches to the citation evidence. Scientometrics, 48(3):307–344, July 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005636420708.

Sangam:2000:PNK

[San00] S. L. Sangam. P. S. Nagpaul, K. C. Garg, B. M. Gupta, editors emerging trends in scientometrics: Essays in honour of Dr. Ashok Jain. Scientometrics, 47(1):165–166, January 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1005630128673.

Sandstrom:2001:SCSa

[San01a] Pamela E. Sandstrom. Scholarly communication as a socioecological system. Scientometrics, 50(3):573–605, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1012786609066.

Sandstrom:2001:SCSbv

[San01b] Pamela E. Sandstrom. Scholarly communication as a socioecological system. Scientometrics, 51(3):573–605, July 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1019655305286.

Sandstrom:2009:RQD

[San09] Ulf Sandström. Research quality and diversity of funding: A model for relating research money to output of research. Scientometrics, 79(2):341–349, May 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0422-2.

Sapa:2007:ICL

[Sap07] Remigiusz Sapa. International contribution to library and information science in Poland: A bibliometric analysis. Scientometrics, 71(3):473–493, June 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1675-2.

Sonderstrup-Andersen:2008:IDR

[SASA08] Eva M. Sønderstrup-Andersen and Hans H. K. Sønderstrup-Andersen. An investigation into diabetes researcher’s perceptions of the Journal Impact Factor — reconsidering evaluating research. Scientometrics, 76(2):391–406, August 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1924-4.

Sternitzke:2009:SMD

[SB09] Christian Sternitzke and Isumo Bergmann. Similarity measures for document mapping: A comparative study on the level of an individual scientist. Scientometrics, 78(1): 113–130, January 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1961-z.

Suarez-Balseiro:2006:PIS

[SBSCOR06] Carlos Suárez-Balseiro, Elias Sanz-Casado, and Laurie Ortiz-Rivera. Patterns of international scientific cooperation in Puerto Rico. Scientometrics, 67(3):335–350, June 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1556/Scient.67.2006.3.1.

Schubert:2000:SMR

[Sch00] András Schubert. Scientometrics in medicine-related fields 1990–1999. Scientometrics, 48(2):251–284, September 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1005600922922.

Schubert:2001:BCB

[Sch01] András Schubert. Scientometrics: A citation based bibliography 1997–2000. Scientometrics, 50(1):99–198, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005646319815.

Schubert:2002:WS

[Sch02] András Schubert. The Web of scientometrics. Scientometrics, 53(1):3–20, January 2002. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1014886202425.

Schummer:2004:MIP

[Sch04] Joachim Schummer. Multidisciplinarity, interdisciplinarity, and patterns of research collaboration in nanoscience and nanotechnology. Scientometrics, 59(3):425–465, March 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000018542.71314.38.

Schneider:2006:CSR

[Sch06] Jesper W. Schneider. Concept symbols revisited: Naming clusters by parsing and filtering of noun phrases from citation contexts of concept symbols. Scientometrics, 68(3): 573–593, September 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0131-z.

Schubert:2007:SI

[Sch07a] András Schubert. Successive h-indices. Scientometrics, 70 (1):201–205, January 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-0112-x.

Schummer:2007:GIN

[Sch07b] Joachim Schummer. The global institutionalization of nanotechnology research: A bibliometric approach to the assessment of science policy. Scientometrics, 70(3):669–692, March 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-0307-1.

Schreiber:2008:ISC

[Sch08] Michael Schreiber. The influence of self-citation corrections on Egghe’s g index. Scientometrics, 76(1):187–200, July 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1886-6.

Schubert:2009:UIA

[Sch09] András Schubert. Using the h-index for assessing single publications. Scientometrics, 78(3):559–565, March 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2208-3.

Shirabe:2000:IEM

[SF00] Masashi Shirabe and Yuko Fujigaki. The introduction of economic methods to scientometrics: The citing–cited table and the autopoietic systems of citations. Scientometrics, 47 (1):117–130, January 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005621910926.

Shelton:2009:DNS

[SFG09] Robert D. Shelton, Patricia Foland, and Roman Gorelskyy. Do new SCI journals have a different national bias? Scientometrics, 79(2):351–363, May 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0423-1.

Sierra-Flores:2009:PMA

[SFGRP09] María Magdalena Sierra-Flores, María V. Guzmán, Alejandro C. Raga, and I. Pérez. The productivity of Mexican astronomers in the field of outflows from young stars. Scientometrics, 81(3):765–777, December 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2264-8.

Schoepflin:2001:TDB

[SG01a] Urs Schoepflin and Wolfgang Glänzel. Two decades of “Scientometrics”. An interdisciplinary field represented by its leading journal. Scientometrics, 50(2):301–312, February 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1010577824449.

Stegmann:2001:CRK

[SG01b] Johannes Stegmann and Guenter Grohmann. Citation rates, knowledge export and international visibility of dermatology journals listed and not listed in the Journal Citation Reports. Scientometrics, 50(3):483–502, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1010562815788.

Stegmann:2003:HGG

[SG03] Johannes Stegmann and Guenter Grohmann. Hypothesis generation guided by co-word clustering. Scientometrics, 56 (1):111–135, January 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1021954808804.

Schubert:2006:CNP

[SG06] András Schubert and Wolfgang Glänzel. Cross-national preference in co-authorship, references and citations. Scientometrics, 69(2):409–428, November 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0160-7.

Spiegelhalter:2009:CCS

[SG09] David Spiegelhalter and Harvey Goldstein. Comment: Citation statistics. Statistical Science, 24(1):21–24, February 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.
ss/1255009005. See [Hal09].

Schloegl:2003:ETA

[SGBB03] Christian Schloegl, Juan Gorraiz, Christoph Bart, and Monika Bargmann. Evaluating two Austrian university departments: Lessons learned. Scientometrics, 56(3):289–299, March 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1022318618200.

Schubert:2006:WAS

[SGT06a] András Schubert, Wolfgang Glänzel, and Bart Thijs. The weight of author self-citations. A fractional approach to self-citation counting. Scientometrics, 67(3):503–514, June 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1556/Scient.67.2006.3.11.

Suzuki:2006:APP

[SGT^{+}06b] Jun Suzuki, Kiminori Gemba, Schumpeter Tamada, Yoshihito Yasaki, and Akira Goto. Analysis of propensity to patent and science-dependence of large Japanese manufacturers of electrical machinery. Scientometrics, 68(2): 265–288, August 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0111-3.

Shelton:2004:UER

[SH04] Robert D. Shelton and Geoffrey M. Holdridge. The US–EU race for leadership of science and technology: Qualitative and quantitative indicators. Scientometrics, 60(3):353–363, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000034379.75236.d9.

Sandstrom:2008:PNP

[SH08a] Ulf Sandström and Martin Hällsten. Persistent nepotism in peer-review. Scientometrics, 74(2):175–189, February 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-008-0211-3. See correction [SH21].

Sotudeh:2008:GER

[SH08b] Hajar Sotudeh and Abbas Horri. Great expectations: The role of open access in improving countries’ recognition. Scientometrics, 76(1):69–93, July 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1890-x.

Sotudeh:2009:CPO

[SH09] Hajar Sotudeh and Abbas Horri. Countries positioning in open access journals system: An investigation of citation distribution patterns. Scientometrics, 81(1):7–31, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-1870-4.

Sandstrom:2021:CPN

[SH21] Ulf Sandström and Martin Hällsten. Correction to: Persistent nepotism in peer-review. Scientometrics, 126(2):1863–1865, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03739-4;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03739-4.pdf. See [SH08a].

Shapiro:2001:UG

[Sha01] Samuel I. Shapiro. The universe grasper. Scientometrics, 52 (2):337–344, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017976009321.

Shelton:2008:RBN

[She08] Robert D. Shelton. Relations between national research investment and publication output: Application to an American paradox. Scientometrics, 74(2):191–205, February 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-008-0212-2.

Shama:2000:COA

[SHO00] Gilbert Shama, Klaus Hellgardt, and Charles Oppenheim. Citation ootprint analysis. Part I: UK and US chemical engineering academics. Scientometrics, 49(2):289–305, October 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1010573220173.

Siddiqi:2006:ALS

[Sid06] Ahmed F. Siddiqi. Age likes some years. Scientometrics, 69(2):315–321, November 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
006-0155-4.

Sigogneau:2000:ADT

[Sig00] Anne Sigogneau. An analysis of document types published in journals related to physics: Proceeding papers recorded in the Science Citation Index database. Scientometrics, 47 (3):589–604, March 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005628218890.

Sambunjak:2008:RJF

[SIMM08] Dario Sambunjak, Ana Ivanis, Ana Marusić, and Matko Marusić. Representation of journals from five neighboring European countries in the Journal Citation Reports. Scientometrics, 76(2):261–271, August 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1915-5.

Shan:2004:MWD

[SJJ04] Shi Shan, Guohua Jiang, and Lan Jiang. The multivariate Waring distribution and its application. Scientometrics, 60 (3):523–535, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000034392.

65855.5c. See comment [Xek05].

Sutter:2001:PLR

[SK01a] Matthias Sutter and Martin G. Kochner. Power laws of research output. Evidence for journals of economics. Scientometrics, 51(2):405–414, June 2001. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%

3A1012757802706. See comments [Rou02].

Szydlowski:2001:SCM

[SK01b] Marek Szydlowski and Adam Krawiec. Scientific cycle model with delay. Scientometrics, 52(1):83–95, September 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1012751028630.

Szydlowski:2009:GCK

[SK09] Marek Szydlowski and Adam Krawiec. Growth cycles of knowledge. Scientometrics, 78(1):99–111, January 2009.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-1958-7.

Small:2008:SPS

[SKB08] Henry Small, Ann Kushmerick, and Doug Benson. Scientists’ perceptions of the social and political implications of their research. Scientometrics, 74(2):207–221, February 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-0213-1.

Skilton:2006:CSC

[Ski06] Paul F. Skilton. A comparative study of communal practice: Assessing the effects of taken-for-granted-ness on citation practice in scientific communities. Scientometrics, 68(1):73– 96, July 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0084-2.

Skilton:2009:DHC

[Ski09] Paul F. Skilton. Does the human capital of teams of natural science authors predict citation frequency? Scientometrics, 78(3):525–542, March 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1953-z.

Sidiropoulos:2007:GHI

[SKM07] Antonis Sidiropoulos, Dimitrios Katsaros, and Yannis Manolopoulos. Generalized Hirsch h-index for disclosing latent facts in citation networks. Scientometrics, 72(2): 253–280, August 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1722-z.

Sombatsompop:2006:SEC

[SKMI06] Narongrit Sombatsompop, Apisit Kositchaiyong, Teerasak Markpin, and Sekson Inrit. Scientific evaluations of citation quality of international research articles in the SCI database: Thailand case study. Scientometrics, 66(3): 521–535, March 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0038-8.

Schubert:2009:HTI

[SKT09] András Schubert, András Korn, and András Telcs. Hirschtype indices for characterizing networks. Scientometrics, 78 (2):375–382, February 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-2218-1.

Kim:2004:EPI

[sKtP04] Moon soo Kim and Yong tae Park. The evolving patterns of inter-industrial knowledge structure: Case of Korean manufacturing in the 1980s. Scientometrics, 61(1):43–54, September 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000037361.19404.a5.

Sternberg:2005:PCO

[SL05] Rolf Sternberg and Timo Litzenberger. The publication and citation output of German faculties of economics and social sciences — a comparison of faculties and disciplines based upon SSCI data. Scientometrics, 65(1):29–53, October 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0258-3.

Salaun:2000:DSA

[SLB00] Jean-Michel Salaün, Thierry Lafouge, and Cherifa Boukacem. Demand for scientific articles and citations: An example from the Institut de l’information scientifique et technique (France). Scientometrics, 47(3):561–588, March 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005676002052.

Schneider:2009:CSF

[SLI09] Jesper W. Schneider, Birger Larsen, and Peter Ingwersen. A comparative study of first and all-author co-citation counting, and two different matrix generation approaches applied for author co-citation analyses. Scientometrics, 80(1):103– 130, July 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-2019-y.

Shin:2006:BMP

[SLP06] Juneseuk Shin, Wook Lee, and Yongtae Park. On the benchmarking method of patent-based knowledge flow structure: Comparison of Korea and Taiwan with USA. Scientometrics, 69(3):551–574, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0169-y.

Solari:2000:NAS

[SM00] Aline Solari and Marie-Helene Magri. A new approach to the SCI Journal Citation Reports, a system for evaluating scientific journals. Scientometrics, 47(3):605–625, March 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005680202961.

Schildt:2006:DNS

[SM06] Henri A. Schildt and Juha T. Mattsson. A dense network sub-grouping algorithm for co-citation analysis and its implementation in the software tool sitkis. Scientometrics, 67 (1):143–163, April 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0054-8.

Small:2001:BHa

[Sma01a] Henry Small. Belver and Henry. Scientometrics, 50(3): 487–491, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1012722222270.

Small:2001:BHb

[Sma01b] Henry Small. Belver and Henry. Scientometrics, 51(3):489– 497, July 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1019690918490.

Small:2004:SRM

[Sma04a] Henry Small. On the shoulders of Robert Merton: Towards a normative theory of citation. Scientometrics, 60(1):71– 79, May 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000027310.68393.bc.

Small:2004:WAT

[Sma04b] Henry Small. Why authors think their papers are highly cited. Scientometrics, 60(3):305–316, August 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000034376.55800.18.

Small:2006:TPG

[Sma06] Henry Small. Tracking and predicting growth areas in science. Scientometrics, 68(3):595–610, September 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0132-y.

Sombatsompop:2007:CTM

[SMB^{+}07] N. Sombatsompop, T. Markpin, T. Buranathiti, P. Ratchatahirun, T. Metheenukul, N. Premkamolnetr, and W. Yochai. Categorization and trend of materials science research from Science Citation Index (SCI) database: A case study of ceramics, metallurgy, and polymer subfields. Scientometrics, 71(2):283–302, May 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1667-2.

Smith:2008:BGS

[Smi08] Alastair G. Smith. Benchmarking Google Scholar with the New Zealand PBRF research assessment exercise. Scientometrics, 74(2):309–316, February 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-0219-8.

Sombatsompop:2004:MMC

[SMP04] Narongrit Sombatsompop, T. Markpin, and N. Premkamolnetr. A modified method for calculating the Impact Factors of journals in ISI Journal Citation Reports: Polymer science category in 1997–2001. Scientometrics, 60(2):217–235, June 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000027794.98854.f6.

Sombatsompop:2009:RRP

[SMW^{+}09] Narongrit Sombatsompop, T. Markpin, E. Wimolmala, P. Ratchatahirun, N. Premkamolnetr, B. Boonradsamee, and W. Yochai. Relationship on research publications and productivity-export volumes for natural rubber. Scientometrics, 81(2):393–405, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-2165-0.

Sombatsompop:2005:ERP

[SMYS05] N. Sombatsompop, T. Markpin, W. Yochai, and M. Saechiew. An evaluation of research performance for different subject categories using Impact Factor Point Average (IFPA) index: Thailand case study. Scientometrics, 65(3):293–305, December 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0275-2.

Soler:2006:RIS

[Sol06] José M. Soler. A rational indicator of scientific creativity. arxiv.org, August 1, 2006. URL http://adsabs.harvard.

edu/abs/2006physics...8006S;   http://arxiv.org/
abs/physics/0608006.

Soler:2007:SAA

[Sol07] José M. Soler. Separating the articles of authors with the same name. Scientometrics, 72(2):281–290, August 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-1730-z.

Sommer:2005:BAP

[Som05] Simon Sommer. Bibliometric analysis and private research funding. Scientometrics, 62(1):165–171, January 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0011-y.

Song:2003:IKI

[Son03] Choong-Han Song. Interdisciplinarity and knowledge inflow/outflow structure among science and engineering research in Korea. Scientometrics, 58(1):129–141, September 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1025483625299.

Sooryamoorthy:2009:CPH

[Soo09a] Radhamany Sooryamoorthy. Collaboration and publication: How collaborative are scientists in South Africa? Scientometrics, 80(2):419–439, August 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2074-z.

Sooryamoorthy:2009:DTC

[Soo09b] Radhamany Sooryamoorthy. Do types of collaboration change citation? Collaboration and citation patterns of South African science publications. Scientometrics, 81(1): 177–193, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-2126-z.

Seol:2008:KSI

[SP08] Sung-Soo Seol and Jung-Min Park. Knowledge sources of innovation studies in Korea: A citation analysis. Scientometrics, 75(1):3–20, April 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1826-5.

Sandstrom:2000:KPK

[SPN00] Anna Sandström, Ingrid Pettersson, and Anna Nilsson. Knowledge production and knowledge flows in the Swedish biotechnology innovation system. Scientometrics, 48(2): 179–201, September 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005644705175.

Simkin:2005:SMC

[SR05] M. V. Simkin and V. P. Roychowdhury. Stochastic modeling of citation slips. Scientometrics, 62(3):367–384, March 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-005-0028-2.

Sombatsompop:2002:CRT

[SRS^{+}02] N. Sombatsompop, P. Ratchatahirun, V. Surathanasakul, N. Premkamolnetr, and T. Markpin. A citation report for Thai academic journals published during 1996– 2000. Scientometrics, 55(3):445–462, November 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1020423031578.

Somogyi:2005:SCC

[SS05] Anikó Somogyi and András Schubert. Short communication: Correlation between national bibliometric and health indicators: The case of diabetes. Scientometrics, 62(2): 285–292, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0021-9.

Schmoch:2008:ICP

[SS08] Ulrich Schmoch and Torben Schubert. Are international co-publications an indicator for quality of scientific research? Scientometrics, 74(3):361–377, March 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1818-5.

Schmoch:2009:SIE

[SS09a] Ulrich Schmoch and Torben Schubert. Sustainability of incentives for excellent research — the German case. Scientometrics, 81(1):195–218, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-2127-y.

Schubert:2009:OYN

[SS09b] András Schubert and Mihály Schubert. Outperform your neighbors. Scientometrics, 80(2):553–558, August 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-008-2209-2.

Sussmuth:2006:TEE

[SSG06] Bernd Süssmuth, Martin Steininger, and Stephane Ghio. Towards a European economics of economics: Monitoring a decade of top research and providing some explanation. Scientometrics, 66(3):579–612, March 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0042-z.

Sangam:2007:CCR

[SSM07] S. L. Sangam, Kiran Savanur, and M. Manjunath. Communication and collaborative research pattern of Sivaraj Ramaseshan: A scientometric portrait. Scientometrics, 71 (2):217–230, May 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1670-7.

Sangam:2006:SPP

[SSMV06] S. L. Sangam, Kiran Savanur, M. Manjunath, and R. Vasudevan. Scientometric portrait of Prof. Peter John Wyllie. Scientometrics, 66(1):43–53, January 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0004-5.

Shirabe:2002:LOA

[ST02a] Masashi Shirabe and Hiroyuki Tomizawa. Likelihood of overseas access to international co-authorships. Scientometrics, 53(1):123–129, January 2002. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/A%
3A1014840121989.

Smith:2002:WIF

[ST02b] Alastair Smith and Mike Thelwall. Web Impact Factors for Australasian universities. Scientometrics, 54(3):363–380, July 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1016030415822.

Shirabe:2004:LIO

[ST04] Masashi Shirabe and Hiroyuki Tomizawa. Likelihood of inbound/outbound access to co-authorship. Scientometrics, 59(3):337–344, March 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000018538.
73884.db.

Sharma:2008:ICR

[ST08] Seema Sharma and V. J. Thomas. Inter-country R&D efficiency analysis: An application of data envelopment analysis. Scientometrics, 76(3):483–501, September 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1896-4.

Stefaniak:2001:ICO

[Ste01] Barbara Stefaniak. International co-operation in science and in social sciences as reflected in multinational papers indexed in SCI and SSCI. Scientometrics, 52(2):193–210, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1017959605687.

Stephan:2004:RKM

[Ste04] Paula E. Stephan. Robert K. Merton’s perspective on priority and the provision of the public good knowledge. Scientometrics, 60(1):81–87, May 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000027311.17226.70.

Sternitzke:2009:DTP

[Ste09a] Christian Sternitzke. Defining triadic patent families as a measure of technological strength. Scientometrics, 81(1): 91–109, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-1836-6.

Sternitzke:2009:IPE

[Ste09b] Christian Sternitzke. The international preliminary examination of patent applications filed under the patent cooperation treaty- a proxy for patent value? Scientometrics, 78 (2):189–202, February 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1837-x.

Sternitzke:2009:PPS

[Ste09c] Christian Sternitzke. Patents and publications as sources of novel and inventive knowledge. Scientometrics, 79(3):551– 561, June 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-2041-0.

Sternitzke:2009:TSP

[Ste09d] Christian Sternitzke. Technological specialization and patenting strategies in East Asia — insights from the electronics industry. Scientometrics, 78(1):69–76, January 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-1988-1.

Stigler:2004:RKM

[Sti04] Stephen M. Stigler. Robert K. Merton: Memorial. Scientometrics, 60(1):89–92, May 2004. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/article/10.1023/B%
3ASCIE.0000027312.39478.b8.

Small:2009:CSE

[SU09] Henry Small and Phineas Upham. Citation structure of an emerging research area on the verge of application. Scientometrics, 79(2):365–375, May 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0424-0.

Salzarulo:2001:BSQ

[SV01] Laurent Salzarulo and Markus Von Ins. Bias, structure and quality in citation indexing. Scientometrics, 50(2):289–299, February 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1010525807611.

Schmoch:2009:O

[SV09] Ulrich Schmoch and Anthony F. J. Van Raan. Obituary. Scientometrics, 80(2):303–304, August 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-009-0201-0.

Schwechheimer:2001:MIR

[SW01] Holger Schwechheimer and Matthias Winterhager. Mapping interdisciplinary research fronts in neuroscience: A bibliometric view to retrograde amnesia. Scientometrics, 51(1):311–318, April 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010541418576.

Skalska-Zlat:2001:NPW

[SZ01] Marta Skalska-Zlat. Nalimov and the Polish way towards science of science. Scientometrics, 52(2):211–223, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1017911722525.

Thelwall:2002:EUA

[TBH^{+}02] Mike Thelwall, Ray Binns, Gareth Harries, Teresa Page-Kennedy, Liz Price, and David Wilkinson. European Union associated university websites. Scientometrics, 53 (1):95–111, January 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014836021080.

Thelwall:2006:WLG

[TBK06] Mike Thelwall, Franz Barjak, and Hildrun Kretschmer. Web links and gender in science: An exploratory analysis. Scientometrics, 67(3):373–383, June 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1556/Scient.67.2006.3.3.

Tussen:2000:TRS

[TBvL00] R. J. W. Tussen, R. K. Buter, and Th. N. van Leeuwen. Technological relevance of science: An assessment of citation linkages between patents and research papers. Scientometrics, 47(2):389–412, February 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005603513439.

Trueba:2004:RFC

[TG04] Frank J. Trueba and Héctor Guerrero. A robust formula to credit authors for their publications. Scientometrics, 60 (2):181–204, June 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000027792.
09362.3f.

Thijs:2006:IAS

[TG06] Bart Thijs and Wolfgang Glänzel. The influence of author self-citations on bibliometric meso-indicators. The case of European universities. Scientometrics, 66(1):71–80, January 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0006-3.

Thijs:2008:SAP

[TG08] Bart Thijs and Wolfgang Glänzel. A structural analysis of publication profiles for the classification of European research institutes. Scientometrics, 74(2):223–236, February 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-0214-0.

Thijs:2009:SAB

[TG09] Bart Thijs and Wolfgang Glänzela. A structural analysis of benchmarks on different bibliometrical indicators for European research institutes based on their research profile. Scientometrics, 79(2):377–388, May 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0425-z.

Turner:2001:IMM

[TGJd01] William A. Turner, R. Gherbi, C. Jacquemin, and M. de Saint Leger. Inometric methods and measures for sharing knowledge over internet. Scientometrics, 50(1):33–57, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005638017998.

Thorsteinsdottir:2000:ERC

[Tho00] O. Halla Thorsteinsdóttir. External research collaboration in two small science systems. Scientometrics, 49(1):145–160, August 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005617426557.

Thomas:2001:RBT

[Tho01] Patrick Thomas. A relationship between technology indicators and stock market performance. Scientometrics, 51 (1):319–333, April 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010597502646.

Tonta:2002:CHU

[TI02] YaSar Tonta and Mustafa Ilhan. Contribution of Hacettepe University Faculty of Medicine to the world’s biomedical literature (1988–1997). Scientometrics, 55(1):123– 136, September 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1016007205344.

Tijssen:2007:ACW

[Tij07] Robert J. W. Tijssen. Africa’s contribution to the worldwide research literature: New analytical perspectives, trends, and performance indicators. Scientometrics, 71(2):303–327, May 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-007-1658-3.pdf.

Tsay:2000:BSS

[TJM00] Ming-Yueh Tsay, Shiow-Jen Jou, and Sheau-Shin Ma. A bibliometric study of semiconductor literature, 1978– 1997. Scientometrics, 49(3):491–509, November 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010593824716.

Takeda:2009:OBA

[TK09] Yoshiyuki Takeda and Yuya Kajikawa. Optics: a bibliometric approach to detect emerging research domains and intellectual bases. Scientometrics, 78(3):543–558, March 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-2012-5.

Tseng:2009:CMD

[TLL^{+}09] Yuen-Hsien Tseng, Yu-I Lin, Yi-Yang Lee, Wen-Chi Hung, and Chun-Hsiang Lee. A comparison of methods for detecting hot topics. Scientometrics, 81(1):73–90, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-009-1885-x.

Lin:2007:EPS

[tLtC07] Chin tsai Lin and Chang tzu Chiang. Evaluating the performance of sponsored Chinese herbal medicine research. Scientometrics, 70(1):67–84, January 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0105-9.

Tsay:2003:NRB

[TM03] Ming-Yueh Tsay and Shiao-Shing Ma. The nature and relationship between the productivity of journals and their citations in semiconductor literature. Scientometrics, 56(2): 201–222, January 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1021915127459.

Takeda:2009:NER

[TMKM09] Yoshiyuki Takeda, Shiho Mae, Yuya Kajikawa, and Katsumori Matsushima. Nanobiotechnology as an emerging research domain from nanotechnology: A bibliometric approach. Scientometrics, 80(1):23–38, July 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1897-3.

Torricella-Morales:2000:CAC

[TMVAR00] Raúl G. Torricella-Morales, Guido Van Hooydonk, and Juan Antonio Araujo-Ruiz. Citation analysis of Cuban research. Part 1. A case study: The Cuban Journal of Agricultural Science. Scientometrics, 47(2):413–426, February 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005655530277.

Tamada:2006:SDD

[TNK^{+}06] Schumpeter Tamada, Yusuke Naito, Fumio Kodama, Kiminori Gemba, and Jun Suzuki. Significant difference of dependence upon scientific knowledge among different technologies. Scientometrics, 68(2):289–302, August 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0112-2.

Tol:2009:IAA

[Tol09] Richard S. J. Tol. The h-index and its alternatives: An application to the 100 most prolific economists. Scientometrics, 80(2):317–324, August 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
008-2079-7.

Tonta:2000:CTR

[Ton00] Yasar Tonta. Contribution of Turkish researchers to the World’s biomedical literature (1988–1997). Scientometrics, 48(1):71–84, June 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005632402450.

Trimble:2000:SCY

[Tri00] Virginia Trimble. Some characteristics of Young vs. Established American astronomers: Entering the new century. Scientometrics, 48(3):403–411, July 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005644622525.

Thelwall:2002:IBA

[TS02] Mike Thelwall and Alastair Smith. Interlinking between Asia–Pacific university Web sites. Scientometrics, 55(3): 363–376, November 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1020410728852.

Tsay:2008:BAH

[Tsa08] Ming-Yueh Tsay. A bibliometric analysis of hydrogen energy literature, 1965–2005. Scientometrics, 75(3):421–438, June 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1785-x.

Tsay:2009:ACS

[Tsa09a] Ming-Yueh Tsay. An analysis and comparison of scientometric data between journals of physics, chemistry and engineering. Scientometrics, 78(2):279–293, February 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1996-1.

Tsay:2009:CAT

[Tsa09b] Ming-Yueh Tsay. Citation analysis of Ted Nelson’s works and his influence on hypertext concept. Scientometrics, 79 (3):451–472, June 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-1641-7.

Torres-Salinas:2009:RDR

[TSLCJC09] Daniel Torres-Salinas, Emilio Delgado Lopez-Cózar, and Evaristo Jiménez-Contreras. Ranking of departments and researchers within a university using two different databases: Web of Science versus Scopus. Scientometrics, 80(3):761–774, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2113-9.

Thelwall:2003:DLC

[TT03] Mike Thelwall and Rong Tang. Disciplinary and linguistic considerations for academic Web linking: An exploratory hyperlink mediated study with mainland China and Taiwan. Scientometrics, 58(1):155–181, September 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1025487826207.

Tang:2004:PNI

[TT04] Rong Tang and Mike Thelwall. Patterns of national and international Web inlinks to US academic departments: An analysis of disciplinary variations. Scientometrics, 60(3): 475–485, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000034388.
70594.cc.

Thelwall:2003:LPA

[TTP03] Mike Thelwall, Rong Tang, and Liz Price. Linguistic patterns of academic Web use in Western Europe. Scientometrics, 56(3):417–432, March 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1022387105904.

Tuzi:2005:SSI

[Tuz05] Fabrizio Tuzi. The scientific specialisation of the Italian regions. Scientometrics, 62(1):87–111, January 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0005-9.

Tijssen:2006:MIA

[TV06] Robert J. W. Tijssen and Thed N. Van Leeuwen. Measuring impacts of academic science on industrial research: A citation-based approach. Scientometrics, 66(1):55–69, January 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0005-4.

Tijssen:2002:BIS

[TVvL02] Robert J. W. Tijssen, Martijn S. Visser, and Thed N. van Leeuwen. Benchmarking international scientific excellence: Are highly cited research papers an appropriate frame of reference? Scientometrics, 54(3):381–397, July 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016082432660.

Thelwall:2008:UCE

[TZ08] Mike Thelwall and Alesia Zuccala. A university-centred European Union link analysis. Scientometrics, 75(3):407–420, June 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1831-8.

Ugolini:2003:VIJ

[UC03] Donatella Ugolini and Cristina Casilli. The visibility of Italian journals. Scientometrics, 56(3):345–355, March 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1022326820017.

Ugolini:2001:HEU

[UCCM01] Donatella Ugolini, Marco Amedeo Cimmino, Cristina Casilli, and Giuseppe Sandro Mela. How the European Union writes about ophthalmology. Scientometrics, 52(1): 45–58, September 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1012746927721.

Upadhye:2004:SAS

[UKKP04] Rekha P. Upadhye, V. L. Kalyane, Vijai Kumar, and E. R. Prakasan. Scientometric analysis of synchronous references in the Physics Nobel lectures, 1981–1985: A pilot study. Scientometrics, 61(1):55–68, September 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000037362.11986.42.

Uzun:2002:PRI

[Uzu02] Ali Uzun. Productivity ratings of institutions based on publication in Scientometrics, Informetrics, and Bibliometrics, 1981–2000. Scientometrics, 53(3):297–307, March 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1014864827441.

Uzun:2004:AIS

[Uzu04] Ali Uzun. Assessing internationality of scholarly journals through foreign authorship patterns: the case of major journals in information science, and scientometrics. Scientometrics, 61(3):457–465, November 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000045121.26810.35.

Uzun:2006:STP

[Uzu06] Ali Uzun. Science and technology policy in Turkey. National strategies for innovation and change during the 1983– 2003 period and beyond. Scientometrics, 66(3):551–559, March 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0040-1.

VanDenBesselaar:2001:CSS

[Van01a] Peter Van Den Besselaar. The cognitive and the social structure of STS. Scientometrics, 51(2):441–460, June 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1012714020453.

VanRaan:2001:CAS

[Van01b] Anthony F. J. Van Raan. Competition amongst scientists for publication status:toward a model of scientific publication and citation distributions. Scientometrics, 51(1): 347–357, April 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010501820393.

Vanecek:2008:BAC

[Van08a] Jiri Vanecek. Bibliometric analysis of the Czech research publications from 1994 to 2005. Scientometrics, 77(2):345– 360, November 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1986-3.

Vanecek:2008:PPC

[Van08b] Jirí Vanecek. Patenting propensity in the Czech Republic. Scientometrics, 75(2):381–394, May 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1874-x.

Vanclay:2009:BJI

[Van09] Jerome K. Vanclay. Bias in the journal impact factor. Scientometrics, 78(1):3–12, January 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-1778-4.

Vaughan:2004:EWF

[Vau04] Liwen Vaughan. Exploring website features for business information. Scientometrics, 61(3):467–477, November 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/B%3ASCIE.0000045122.93018.2a.

Villarroya:2008:PTS

[VBBF08] Anna Villarroya, Maite Barrios, Angel Borrego, and Amparo Frías. PhD theses in Spain: A gender study covering the years 1990–2004. Scientometrics, 77(3):469–483, December 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1965-8.

VanBorm:2000:PMB

[VCP00] Julien Van Borm, Jan Corthouts, and Richard Philips. Performance measurement in the Belgian document ordering and delivery system impala. Scientometrics, 47(2):207–225, February 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005630824825.

vanCampenhout:2008:COS

[vCvCvU08] Geert van Campenhout, Tom van Caneghem, and Steve van Uytbergen. A comparison of overall and sub-area journal influence: The case of the accounting literature. Scientometrics, 77(1):61–90, October 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1907-5.

Verbeek:2006:PER

[VD06] Arnold Verbeek and Koenraad Debackere. Patent evolution in relation to public/private R&D investment and corporate profitability: Evidence from the United States. Scientometrics, 66(2):279–294, February 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0021-4.

vandenBesselaar:2000:CBS

[vdB00] Peter van den Besselaar. Communication between science and technology studies journals: A case study in differentiation and integration in scientific fields. Scientometrics, 47 (2):169–193, February 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005686123917.

vandenBesselaar:2006:MRT

[vdBH06] Peter van den Besselaar and Gaston Heimeriks. Mapping research topics using word-reference co-occurrences: A method and an exploratory case study. Scientometrics, 68(3):377–393, September 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
006-0118-9.

VanLooy:2006:SCT

[VDC^{+}06] Bart Van Looy, Koenraad Debackere, Julie Callaert, Robert Tijssen, and Thed van Leeuwen. Scientific capabilities and technological performance of national innovation systems: An exploration of emerging industrial relevant research domains. Scientometrics, 66(2):295–310, February 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0030-3.

vanDalen:2005:SSI

[vDH05] Hendrik P. van Dalen and Kène Henkens. Signals in science — on the importance of signaling in gaining attention in science. Scientometrics, 64(2):209–233, August 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0248-5.

Verbeek:2002:LST

[VDL^{+}02] Arnold Verbeek, Koenraad Debackere, Marc Luwel, Petra Andries, Edwin Zimmermann, and Filip Deleus. Linking science to technology: Using bibliographic references in patents to build linkage schemes. Scientometrics, 54(3):399– 420, July 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1016034516731.

Verbeek:2003:SCP

[VDL03] Arnold Verbeek, Koenraad Debackere, and Marc Luwel. Science cited in patents: A geographic “flow” analysis of bibliographic citation patterns in patents. Scientometrics, 58 (2):241–263, October 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026232526034.

vanderPanne:2007:IMI

[vdP07] Gerben van der Panne. Issues in measuring innovation. Scientometrics, 71(3):495–507, June 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-007-1691-2.pdf.

vanderWal:2009:BNB

[vdWFM^{+}09] René van der Wal, Anke Fischer, Mick Marquiss, Steve Redpath, and Sarah Wanless. Is bigger necessarily better for environmental research? Scientometrics, 78(2):317–322, February 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-2017-0.

Verspagen:2000:RLM

[Ver00] Bart Verspagen. The role of large multinationals in the Dutch technology infrastructure. A patent citation analysis. Scientometrics, 47(2):427–448, February 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005607614347.

Vaughan:2006:P

[VG06] Liwen Vaughan and Wolfgang Glänzel. Preface. Scientometrics, 67(2):261–262, May 2006. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-006-0097-x.

Vieira:2009:CSW

[VG09] Elizabeth S. Vieira and José A. N. F. Gomes. A comparison of Scopus and Web of Science for a typical university. Scientometrics, 81(2):587–600, November 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-2178-0.

Vaughan:2006:WHB

[VGK06] Liwen Vaughan, Yijun Gao, and Margaret Kipp. Why are hyperlinks to business Websites created? A content analysis. Scientometrics, 67(2):291–300, May 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0100-6.

VanDalen:2001:WMS

[VH01] Hendrik P. Van Dalen and Kène Henkens. What makes a scientific article influential? The case of demographers. Scientometrics, 50(3):455–482, January 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010510831718.

Vinkler:2001:ADS

[Vin01] Péter Vinkler. An attempt for defining some basic categories of scientometrics and classifying the indicators of evaluative scientometrics. Scientometrics, 50(3):539–544, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1010519000767.

Vinkler:2002:DCC

[Vin02a] Péter Vinkler. Dynamic changes in the chance for citedness. Scientometrics, 54(3):421–434, July 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1016086500801.

Vinkler:2002:SPA

[Vin02b] Péter Vinkler. Subfield problems in applying the Garfield (Impact) Factors in practice. Scientometrics, 53(2):267–279, February 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1014860726532.

Vinkler:2003:RRS

[Vin03] Péter Vinkler. Relations of relative scientometric indicators. Scientometrics, 58(3):687–694, November 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000006888.69146.24.

Vinkler:2006:CSI

[Vin06] Péter Vinkler. Composite scientometric indicators for evaluating publications of research institutes. Scientometrics, 68(3):629–642, September 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
006-0123-z.

Vinkler:2008:CBS

[Vin08] Peter Vinkler. Correlation between the structure of scientific research, scientometric indicators and GDP in EU and non-EU countries. Scientometrics, 74(2):237–254, February 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-0215-z.

Vinkler:2009:ICC

[Vin09] Péter Vinkler. Introducing the current contribution index for characterizing the recent, relevant impact of journals. Scientometrics, 79(2):409–420, May 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0427-x.

Vaughan:2007:WWC

[VKG07] Liwen Vaughan, Margaret E. I. Kipp, and Yijun Gao. Why are Websites co-linked? The case of Canadian universities. Scientometrics, 72(1):81–92, July 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1707-y.

Valkimadi:2009:IDE

[VKVF09] Politimi Eleni Valkimadi, Drosos E. Karageorgopoulos, Harissios Vliagoftis, and Matthew E. Falagas. Increasing dominance of English in publications archived by PubMed. Scientometrics, 81(1):219–223, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-008-2139-z.

Vilanova:2001:WCC

[VL01] Marta Riba Vilanova and Loet Leydesdorff. Why Catalonia cannot be considered as a regional innovation system. Scientometrics, 50(2):215–240, February 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010517505793.

vanLeeuwen:2006:ABA

[vL06] Thed van Leeuwen. The application of bibliometric analyses in the evaluation of social science research. Who benefits from it, and why it is still feasible. Scientometrics, 66(1): 133–154, January 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0010-7.

vanLeeuwen:2009:SWN

[vL09] Thed N. van Leeuwen. Strength and weakness of national science systems: A bibliometric analysis through cooperation patterns. Scientometrics, 79(2):389–408, May 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-009-0426-y.

Velloso:2004:CSB

[VLdM04] Andréa Velloso, Denise Lannes, and Leopoldo de Meis. Concentration of science in Brazilian governmental universities. Scientometrics, 61(2):207–220, October 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000041649.24713.ca.

vanLeeuwen:2002:DAJ

[vLM02] Thed N. van Leeuwen and Henk F. Moed. Development and application of journal impact measures in the Dutch science system. Scientometrics, 53(2):249–266, February 2002. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1014808709694.

vanLeeuwen:2005:CJI

[vLM05] Thed N. van Leeuwen and Henk F. Moed. Characteristics of journal impact factors: The effects of uncitedness and citation distribution on the understanding of journal impact factors. Scientometrics, 63(2):357–371, April 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0217-z.

Ventura:2006:UBI

[VM06] Oscar N. Ventura and Alvaro W. Mombrú. Use of bibliometric information to assist research policy making. A comparison of publication and citation profiles of Full and Associate Professors at a School of Chemistry in Uruguay. Scientometrics, 69(2):287–313, November 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-006-0154-5.

VanLooy:2007:DTV

[VMD07] Bart Van Looy, Tom Magerman, and Koenraad Debackere. Developing technology in the vicinity of science: An examination of the relationship between science intensity (of patents) and technological productivity within the field of biotechnology. Scientometrics, 70(2):441–458, February 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-0211-8.

VanLeeuwen:2001:LBC

[VMT^{+}01] Thed N. Van Leeuwen, Henk F. Moed, Robert J. W. Tijssen, Martijn S. Visser, and Anthony F. J. Van Raan. Language biases in the coverage of the Science Citation Index and its consequencesfor international comparisons of national research performance. Scientometrics, 51(1):335–346, April 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1010549719484.

delaPotterie:2008:BHS

[vPdlPvZ08] Bruno van Pottelsberghe de la Potterie and Nicolas van Zeebroeck. A brief history of space and time: The scope-year index as a patent value indicator based on families and renewals. Scientometrics, 75(2):319–338, May 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1864-z.

vanRaan:2000:GAF

[vR00] Anthony F. J. van Raan. On growth, ageing, and fractal differentiation of science. Scientometrics, 47(2):347–362, February 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005647328460.

vanRaan:2001:BIS

[vR01] Anthony F. J. van Raan. Bibliometrics and internet: Some observations and expectations. Scientometrics, 50(1):59–63, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005690102069.

vanRaan:2004:SBS

[vR04] Anthony F. J. van Raan. Sleeping beauties in science. Scientometrics, 59(3):467–472, March 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000018543.82441.f1.

vanRaan:2005:FAC

[vR05a] Anthony F. J. van Raan. Fatal attraction: Conceptual and methodological problems in the ranking of universities by bibliometric methods. Scientometrics, 62(1):133–143, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0008-6.

vanRaan:2005:RBP

[vR05b] Anthony F. J. van Raan. Reference-based publication networks with episodic memories. Scientometrics, 63(3):549– 566, June 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0227-x.

vanRaan:2005:RCL

[vR05c] Anthony F. J. van Raan. Reply to the comments of Liu et al. Scientometrics, 64(1):111–112, July 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-005-0242-y.

vanRaan:2006:CHI

[vR06] Anthony F. J. van Raan. Comparison of the Hirsch-index with standard bibliometric indicators and with peer judgment for 147 chemistry research groups. Scientometrics, 67 (3):491–502, June 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1556/Scient.67.2006.3.10.

Vaughan:2008:NLE

[VS08] Liwen Vaughan and Debora Shaw. A new look at evidence of scholarly citation in citation indexes and from web sources. Scientometrics, 74(2):317–330, February 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-0220-2.

Vasconcelos:2009:NII

[VSL09] Sonia M. R. Vasconcelos, Martha M. Sorenson, and Jacqueline Leta. A new input indicator for the assessment of science & technology research? Scientometrics, 80(1):217–230, July 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2082-z.

vonUngern-Sternberg:2000:BLC

[vUS00] Sara von Ungern-Sternberg. Bradford’s Law in the context of information provision. Scientometrics, 49(1):161–186, August 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005669410627.

VanLeeuwen:2003:HGS

[VVM^{+}03] Thed N. Van Leeuwen, Martijn S. Visser, Henk F. Moed, Ton J. Nederhof, and Anthony F. J. Van Raan. The Holy Grail of science policy: Exploring and combining bibliometric tools in search of scientific excellence. Scientometrics, 57(2):257–280, June 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1024141819302.

Valles-Valenzuela:2009:QAS

[VVPCOL09] Javier Valles-Valenzuela, María D. Pérez-Cárceles, Eduardo Osuna, and Aurelio Luna. Quantitative analysis of Spanish university scientific output in the area of legal and forensic medicine: International exposure. Scientometrics, 78 (3):383–395, March 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1808-2.

Vaughan:2004:LCW

[VW04] Liwen Vaughan and Guozhu Wu. Links to commercial websites as a source of business information. Scientometrics, 60 (3):487–496, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000034389.
14825.bc.

VonElm:2009:RCS

[VWJ09] Erik Von Elm, Simon Wandel, and Peter Jüni. The role of correspondence sections in post-publication peer review: A bibliometric study of general and internal medicine journals. Scientometrics, 81(3):747–755, December 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-2236-0.

Vaughan:2006:CBC

[VY06] Liwen Vaughan and Justin You. Comparing business competition positions based on Web co-link data: The global market vs. the Chinese market. Scientometrics, 68(3):611– 628, September 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0133-x.

Vaughan:2008:CAW

[VY08] Liwen Vaughan and Justin You. Content assisted web co-link analysis for competitive intelligence. Scientometrics, 77(3):433–444, December 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-1999-y.

VanLooy:2003:DST

[VZV^{+}03] Bart Van Looy, Edwin Zimmermann, Reinhilde Veugelers, Arnold Verbeek, Johanna Mello, and Koenraad Debackere. Do science-technology interactions pay off when developing technology? Scientometrics, 57(3):355–367, July 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1025052617678.

vanZeebroeck:2006:IMD

[vZvPdlPH06] Nicolas van Zeebroeck, Bruno van Pottelsberghe de la Potterie, and Wook Han. Issues in measuring the degree of technological specialisation with patent data. Scientometrics, 66 (3):481–492, March 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0035-y.

Wagner:2005:SCS

[Wag05] Caroline S. Wagner. Six case studies of international collaboration in science. Scientometrics, 62(1):3–26, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0001-0.

Walters:2006:PSC

[Wal06] Glenn D. Walters. Predicting subsequent citations to articles published in twelve crime-psychology journals: Author impact versus journal impact. Scientometrics, 69(3): 499–510, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0166-1.

Wang:2007:FEP

[Wan07] Shyh-Jen Wang. Factors to evaluate a patent in addition to citations. Scientometrics, 71(3):509–522, June 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1698-8.

Wolfgang:2004:BAR

[WBB04] Glänzel Wolfgang, Thijs Bart, and Schlemmer Balázs. A bibliometric approach to the role of author self-citations in scientific communication. Scientometrics, 59(1):63–77, January 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000013299.38210.74.

Wagner-Dobler:2001:CDC

[WD01a] Roland Wagner-Döbler. Continuity and discontinuity of collaboration behaviour since 1800 — from a bibliometric point of view. Scientometrics, 52(3):503–517, November 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1014208219788.

Wagner-Dobler:2001:RPD

[WD01b] Roland Wagner-Döbler. Rescher’s principle of decreasing marginal returns of scientific research. Scientometrics, 50 (3):419–436, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010506730809.

Wagner-Dobler:2005:SRD

[WD05] Roland Wagner-Döbler. The system of research and development indicators: Entry points for information agents. Scientometrics, 62(1):145–153, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0009-5.

Weingart:2005:IBU

[Wei05] Peter Weingart. Impact of bibliometrics upon the science system: Inadvertent consequences? Scientometrics, 62(1): 117–131, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0007-7.

Wenzel:2001:CSN

[Wen01] Vera Wenzel. Complex systems in natural science and humanities. Scientometrics, 52(3):525–529, November 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1014216520697.

Wetterer:2006:QEC

[Wet06] James K. Wetterer. Quotation error, citation copying, and ant extinctions in Madeira. Scientometrics, 67(3):351–372, June 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1556/Scient.67.2006.3.2.

Wallner:2003:APP

[WFI03] Bernard Wallner, Martin Fieder, and Karoline Iber. Age profile, personnel costs and scientific productivity at the University of Vienna. Scientometrics, 58(1):143– 153, September 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025435709369.

Wong:2007:KSI

[WH07] Poh-Kam Wong and Yuen-Ping Ho. Knowledge sources of innovation in a small open economy: The case of Singapore. Scientometrics, 70(2):223–249, February 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0201-x.

Wong:2007:IEA

[WHC07] Poh Kam Wong, Yuen Ping Ho, and Casey K. Chan. Internationalization and evolution of application areas of an emerging technology: The case of nanotechnology. Scientometrics, 70(3):715–737, March 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-0309-z.

White:2001:ACBa

[Whi01a] Howard D. White. Author-centered bibliometrics through CAMEOs: Characterizations automatically made and edited online. Scientometrics, 50(3):607–637, January 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1012738725904.

White:2001:ACBb

[Whi01b] Howard D. White. Author-centered bibliometrics through CAMEOs: Characterizations automatically made and edited online. Scientometrics, 51(3):607–637, July 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1023/A%3A1019607522125.

White:2004:RPS

[Whi04] Howard D. White. Reward, persuasion, and the Sokal Hoax: A study in citation identities. Scientometrics, 60(1):93–120, May 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000027313.91401.9b.

White:2008:KMR

[Whi08] Howard D. White. Katherine McCain: Recipient of the 2007 Derek de Solla Price Award of the journal Scientometrics. Scientometrics, 74(1):7–10, January 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-0100-9.

Wagner:2003:SDD

[WL03] Caroline S. Wagner and Loet Leydesdorff. Seismology as a dynamic, distributed area of scientific research. Scientometrics, 58(1):91–114, September 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1025479524390.

Wilson:2004:CSO

[WM04] Concepción S. Wilson and Valentina A. Markusova. Changes in the scientific output of Russia from 1980 to 2000, as reflected in the Science Citation Index, in relation to national politico-economic changes. Scientometrics, 59 (3):345–389, March 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000018539.
93567.d8.

Wohlin:2009:NIC

[Woh09] Claes Wohlin. A new index for the citation curve of researchers. Scientometrics, 81(2):521–533, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-008-2155-z.

Wolfram:2006:ASI

[Wol06] Dietmar Wolfram. Applications of SQL for informetric frequency distribution processing. Scientometrics, 67(2):301– 313, May 2006. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-006-0101-5.

Wormell:2000:BAW

[Wor00a] Irene Wormell. Bibliometric analysis of the welfare topic. Scientometrics, 48(2):203–236, September 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005696722014.

Wormell:2000:CAD

[Wor00b] Irene Wormell. Critical aspects of the Danish welfare state — as revealed by issue tracking. Scientometrics, 48(2): 237–250, September 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005648806084.

Wormell:2000:PNW

[Wor00c] Irene Wormell. Proceedings of the 4th Nordic Workshop in Bibliometrics Copenhagen (Denmark), August 27–28, 1999. Scientometrics, 48(2):117–120, September 2000. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1005688520197.

Wu:2004:CST

[WPZ^{+}04] Yishan Wu, Yuntao Pan, Yuhua Zhang, Zheng Ma, Jingan Pang, Hong Guo, Bo Xu, and Zhiqing Yang. China scientific and technical papers and citations (CSTPC): History, impact and outlook. Scientometrics, 60(3):385–397, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000034381.64865.2b.

Wray:2004:ECY

[Wra04] K. Brad Wray. An examination of the contributions of young scientists in new fields. Scientometrics, 61(1):117– 128, September 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000037367.
99073.bb.

Wray:2009:DPA

[Wra09a] K. Brad Wray. Did professionalization afford better opportunities for young scientists? Scientometrics, 81(3): 757–764, December 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-2254-x.

Wray:2009:SIR

[Wra09b] K. Brad Wray. The salaries of Italian Renaissance professors. Scientometrics, 80(2):351–357, August 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2073-0.

Wu:2009:ESA

[Wu09] Zhiqiang Wu. An empirical study of the accessibility of web references in two Chinese academic journals. Scientometrics, 78(3):481–503, March 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1951-1.

Wooding:2006:CAI

[WWJLG06] Steven Wooding, Kate Wilcox-Jay, Grant Lewison, and Jonathan Grant. Co-author inclusion: A novel recursive algorithmic method for dealing with homonyms in bibliometric analysis. Scientometrics, 66(1):11–21, January 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-006-0002-7.

Wang:2005:SCC

[WWP^{+}05] Yan Wang, Yishan Wu, Yuntao Pan, Zheng Ma, and Ronald Rousseau. Scientific collaboration in China as reflected in co-authorship. Scientometrics, 62(2):183–198, January 2005. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-005-0013-9.

Wang:2007:BAE

[WWW07] Shuhua Wang, Hengjun Wang, and Paul R. Weldon. Bibliometric analysis of English-language academic journals of China and their internationalization. Scientometrics, 73(3): 331–343, December 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1775-z.

Wainer:2009:SPC

[WXB09] Jacques Wainer, Eduardo C. Xavier, and Fabio Bezerra. Scientific production in computer science: A comparative study of Brazil and other countries. Scientometrics, 81(2): 535–547, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-2156-y.

Xu:2003:NOC

[XCS03] Wei Xu, Yi-Zhang Chen, and Zhi-Chao Shen. Neuroscience output of China: A MEDLINE-based bibliometric study. Scientometrics, 57(3):399–409, July 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1025060819495.

Xekalaki:2005:LEC

[Xek05] Evdokia Xekalaki. Letter to the Editor; comments on the paper of Shan et al.: The multivariate Waring distribution. Scientometrics, 62(2):293–296, January 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0022-8. See [SJJ04].

Xie:2008:AWA

[XZH08] Shaodong Xie, Jing Zhang, and Yuh-Shan Ho. Assessment of world aerosol research trends by bibliometric analysis. Scientometrics, 77(1):113–130, October 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1928-0.

Yaman:2007:PTT

[YA07] Hakan Yaman and E. Atay. PhD theses in Turkish sports sciences: A study covering the years 1988– 2002. Scientometrics, 71(3):415–421, June 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1679-y.

Yegros:2009:EDF

[YA09] Alfredo Yegros Yegros and Carlos B. Amat. Editorial delay of food research papers is influenced by authors’ experience but not by country of origin of the manuscripts. Scientometrics, 81(2):367–380, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-2164-y.

Yi:2008:UCP

[YAH08] Huang Yi, Xiaolan Ao, and Yuh-Shan Ho. Use of citation per publication as an indicator to evaluate pentachlorophenol research. Scientometrics, 75(1):67–80, April 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1849-y.

Yang:2009:TCC

[Yan09] Hua Yang. The top 40 citation classics in the Journal of the American Society for Information Science and Technology. Scientometrics, 78(3):421–426, March 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-2011-6.

Ynalvez:2005:WDS

[YDM^{+}05] Marcus Ynalvez, Ricardo B. Duque, Paul Mbatia, R. Sooryamoorthy, Antony Palackal, and Wesley Shrum. When do scientists “adopt” the Internet? Dimensions of connectivity in developing areas. Scientometrics, 63(1):39–67, March 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-005-0203-5.

Ye:2007:QRB

[Ye07] Fred Y. Ye. A quantitative relationship between per capita GDP and scientometric criteria. Scientometrics, 71(3):407– 413, June 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1678-z.

Ye:2009:IMM

[Ye09] Fred Y. Ye. An investigation on mathematical models of the h-index. Scientometrics, 81(2):493–498, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-008-2169-6.

Yglesias:2003:PVP

[Ygl03] Elmer Yglesias. Porter vs. Porter: Modeling the technological competitiveness of nations. Scientometrics, 57(2):281– 293, June 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1024193803372.

Yu:2006:IPDb

[YGL06] Guang Yu, Rui Guo, and Yi-Jun Li. The influence of publication delays on three ISI indicators. Scientometrics, 69(3): 511–527, December 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0167-0.

Yu:2006:IPDa

[YGY06] Guang Yu, Rui Guo, and Da-Ren Yu. The influence of the publication delay on journal rankings according to the impact factor. Scientometrics, 67(2):201–211, May 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0094-0.

Yalpani:2005:ASM

[YHM05] Mohamed Yalpani, Akbar Heydari, and Morteza Mehrdad. Application of scientometric methods to chemical research in Iran: Reflections on Iran’s current science policy. Scientometrics, 63(3):531–547, June 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0226-y.

Yitzhaki:2002:RTL

[Yit02] Moshe Yitzhaki. Relation of the title length of a journal article to the length of the article. Scientometrics, 54(3):435– 447, July 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1016038617639.

Yoshikane:2004:CAC

[YK04] Fuyuki Yoshikane and Kyo Kageura. Comparative analysis of coauthorship networks of different domains: The growth and change of networks. Scientometrics, 60(3):433–444, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/B%3ASCIE.0000034385.05897.46.

Yu:2007:PIO

[YL07] Guang Yu and Yi-Jun Li. Parameter identification of the observed citation distribution. Scientometrics, 71(2):339– 348, May 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1662-7.

Yoo:2006:SPM

[YM06] Seung-Hoon Yoo and Hye-Seon Moon. A semi-parametric modeling of firms’ R&D expenditures with zero values. Scientometrics, 69(1):57–67, October 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0138-5.

Yoo:2008:NAD

[YM08] Seung-Hoon Yoo and Hye-Seon Moon. A note on approximation of distribution function for the number of innovation activities. Scientometrics, 77(3):445–452, December 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-1802-8.

Yang:2009:MIT

[YMB09] Liying Yang, Steven A. Morris, and Elizabeth M. Barden. Mapping institutions and their weak ties in a specialty: A case study of cystic fibrosis body composition research. Scientometrics, 79(2):421–434, May 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0428-9.

Yoshikane:2009:ACB

[YNSS09] Fuyuki Yoshikane, Takayuki Nozawa, Susumu Shibui, and Takafumi Suzuki. An analysis of the connection between researchers’ productivity and their co-authors’ past attributions, including the importance in collaboration networks. Scientometrics, 79(2):435–449, May 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-008-0429-8.

Yoshikane:2006:CAC

[YNT06] Fuyuki Yoshikane, Takayuki Nozawa, and Keita Tsuji. Comparative analysis of co-authorship networks considering authors’ roles in collaboration: Differences between the theoretical and application areas. Scientometrics, 68(3):643– 655, September 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-0113-1.

Yamashita:2006:PSC

[YO06] Yasuhiro Yamashita and Yoshiko Okubo. Patterns of scientific collaboration between Japan and France: Intersectoral analysis using probabilistic partnership index (PPI). Scientometrics, 68(2):303–324, August 2006. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-006-0105-1.

Yong:2001:LCN

[YR01] Fang Yong and Ronald Rousseau. Lattices in citation networks: An investigation into the structure of citation graphs. Scientometrics, 50(2):273–287, February 2001. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/A%3A1010573723540.

Yu:2003:MMD

[YRL03] Guang Yu, Yihong Rong, and Yijun Li. Mathematical model of delay in the secondary literature publishing process. Scientometrics, 58(3):673–685, November 2003. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000006887.31509.7a.

Yuh-Shan:2004:CRL

[YS04] Ho Yuh-Shan. Citation review of Lagergren kinetic rate equation on adsorption reactions. Scientometrics, 59(1): 171–177, January 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000013305.
99473.cf.

Ynalvez:2008:IGT

[YS08] Marcus Antonius Ynalvez and Wesley Shrum. International graduate training, digital inequality and professional network structure: An ego-centric social network analysis of knowledge producers at the “global south”. Scientometrics, 76(2):343–368, August 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1936-0.

Yanagisawa:2009:SEE

[YT09] Kazuaki Yanagisawa and Shoji Takahashi. Socio-economic effects of the material science in JAERI. Scientometrics, 78 (3):505–524, March 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-1636-4.

Tsay:2005:JGI

[yTlC05] Ming yueh Tsay and Yi ling Chen. Journals of general & internal medicine and surgery: An analysis and comparison of citation. Scientometrics, 64(1):17–30, July 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0235-x.

Tsay:2003:ACC

[yTXwW03a] Ming yueh Tsay, Hong Xu, and Chia wen Wu. Author co-citation analysis of semiconductor literature. Scientometrics, 58(3):529–545, November 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000006878.83104.61.

Tsay:2003:JCC

[yTXwW03b] Ming yueh Tsay, Hong Xu, and Chia wen Wu. Journal cocitation analysis of semiconductor literature. Scientometrics, 57(1):7–25, May 2003. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023667318934.

Yue:2004:MCI

[YW04] Weiping Yue and Concepción S. Wilson. Measuring the citation impact of research journals in clinical neurology: A structural equation modelling analysis. Scientometrics, 60 (3):317–332, August 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000034377.
93437.18.

Yu:2007:SCR

[YW07] Guang Yu and Liang Wang. The self-cited rate of scientific journals and the manipulation of their impact factors. Scientometrics, 73(3):321–330, December 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1779-8.

Yu:2005:IPD

[YWY05] Guang Yu, Xiao-Hong Wang, and Da-Ren Yu. The influence of publication delays on impact factors. Scientometrics, 64 (2):235–246, August 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-005-0249-4.

Yi:2008:TDR

[YX08] Huang Yi and Zhao Xi. Trends of DDT research during the period of 1991 to 2005. Scientometrics, 75(1):111–122, April 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1828-3.

Yoon:2008:CAB

[YY08a] Yoon Hee Yoon and Kim Sin Young. Correlation analysis between university research competitiveness and library’s scholarly information in OECD nations and Korea. Scientometrics, 74(3):345–360, March 2008. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1753-5.

Yu:2008:DSP

[YY08b] Guang Yu and Da-Ren Yu. Design and simulation on the publication delay control system. Scientometrics, 76(3): 407–427, September 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-1748-2.

Yu:2004:UEP

[YYL04] Guang Yu, Daren Yu, and Yijun Li. The universal expression of periodical average publication delay at steady state. Scientometrics, 60(2):121–129, June 2004. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1023/B%3ASCIE.0000027675.69901.5b.

Yu:2005:SSP

[YYL05] Guang Yu, Daren Yu, and Yijun Li. A simulation study of the periodicals’ publication delay control process. Scientometrics, 63(1):25–38, March 2005. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-005-0202-6.

Yan:2008:HAG

[YZ08] Erjia Yan and Qinghua Zhu. Hyperlink analysis for government websites of Chinese provincial capitals. Scientometrics, 76(2):315–326, August 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-1923-5.

Zanotto:2002:STD

[Zan02] Edgar Dutra Zanotto. Scientific and technological development in Brazil. The widening gap. Scientometrics, 55(3): 411–419, November 2002. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1020466813831.

Zanotto:2006:SP

[Zan06] Edgar Dutra Zanotto. The scientists pyramid. Scientometrics, 69(1):175–181, October 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
006-0134-9.

Zitt:2000:SPI

[ZBO00] Michel Zitt, Elise Bassecoulard, and Yoshiko Okubo. Shadows of the past in international cooperation: Collaboration profiles of the top five producers of science. Scientometrics, 47(3):627–657, March 2000. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1005632319799.

Zhao:2009:ERC

[ZCY09] Yueyang Zhao, Lei Cui, and Hua Yang. Evaluating reliability of co-citation clustering analysis in representing the research history of subject. Scientometrics, 80(1):91–102, July 2009. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-2056-1.

Zimmerman:2009:SCB

[ZGBI09] Eric Zimmerman, Wolfgang Glänzel, and Judit Bar-Ilan. Scholarly collaboration between Europe and Israel: A scientometric examination of a changing landscape. Scientometrics, 78(3):427–446, March 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
007-2044-x.

Zhou:2007:SAG

[ZGHW07] Feng Zhou, Huai-Cheng Guo, Yuh-Shan Ho, and Chao-Zhong Wu. Scientometric analysis of geostatistics using multivariate methods. Scientometrics, 73(3):265–279, December 2007. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-007-1798-5.

Zhang:2009:TRI

[ZGL09] Lin Zhang, Wolfgang Glänzel, and Liming Liang. Tracing the role of individual journals in a cross-citation network based on different indicators. Scientometrics, 81(3): 821–838, December 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-2245-y.

Zabala-Iturriagagoitia:2007:WID

[ZIJSCMGG07] Jon Mikel Zabala-Iturriagagoitia, Fernando Jiménez-Sáez, Elena Castro-Martínez, and Antonio Gutiérrez-Gracia. What indicators do (or do not) tell us about regional innovation systems. Scientometrics, 70(1):85–106, January 2007. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-007-0106-8.

Zhivotovsky:2008:SCC

[ZK08] Lev A. Zhivotovsky and Konstantin V. Krutovsky. Selfcitation can inflate h-index. Scientometrics, 77(2):373– 375, November 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-006-1716-2.

Zelman:2000:TEM

[ZL00] Andrés Zelman and Loet Leydesdorff. Threaded email messages in self-organization and science & technology studies oriented mailing lists. Scientometrics, 48(3):361–380, July 2000. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1005640521616.

Zhao:2002:CAU

[ZL02] Dangzhi Zhao and Elisabeth Logan. Citation analysis using scientific publications on the Web as data source: A case study in the XML research area. Scientometrics, 54(3):449– 472, July 2002. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1016090601710.

Zhou:2007:CIC

[ZL07] Ping Zhou and Loet Leydesdorff. The citation impacts and citation environments of Chinese journals in mathematics. Scientometrics, 72(2):185–200, August 2007. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1713-0.

Zbikowska-Migon:2001:KHF

[ZM01] Anna Zbikowska-Migoń. Karl Heinrich Frömmichen (1736– 1783) and Adrian Balbi (1782–1848) — the pioneers of biblio- and scientometrics. Scientometrics, 52(2):225–233, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1017963706595.

Zorin:2001:FAP

[ZNK01] Nikita A. Zorin, Alexander V. Nemtsov, and Vladimir V. Kalinin. Formalised assessment of publication quality in Russian psychiatry. Scientometrics, 52(2):315–322, October 2001. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1017971908413.

Zhang:2009:BAR

[ZQH09] Weiwei Zhang, Weihong Qian, and Yuh-Shan Ho. A bibliometric analysis of research related to ocean circulation. Scientometrics, 80(2):305–316, August 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-007-1863-0.

Zitt:2003:BCR

[ZRRB03a] Michel Zitt, Suzy Ramanana-Rahary, and Elise Bassecoulard. Bridging citation and reference distributions: Part I — the referencing-structure function and its application to co-citation and co-item studies. Scientometrics, 57(1):93– 118, May 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1023675620752.

Zitt:2003:CGH

[ZRRB03b] Michel Zitt, Suzy Ramanana-Rahary, and Elise Bassecoulard. Correcting glasses help fair comparisons in international science landscape: Country indicators as a function of ISI database delineation. Scientometrics, 56(2):259–282, January 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1021923329277.

Zitt:2005:RCP

[ZRRB05] Michel Zitt, Suzy Ramanana-Rahary, and Elise Bassecoulard. Relativity of citation performance and excellence measures: From cross-field to cross-scale effects of fieldnormalisation. Scientometrics, 63(2):373–401, April 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-005-0218-y.

Zitt:2003:PST

[ZRRBL03] Michel Zitt, Suzy Ramanana-Rahary, Elise Bassecoulard, and Françoise Laville. Potential science-technology spillovers in regions: An insight on geographic co-location of knowledge activities in the EU. Scientometrics, 57(2):295–320, June 2003. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1024145920210.

Zsindely:2008:VFS

[Zsi08] Sándor Zsindely. From vanity fair to scientific research: The place of genealogy in contemporary science. A scientometric approach. Scientometrics, 77(1):197–206, October 2008. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-008-1010-6.

Zhou:2009:CBG

[ZTG09a] Ping Zhou, Bart Thijs, and Wolfgang Glänzel. Is China also becoming a giant in social sciences? Scientometrics, 79(3):593–621, June 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-007-2068-x.

Zhou:2009:RAC

[ZTG09b] Ping Zhou, Bart Thijs, and Wolfgang Glänzel. Regional analysis on Chinese scientific output. Scientometrics, 81(3): 839–857, December 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-008-2255-9.

Zuccala:2009:MRN

[ZV09] Alesia Zuccala and Peter Van Den Besselaar. Mapping review networks: Exploring research community roles and contributions. Scientometrics, 81(1):111–122, October 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-008-2136-2.

Zou:2009:OAO

[ZWW09] Feng Zou, Mingxing Wu, and Kaili Wu. Outcomes associated with ophthalmology, optometry and visual science literature in the Science Citation Index from mainland China, 2000–2007. Scientometrics, 81(3):671–682, December 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-009-2191-3.

Zhu:2004:HCR

[ZWZM04] Xing Zhu, Qi Wu, Yingzi Zheng, and Xin Ma. Highly cited research papers and the evaluation of a research university: A case study: Peking University 1974–2003. Scientometrics, 60(2):237–247, June 2004. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1023/B%3ASCIE.0000027795.
69665.09.